### v1

# Accelerated Convergence of Frank–Wolfe Algorithms with Adaptive Bregman Step-Size Strategy

Abstract
We propose a Frank–Wolfe (FW) algorithm with an adaptive Bregman step-size strategy for smooth adaptable (also called: relatively smooth) (weakly-) convex functions. This means that the gradient of the objective function is not necessarily Lipschitz continuous, and we only require the smooth adaptable property. Compared to existing FW algorithms, our assumptions are less restrictive. We establish convergence guarantees in various settings, such as sublinear to linear convergence rates, depending on the assumptions for convex and nonconvex objective functions. Assuming that the objective function is weakly convex and satisfies the local quadratic growth condition, we provide both local sublinear and local linear convergence regarding the primal gap. We also propose a variant of the away-step FW algorithm using Bregman distances. We establish global accelerated (up to linear) convergence for convex optimization under the Hölder error bound condition and its local linear convergence for nonconvex optimization under the local quadratic growth condition over polytopes. Numerical experiments on nonnegative linear inverse problems, $\ell_{p}$ loss problems, phase retrieval, low-rank minimization, and nonnegative matrix factorization demonstrate that our proposed FW algorithms outperform existing methods.
1 Introduction
In this paper, we consider constrained optimization problems of the form
$\min_{x\in P}\quad f(x),$ (1.1)
where $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is a continuously differentiable function and $P\subset\mathbb{R}^{n}$ is a compact convex set. We are interested in both convex and nonconvex $f$ and assume that we have first-order oracle access to $f$, i.e., given $x\in\mathbb{R}^{n}$, we can compute $\nabla f(x)$. Usually, Problem (1.1) is solved by a variant of the projected gradient method (see, for example, [Nesterov2018-gj, Nocedal2006-sb]). However, even if $P$ is convex, it is not always possible to access the projection operator of $P$ or compute the projection efficiently. While interior point methods are a potential alternative approach to constrained optimization problems, their updates often require second-order information, which can be computationally expensive for large-scale problems. To alleviate the situation, Frank and Wolfe [Frank1956-cq] proposed the Frank–Wolfe (FW) algorithm (also known as the conditional gradient method), which is a projection-free first-order method. Instead of requiring access to a projection oracle, the FW algorithm requires only access to a so-called linear minimization oracle (LMO), which for a given linear function $a\in\mathbb{R}^{n}$ computes $y\in\operatorname*{argmin}_{v\in P}\langle a,v\rangle$. Linear minimization oracles are often much cheaper than projection oracles, as shown in [CP2021] (see also summary in [Braun2022-ew, Table 1.1]), so that, in practice, FW algorithms are often faster than projected gradient methods when the projection operation is not trivial. Additionally, FW algorithms tend to be numerically quite robust and stable due to the affine-invariance of the algorithm and can be also used, e.g., to provide theoretical guarantees for the approximate Carathéodory problem [cp2019approxCara].
The original FW algorithm is given in Algorithm 1.
Input: Initial point $x_{0}\in P$, objective function $f$, step-size strategy $\gamma_{t}\in[0,1]$
1
2for $t=0,\ldots$ do
3      $v_{t}\leftarrow\operatorname*{argmin}_{v\in P}\langle\nabla f(x_{t}),v\rangle$
4     $x_{t+1}\leftarrow(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$
Algorithm 1 Frank–Wolfe algorithm
Related Work
In this section, we briefly review the work most closely related to this paper. The FW algorithm was originally proposed by Frank and Wolfe [Frank1956-cq] and was independently rediscovered and extended by Levitin and Polyak [Levitin1966-sm] as the conditional gradient method; we will use these terms interchangeably. Canon and Cullum [Canon1968-sr] established an initial lower bound for the convergence rate of the FW algorithm, which was later improved by GuéLat and Marcotte  [GueLat1986-bg], which also provided the first analysis of the away-step FW algorithm by Wolfe [Wolfe1970-ax]. Jaggi [Jaggi2013-oz] provided a more detailed convergence analysis of the FW algorithm, establishing a new lower bound that demonstrated a trade-off between sparsity and error. Concurrently, Lan [lan2013complexity] examined the complexity of linear programming-based first-order methods, establishing a similar complexity lower bound.
The classical FW algorithm is known to zigzag when approaching the optimal face containing the optimal solution $x^{*}$. This behavior inspired Wolfe [Wolfe1970-ax] to propose the away-step FW algorithm in 1970 that shortcuts the zigzagging by removing atoms that slow down the iterate sequence. Lacoste-Julien and Jaggi [Lacoste-Julien2015-gb] showed the linear convergence of the away-step FW algorithm. Their analysis introduced a geometric constant, the pyramidal width, that measures the conditioning of the polytope $P$, representing the feasible region. Similar analyses based on this constant have been developed for other advanced variants, such as the pairwise FW algorithm [Lacoste-Julien2015-gb], Wolfe’s algorithm [Lacoste-Julien2015-gb], the blended conditional gradient method [BPTW2018], and blended pairwise conditional gradient methods and similar variants [CP2020boost, TTP2021]. Even Nesterov-style acceleration is possible under weak assumptions [CDP2019] building upon this analysis.
Pedregosa et al. [Pedregosa2020-ay] introduced a powerful adaptive step-size strategy that dynamically approximates the local Lipschitz continuity of the function $f$. They proved that this strategy is at least as effective as the traditional short-step strategy, and it was later enhanced by Pokutta [Pokutta2024-qu] to improve numerical stability. For nonconvex functions, Lacoste-Julien [Lacoste-Julien2016-mj] established sublinear convergence, and Maskan et al. [Maskan2025-oa] incorporated the difference of convex functions (DC) into FW algorithms. For a comprehensive review of FW algorithms, we refer the interested reader to the survey by Braun et al. [Braun2022-ew] and the brief introduction by Pokutta [Pokutta2024-qu].
Establishing the convergence rate of FW algorithms typically requires that $\nabla f$ is Lipschitz continuous, although exceptions exist, such as in the case of generalized self-concordant functions [CBP2021, SelfConcordantFW_2020]. Some first-order methods using the Bregman distance do not require the global Lipschitz continuity of $\nabla f$. Bolte et al. [Bolte2018-zt] proposed the Bregman proximal gradient algorithm and established its global convergence under the smooth adaptable property (see Definition 2.6). Hanzely et al. [Hanzely2021-sz] introduced accelerated Bregman proximal gradient algorithms. Takahashi et al. [Takahashi2022-ml] developed a Bregman proximal gradient algorithm for the DC optimization and its accelerated version. Subproblems of Bregman proximal gradient algorithms are often not solvable in closed form. Takahashi and Takeda [Takahashi2024-wz] used approximations to Bregman distances to solve subproblems in closed form. Rebegoldi et al. [Rebegoldi2018-pa] proposed an inexact version of the Bregman proximal gradient algorithm.
Contribution
Our contributions can be roughly summarized as follows:
Frank–Wolfe for Smooth Adaptable functions
We study Frank–Wolfe algorithms for smooth adaptable functions (also called: relatively smooth) functions. Some functions whose gradients are not Lipschitz continuous satisfy the smooth adaptable property with kernel generating distances (see Definition 2.1). The class of smooth adaptable functions appears in many applications, such as nonnegative linear inverse problems [Bauschke1997-vk, Takahashi2024-wz], $\ell_{p}$ loss problems [Kyng2015-uo, Maddison2021-nt] phase retrieval [Bolte2018-zt, Takahashi2022-ml], nonnegative matrix factorization (NMF) [Mukkamala2019-mk, Takahashi2024-cu], and blind deconvolution [Takahashi2023-uh]. While the gradient of any $\mathcal{C}^{2}$ function over compact sets is Lipschitz continuous, an $\ell_{p}$ loss function is not $\mathcal{C}^{2}$ when $21<p<$ 2, i.e., its gradient is not Lipschitz continuous over compact sets. Moreover, functions in these applications are weakly convex on compact sets. We establish sublinear to linear convergence rates for the respective variants, which are compatible and in line with existing results.
Adaptive Bregman Step-Size Strategy
We propose an adaptive Bregman step-size strategy and establish its convergence results for convex and nonconvex optimization. Table 1 summarizes our contributions for the various cases that we consider. While Vyguzov and Stonyakin [Vyguzov2024-tp] also proposed an adaptive FW algorithm with the Bregman distance similar to ours for convex optimization problems. They established a linear convergence rate assuming that the objective function is relatively strongly convex, the optimal solution belongs to $\operatorname{int}P$, and the angle condition holds; in contrast, we establish linear convergence under much weaker assumptions. Beyond that, we also establish sublinear convergence for convex optimization for the case of general relatively smooth functions. Assuming that $f$ is convex, satisfies the Hölder error bound condition, and the optimal solution $x^{*}\in\operatorname{int}P$, we also provide the accelerated (up to linear) convergence, however, in a different way than [Vyguzov2024-tp] with fewer assumptions. For nonconvex optimization, assuming that $f$ is weakly convex and satisfies the local quadratic growth condition, we establish local linear convergence for our step-size strategy. These results include a special case that $\nabla f$ is Lipschitz continuous, i.e., $L$-smoothness. Without weak convexity, we prove the global convergence rate of the FW algorithm to a stationary point of (1.1), similar to [Lacoste-Julien2016-mj].
Away-Step FW Algorithm
We propose a variant of the away-step FW algorithm with the Bregman distance and establish its linear convergence under the Hölder error bound condition for convex optimization and under the local quadratic growth condition for weakly convex optimization.
Numerical Experiments
In numerical experiments, we have applied the FW algorithm with the adaptive Bregman step-size strategy to nonnegative linear systems [Bauschke2017-hg], $\ell_{p}$ loss problems [Kyng2015-uo, Maddison2021-nt], phase retrieval [Patterson1934-sg, Patterson1944-nm], low-rank minimization [Dragomir2021-rv], and NMF [Gillis2020-dh]. Bauschke et al. [Bauschke2017-hg] showed applications to linear inverse problems to solve nonnegative linear systems. Maddison et al. [Maddison2021-nt] considered $\ell_{p}$ loss problems, and Kyng et al. [Kyng2015-uo] showed its application as Lipschitz learning on graphs. Bolte et al. [Bolte2018-zt] and Takahashi et al. [Takahashi2022-ml] applied Bregman proximal algorithms to phase retrieval, Dragomir et al. [Dragomir2021-rv] applied them to low-rank minimization, and Mukkamala and Ochs [Mukkamala2019-mk] and Takahashi et al. [Takahashi2024-cu] applied them to NMF. Throughout these applications, we compare our proposed algorithms with existing FW algorithms and the mirror descent algorithm [Nemirovskij1983-sm] as well as away-step FW algorithms (see e.g. [Braun2022-ew, GueLat1986-bg, Lacoste-Julien2015-gb, Wolfe1970-ax]).
We would also like to stress that while we formulate some of the results for the case that $x^{*}\in\operatorname{int}P$, the results can be extended to the case that $x^{*}\in\operatorname{ri}P$, i.e., the relative interior of $P$, which we did not do for the sake of clarity. Basically, the analysis is performed in the affine space spanned by the optimal face of $P$ in that case; the interested reader is referred to [Braun2022-ew] for details on how to extend the results.
Table 1: Convergence rates for the Frank–Wolfe algorithm. For nonconvex functions, convergence is measured using the Frank-Wolfe gap, $\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\leq\epsilon$, instead of the primal gap, $f(x_{t})-f^{*}\leq\epsilon$. Weak convexity can be replaced by twice continuous differentiability (see Proposition 2.10). Convergence rates for the weakly convex optimization locally hold. AFW, HEB, and quad. denote the away-step Frank–Wolfe algorithm, the Hölder error bound condition with $q\geq 1$, and the quadratic growth condition, respectively.
FW Assumptions Convergence rate
$f$ convexity $f$ growth $x^{*}\in\operatorname{int}P$ $P$ polytope $L$-smooth $L$-smad ($\nu\in(0,1]$)
any convex $\mathcal{O}(\epsilon^{-1})$ [Jaggi2013-oz] $\mathcal{O}(\epsilon^{-1/\nu})$ ([Vyguzov2024-tp], Thm. 4.2)
[…]
any weak quad. [^3] We assume that $f$ is local quadratic growth. $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(\epsilon^{-1/\nu})$ (Thm. 5.5)
[…]
any [^5] Its rate is established for the Frank–Wolfe gap. nonconvex $\mathcal{O}(\epsilon^{-2})$ [Lacoste-Julien2016-mj] $\mathcal{O}(\epsilon^{-1-1/\nu})$ (Thm. 5.1)
Outline
The structure of this paper is as follows. Section 2 introduces essential notations, including Bregman distances, the smooth adaptable property, weak convexity, growth conditions, and scaling inequalities, which are crucial for the convergence analyses that come later. In Section 3, we present the FW algorithm with an adaptive Bregman step-size strategy and the away-step FW algorithm utilizing the Bregman distance. Section 4 details the convergence results for convex optimization, demonstrating both sublinear and linear convergence rates. Section 5 addresses convergence results for the nonconvex case, showing that our algorithm achieves global convergence to a stationary point of (1.1). Under the assumption that $f$ is weakly convex and local quadratic growth, we establish local linear convergence to a minimizer of (1.1), indicating that the proposed algorithms converge to a minimizer when the initial point is sufficiently close. Section 6 presents applications to nonnegative linear inverse problems, phase retrieval, low-rank minimization, and nonnegative matrix factorization (NMF).
2 Preliminaries
[…]
Let $\operatorname{int}{C}$, $\operatorname{ri}{C}$, and $\operatorname{cl}{C}$ be the interior, the relative interior, and the closure of a set $C\subset\mathbb{R}^{n}$, respectively. The set $\operatorname{conv}C$ denotes the convex hull of $C$, and the set $\operatorname{Vert}C$ denotes the set of vertices of $C$. We also define the distance from a point $x\in\mathbb{R}^{n}$ to $C$ by $\operatorname{dist}(x,C):=\inf_{y\in C}\|x-y\|$.
For an extended-real-valued function $f:\mathbb{R}^{n}\to[-\infty,+\infty]$, the effective domain of $f$ is defined by $\operatorname{dom}f:=\{x\in\mathbb{R}^{n}\,|\,f(x)<+\infty\}$. The function $f$ is said to be proper if $f(x)>-\infty$ for all $x\in\mathbb{R}^{n}$ and $\operatorname{dom}f\neq\emptyset$. Let $[f\leq\zeta]:=\{x\in\mathbb{R}^{n}\,|\,f(x)\leq\zeta\}$ be a $\zeta$-sublevel set of $f$ for some $\zeta\in\mathbb{R}$. Let $\mathcal{C}^{k}$ be the class of $k$-times continuously differentiable functions for $k\geq 0$. The sign function $\operatorname{sgn}(x)$ is defined by $\operatorname{sgn}(x)=-1$ for $x<0$, $\operatorname{sgn}(x)=0$ for $x=0$, and $\operatorname{sgn}(x)=1$ for $x>0$.
2.1 Bregman Distances
Let $C$ be a nonempty open convex subset of $\mathbb{R}^{n}$.
Definition 2.1 (Kernel Generating Distance [Bolte2018-zt, Definition 2.1]).
A function $\phi:\mathbb{R}^{n}\to(-\infty,+\infty]$ is called a kernel generating distance associated with $C$ if it satisfies the following conditions:
  1. (i)
    $\phi$ is proper, lower semicontinuous, and convex, with $\operatorname{dom}\phi\subset\operatorname{cl}C$ and $\operatorname{dom}\partial\phi=C$.
  2. (ii)
    $\phi$ is $\mathcal{C}^{1}$ on $\operatorname{int}\operatorname{dom}\phi\equiv C$.
We denote the class of kernel generating distances associated with $C$ by $\mathcal{G}(C)$.
Definition 2.2 (Bregman Distance [Bregman1967-rf]).
Given a kernel generating distance $\phi\in\mathcal{G}(C)$, a Bregman distance $D_{\phi}:\operatorname{dom}\phi\times\operatorname{int}\operatorname{dom}\phi\to\mathbb{R}_{+}$ associated with $\phi$ is defined by
$D_{\phi}(x,y)=\phi(x)-\phi(y)-\langle\nabla\phi(y),x-y\rangle.$ (2.1)
The Bregman distance $D_{\phi}(x,y)$ measures the proximity between $x\in\operatorname{dom}\phi$ and $y\in\operatorname{int}\operatorname{dom}\phi$. Indeed, since $\phi$ is convex, it holds that $D_{\phi}(x,y)\geq 0$ for all $x\in\operatorname{dom}\phi$ and $y\in\operatorname{int}\operatorname{dom}\phi$. Moreover, when $\phi$ is strictly convex, $D_{\phi}(x,y)=0$ holds if and only if $x=y$. However, the Bregman distance is not always symmetric and does not have to satisfy the triangle inequality. The Bregman distance is also called the Bregman divergence.
Example 2.3.
Well-known choices for $\phi$ and $D_{\phi}$ are listed below; for more examples, see, e.g., [Bauschke2017-hg, Example 1], [Bauschke1997-vk, Section 6], and [Dhillon2008-zz, Table 2.1].
  1. (i)
    Let $\phi(x)=\frac{1}{2}\|x\|^{2}$ and $\operatorname{dom}\phi=\mathbb{R}^{n}$. Then, the Bregman distance corresponds to the squared Euclidean distance, i.e., $D_{\phi}(x,y)=\frac{1}{2}\|x-y\|^{2}$.
  2. (ii)
    The Boltzmann–Shannon entropy $\phi(x)=\sum_{i=1}^{n}x_{i}\log x_{i}$ with $00\log 0=$ 0 and $\operatorname{dom}\phi=\mathbb{R}_{+}^{n}$. Then, $D_{\phi}(x,y)=\sum_{i=1}^{n}(x_{i}\log\frac{x_{i}}{y_{i}}-x_{i}+y_{i})$ is called the Kullback–Leibler divergence [Kullback1951-yv].
  3. (iii)
    The Burg entropy $\phi(x)=-\sum_{i=1}^{n}\log x_{i}$ and $\operatorname{dom}\phi=\mathbb{R}_{++}^{n}$. Then, $D_{\phi}(x,y)=\sum_{i=1}^{n}(\frac{x_{i}}{y_{i}}-\log\frac{x_{i}}{y_{i}}-1)$ is called the Itakura–Saito divergence [Itakura1968-en].
  4. (iv)
    Let $\phi(x)=\frac{1}{4}\|x\|^{4}+\frac{1}{2}\|x\|^{2}$ and $\operatorname{dom}\phi=\mathbb{R}^{n}$. The Bregman distance $D_{\phi}$ is used in phase retrieval [Bolte2018-zt], low-rank minimization [Dragomir2021-rv], NMF [Mukkamala2019-mk], and blind deconvolution [Takahashi2023-uh].
We recall the triangle scaling property for Bregman distances from Hanzely et al. [Hanzely2021-sz, Section 2], where several properties of the triangle scaling property are also shown.
Definition 2.4 (Triangle Scaling Property [Hanzely2021-sz, Definition 2]).
Given a kernel generating distance $\phi\in\mathcal{G}(C)$, the Bregman distance $D_{\phi}$ has the triangle scaling property if there exists a constant $\nu>0$ such that, for all $x,y,z\in\operatorname{int}\operatorname{dom}\phi$ and all $\gamma\in[0,1]$, it holds that
$D_{\phi}((1-\gamma)x+\gamma y,(1-\gamma)x+\gamma z)\leq\gamma^{\nu}D_{\phi}(y,z).$ (2.2)
Now, substituting $z\leftarrow x$ on the left-hand side of (2.2), we obtain
$D_{\phi}((1-\gamma)x+\gamma y,x)$ $=\phi((1-\gamma)x+\gamma y)-\phi(x)-\gamma\langle\nabla\phi(x),y-x\rangle$
$\leq(1-\gamma)\phi(x)+\gamma\phi(y)-\phi(x)-\gamma\langle\nabla\phi(x),y-x\rangle=\gamma D_{\phi}(y,x),$
where the inequality holds because of the convexity of $\phi$. Therefore, there exists $\nu>0$ such that $D_{\phi}((1-\gamma)x+\gamma y,x)\leq\gamma^{\nu}D_{\phi}(y,x)$ holds for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and all $\gamma\in[0,1]$.
We will now show that a stronger version can be obtained if $\phi$ is strictly convex, so that we can rephrase $\nu$ with $1+\nu$ where $\nu\in(0,1]$, i.e., the right-hand side is superlinear, which will be crucial for the convergence analysis later.
Lemma 2.5.
Given a kernel generating distance $\phi\in\mathcal{G}(C)$, if $\phi$ is strictly convex, then there exists $\nu\in(0,1]$ such that, for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and all $\gamma\in[0,1]$, it holds that
$D_{\phi}((1-\gamma)x+\gamma y,x)\leq\gamma^{1+\nu}D_{\phi}(y,x).$ (2.3)
Proof.
Obviously, (2.3) holds if $y=x$ or $\gamma\in\{0,1\}$. In what follows, we assume $y\neq x$ and $10<\gamma<$ 1. Let $g(\nu):=\gamma^{1+\nu}D_{\phi}(y,x)-D_{\phi}((1-\gamma)x+\gamma y,x)$ for all $x,y\in\operatorname{int}\operatorname{dom}\phi$, $y\neq x$ and all $\gamma\in(0,1)$. Using $D_{\phi}(y,x)>0$ due to the strict convexity of $\phi$, we have, for any $\nu\geq 0$,
$g^{\prime}(\nu)=\gamma^{1+\nu}D_{\phi}(y,x)\log\gamma<0,$
which implies $g$ monotonically decreases. In addition, it holds that
$g(0)$ $=\gamma D_{\phi}(y,x)-D_{\phi}((1-\gamma)x+\gamma y,x)$
$=\gamma(\phi(y)-\phi(x)-\langle\nabla\phi(x),y-x\rangle)-\phi((1-\gamma)x+\gamma y)+\phi(x)+\gamma\langle\nabla\phi(x),y-x\rangle$
$=(1-\gamma)\phi(x)+\gamma\phi(y)-\phi((1-\gamma)x+\gamma y)>0,$
where the last inequality holds because $\phi$ is strictly convex, i.e., $\phi((1-\gamma)x+\gamma y)<(1-\gamma)\phi(x)+\gamma\phi(y)$. Therefore, there exists $\nu>0$ such that $g(\nu)\geq 0$ by the intermediate value theorem. ∎
Because $\phi$ is quadratic when $D_{\phi}$ is symmetric [Bauschke2001-ry, Lemma 3.16], it always holds that $D_{\phi}((1-\gamma)x+\gamma y,x)=\gamma^{1+\nu}D_{\phi}(y,x)$ with $\nu=1$.
In the remainder of the paper, if not stated otherwise, we will assume that $\phi$ is strictly convex and as such for a given $\phi$ and $C$ there exists a $\nu\in(0,1]$ such that $D_{\phi}((1-\gamma)x+\gamma y,x)\leq\gamma^{1+\nu}D_{\phi}(y,x)$ holds for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and all $\gamma\in[0,1]$.
2.2 Smooth Adaptable Property
We will now recall the smooth adaptable property, which is a generalization of $L$-smoothness and was first introduced by [Bauschke2017-hg]. The smooth adaptable property is also called relative smoothness [Lu2018-ii].
Definition 2.6 ($L$-smooth Adaptable Property [Bolte2018-zt, Definition 2.2]).
Consider a pair of functions $(f,\phi)$ satisfying the following conditions:
  1. (i)
    $\phi\in\mathcal{G}(C)$,
  2. (ii)
    $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is proper and lower semicontinuous with $\operatorname{dom}\phi\subset\operatorname{dom}f$, which is $\mathcal{C}^{1}$ on $C\equiv\operatorname{int}\operatorname{dom}\phi$.
The pair $(f,\phi)$ is said to be $L$-smooth adaptable (for short: $L$-smad) on $C$ if there exists $L>0$ such that $L\phi-f$ and $L\phi+f$ are convex on $C$.
The convexity of $L\phi-f$ and $L\phi+f$ plays a central role in developing and analyzing algorithms, and the smooth adaptable property implies the extended descent lemma.
Lemma 2.7 (Extended Descent Lemma [Bolte2018-zt, Lemma 2.1]).
The pair of functions $(f,\phi)$ is $L$-smad on $C$ if and only if for all $x,y\in\operatorname{int}\operatorname{dom}\phi$,
$|f(x)-f(y)-\langle\nabla f(y),x-y\rangle|\leq LD_{\phi}(x,y).$
From this, it can be seen that the $L$-smad property for $(f,\phi)$ provides upper and lower approximations for $f$ majorized by $\phi$ with $L>0$. In addition, if $\phi(x)=\frac{1}{2}\|x\|^{2}$ on $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$, Lemma 2.7 corresponds to the classical descent lemma. While the $L$-smad property might seem unfamiliar at first, it is a natural generalization of $L$-smoothness, and examples of functions $f$ and $\phi$ satisfying the $L$-smad property are given, e.g., in [Bauschke2017-hg, Lemmas 7 and 8][Bolte2018-zt, Lemma 5.1], [Dragomir2021-rv, Propositions 2.1 and 2.3], [Mukkamala2019-mk, Proposition 2.1], [Takahashi2024-wz, Proposition 24], [Takahashi2023-uh, Theorem 1], and [Takahashi2024-cu, Theorem 15].
The extended descent lemma immediately implies primal progress of FW algorithms under the $L$-smad property as a straightforward generalization of the $L$-smooth case.
Lemma 2.8 (Primal progress from the smooth adaptable property).
Let the pair of functions $(f,\phi)$ be $L$-smad with a strictly convex function $\phi$ and let $x_{+}=(1-\gamma)x+\gamma v$ with $x,v\in\operatorname{int}\operatorname{dom}\phi$ and $\gamma\in[0,1]$. Then it holds:
$f(x)-f(x_{+})\geq\gamma\langle\nabla f(x),x-v\rangle-L\gamma^{1+\nu}D_{\phi}(v,x).$
Proof.
Using Lemma 2.7 and substituting $x_{+}$ for $x$ and $x$ for $y$, we have
$f(x)-f(x_{+})$ $\geq\gamma\langle\nabla f(x),x-v\rangle-LD_{\phi}(x_{+},x).$
It holds that $D_{\phi}(x_{+},x)\leq\gamma^{1+\nu}D_{\phi}(v,x)$ for some $\nu\in(0,1]$ by Lemma 2.5. Therefore, this provides the desired inequality. ∎
Often the point $v\in P\subset\operatorname{int}\operatorname{dom}\phi$ in Lemma 2.8 is chosen as a Frank–Wolfe vertex, i.e.,
$v\in\operatorname*{argmax}_{u\in P}\langle\nabla f(x),x-u\rangle,$
but other choices, e.g., those arising from away-directions from the away-step FW method, are also possible, as we will see in Section 3.2.
2.3 Weakly Convex Functions
A function $f$ is called weakly convex if it becomes convex upon adding a quadratic perturbation.
Definition 2.9 (Weakly Convex Function [Davis2018-bb, Nurminskii1975-vp]).
A function $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is said to be $\rho$-weakly convex for some $\rho>0$ if the function $f+\frac{\rho}{2}\|\cdot\|^{2}$ is convex.
Obviously, $f$ is $\rho$-weakly convex and differentiable if and only if, for all $x,y\in\operatorname{dom}f$,
$f(y)-f(x)\geq\langle\nabla f(x),y-x\rangle-\frac{\rho}{2}\|y-x\|^{2}.$
For examples of such $f$, see [Davis2018-bb, Examples 3.1 and 3.2]. Moreover, it is easy to see that any $\mathcal{C}^{2}$ function defined on a compact set is weakly convex [Vial1983-oo, Proposition 4.11] and [Wu2007-dt], the rationale being that we can simply shift the function by the quadratic belonging to its smallest eigenvalue. Since the proof of this fact is omitted therein, we provide the argument for the sake of completeness below.
Proposition 2.10.
Let $P\subset\mathbb{R}^{n}$ be a compact set and $f$ be a proper and $\mathcal{C}^{2}$ function on $P$. Then, $f$ is weakly convex on $P$.
Proof.
Let $g=\lambda_{\min}(\nabla^{2}f(\cdot))$. Obviously, $g$ is continuous because $\lambda_{\min}(\cdot)$ and $\nabla^{2}f(\cdot)$ are continuous. Therefore, there exists the minimum value of $g$ on $P$ due to the continuity of $g$ and the compactness of $P$. For $\rho=|\min_{x\in P}g(x)|$, we have $\nabla^{2}f(x)+\rho I\succeq O$ for any $x\in P$, which implies $f+\frac{\rho}{2}\|\cdot\|^{2}$ is convex on $P$, i.e., $f$ is $\rho$-weakly convex on $P$. ∎
Minimizing a nonconvex $\mathcal{C}^{2}$-function over a compact set $P$ is equivalent to minimizing a weakly convex function over $P$. However, it is important to note that when $\rho$ becomes large, the assumptions of algorithms may not be satisfied (see also Theorems 5.5 and 5.9).
2.4 Growth Conditions
In [garber2015faster] (see also [Garber2020-tr]) the linear convergence of the FW algorithm for convex optimization under the quadratic growth condition, which is a weaker assumption than assuming strong convexity was established over strongly convex sets and later generalized to uniformly convex sets in [KAP2020] as well as to conditions weaker than quadratic growth in [KDP2018, KDP2018jour]. Local variants of these notions, as necessary, e.g., for the nonconvex case, have been studied in [KAP2021] in the context of Frank–Wolfe methods.
Definition 2.11 (Hölder Error Bound [Braun2022-ew, Definition 3.25], [alex17sharp, Definition 1.2] and Quadratic Growth Conditions [garber2015faster, Garber2020-tr, Liao2024-vd]).
Let $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ be a proper lower semicontinuous function and $P\subset\mathbb{R}^{n}$ be a compact convex set. Let $\mathcal{X}^{*}\neq\emptyset$ be the set of optimal solutions, i.e., $\mathcal{X}^{*}:=\operatorname*{argmin}_{x\in P}f(x)$, and let $f^{*}=\min_{x\in P}f(x)$ and $\zeta>0$. The function $f$ satisfies the Hölder error bound condition on $P$ if there exists constants $q\geq 1$ and $\mu>0$ such that, for any $x\in P$,
$\operatorname{dist}(x,\mathcal{X}^{*})^{q}\leq\frac{q}{\mu}(f(x)-f^{*}).$ (2.4)
Particularly, when $q=2$, it is called the $\mu$-quadratic growth condition. Moreover, the function $f$ is said to be local $\mu$-quadratic growth (with $\zeta$) if there exists a constant $\mu>0$ such that, for any $x\in[f\leq f^{*}+\zeta]\cap P$, (2.4) holds with $q=2$.
The Hölder error bound condition shows sharpness bounds on the primal gap (see, e.g., [bolte17, alex17sharp]), which has also been extensively analyzed for Frank–Wolfe methods in [KDP2018, KDP2018jour]. The convergence analysis for nonconvex optimization under the quadratic growth or Hölder error bound condition was established in [Davis2024-rm, Davis2018-bb, Davis2024-og]. We distinguish between the global condition and the local condition of Definition 2.11. We employ the Hölder error bound condition for convex optimization while we use the local quadratic growth for nonconvex optimization.
Next, we show in the following lemma that the Hölder error bound condition immediately provides a bound on the primal optimality gap, which will be useful for establishing convergence rates of the proposed algorithms as well as the Łojasiewicz inequality [Bolte2007-xr, Lojasiewicz1963-sm, Lojasiewicz1965-ks], provided $f$ is convex.
Lemma 2.12 (Primal gap bound from Hölder error bound).
Let $f$ be a convex function and satisfy the Hölder error bound condition with $q\geq 1$ and $\mu>0$. Let $x^{*}$ be the unique minimizer of $f$ over $P$. Then, the following argument holds in general, for all $x\in P$:
$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{\frac{q}{q-1}},$ (2.5)
or equivalently,
$\left(\frac{\mu}{q}\right)^{1/q}(f(x)-f^{*})^{1-1/q}\leq\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}.$
Proof.
By first applying convexity and then the Hölder error bound condition for any $x^{*}\in\mathcal{X}^{*}$ with $f^{*}=f(x^{*})$ it holds:
$f(x)-f^{*}$ $\leq\langle\nabla f(x),x-x^{*}\rangle$
$=\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\|x-x^{*}\|$
$\leq\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\left(\frac{q}{\mu}(f(x)-f^{*})\right)^{1/q},$
which implies
$\left(\frac{\mu}{q}\right)^{1/q}(f(x)-f^{*})^{1-1/q}\leq\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|},$
or equivalently
$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{\frac{q}{q-1}}.$
Remark 2.13.
With the remark from above, we can immediately relate the Hölder error bound condition to the Łojasiewicz inequality [Bolte2007-xr, Lojasiewicz1963-sm, Lojasiewicz1965-ks]; the Łojasiewicz inequality is used to establish convergence analysis [Bolte2007-xr]. To this end, using (2.5), we estimate $\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\leq\frac{\|\nabla f(x)\|\|x-x^{*}\|}{\|x-x^{*}\|}=\|\nabla f(x)\|$ to then obtain the weaker condition:
$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\|\nabla f(x)\|^{\frac{q}{q-1}}=c^{\frac{1}{\theta}}\|\nabla f(x)\|^{\frac{1}{\theta}},$ (2.6)
where $c=\left(\frac{q}{\mu}\right)^{\frac{1}{q}}$ and $\theta=\frac{q-1}{q}$. Inequality (2.6) is called the $c$-Łojasiewicz inequality with $\theta\in[0,1)$. If $\mathcal{X}^{*}\subseteq\operatorname{ri}P$, then the two conditions are equivalent. However, if the optimal solution(s) are on the boundary of $P$ as is not infrequently the case, then the two conditions are not equivalent as $\|\nabla f(x)\|$ might not vanish for $x\in\mathcal{X}^{*}$, whereas $\langle\nabla f(x),x-x^{*}\rangle$ does, i.e., the Hölder error bound condition is tighter than the one induced by the Łojasiewicz inequality.
The next lemma shows that we can also obtain primal gap bounds in the weakly convex case together with (local) quadratic growth. Its proof can be found in Appendix A.1.
Lemma 2.14 (Primal gap bound from the quadratic growth).
Let $f$ be a $\rho$-weakly convex function that satisfies the local $\mu$-quadratic growth condition such that $\rho\leq\mu$. Let $x^{*}$ be the unique minimizer of $f$ over $P$ and let $\zeta>0$. Then, the following holds: for all $x\in[f\leq f^{*}+\zeta]\cap P$,
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{2},$ (2.7)
or equivalently,
$\left(\frac{\mu}{2}\right)^{1/2}\left(1-\frac{\rho}{\mu}\right)(f(x)-f^{*})^{1/2}\leq\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}.$
Remark 2.15.
In the same vein as the discussion in Remark 2.13, we can immediately relate the local $\mu$-quadratic growth condition with $\zeta>0$ to the local Polyak–Łojasiewicz (PL) inequality. Using (2.7), for all $x\in[f\leq f^{*}+\zeta]\cap P$, we have
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\|\nabla f(x)\|^{2},$ (2.8)
which is equivalent to the PL inequality [Polyak1963-dr], also called the gradient dominance property [Combettes2020-jj], with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$. The PL inequality is also equivalent to the Łojasiewicz inequality with $\theta=\frac{1}{2}$. Strongly convex functions satisfy the PL inequality [Braun2022-ew, Lemma 2.13]. Note that we will not have the primal gap bound under the Hölder error bound condition and weak convexity because an inequality like (A.1) does not hold from the Hölder error bound condition.
2.5 Scaling Inequalities, Geometric Hölder Error Bounds, and Contractions
In the following, we will now bring things together to derive tools that will be helpful in establishing convergence rates.
Scaling inequalities
Scaling inequalities are a key tool in establishing convergence rates of FW algorithms. We will introduce two such inequalities that we will use in the following. The first scaling inequality is useful for analyzing the case where the optimal solution lies in the relative interior of the feasible region. While its formulation in [Braun2022-ew, Proposition 2.16] required $L$-smoothness of $f$, it is actually not used in the proof, and the results hold more broadly. We restate it here for the sake of completeness.
Proposition 2.16 (Scaling inequality for inner optima from convexity [Braun2022-ew, Proposition 2.16]).
Let $P\subset\mathbb{R}^{n}$ be a nonempty compact convex set. Let $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ be $\mathcal{C}^{1}$ and convex on $P$. If there exists $r>0$ so that $B(x^{*},r)\subset P$ for a minimizer $x^{*}$ of $f$, then for all $x\in P$, we have
$\langle\nabla f(x),x-v\rangle\geq r\|\nabla f(x)\|\geq\frac{r\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|},$
where $v\in\operatorname*{argmax}_{u\in P}\langle\nabla f(x),x-u\rangle$.
When $f$ is not convex, assuming that $f$ is weakly convex and local quadratic growth, we have the scaling inequality for a nonconvex objective function.
Proposition 2.17 (Scaling inequality for inner optima from weak convexity).
Let $P\subset\mathbb{R}^{n}$ be a nonempty compact convex set. Let $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ be $\mathcal{C}^{1}$, $\rho$-weakly convex, local $\mu$-quadratic growth with $\zeta>0$ on $P$. If there exists $r>0$ so that $B(x^{*},r)\subset P$ for a minimizer $x^{*}$ of $f$ and $\rho\leq\mu$, then for all $x\in[f\leq f^{*}+\zeta]\cap P$, we have
$\langle\nabla f(x),x-v\rangle\geq r\|\nabla f(x)\|\geq\frac{r\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|},$
where $v\in\operatorname*{argmax}_{u\in P}\langle\nabla f(x),x-u\rangle$.
Proof.
We consider $x^{*}-rz$, where $z$ is a point with $\|z\|=1$ and $\langle\nabla f(x),z\rangle=\|\nabla f(x)\|$. It holds that
$\langle\nabla f(x),v\rangle\leq\langle\nabla f(x),x^{*}-rz\rangle=\langle\nabla f(x),x^{*}\rangle-r\|\nabla f(x)\|.$ (2.9)
From weak convexity, for all $x\in[f\leq f^{*}+\zeta]\cap P$, we have
$f^{*}-f(x)$ $\geq\langle\nabla f(x),x^{*}-x\rangle-\frac{\rho}{2}\|x-x^{*}\|^{2}$
$\geq\langle\nabla f(x),x^{*}-x\rangle-\frac{\rho}{\mu}(f(x)-f^{*}),$
where the last inequality holds due to the local quadratic growth condition. Because of $\rho\leq\mu$, this inequality implies
$\langle\nabla f(x),x-x^{*}\rangle\geq\left(1-\frac{\rho}{\mu}\right)(f(x)-f^{*})\geq 0.$
By rearranging (2.9) and using $\langle\nabla f(x),x-x^{*}\rangle\geq 0$, we obtain
$\langle\nabla f(x),x-v\rangle\geq\langle\nabla f(x),x-x^{*}\rangle+r\|\nabla f(x)\|\geq r\|\nabla f(x)\|.$
In addition, it holds that
$\frac{r\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\leq\frac{r\|\nabla f(x)\|\|x-x^{*}\|}{\|x-x^{*}\|}=r\|\nabla f(x)\|,$
where the first inequality holds because of the Cauchy–Schwarz inequality. ∎
The authors in [Lacoste-Julien2015-gb] defined a geometric distance-like constant of a polytope, known as the pyramidal width, to analyze the convergence of the away-step Frank–Wolfe algorithm (and other variants that use the pyramidal width) over polytopes. It can be interpreted as the minimal $\delta>0$ satisfying the following scaling inequality (2.10), which plays a central role in establishing convergence rates for the away-step FW algorithm; see [Braun2022-ew] for an in-depth discussion.
Lemma 2.18 (Scaling inequality via pyramidal width [Braun2022-ew, Theorem 2.26][Lacoste-Julien2015-gb, Theorem 3]).
Let $P\subset\mathbb{R}^{n}$ be a polytope and let $\delta$ denote the pyramidal width of $P$. Let $x\in P$, and let $\mathcal{S}$ denote any set of vertices of $P$ with $x\in\operatorname{conv}\mathcal{S}$. Let $\psi$ be any vector, so that we define $v^{\textup{FW}}=\operatorname*{argmin}_{v\in P}\langle\psi,v\rangle$ and $v^{\textup{A}}=\operatorname*{argmax}_{v\in\mathcal{S}}\langle\psi,v\rangle$. Then for any $y\in P$
$\langle\psi,v^{\textup{A}}-v^{\textup{FW}}\rangle\geq\delta\frac{\langle\psi,x-y\rangle}{\|x-y\|}.$ (2.10)
Note that Lemma 2.18 does not require the convexity of $f$, and we will use it for nonconvex optimization. When $\phi$ is $\sigma$-strongly convex and $\langle\psi,x-y\rangle\geq 0$, Lemma 2.18 implies
$\langle\psi,v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}\geq\delta^{2}\frac{\langle\psi,x-y\rangle^{2}}{\|x-y\|^{2}}\geq\delta^{2}\sigma\frac{\langle\psi,x-y\rangle^{2}}{2D_{\phi}(x,y)}.$
Note that the last inequality of the above also holds for $\delta^{2}\sigma\frac{\langle\psi,x-y\rangle^{2}}{2D_{\phi}(y,x)}$, i.e., with $x$ and $y$ swapped in the divergence.
Geometric Hölder Error Bound Condition
We will now introduce the more compact notion of geometric Hölder error bound condition, which simply combines the pyramidal width and the Hölder error bound condition of the function $f$; see also [Braun2022-ew, Lemma 2.27] for details when $q=2$.
Lemma 2.19 (Geometric Hölder Error Bound).
Let $P$ be a polytope with pyramidal width $\delta>0$ and let $f$ be a convex function and satisfy the Hölder error bound condition with $v^{\textup{FW}}=\operatorname*{argmin}_{v\in P}\langle\nabla f(x),v\rangle$ and $v^{\textup{A}}=\operatorname*{argmax}_{v\in\mathcal{S}}\langle\nabla f(x),v\rangle$ with $\mathcal{S}\subseteq\operatorname{Vert}{P}$, so that $x\in\operatorname{conv}{\mathcal{S}}$, we have
$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle}{\delta}\right)^{\frac{q}{q-1}}.$
Proof.
Combining (2.10) with Lemma 2.12 for $\psi=\nabla f(x)$, we have
$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{\frac{q}{q-1}}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle}{\delta}\right)^{\frac{q}{q-1}}.$
From Contractions to Convergence Rates
Besides scaling inequalities and the geometric Hölder error bound, we also utilize the following lemma for convex and nonconvex optimization, which allows us to turn a contraction into a convergence rate.
Lemma 2.20 (From contractions to convergence rates [Braun2022-ew, Lemma 2.21]).
Let $\{h_{t}\}_{t}$ be a decreasing sequence of positive numbers and $c_{0}$, $c_{1}$, $c_{2}$, $\theta_{0}$ be positive numbers with $c_{1}<1$ such that $h_{1}\leq c_{0}$ and $h_{t}-h_{t+1}\geq h_{t}\min\{c_{1},c_{2}h_{t}^{\theta_{0}}\}$ for $t\geq 1$, then
$h_{t}\leq\begin{cases}c_{0}(1-c_{1})^{t-1}&\text{if}\ 1\leq t\leq t_{0},\\ \frac{(c_{1}/c_{2})^{1/\theta_{0}}}{(1+c_{1}\theta_{0}(t-t_{0}))^{1/\theta_{0}}}=\mathcal{O}(1/t^{1/\theta_{0}})&\text{if}\ t\geq t_{0},\end{cases}$
where
$t_{0}:=\max\left\{\left\lfloor\log_{1-c_{1}}\left(\frac{(c_{1}/c_{2})^{1/\theta_{0}}}{c_{0}}\right)\right\rfloor+2,1\right\}.$
In particular, we have $h_{t}\leq\epsilon$ if $t\geq t_{0}+\frac{1}{\theta_{0}c_{2}\epsilon^{\theta_{0}}}-\frac{1}{\theta_{0}c_{1}}$ and $\epsilon\leq(c_{1}/c_{2})^{1/\theta_{0}}$.
3 Proposed Algorithms
Throughout this paper, we make the following assumptions.
Assumption 3.1.
  1. (i)
    $\phi\in\mathcal{G}(C)$ with $\operatorname{cl}{C}=\operatorname{cl}{\operatorname{dom}\phi}$ is strictly convex on $C\equiv\operatorname{int}\operatorname{dom}\phi$.
  2. (ii)
    $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is a proper and lower semicontinuous function with $\operatorname{dom}\phi\subset\operatorname{dom}f$, which is $\mathcal{C}^{1}$ on $C$.
  3. (iii)
    The pair $(f,\phi)$ is $L$-smad on $P$.
  4. (iv)
    $P\subset\mathbb{R}^{n}$ is a nonempty compact convex set with $P\subset C$.
Assumption 3.1(i)-(iii) are standard for Bregman-type algorithms [Bolte2018-zt, Takahashi2024-wz]. The strict convexity ensures the existence of $\nu$ by Lemma 2.5 and is satisfied by many kernel generating distances (see Example 2.3). Assumption 3.1(iv) is also standard for FW algorithms, and the condition $P\subset C$ is natural ensuring that the Bregman distance $D_{\phi}$ is well-defined on $P$.
In what follows, let $x^{*}\in\operatorname*{argmin}_{x\in P}f(x)$ and let $f^{*}=f(x^{*})$. We recall a key property for analyzing the convergence rates of FW algorithms in convex optimization.
Lemma 3.2 (Primal gap, dual gap, and Frank–Wolfe gap [Pokutta2024-qu, Lemma 4.1]).
Let $x^{*}\in P$ be an optimal solution of (1.1) and let $f^{*}=f(x^{*})$. Suppose that Assumptions 3.1(ii), (iv) and that $f$ is convex. Then for all $x\in P$, it holds that
$f(x)-f^{*}\leq\langle\nabla f(x),x-x^{*}\rangle\leq\max_{v\in P}\langle\nabla f(x),x-v\rangle.$ (3.1)
3.1 Adaptive Bregman Step-Size Strategy
Assume that $(f,\phi)$ is $L$-smad. Substituting $x_{t+1}=(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$ with $x_{t},v_{t}\in P$ for $x_{+}$ in Lemma 2.8 provides
$f(x_{t})-f(x_{t+1})\geq\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle-L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t}).$ (3.2)
Using the optimality condition of the right-hand side of (3.2) in terms of $\gamma_{t}$, we find
$\gamma_{t}=\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}}\quad\text{and}\quad f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle^{1+1/\nu}}{(L(1+\nu)D_{\phi}(v_{t},x_{t}))^{1/\nu}}.$
Theoretically, when $\gamma_{t}\in[0,1]$, the Frank–Wolfe step $x_{t+1}=(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}\in P$ is well-defined. We update $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},\gamma_{\max}\right\}$ with some $\gamma_{\max}\in\mathbb{R}$ (usually set $\gamma_{\max}=1$). This step-size strategy provides the progress lemma. Note that the following lemma does not require the convexity of $f$.
Lemma 3.3 (Progress lemma from the Bregman short step-size).
Suppose that Assumption 3.1 holds. Define $x_{t+1}=(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$. Then if $x_{t+1}\in\operatorname{dom}f$,
$f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\quad\text{for}\quad 0\leq\gamma_{t}\leq\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}}.$
Proof.
Recalling (3.2), we have
$f(x_{t})-f(x_{t+1})$ $\geq\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle-L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t})$
$\geq\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle-\frac{\gamma_{t}}{1+\nu}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle$
$=\frac{\nu}{1+\nu}\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle,$
where the second inequality holds because of an upper bound of $\gamma_{t}$. ∎
If $f$ is convex, using Lemma 3.3, we have
$f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq\frac{\nu}{1+\nu}\gamma_{t}(f(x_{t})-f^{*}),$
which implies that the primal progress is at least $\frac{\nu}{1+\nu}$ of the Frank–Wolfe gap and of the primal gap multiplied by $\gamma_{t}$. If $\phi=\frac{1}{2}\|\cdot\|^{2}$, we have $\nu=1$ and
$\gamma_{t}=\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L\|v_{t}-x_{t}\|^{2}},$
which is often called the (Euclidean) short step-size. Although the short step-size strategy does not require line searches, it requires knowledge of the value of $L$.
The exact value or the tight upper bound of $L$ is often unknown, however, the algorithm’s performance heavily depends on it; an underestimation of $L$ might lead to non-convergence, and an overestimation of $L$ might lead to slow convergence. Moreover, a worst-case $L$ might be too conversative for regimes where the function is better behaved. Due to all these reasons, Pedregosa et al. [Pedregosa2020-ay] proposed an adaptive step-size strategy for FW algorithms in the case where $\phi=\frac{1}{2}\|\cdot\|^{2}$. Vyguzov and Stonyakin [Vyguzov2024-tp] proposed a variant using Bregman distances. We have concurrently and independently formulated an improved version similar (but different in details) to [Vyguzov2024-tp], which includes updates to $\nu$ and uses different coefficients for $\gamma_{t}$. We present our algorithm in Algorithm 2.
Output : Updated estimates $\tilde{L}^{*}$ and $\tilde{\nu}^{*}$, step-size $\gamma$
[…]
2      Choose $\beta<1$, $\eta\leq 1$, and $\tau>1$
3     $\kappa\leftarrow 1$
4     $M\leftarrow\eta\tilde{L}$
5     loop
6           $\gamma\leftarrow\min\left\{\left(\frac{\langle\nabla f(x),x-v\rangle}{M(1+\kappa)D_{\phi}(v,x)}\right)^{\frac{1}{\kappa}},\gamma_{\max}\right\}$
7          if $f((1-\gamma)x+\gamma v)-f(x)-\gamma\langle\nabla f(x),v-x\rangle\leq M\gamma^{1+\kappa}D_{\phi}(v,x)$ then
8                $\tilde{L}^{*}\leftarrow M$
9               $\tilde{\nu}^{*}\leftarrow\kappa$
10               return $\tilde{L}^{*}$, $\tilde{\nu}^{*}$, $\gamma$
11                $M\leftarrow\tau M$
12               $\kappa\leftarrow\beta\kappa$
13
Algorithm 2 Adaptive Bregman step-size strategy
This step-size strategy can be used as a drop-in replacement. For example, inserting $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},1)$ between lines 1 and 1 in Algorithm 1, we obtain that Algorithm 2 searches $L$ and $\nu$ until
$f((1-\gamma)x+\gamma v)-f(x)-\gamma\langle\nabla f(x),v-x\rangle\leq M\gamma^{1+\kappa}D_{\phi}(v,x)$ (3.3)
holds for the approximation values $M$ and $\kappa$ of $L$ and $\nu$, respectively.
Remark 3.4 (Well-definedness and termination of Algorithm 2).
By the $L$-smad property of $(f,\phi)$ and Lemma 2.8, we know that (3.3) holds for all $M\geq L$ and $\nu\geq\kappa>0$. Therefore, the loop in line 2 in Algorithm 2 is well-defined and guaranteed to terminate. Thus it suffices to update $\kappa$ in line 2 only if $D_{\phi}((1-\gamma)x+\gamma v,x)>\gamma^{1+\kappa}D_{\phi}(v,x)$ holds. Thus, $\kappa$ cannot become too small.
3.2 The Away-Step Frank–Wolfe Algorithm
In the case where $P$ is a polytope, the classical FW algorithm might zigzag when approaching the optimal face and, in consequence, converge slowly. To overcome this drawback, Wolfe [Wolfe1970-ax] introduced the away-step Frank–Wolfe algorithm and away steps, which allow the algorithm to move away from vertices in a convex combination of $x_{t}$, which can effectively shortcircuit the zigzagging. The convergence properties of this algorithm were unknown for a long time, and it was only quite recently that Lacoste-Julien and Jaggi [Lacoste-Julien2015-gb] established the linear convergence of the away-step Frank-Wolfe algorithm. Inspired by [GueLat1986-bg, Lacoste-Julien2015-gb, Wolfe1970-ax], we propose a variant of the away-step Frank–Wolfe algorithm utilizing Bregman distances as given in Algorithm 3.
For the following discussion, we introduce the following notions. In the same way [Braun2022-ew], we define an active set $\mathcal{S}\subset\operatorname{Vert}P$ and the away vertex as $v_{t}^{\textup{A}}\in\operatorname*{argmax}_{v\in\mathcal{S}}\langle\nabla f(x_{t}),v\rangle$, where $x_{t}$ is a (strict) convex combination of elements in $\mathcal{S}$, i.e., $x_{t}=\sum_{v\in\mathcal{S}}\lambda_{v}x_{t}$ with $\lambda_{v}>0$ for all $v\in\mathcal{S}$ and $\sum_{v\in\mathcal{S}}\lambda_{v}=1$. Algorithm 3 updates $\nu$ until $D_{\phi}((1-\gamma)x_{t}+\gamma v_{t},x_{t})\leq\gamma^{1+\nu}D_{\phi}(v,x)$ holds in lines 3-3. If $\nu$ is known, one only updates $\gamma_{t}$. If $\nu$ and $L$ are unknown, we can use $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},\gamma_{t,\max})$ with $\gamma\leftarrow\min\left\{\left(\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{M(1+\kappa)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\kappa}},\gamma_{t,\max}\right\}$ instead of line 2 in Algorithm 2. When $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$, Algorithm 3 corresponds to the classical away-step FW algorithm.
Input: Initial point $x_{0}\in\operatorname*{argmin}_{v\in P}\langle\nabla f(x),v\rangle$ for $x\in P$, $\beta<1$
1 $\mathcal{S}_{0}\leftarrow\{x_{0}\}$, $\lambda_{x_{0},0}\leftarrow 1$
2for $t=0,\ldots$ do
3      $v_{t}^{\textup{FW}}\leftarrow\operatorname*{argmin}_{v\in P}\langle\nabla f(x_{t}),v\rangle$
4     $v_{t}^{\textup{A}}\leftarrow\operatorname*{argmax}_{v\in\mathcal{S}_{t}}\langle\nabla f(x_{t}),v\rangle$
5     if $\langle\nabla f(x_{t}),x_{t}-v_{t}^{\textup{FW}}\rangle\geq\langle\nabla f(x_{t}),v_{t}^{\textup{A}}-x_{t}\rangle$ then
6           $v_{t}\leftarrow v_{t}^{\textup{FW}}$, $d_{t}\leftarrow x_{t}-v_{t}^{\textup{FW}}$, $\gamma_{t,\max}\leftarrow 1$
7           else
8                $v_{t}\leftarrow v_{t}^{\textup{A}}$, $d_{t}\leftarrow v_{t}^{\textup{A}}-x_{t}$, $\gamma_{t,\max}\leftarrow\frac{\lambda_{v_{t}^{\textup{A}},t}}{1-\lambda_{v_{t}^{\textup{A}},t}}$
9                $\nu\leftarrow 1$
10               while $D_{\phi}((1-\gamma)x_{t}+\gamma v_{t},x_{t})\leq\gamma^{1+\nu}D_{\phi}(v_{t},x_{t})$ do
11                     $\gamma\leftarrow\min\left\{\left(\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},\gamma_{t,\max}\right\}$, $\nu\leftarrow\beta\nu$
12
13                    $\gamma_{t}\leftarrow\gamma$, $x_{t+1}\leftarrow x_{t}-\gamma_{t}d_{t}$
14                    if $\langle\nabla f(x_{t}),x_{t}-v_{t}^{\textup{FW}}\rangle\geq\langle\nabla f(x_{t}),v_{t}^{\textup{A}}-x_{t}\rangle$ then
15                          $\lambda_{v,t+1}\leftarrow(1-\gamma_{t})\lambda_{v,t}$ for all $v_{t}\in\mathcal{S}_{t}\setminus\{v_{t}^{\textup{FW}}\}$
16                         $\lambda_{v_{t}^{\textup{FW}},t+1}\leftarrow\begin{cases}\gamma_{t}&\text{if}\ v_{t}^{\textup{FW}}\notin\mathcal{S}_{t}\\ (1-\gamma_{t})\lambda_{v_{t}^{\textup{FW}},t}+\gamma_{t}&\text{if}\ v_{t}^{\textup{FW}}\in\mathcal{S}_{t}\end{cases}$
17                         $\mathcal{S}_{t+1}\leftarrow\begin{cases}\mathcal{S}_{t}\cup\{v_{t}^{\textup{FW}}\}&\text{if}\ \gamma_{t}<1\\ \{v_{t}^{\textup{FW}}\}&\text{if}\ \gamma_{t}=1\end{cases}$
18                          else
19                               $\lambda_{v,t+1}\leftarrow(1+\gamma_{t})\lambda_{v,t}$ for all $v_{t}\in\mathcal{S}_{t}\setminus\{v_{t}^{\textup{A}}\}$
20                              $\lambda_{v_{t}^{\textup{A}},t+1}\leftarrow(1+\gamma_{t})\lambda_{v_{t}^{\textup{A}},t}-\gamma_{t}$
21                              $\mathcal{S}_{t+1}\leftarrow\begin{cases}\mathcal{S}_{t}\setminus\{v_{t}^{\textup{A}}\}&\text{if}\ \lambda_{v_{t}^{\textup{A}},t+1}=0\\ \mathcal{S}_{t}&\text{if}\ \lambda_{v_{t}^{\textup{A}},t+1}>0\end{cases}$
22
Algorithm 3 Away-Step Frank–Wolfe algorithm with the Bregman distance
4 Convergence Analysis for Convex Optimization
In this section, we assume that $f$ is convex.
Assumption 4.1.
The objective function $f$ is convex.
4.1 Sublinear Convergence
We establish a sublinear convergence rate of the FW algorithm under the smooth adaptable property. It is a similar result to [Vyguzov2024-tp, Theorem 1]. We conducted its proof following [Braun2022-ew, Jaggi2013-oz, Pokutta2024-qu]. The FW algorithm uses the open loop step-size, i.e., $\gamma_{t}=\frac{2}{2+t}$ in the following theorem.
Theorem 4.2 (Primal convergence of the Frank–Wolfe algorithm).
Suppose that Assumptions 3.1 and 4.1 hold. Let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$ characterized by the Bregman distance $D_{\phi}$. Consider the iterates of Algorithm 1 with the open loop step-size, i.e., $\gamma_{t}=\frac{2}{2+t}$. Then, it holds that, for all $t\geq 1$,
$f(x_{t})-f^{*}\leq\frac{2^{1+\nu}LD^{2}}{(t+2)^{\nu}},$ (4.1)
and hence for any accuracy $\epsilon>0$ we have $f(x_{t})-f^{*}\leq\epsilon$ for all $t\geq\left(\frac{2^{1+\nu}LD^{2}}{\epsilon}\right)^{\frac{1}{\nu}}$.
Proof.
For some $\nu\in(0,1]$, we have
$f(x_{t})-f(x_{t+1})$ $\geq\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle-L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t})$
$\geq\gamma_{t}(f(x_{t})-f^{*})-L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t}),$
where the first inequality holds because of Lemma 2.8 and the last inequality holds because of Lemma 3.2. Subtracting $f^{*}$ on both sides, using $D_{\phi}(v_{t},x_{t})\leq D^{2}$, and rearranging leads to
$f(x_{t+1})-f^{*}\leq(1-\gamma_{t})(f(x_{t})-f^{*})+L\gamma_{t}^{1+\nu}D^{2}.$
When $t=0$, it follows $f(x_{1})-f^{*}\leq LD^{2}\leq 2LD^{2}$. Now, we consider $t\geq 1$ and obtain
$f(x_{t+1})-f^{*}$ $\leq(1-\gamma_{t})(f(x_{t})-f^{*})+L\gamma_{t}^{1+\nu}D^{2}$
$\leq\frac{t}{2+t}(f(x_{t})-f^{*})+\frac{2^{1+\nu}}{(2+t)^{1+\nu}}LD^{2}$
$\leq\frac{t}{2+t}\frac{2^{1+\nu}LD^{2}}{(2+t)^{\nu}}+\frac{2^{1+\nu}}{(2+t)^{1+\nu}}LD^{2}$
$=\frac{2^{1+\nu}LD^{2}}{(3+t)^{\nu}}\left(\frac{(3+t)^{\nu}(1+t)}{(2+t)^{1+\nu}}\right)\leq\frac{2^{1+\nu}LD^{2}}{(3+t)^{\nu}},$
where the last inequality holds due to $(3+t)^{\nu}(1+t)\leq(2+t)^{1+\nu}$ with $0<\nu\leq 1$. ∎
If $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$ in (4.1), we have
$f(x_{t})-f^{*}\leq\frac{4LD^{2}}{t+2},$
which is the same as a sublinear convergence rate of the classical FW algorithm.
Remark 4.3.
While the convergence rate (4.1) is the same as [Vyguzov2024-tp], Vyguzov and Stonyakin assume that the triangle scaling property holds for $D_{\phi}$. In contrast, we require significantly weaker assumptions: it is enough to assume that $\phi$ is strictly convex due to Lemma 2.5 in order to establish Theorem 4.2.
4.2 Accelerated Convergence
We will now establish accelerated convergence rates better than $O(1/t)$ up to linear convergence depending on the choice of parameters. First, we establish the accelerated convergence of Algorithm 1 with the Bregman short step-size, i.e.,
$\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},1\right\},$
in the case where the optimal solution lies in the relative interior.
Theorem 4.4 (Accelerated convergence for inner optima).
Suppose that Assumptions 3.1 and 4.1 hold. Let $f$ satisfy the Hölder error bound condition with $q\geq 1+\nu$ and $\mu>0$ and $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$. Assume that there exists a minimizer $x^{*}\in\operatorname{int}P$, i.e., there exists an $r>0$ with $B(x^{*},r)\subset P$. Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},1\right\}$. Then, it holds that
$f(x_{t})-f^{*}\leq\begin{cases}h_{t}\leq\max\left\{\frac{1}{1+\nu},1-\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{c^{1+1/\nu}D^{2/\nu}}\right\}^{t-1}LD^{2}&\text{if}\ q=1+\nu,\\ \frac{LD^{2}}{(1+\nu)^{t-1}}&\text{if}\ 1\leq t\leq t_{0},q>1+\nu,\\ \frac{(L(1+\nu)(c/r)^{1+\nu}D^{2})^{q/(q-1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}(t-t_{0})\right)^{\nu q/(q-1-\nu)}}=\mathcal{O}(1/t^{\nu q/(q-1-\nu)})&\text{if}\ t\geq t_{0},q>1+\nu,\end{cases}$
for all $t\geq 1$ where $c=(q/\mu)^{1/q}$ and
$t_{0}:=\max\left\{\left\lfloor\log_{\frac{1}{1+\nu}}\left(\frac{(L(1+\nu(c/r)^{1+\nu}D^{2})^{q/(q-1-\nu)}}{LD^{2}}\right)\right\rfloor+2,1\right\}.$
Proof.
Let $h_{t}:=f(x_{t})-f^{*}$. Using Lemma 3.3, we have
$h_{t}-h_{t+1}=f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\gamma_{t},$
where $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},1\right\}$. We consider two cases: (i) $\gamma_{t}<1$ and (ii) $\gamma_{t}=1$.
(i) $\gamma_{t}<1$: Using $\gamma_{t}=\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}}$, we have
$h_{t}-h_{t+1}$ $\geq\frac{\nu}{1+\nu}\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle^{1+1/\nu}}{(L(1+\nu)D_{\phi}(v_{t},x_{t}))^{1/\nu}}$
$\geq\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}\|\nabla f(x_{t})\|^{1+1/\nu}}{(L(1+\nu))^{1/\nu}D^{2/\nu}}$
$\geq\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}h_{t}^{\frac{(1+\nu)(q-1)}{\nu q}},$
where the second inequality holds from $\langle f(x_{t}),x_{t}-v_{t}\rangle\geq r\|\nabla f(x_{t})\|$ in Proposition 2.16, and the last inequality holds because $f$ satisfies the $c$-Łojasiewicz inequality with $c=(q/\mu)^{1/q}$ (see Remark 2.13), and $M:=(L(1+\nu))^{1/\nu}$.
(ii) $\gamma_{t}=1$: Using (3.1), we have
$h_{t}-h_{t+1}=f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq\frac{\nu}{1+\nu}(f(x_{t})-f^{*})=\frac{\nu}{1+\nu}h_{t},$
where the second inequality holds due to the convexity of $f$.
From (i) and (ii), we have
$h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu}h_{t},\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}h_{t}^{\frac{(1+\nu)(q-1)}{\nu q}}\right\}.$
When $q=1+\nu$, we have
$h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu},\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}\right\}\cdot h_{t}.$
This inequality and the initial bound $f(x_{1})-f^{*}\leq LD^{2}$ due to Lemma 2.7 imply
$h_{t}\leq\max\left\{\frac{1}{1+\nu},1-\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}\right\}^{t-1}LD^{2}.$
On the other hand, when $q>1+\nu$, we have
$h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu},\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}h_{t}^{\frac{q-1-\nu}{\nu q}}\right\}\cdot h_{t}.$
Using $f(x_{1})-f^{*}\leq LD^{2}$ and Lemma 2.20 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}$, $c_{2}=\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}$, and $\theta_{0}=\frac{q-1-\nu}{\nu q}>0$ from $q>1+\nu$, we have the claim. ∎
Because Algorithm 2 is well-defined (see Remark 3.4), the convergence result of the FW algorithm using the adaptive step-size strategy (Algorithm 2) is essentially the same as the one from Theorem 4.4 which uses Bregman short steps; up to small errors arising from the approximation whose precise analysis we skip for the sake of brevity. Note that Vyguzov and Stonyakin [Vyguzov2024-tp, Theorem 2] also established linear convergence, assuming that $f$ is relatively strongly convex, the optimal solution belongs to int $P$, the angle condition holds, and the triangle scaling property holds. However, these requirements are stronger than ours here, and they do not establish sublinear but accelerated rates for intermediate parameter regimes.
Next, we establish the linear convergence of Algorithm 3. Recall that the pyramidal width is the minimal $\delta>0$ satisfying Lemma 2.18 (see [Braun2022-ew, Lemma 2.26] or [Lacoste-Julien2015-gb, Theorem 3] for an in-depth discussion). We make the following assumption:
Assumption 4.5.
The kernel generating distance $\phi$ is $\sigma$-strongly convex.
Under Assumption 4.5, it holds that $\delta\leq D$, and this will be important in establishing convergence rates of Algorithm 3.
Theorem 4.6 (Accelerated convergence of the away-step FW algorithm).
Suppose that Assumptions 3.14.1, and 4.5 hold. Let $P\subset\mathbb{R}^{n}$ be a polytope and $f$ satisfy the Hölder error bound condition with $q>1+\nu$ or $(\nu,q)=(1,2)$. The convergence rate of Algorithm 3 is linear: for all $t\geq 1$
$f(x_{t})-f^{*}\leq\begin{cases}\left(1-\frac{\mu}{32L}\frac{\delta^{2}}{D^{2}}\right)^{\lceil(t-1)/2\rceil}LD^{2}&\text{if}\ (\nu,q)=(1,2),\\ \frac{1}{(1+\nu)^{\lceil(t-1)/2\rceil}}LD^{2}&\text{if}\ 1\leq t\leq t_{0},q>1+\nu,\\ \frac{(L(1+\nu)D^{2}/((\mu/q)^{1/q}\delta/2)^{1+\nu})^{q/(q-1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}\lceil(t-t_{0})/2\rceil\right)^{\nu q/(q-1-\nu)}}=\mathcal{O}(1/t^{\nu q/(q-1-\nu)})&\text{if}\ t\geq t_{0},q>1+\nu,\end{cases}$
where $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ and $\delta$ are the diameter and the pyramidal width of the polytope $P$, respectively, and
$t_{0}:=\max\left\{\left\lfloor\log_{\frac{1}{1+\nu}}\frac{(L(1+\nu)D^{2}/((\mu/q)^{1/q}\delta/2)^{1+\nu})^{q/(q-1-\nu)}}{LD^{2}}\right\rfloor+2,1\right\}.$
Proof.
By using the induced guarantee on the primal gap via Lemma 2.19, we have
$h_{t}=f(x_{t})-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x_{t}),v_{t}^{\textup{A}}-v_{t}^{\textup{FW}}\rangle}{\delta}\right)^{\frac{q}{q-1}}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{2\langle\nabla f(x_{t}),d_{t}\rangle}{\delta}\right)^{\frac{q}{q-1}},$ (4.2)
where the last inequality holds because $d_{t}$ is either $x_{t}-v_{t}^{\textup{FW}}$ or $v_{t}^{\textup{A}}-x_{t}$ with $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle/2$ in Lines 3 and 3 of Algorithm 3. We obtain
$h_{t}-h_{t+1}$ $\geq\frac{\nu}{1+\nu}\langle\nabla f(x_{t}),d_{t}\rangle\min\left\{\gamma_{t,\max},\left(\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}}\right\}$
$=\min\left\{\gamma_{t,\max}\frac{\nu}{1+\nu}\langle\nabla f(x_{t}),d_{t}\rangle,\frac{\nu}{1+\nu}\frac{\langle\nabla f(x_{t}),d_{t}\rangle^{1+1/\nu}}{(L(1+\nu)D_{\phi}(v_{t},x_{t}))^{1/\nu}}\right\}$
$\geq\min\left\{\gamma_{t,\max}\frac{\nu h_{t}}{1+\nu},\frac{\nu}{1+\nu}\frac{((\mu/q)^{1/q}h_{t}^{1-1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}\right\}$
$=\min\left\{\frac{\nu h_{t}}{1+\nu}\gamma_{t,\max},\frac{\nu}{1+\nu}\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}h_{t}^{\frac{(1+\nu)(q-1)}{\nu q}}\right\}$
where the first inequality holds due to Lemma 3.3 with $\gamma_{t,\max}=1$ (Frank–Wolfe steps) and $\gamma_{t,\max}=\frac{\lambda_{v_{t}^{\textup{A}},t}}{1-\lambda_{v_{t}^{\textup{A}},t}}$ (away steps), and the second inequality holds because of (4.2) and $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq h_{t}$.
In $(\nu,q)=(1,2)$, for Frank–Wolfe steps, $\gamma_{t,\max}=1\geq\mu\delta^{2}/LD^{2}\geq\mu\delta^{2}/32LD^{2}$. For away steps, we only rely on monotone progress $h_{t+1}<h_{t}$ because it is difficult to estimate $\gamma_{t,\max}$ below. However, $\gamma_{t,\max}$ cannot be small too often, which is the key point. Let us consider $\gamma_{t}=\gamma_{t,\max}$ in an away step. In that case, $v^{\textup{A}}_{t}$ is removed from the active set $\mathcal{S}_{t+1}$ in line 3. Moreover, the active set can only grow in Frank–Wolfe steps (line 3). It is impossible to remove more vertices from $\mathcal{S}_{t+1}$ than have been added in Frank–Wolfe steps. Therefore, at most half of iterations until $t$ iterations are in away steps, i.e., in all other steps, we have $\gamma_{t}=\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{2LD_{\phi}(v_{t},x_{t})}<\gamma_{t,\max}$ and then $h_{t}-h_{t+1}\geq\frac{\mu\delta^{2}}{32LD^{2}}h_{t}$. Because we have $h_{t+1}\leq\left(1-\frac{\mu\delta^{2}}{32LD^{2}}\right)h_{t}$ for at least half of the iterations and $h_{t+1}\leq h_{t}$ for the rest, we obtain
$f(x_{t})-f^{*}\leq\left(1-\frac{\mu}{32L}\frac{\delta^{2}}{D^{2}}\right)^{\lceil(t-1)/2\rceil}LD^{2}.$
In $q>1+\nu$, we have $h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu},\frac{\nu}{1+\nu}\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}h_{t}^{\frac{q-1-\nu}{\nu q}}\right\}\cdot h_{t}$ for at least half of the iterations (Frank–Wolfe steps) and $h_{t+1}\leq h_{t}$ for the rest (away steps). The initial bound $f(x_{1})-f^{*}\leq LD^{2}$ holds generally for the Frank–Wolfe algorithm (see the proof of Theorem 4.2 and [Braun2022-ew, Remark 2.4]). We use Lemma 2.20 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}$, $c_{2}=c_{1}\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}$, and $\theta_{0}=\frac{q-1-\nu}{\nu q}>0$ from $q>1+\nu$ and obtain the claim. ∎
Remark 4.7 (Compatibility of parameters).
The condition $q>1+\nu$ is necessary in case that $(\nu,q)\neq(1,2)$. The reason is as follows. In the case $q=1+\nu$, for Frank–Wolfe steps, we have
$\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}=\left(\frac{\mu}{L}\right)^{1/\nu}\left(\frac{\delta}{D}\right)^{\frac{1+\nu}{\nu}}\frac{1}{2^{\frac{1+\nu}{\nu}}(1+\nu)^{2/\nu}D^{(1-\nu)/\nu}},$ (4.3)
which might be greater than 1 because $\frac{1}{D^{(1-\nu)/\nu}}\geq 1$ if $D<1$ and $\nu$ are small enough. In order to make (4.3) smaller than 1, $\nu$ should be 1, i.e., $(\nu,q)=(1,2)$. When $\nu=1$ and $q\neq 2$, it reduces to the $q>1+\nu$ case.
5 Convergence Analysis for Nonconvex Optimization
In this section, we consider a nonconvex objective function. We establish global sublinear convergence to a stationary point of (1.1), i.e., $x\in P$ such that $\max_{v\in P}\langle\nabla f(x),x-v\rangle=0$. We also obtain local sublinear and local linear convergence to a minimizer of (1.1).
5.1 Global Convergence
We show that Algorithm 1 with $\gamma_{t}=\gamma:=1/(T+1)^{\frac{1}{1+\nu}}$ globally converges to a stationary point, where $T\in\mathbb{N}$ is the number of iterations. Its proof is inspired by [Lacoste-Julien2016-mj, Theorem 1] and [Pokutta2024-qu, Theorem 4.7] and identical to those for the case when $\phi=\frac{1}{2}\|\cdot\|^{2}$.
Theorem 5.1 (Global sublinear convergence for nonconvex optimization).
Suppose that Assumption 3.1 holds. Let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$ characterized by $D_{\phi}$ and let $T\in\mathbb{N}$. Then, the iterates of the FW algorithm with $\gamma_{t}=\gamma:=1/(T+1)^{\frac{1}{1+\nu}}$ satisfy
$G_{T}:=\min_{0\leq t\leq T}\max_{v_{t}\in P}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\leq\frac{2\max\{h_{0},LD^{2}\}}{(T+1)^{\frac{\nu}{1+\nu}}},$
where $h_{0}=f(x_{0})-f^{*}$ is the primal gap at $x_{0}$.
Proof.
Substituting $x_{t+1}$ for $x_{+}$ and $x_{t}$ for $x$ in Lemma 2.8, we have the following inequality:
$f(x_{t})-f(x_{t+1})$ $\geq\gamma\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle-L\gamma^{1+\nu}D_{\phi}(v_{t},x_{t}).$
Summing up the above inequality along $t=1,\ldots,T$ and rearranging provides
$\gamma\sum_{t=0}^{T}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle$ $\leq f(x_{0})-f(x_{T+1})+\gamma^{1+\nu}\sum_{t=0}^{T}LD_{\phi}(x_{t},v_{t})$
$\leq f(x_{0})-f^{*}+\gamma^{1+\nu}\sum_{t=0}^{T}LD^{2}=h_{0}+\gamma^{1+\nu}(T+1)LD^{2}.$
Dividing by $\gamma(T+1)$ on the both sides, we obtain
$G_{T}\leq\frac{1}{T+1}\sum_{t=0}^{T}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\leq\frac{h_{0}}{\gamma(T+1)}+\gamma^{\nu}LD^{2},$
which, for $\gamma=1/(T+1)^{\frac{1}{1+\nu}}$, implies
$G_{T}\leq\frac{1}{T+1}\sum_{t=0}^{T}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\leq(h_{0}+LD^{2})(T+1)^{-\frac{\nu}{1+\nu}}\leq 2\max\{h_{0},LD^{2}\}(T+1)^{-\frac{\nu}{1+\nu}}.$
This is the desired claim. ∎
As mentioned above, we generalize prior similar results. In fact, in the case where $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$, we have $\gamma_{t}=\frac{1}{\sqrt{T+1}}$ and obtain as guarantee
$\min_{0\leq t\leq T}\max_{v_{t}\in P}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\leq\frac{2\max\{h_{0},LD^{2}\}}{\sqrt{T+1}},$
which is the same rate as [Lacoste-Julien2016-mj, Theorem 1].
5.2 Local Convergence
Next, we show that the FW algorithm converges to a minimizer $x^{*}\in\operatorname*{argmin}_{x\in P}f(x)$ when an initial point is close enough to $x^{*}$. We need the weak convexity of $f$ for its proof.
Assumption 5.2.
  1. (i)
    $f$ is $\rho$-weakly convex.
  2. (ii)
    $f$ is local $\mu$-quadratic growth with $\zeta>0$.
Under Assumption 5.2(i), the following primal gap lemma holds, whose proof is immediate.
Lemma 5.3 (Primal gap, dual gap, and Frank–Wolfe gap for weakly convex functions).
Suppose that Assumptions 3.1(ii)(iv), and 5.2(i). Let $x^{*}\in P$ an optimal solution of (1.1) and let $f^{*}=f(x^{*})$. For all $x\in P$, it holds:
$f(x)-f^{*}$ $\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}$
$\leq\max_{v\in P}\langle\nabla f(x),x-v\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}.$ (5.1)
Proof.
The first inequality follows from the weak convexity of $f$ and the second follows from the maximality of $\langle\nabla f(x),x-v\rangle$. ∎
Note that Lemma 5.3 differs from Lemma 3.2 in the additional term $\frac{\rho}{2}\|x-x^{*}\|^{2}$. Moreover, we prove a lemma used in the proof of convergence analysis.
Lemma 5.4.
For any $\nu\in(0,1]$ and $t\geq 0$, it holds that
$h(t):=\frac{(t+3)^{\nu}(t+2-\nu)}{(t+2)^{1+\nu}}\leq 1.$ (5.2)
Proof.
We have
$\lim_{t\to\infty}h(t)=\lim_{t\to\infty}\frac{(t+3)^{\nu}(t+2-\nu)}{(t+2)^{1+\nu}}=\lim_{t\to\infty}\left(1+\frac{1}{t+2}\right)^{\nu}\left(1-\frac{\nu}{t+2}\right)=1.$
We take the logarithm of $h(t)$, that is, $\log h(t)=\nu\log(t+3)+\log(t+2-\nu)-(1+\nu)\log(t+2)$ and obtain
$\frac{h^{\prime}(t)}{h(t)}$ $=\frac{\nu}{t+3}+\frac{1}{t+2-\nu}-\frac{1+\nu}{t+2}=\frac{-2\nu^{2}+10\nu}{(t+2)(t+3)(t+2-\nu)}>0,$
where the last inequality holds for $\nu\in(0,1]$ and $t\geq 0$. Since $h(t)>0$ for $t\geq 0$, the above inequality implies $h^{\prime}(t)>0$. Therefore, we have $\sup h(t)=1$, which implies $h(t)\leq 1$. ∎
Now we show sublinear convergence with $\gamma_{t}=\frac{2}{2+t}$. The proof is a modified version of Theorem 4.2.
Theorem 5.5 (Local sublinear convergence).
Suppose that Assumptions 3.14.5, and 5.2 hold. Let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$ characterized by the Bregman distance $D_{\phi}$. Consider the iterates of Algorithm 1 with the open loop step-size, i.e., $\gamma_{t}=\frac{2}{2+t}$. Then, if $\rho/\sigma\leq L$ and $\frac{3\rho}{\mu}\leq 2-\nu$ hold, it holds that, for all $t\geq 1$,
$f(x_{t})-f^{*}\leq\frac{2^{1+\nu}\mu LD^{2}}{\rho(t+2)^{\nu}},$ (5.3)
and hence for any accuracy $\epsilon>0$ we have $f(x_{t})-f^{*}\leq\epsilon$ for all $t\geq\left(\frac{2^{1+\nu}\mu LD^{2}}{\rho\epsilon}\right)^{\frac{1}{\nu}}$.
Proof.
Using Lemma 2.8, we have
$f(x_{t})-f(x_{t+1})$ $\geq\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle-L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t})$
$\geq\gamma_{t}\left(f(x_{t})-f^{*}-\frac{\rho}{2}\|x_{t}-x^{*}\|^{2}\right)-L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t}),$
where the last inequality holds because of (5.1) in Lemma 5.3. Subtracting $f^{*}$ and rearranging provides
$f(x_{t+1})-f^{*}$ $\leq(1-\gamma_{t})(f(x_{t})-f^{*})+\frac{\rho\gamma_{t}}{2}\|x_{t}-x^{*}\|^{2}+L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t})$ (5.4)
$\leq\left(1-\left(1-\frac{\rho}{\mu}\right)\gamma_{t}\right)(f(x_{t})-f^{*})+L\gamma_{t}^{1+\nu}D^{2},$ (5.5)
where the last inequality holds due to the quadratic growth condition. For $t=0$, using (5.4), we have
$f(x_{1})-f^{*}\leq\frac{\rho}{2}\|x_{0}-x^{*}\|^{2}+LD_{\phi}(v_{0},x_{0})\leq\left(\frac{\rho}{\sigma}+L\right)D^{2}\leq 2LD^{2}\leq\frac{2\mu}{\rho}LD^{2},$ (5.6)
where the second inequality holds because $\phi$ is $\sigma$-strongly convex and the last inequality holds because of $3<6/(2-\nu)\leq 2\mu/\rho$. Now we consider $t\geq 1$. Using (5.5) and $\gamma_{t}=\frac{2}{2+t}$, we have
$f(x_{t+1})-f^{*}$ $\leq\left(t+\frac{2\rho}{\mu}\right)\frac{2^{1+\nu}\mu LD^{2}}{\rho(t+2)^{1+\nu}}+\frac{2^{1+\nu}LD^{2}}{(t+2)^{1+\nu}}$
$=\left(t+\frac{2\rho}{\mu}+\frac{\rho}{\mu}\right)\frac{\mu}{\rho}\frac{2^{1+\nu}LD^{2}}{(t+2)^{1+\nu}}$
$=\frac{\mu}{\rho}\frac{2^{1+\nu}LD^{2}}{(t+3)^{\nu}}\left(\frac{(t+3)^{\nu}(t+3\rho/\mu)}{(t+2)^{1+\nu}}\right)$
$\leq\frac{2^{1+\nu}\mu LD^{2}}{\rho(t+3)^{\nu}},$
where the last inequality holds because of $(t+3)^{\nu}(t+3\rho/\mu)\leq(t+3)^{\nu}(t+2-\nu)\leq(t+2)^{1+\nu}$ from (5.2). ∎
We can apply the quadratic growth condition again as before and obtain
$\operatorname{dist}(x_{t},\mathcal{X}^{*})^{2}\leq\frac{2}{\mu}(f(x_{t})-f^{*})\leq\frac{2^{2+\nu}LD^{2}}{\rho(t+2)^{\nu}},$
and hence
$\operatorname{dist}(x_{t},\mathcal{X}^{*})\leq\frac{2^{1+\nu/2}D\sqrt{L}}{\sqrt{\rho}(t+2)^{\nu/2}}.$
Note that Theorem 5.5 does not require knowledge of the number of iterations $T$ ahead of time compared to Theorem 5.1. We stress, nonetheless, that the latter can also be adjusted using a different step-size strategy to obtain an any-time guarantee; see [Braun2022-ew] for details for the standard Euclidean case, which can be generalized to our setup. Moreover, we can apply Theorem 5.5 to the classical FW algorithm, i.e., $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$. Using (5.3), $D_{\textup{Euc}}:=\sup_{x,y\in P}\|x-y\|^{2}$, and $D_{\textup{Euc}}=\sqrt{2}D$, we obtain
$f(x_{t})-f^{*}\leq\frac{2\mu LD_{\textup{Euc}}^{2}}{\rho(t+2)},$
Theorem 5.5 requires $\rho/\sigma\leq L$ and $\frac{3\rho}{\mu}\leq 2-\nu$. These assumptions are easy to satisfy.
Example 5.6 (Example 3.1 in the arXiv version of [Liao2024-vd]).
Let us consider
$f(x)=\begin{cases}-x^{2}+1,&\text{if}\ -1<x<-0.5,\\ 3(x+1)^{2},&\text{otherwise}.\end{cases}$
The function $f$ is not convex but $\rho$-weakly convex with $\rho=2$. A global optimal solution of $f$ is $x^{*}=-1$ and its value is $f(x^{*})=0$. Moreover, $f$ has the quadratic growth property with $0<\mu\leq 6$. Because $f$ is a quadratic function, we have $L\geq 6$. It holds that $2=\rho\leq L$ and $1=\frac{3\rho}{\mu}\leq 1<2-\nu$ (set $\mu=6$). Therefore, the assumption of Theorem 5.5 holds.
In order to verify $\frac{3\rho}{\mu}\leq 2-\nu$, it is easier to examine the sufficient condition $\frac{3\rho}{\mu}\leq 1$ instead because the exact value of $\nu$ is difficult to estimate. On the other hand, without loss of generality, we can assume $\sigma=1$. When $\sigma>1$, $\rho/\sigma<\rho\leq L$ holds. When $\sigma<1$, we can use $\phi_{1}=\phi+\frac{1-\sigma}{2}\|\cdot\|^{2}$, which is 1-strongly convex. When $\phi$ is convex but not strongly convex, we can use $\phi_{2}=\phi+\frac{1}{2}\|\cdot\|^{2}$. Therefore, it suffices to verify $\rho/\sigma\leq\rho\leq L$, which often holds (for example, see an example of phase retrieval in Section 6.3). Next we establish local accelerated convergence with the short step step-size, i.e., $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},1\right\}$.
Theorem 5.7 (Local accelerated convergence for inner optima).
Suppose that Assumptions 3.1 and 5.2 hold. Let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter. Assume that there exists a minimizer $x^{*}\in\operatorname{int}P$, i.e., there exists an $r>0$ with $B(x^{*},r)\subset P$. Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},1\right\}$. Then, if $\rho/\mu<1$, it holds that, for all $t\geq 1$,
$f(x_{t})-f^{*}\leq\begin{cases}\max\left\{\frac{1}{2}\left(1+\frac{\rho}{\mu}\right),1-\frac{r^{2}}{2Mc^{2}D^{2}}\right\}^{t-1}LD^{2}&\text{if}\ \nu=1,\\ \left(\frac{1}{1+\nu}\left(1+\frac{\nu\rho}{\mu}\right)\right)^{t-1}LD^{2}&\text{if}\ 1\leq t\leq t_{0},\nu\in(0,1),\\ \frac{(L(1+\nu)(1-\rho/\mu)^{\nu}c^{1+\nu}D^{2}/r^{1+\nu})^{2/(1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}\left(1-\frac{\rho}{\mu}\right)(t-t_{0})\right)^{2\nu/(1-\nu)}}=\mathcal{O}(1/t^{2\nu/(1-\nu)})&\text{if}\ t\geq t_{0},\nu\in(0,1),\end{cases}$
where $c=\frac{\sqrt{2\mu}}{\mu-\rho}$, $M:=(L(1+\nu))^{1/\nu}$, and
$t_{0}:=\max\left\{\left\lfloor\log_{\frac{1}{1+\nu}\left(1+\frac{\nu\rho}{\mu}\right)}\left(\frac{(L(1+\nu)(1-\rho/\mu)^{\nu}c^{1+\nu}D^{2}/r^{1+\nu})^{2/(1-\nu)}}{LD^{2}}\right)\right\rfloor+2,1\right\}.$
Proof.
Let $h_{t}:=f(x_{t})-f^{*}$. Using Lemma 3.3, we have
$h_{t}-h_{t+1}=f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\gamma_{t},$
where $\gamma_{t}=\min\left\{1,\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}}\right\}$. We consider two cases: (i) $\gamma_{t}<1$ and (ii) $\gamma_{t}=1$.
(i) $\gamma_{t}<1$: We have
$h_{t}-h_{t+1}$ $\geq\frac{\nu\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{1+\nu}\min\left\{1,\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu}\right\}$
$\geq\frac{\nu}{1+\nu}\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle^{1+1/\nu}}{(L(1+\nu)D_{\phi}(v_{t},x_{t}))^{1/\nu}}$
$\geq\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}\|\nabla f(x_{t})\|^{1+1/\nu}}{(L(1+\nu))^{1/\nu}D^{2/\nu}}$
$\geq\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}h_{t}^{\frac{1+\nu}{2\nu}},$
where the third inequality holds due to Proposition 2.17 and the definition of $D$ and the last inequality holds due to the local PL inequality from Remark 2.15 with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$ and $M:=(L(1+\nu))^{1/\nu}$.
(ii) In the case where $\gamma_{t}=1$, i.e., $\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}}\geq 1$, this implies
$\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq L(1+\nu)D_{\phi}(v_{t},x_{t}),$ (5.7)
because of $11/\nu>$ 1. Using Lemma 2.8 and (5.7), we have
$h_{t+1}-h_{t}$ $\leq LD_{\phi}(v_{t},x_{t})-\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle$
$\leq-\frac{\nu}{1+\nu}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle$
$\leq-\frac{\nu}{1+\nu}\left(h_{t}-\frac{\rho}{2}\|x_{t}-x^{*}\|^{2}\right)$
$\leq-\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right)h_{t},$
where the third inequality holds because of (5.1) in Lemma 5.3, and the last inequality holds because of the local quadratic growth condition of $f$ with $\mu>0$. Therefore, we have
$h_{t+1}\leq\frac{1}{1+\nu}\left(1+\frac{\rho\nu}{\mu}\right)h_{t}.$
From (i) and (ii), we have
$h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right)h_{t},\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}h_{t}^{\frac{1+\nu}{2\nu}}\right\}.$
When $\nu=1$, we have $h_{t}-h_{t+1}\geq\min\left\{\frac{1}{2}\left(1-\frac{\rho}{\mu}\right),\frac{r^{2}}{2Mc^{2}D^{2}}\right\}\cdot h_{t}$. This inequality and the initial bound $f(x_{1})-f^{*}\leq LD^{2}$ due to Lemma 2.7 imply
$h_{t}\leq\max\left\{\frac{1}{2}\left(1+\frac{\rho}{\mu}\right),1-\frac{r^{2}}{2Mc^{2}D^{2}}\right\}^{t-1}LD^{2}.$
On the other hand, when $\nu\in(0,1)$,
$h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right),\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}h_{t}^{\frac{1-\nu}{2\nu}}\right\}\cdot h_{t}.$
Using $f(x_{1})-f^{*}\leq LD^{2}$ and Lemma 2.20 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}$, and $\theta_{0}=\frac{1-\nu}{2\nu}>0$, we have the claim. ∎
When $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$, we have local linear convergence with $\gamma_{t}=\min\left\{\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L\|v_{t}-x_{t}\|^{2}},1\right\}$ from Theorem 5.7. When $\phi=\frac{1}{2}\|\cdot\|$, i.e., $D_{\phi}(x,y)=\frac{1}{2}\|x-y\|^{2}$, $D_{\textup{Euc}}=\sqrt{2}D$ provides
$f(x_{t})-f^{*}\leq\max\left\{\frac{1}{2}\left(1+\frac{\rho}{\mu}\right),1-\frac{r^{2}}{2Lc^{2}D_{\textup{Euc}}^{2}}\right\}^{t-1}\frac{LD_{\textup{Euc}}^{2}}{2}.$
Because Algorithm 2 is well-defined (see also Remark 3.4), the convergence result of the FW algorithm with Algorithm 2 is the same as Theorem 5.7.
Finally, we establish the local linear convergence of Algorithm 3. In the same way as Theorem 4.6, the pyramidal width is the minimal $\delta>0$ satisfying Lemma 2.18 (see also [Braun2022-ew, Lemma 2.26] or [Lacoste-Julien2015-gb, Theorem 3]). An upper bound exists on the primal gap for weakly convex functions. Its proof can be found in Appendix A.1.
Lemma 5.8 (Upper bound on primal gap for weakly convex functions).
Suppose that Assumptions 3.1 and 5.2 hold. Let $P$ be a polytope with the pyramidal width $\delta>0$. Let $\mathcal{S}$ denote any set of vertices of $P$ with $x\in\operatorname{conv}\mathcal{S}$. Let $\psi$ be any vector, so that we define $v^{\textup{FW}}=\operatorname*{argmin}_{v\in P}\langle\psi,v\rangle$ and $v^{\textup{A}}=\operatorname*{argmax}_{v\in\mathcal{S}}\langle\psi,v\rangle$. If $\rho/\mu<1$, then it holds that, for all $x\in[f\leq f^{*}+\zeta]\cap P$,
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}\delta^{2}}\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}.$ (5.8)
We are ready to prove the local linear convergence of Algorithm 3 for nonconvex optimization. We also assume $\rho\leq L$, which is not restrictive (see also the discussion of Theorem 5.5 and Example 5.6).
Theorem 5.9 (Local linear convergence by the away-step FW algorithm).
Suppose that Assumptions 3.14.5, and 5.2 hold. Let $P\subset\mathbb{R}^{n}$ be a polytope. The convergence rate of Algorithm 3 with $f$ is linear: if $\rho<\mu\leq L$, for all $t\geq 1$
$f(x_{t})-f^{*}\leq\begin{cases}2\left(1-\frac{\omega}{4L}\frac{\delta^{2}}{D^{2}}\right)^{\lceil(t-1)/2\rceil}LD^{2}&\text{if}\ \nu=1,\\ 2\left(\frac{1}{1+\nu}\left(1+\frac{\nu\rho}{\mu}\right)\right)^{\lceil(t-1)/2\rceil}LD^{2}&\text{if}\ 1\leq t\leq t_{0},\nu\in(0,1),\\ \frac{(L^{2}(1+\nu)^{2}D^{4}(1-\rho/\mu)^{2\nu}/(\omega\delta^{2})^{1+\nu})^{1/(1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}\left(1-\frac{\rho}{\mu}\right)\lceil(t-t_{0})/2\rceil\right)^{2\nu/(1-\nu)}}=\mathcal{O}(1/t^{2\nu/(1-\nu)})&\text{if}\ t\geq t_{0},\nu\in(0,1),\end{cases}$
where $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ and $\delta$ are the diameter and the pyramidal width of the polytope $P$, respectively, and $\omega:=\frac{(\mu-\rho)^{2}}{8\mu}$, and
$t_{0}:=\max\left\{\left\lfloor\log_{\frac{1}{1+\nu}\left(1+\frac{\nu\rho}{\mu}\right)}\frac{(L^{2}(1+\nu)^{2}D^{4}(1-\rho/\mu)^{2\nu}/(\omega\delta^{2})^{1+\nu})^{1/(1-\nu)}}{2LD^{2}}\right\rfloor+2,1\right\}.$
Proof.
By letting $h_{t}=f(x_{t})-f^{*}$ and using Lemma 5.8, we have
$h_{t}\leq\frac{2\mu}{(\mu-\rho)^{2}\delta^{2}}\langle\nabla f(x_{t}),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}\leq\frac{8\mu\langle\nabla f(x_{t}),d_{t}\rangle^{2}}{(\mu-\rho)^{2}\delta^{2}}=\frac{\langle\nabla f(x_{t}),d_{t}\rangle^{2}}{\omega\delta^{2}},$ (5.9)
where the second inequality holds because $d_{t}$ is either $x_{t}-v_{t}^{\textup{FW}}$ or $v_{t}^{\textup{A}}-x_{t}$ with $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle/2$ in Lines 3 and 3 of Algorithm 3. From Lemma 3.3 with $\gamma_{t,\max}=1$ (Frank–Wolfe steps) and $\gamma_{t,\max}=\frac{\lambda_{v_{t}^{\textup{A}},t}}{1-\lambda_{v_{t}^{\textup{A}},t}}$ (away steps), we obtain
$h_{t}-h_{t+1}$ $\geq\frac{\nu}{1+\nu}\langle\nabla f(x_{t}),d_{t}\rangle\min\left\{\gamma_{t,\max},\left(\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}}\right\}$
$=\min\left\{\gamma_{t,\max}\frac{\nu}{1+\nu}\langle\nabla f(x_{t}),d_{t}\rangle,\frac{\nu}{1+\nu}\frac{\langle\nabla f(x_{t}),d_{t}\rangle^{1+1/\nu}}{(L(1+\nu)D_{\phi}(v_{t},x_{t}))^{1/\nu}}\right\}$
$\geq\min\left\{\frac{\nu\gamma_{t,\max}}{1+\nu}\left(1-\frac{\rho}{\mu}\right)h_{t},\frac{\nu}{1+\nu}\frac{(h_{t}\omega\delta^{2})^{\frac{1+\nu}{2\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}\right\}$
$=\min\left\{\frac{\nu\gamma_{t,\max}}{1+\nu}\left(1-\frac{\rho}{\mu}\right)h_{t},\frac{\nu}{1+\nu}\frac{(\omega\delta^{2})^{\frac{1+\nu}{2\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}h_{t}^{\frac{1+\nu}{2\nu}}\right\},$
where the second inequality holds because of $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq h_{t}-\frac{\rho}{2}\|x_{t}-v_{t}\|^{2}\geq\left(1-\frac{\rho}{\mu}\right)h_{t}$ (from Lemma 5.3 and the quadratic growth condition of $f$) and (5.9).
In $\nu=1$, for Frank–Wolfe steps, we have $\gamma_{t,\max}\left(1-\frac{\rho}{\mu}\right)=\left(1-\frac{\rho}{\mu}\right)\geq\frac{1}{8}\left(1-\frac{\rho}{\mu}\right)^{2}\frac{\delta^{2}}{D^{2}}=\frac{\omega\delta^{2}}{\mu D^{2}}\geq\frac{\omega\delta^{2}}{2LD^{2}}$ by $0<\rho<\mu\leq L$ and $\delta\leq D$. For away steps, it seems that we only obtain a monotone progress $h_{t+1}<h_{t}$ because it is difficult to estimate $\gamma_{t,\max}$ below. However, $\gamma_{t,\max}$ cannot be small too often. Let us consider $\gamma_{t}=\gamma_{t,\max}$ in an away step. In that case, $v^{\textup{A}}_{t}$ is removed from the active set $\mathcal{S}_{t+1}$ in line 3. Moreover, the active set can only grow in Frank–Wolfe steps (line 3). It is impossible to remove more vertices from $\mathcal{S}_{t+1}$ than have been added in Frank–Wolfe steps. Therefore, at most half of iterations until $t$ iterations are in away steps, i.e., in all other steps, we have $\gamma_{t}=\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{2LD_{\phi}(v_{t},x_{t})}<\gamma_{t,\max}$ and then $h_{t+1}\leq\left(1-\frac{\omega\delta^{2}}{4LD^{2}}\right)h_{t}$. Because we have $h_{t+1}\leq\left(1-\frac{\omega\delta^{2}}{4LD^{2}}\right)h_{t}$ for at least half of the iterations and $h_{t+1}\leq h_{t}$ for the rest, using $h_{1}\leq 2LD^{2}$ from (5.6) and $\rho<L$, we obtain
$f(x_{t})-f^{*}\leq 2\left(1-\frac{\omega}{4L}\frac{\delta^{2}}{D^{2}}\right)^{\lceil(t-1)/2\rceil}LD^{2}.$
In $\nu\in(0,1)$, we have $h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right),\frac{\nu}{1+\nu}\frac{(\omega\delta^{2})^{\frac{1+\nu}{2\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}h_{t}^{\frac{1-\nu}{2\nu}}\right\}\cdot h_{t}$ for at least half of the iterations and $h_{t+1}\leq h_{t}$ for the rest. The initial bound $f(x_{1})-f^{*}\leq(\rho+L)D^{2}\leq 2LD^{2}$ holds from (5.6) and $\rho<L$. We use Lemma 2.20 with $c_{0}=2LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=\frac{\nu}{1+\nu}\frac{(\omega\delta^{2})^{\frac{1+\nu}{2\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}$, and $\theta_{0}=\frac{1-\nu}{2\nu}$ and obtain the claim. ∎
When $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$, the local linear convergence of Algorithm 3 is equivalent to
$f(x_{t})-f(x^{*})\leq\left(1-\frac{\omega}{2L}\frac{\delta^{2}}{D_{\textup{Euc}}^{2}}\right)^{\lceil(t-1)/2\rceil}LD_{\textup{Euc}}^{2}.$
Without loss of generality, we set $\sigma=1$. If $\rho\leq L$ does not hold, we can use $h_{1}\leq(\rho+L)D^{2}$ as the initial bound.
Moreover, when $P$ is $\alpha$-strongly convex set and $\phi=\frac{1}{2}\|\cdot\|^{2}$, we also establish local linear convergence. Its proof can be found in Appendix A.2.
6 Numerical Experiments
In this section, we conducted numerical experiments to examine the performance of our proposed algorithms. All numerical experiments were performed in Julia 1.11 using the FrankWolfe.jl package [Besancon2022-lx] [^2] https://github.com/ZIB-IOL/FrankWolfe.jl on a MacBook Pro with an Apple M2 Max and 64GB LPDDR5 memory.
We compare our algorithm with the following algorithms and use the following notation:
  • BregFW: the FW algorithm with the adaptive Bregman step-size strategy (Algorithm 2, our proposed)
  • BregAFW: the away-step FW algorithm with the adaptive Bregman step-size strategy (Algorithm 3 using Algorithm 2, i.e., $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},\gamma_{t,\max})$, our proposed update)
  • EucFW: the FW algorithm with the adaptive (Euclidean) step-size strategy [Pedregosa2020-ay]
  • EucAFW: the away-step FW algorithm with the adaptive (Euclidean) step-size strategy [Pedregosa2020-ay]
  • ShortFW: the FW algorithm with the (Euclidean) short step
  • ShortAFW: the away-step FW algorithm with the (Euclidean) short step
  • OpenFW: the FW algorithm with the open loop with $\gamma_{t}=\frac{2}{2+t}$
  • OpenAFW: the away-step FW algorithm with the open loop with $\gamma_{t}=\frac{2}{2+t}$
  • MD: the mirror descent [Nemirovskij1983-sm]
Note that we included OpenAFW only for comparison purposes; there is no established convergence theory for AFW with open-loop strategies. In particular, there are no proper drop steps, and the favorable properties of the away-step Frank-Wolfe algorithm are lost; see [Braun2022-ew]. We use $\beta=0.9$, $\eta=0.9$, $\tau=2$, and $\gamma_{\max}=1$ throughout all numerical experiments.
6.1 Nonnegative Linear Inverse Problems
Given a nonnegative matrix $A\in\mathbb{R}_{+}^{m\times n}$ and a nonnegative vector $b\in\mathbb{R}_{+}^{m}$, the goal of nonnegative (Poisson) linear inverse problems is to recover a signal $x\in\mathbb{R}_{+}^{n}$ such that $Ax\simeq b$. This class of problems has been studied in image deblurring [Bertero2009-jq] and positron emission tomography [Vardi1985-de] as well as optimization [Bauschke2017-hg, Takahashi2024-wz]. Since the dimension of $x$ is often larger than the number of observations $m$, the system is indeterminate. From this point of view, we consider the constraint $\Delta_{n}:=\{x\in\mathbb{R}_{+}^{n}\,|\,\sum_{j=1}^{n}x_{j}\leq 1\}$. Recovering $x$ can be formulated as a minimization problem:
$\min_{x\in\Delta_{n}}\quad f(x):=d(Ax,b),$ (6.1)
where $d(x,y):=\sum_{i=1}^{m}\left(x_{i}\log\frac{x_{i}}{y_{i}}+y_{i}-x_{i}\right)$ is the Kullback–Leibler divergence (see also Example 2.3). Problem (6.1) is convex, while $\nabla f$ is not Lipschitz continuous on $\mathbb{R}_{+}^{n}$. The pair $(f,\phi)$ is $L$-smad on $\mathbb{R}_{+}^{n}$ with $\phi(x)=\sum_{j=1}^{n}x_{j}\log x_{j}$ and $L\geq\max_{1\leq j\leq n}\sum_{i=1}^{m}a_{ij}$ from [Bauschke2017-hg, Lemma 8].
We compared BregFW with EucFW, ShortFW, OpenFW, and the mirror descent algorithm (MD[Nemirovskij1983-sm]. The subproblem of MD can be solved in closed-form for $\{x\in\mathbb{R}_{+}^{n}\,|\,\sum_{j=1}^{n}x_{j}=1\}$ by [Beck2017-qc, Example 3.71], and this can be readily extended to $\Delta_{n}$. We used $1000$ as maximum iteration limit and we generated $\tilde{A}$ from an i.i.d. normal distribution and set $a_{ij}=|\tilde{a}_{ij}|/\sum_{i=1}^{m}|\tilde{a}_{ij}|$. We also generated $\tilde{x}$ from an i.i.d. uniform distribution in $[0,1]$ and set the ground truth $x^{*}=0.8\tilde{x}/\sum_{j=1}^{n}\tilde{x}_{j}$ so that $x^{*}\in\operatorname{int}\Delta_{n}$. All components of the initial point $x_{0}$ were $1/n$. For $(m,n)=(100,1000)$, Figure 1 shows the primal gap $f(x_{t})-f^{*}$ and the FW gap $\max_{v\in P}\langle\nabla f(x_{t}),x_{t}-v\rangle$ per iteration (left) and the primal gap per second (right). Table 2 shows average numbers of the primal gap, the FW gap, and computational time over 20 different instances for $(m,n)=(100,1000)$. Here, BregFW outperformed other algorithms, both in iterations and time.
Figure 1: Nonnegative linear inverse problem for $(m,n)=(100,1000)$.
Table 2: Average numbers of the primal gap, the FW gap, and computational time (s) over 20 different instances of nonnegative linear inverse problems for $(m,n)=(100,1000)$.
algorithm primal gap FW gap time (s)
BregFW 6.963691e-08 1.145520e-05 1.457719e-01
EucFW 3.028696e-07 2.922331e-05 1.077826e-01
ShortFW 4.747044e-05 4.613479e-04 6.251837e-02
OpenFW 4.957628e-07 8.538245e-05 6.503594e-02
MD 2.249368e-06 1.721359e-04 1.888499e-01
6.2 $\ell_{p}$oss Problem
We consider the $\ell_{p}$ loss problem [Kyng2015-uo, Maddison2021-nt] to find $x\in P$ such that $Ax\simeq b$, defined by
$\min_{x\in P}\quad\|Ax-b\|_{p}^{p},$ (6.2)
where $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, and $p>1$. Defining $f(x)=\|Ax-b\|_{p}^{p}$, we have
$\nabla f(x)$ $=p\sum_{i=1}^{m}|\langle a_{i},x\rangle-b_{i}|^{p-1}\operatorname{sgn}(\langle a_{i},x\rangle-b_{i}),$
$\nabla^{2}f(x)$ $=p(p-1)\sum_{i=1}^{m}|\langle a_{i},x\rangle-b_{i}|^{p-2}a_{i}a_{i}^{\mathsf{T}},$
where $a_{i}$ is the $i$th row vector of $A$. Problem (6.2) is convex, but $\nabla f$ is not Lipschitz continuous on $\mathbb{R}^{n}$ when $p\neq 2$. Furthermore, when $p<2$, $f$ is not $\mathcal{C}^{2}$ but $\mathcal{C}^{1}$. Therefore, for $21<p<$ 2, $\nabla f$ is also not Lipschitz continuous over compact sets. Since $f$ is convex, the pair $(f,\phi)$ is 1-smad with $\phi=f$.
We use an $\ell_{2}$ norm ball as a constraint, i.e., $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|_{2}\leq 1\}$. We compared BregFW with EucFW, ShortFW, and OpenFW. We generated $A$ from an i.i.d. normal distribution and normalized it so that $\|a_{i}\|=1$. We also generated $\tilde{x}$ from an i.i.d. normal distribution and set $x^{*}=0.8\tilde{x}/\|\tilde{x}\|$ to ensure $x^{*}\in\operatorname{int}P$. The initial point $x_{0}$ was generated by computing an extreme point of $P$ that minimizes the linear approximation of $f$. For $(n,m)=(100,100)$ and $p=1.1$, Figure 2 shows the primal gap $f(x_{t})-f^{*}$ and the FW gap $\max_{v\in P}\langle\nabla f(x_{t}),x_{t}-v\rangle$ per iteration and those gaps per second up to 1000 iterations. Table 3 also shows the average performance over 20 different instances. BregFW outperformed the other algorithms. Since $\nabla f$ is not Lipschitz continuous, ShortFW did not reduce the primal and FW gap.
Figure 2: The $\ell_{p}$ loss problem for $(m,n)=(1000,100)$.
Table 3: Average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of $\ell_{p}$ loss problems for $(m,n)=(1000,100)$.
algorithm primal gap FW gap time (s)
BregFW 1.056988e-13 3.764084e-01 5.600876e-01
EucFW 6.341301e-10 8.576574e-01 3.138045e-01
ShortFW 2.465054e+01 4.754413e+01 9.198800e-02
OpenFW 1.698968e-02 4.007941e+00 8.713342e-02
6.3 Phase Retrieval
We are interested in phase retrieval, which involves finding a signal $x\in\mathbb{R}^{n}$ such that
$|\langle a_{i},x\rangle|^{2}\simeq b_{i},\quad i=1,\ldots,m,$
where $a_{i}\in\mathbb{R}^{n}$ describes the model and $b\in\mathbb{R}^{m}$ is a vector of measurements. Phase retrieval has a long history. Patterson studied phase retrieval in X-ray crystallography in 1934 [Patterson1934-sg] and 1944 [Patterson1944-nm]. Phase retrieval arises in many fields of science and engineering, such as image processing [Candes2015-mk], X-ray crystallography [Patterson1934-sg, Patterson1944-nm], and optics [Shechtman2015-ic]. Bolte et al. [Bolte2018-zt] applied the Bregman proximal gradient algorithm to phase retrieval.
In applications, $a_{i}$ and $x$ are often complex vectors. Now, we consider real vectors. To achieve the goal of phase retrieval, we focus on the following nonconvex optimization problem:
$\min_{x\in P}\quad f(x):=\frac{1}{4}\sum_{i=1}^{m}\left(|\langle a_{i},x\rangle|^{2}-b_{i}\right)^{2},$
where $P\subset\mathbb{R}^{n}$ is a compact convex set. We introduce some properties of $f$. Let $\phi$ be defined by
$\phi(x)=\frac{1}{4}\|x\|^{4}+\frac{1}{2}\|x\|^{2},$
which is $\sigma$-strongly convex for $\sigma\leq 1$. For any $L$ satisfying
$L\geq\sum_{i=1}^{m}\left(3\|a_{i}\|^{4}+\|a_{i}\|^{2}|b_{i}|\right),$ (6.3)
the pair $(f,\phi)$ is $L$-smad on $\mathbb{R}^{n}$ [Bolte2018-zt, Lemma 5.1]. The right-hand side of (6.3) is relatively large in applications, while Takahashi et al. derived a smaller lower bound of $L$ [Takahashi2023-uh, Propositions 5 and 6]. In addition, because we have
$\nabla^{2}f(x)=\sum_{i=1}^{m}\left((3|\langle a_{i},x\rangle|^{2}-b_{i})a_{i}a_{i}^{\mathsf{T}}\right),$
the matrix $\nabla^{2}f(x)+\rho I$ is positive semidefinite for $\rho\geq\sum_{i=1}^{m}\|a_{i}\|^{2}|b_{i}|$, i.e., $f$ is $\rho$-weakly convex. For example, $L=\sum_{i=1}^{m}\left(3\|a_{i}\|^{4}+\|a_{i}\|^{2}|b_{i}|\right)\geq\sum_{i=1}^{m}\|a_{i}\|^{2}|b_{i}|=\rho/\sigma$ holds with $\sigma=1$.
We use a $K$-sparse polytope as a constraint, i.e., $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|_{1}\leq K,\|x\|_{\infty}\leq 1\}$. We compared BregFW with EucFW, ShortFW, and OpenFW. The maximum iteration is 1000. We generated $a_{i}$, $i=1,\ldots,m$ from an i.i.d. normal distribution and normalized it to have norm 1. We also generated $x^{*}$ from an i.i.d. uniform distribution in $[0,1]$ and normalized $x^{*}$ to have sum 1. The initial point $x_{0}$ was generated by computing an extreme point of $P$ that minimizes the linear approximation of $f$. For $(m,n)=(100,2000)$ and $K=200$, Figure 3 shows the primal and FW gaps per iteration and gaps per second. Because $\nabla f$ is not Lipschitz continuous, ShortFW does not converge, and EucFW performs slowly. The primal gap and the FW gap by the Bregman adaptive step-size strategy are the smallest among these step-size strategies. Table 4 shows the average performance over 20 different instances of $(m,n)=(100,2000)$. We compared BregAFW with EucAFW, ShortAFW, and OpenAFW. In only this setting, we generated $x^{*}$ from an i.i.d. uniform distribution in $[0,1]$ and did not normalize it; that is, $x^{*}$ might be in the face of $P$. Figure 4 shows another setting’s results for $(m,n)=(200,200)$ and $K=110$. Table 5 shows the average performance over 20 different instances for $(m,n)=(200,200)$ and $K=110$. The primal gap by BregAFW is the smallest among these algorithms, while ShortAFW has the smallest value of the FW gap.
Figure 3: Phase retrieval for $(m,n)=(100,2000)$ and $K=200$.
Table 4: Average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of phase retrieval for $(m,n)=(100,2000)$ and $K=200$.
algorithm primal gap FW gap time
BregFW 3.307714e-09 4.372800e-06 9.532406e-01
EucFW 1.419221e-04 1.271594e+01 8.291155e-01
ShortFW 5.488362e+04 5.389255e+03 1.822046e-01
OpenFW 4.710460e-08 2.747374e+00 1.945798e-01
Figure 4: Phase retrieval for $(m,n)=(200,200)$ and $K=110$ via away-step FW algorithms.
Table 5: Average numbers of the primal gap, FW gap, and computational time (sec) over 20 different instances of phase retrieval for $(m,n)=(200,200)$ and $K=110$ via away-step FW algorithms.
algorithm primal gap FW gap time (s)
BregAFW 4.343027e+00 9.757773e-01 2.333602e-01
EucAFW 4.396803e+00 1.221226e+00 3.942057e-01
ShortAFW 4.409245e+00 2.597756e-02 7.353968e-02
OpenAFW 4.409352e+00 7.199201e-02 5.899826e-02
6.4 Low-Rank Minimization
Given a symmetric matrix $M\in\mathbb{R}^{n\times n}$, our goal is to find $X\in\mathbb{R}^{n\times r}$ such that $M\simeq XX^{\mathsf{T}}$. This is accomplished by minimizing the function
$\min_{X\in P}\quad f(X):=\frac{1}{2}\|XX^{\mathsf{T}}-M\|_{F}^{2},$
where $P\subset\mathbb{R}^{n\times r}$. We assume that $r\leq n$. This problem is known as low-rank minimization [Dragomir2021-rv]. In this paper, we define $P=\{X\in\mathbb{R}^{n\times r}\,|\,\|X\|_{*}\leq\xi\}$, where $\|\cdot\|_{*}$ denotes the nuclear norm and $r\in\mathbb{R}_{+}$ for the low-rank assumption.
We define
$\phi(X)=\frac{1}{4}\|X\|_{F}^{4}+\frac{1}{2}\|X\|_{F}^{2}.$
There exists a constant $L$ such that the pair $(f,\phi)$ is $L$-smad on $\mathbb{R}^{n}$ [Dragomir2021-rv]. Additionally, $f(X)$ is weakly convex on any compact set due to Proposition 2.10, which follows from the twice continuous differentiability of $f$.
We also compared BregFW with EucFW, ShortFW, and OpenFW. The parameter settings are the same as those in the previous subsection. We generated $X^{*}$ from an i.i.d. uniform distribution in $[0,1]$, normalized each column of $X^{*}$, and set $M=X^{*}(X^{*})^{\mathsf{T}}$. The initial point $X_{0}$ was generated from an i.i.d. uniform distribution in $[0,1]$. We set $\xi=10\lambda_{\max}(M)$ for $P$. Figure 5 shows the primal and FW gaps per iteration and gaps per second for $(n,r)=(1000,20)$ up to the 1000th iteration. Table 6 presents the average performance over 20 different instances for $(n,r)=(1000,20)$. BregFW performs slightly better than EucFW. OpenFW also performed as fast as BregFW and EucFW, but its performance was unstable. ShortFW did not converge due to the lack of Lipschitz continuity of $\nabla f$.
Figure 5: Low-rank minimization for $(n,r)=(1000,20)$.
Table 6: Average values of the primal gap, FW gap, and computational time (s) over 20 different instances of low-rank minimization for $(n,r)=(1000,20)$.
algorithm primal gap FW gap time (s)
BregFW 5.651947e-01 1.472666e+01 1.260218e+01
EucFW 5.776476e-01 1.723729e+01 1.228013e+01
ShortFW 2.710108e+08 2.169351e+09 4.137911e+00
OpenFW 1.978179e+00 2.121183e+03 4.176979e+00
6.5 Nonnegative Matrix Factorization
Given a nonnegative matrix $V\in\mathbb{R}_{+}^{m\times n}$, nonnegative matrix factorization (NMF) aims to find nonnegative matrices $W\in\mathbb{R}_{+}^{m\times r}$ and $H\in\mathbb{R}_{+}^{r\times n}$ such that $V\simeq WH$. NMF can be formulated as a minimization problem of the loss function that measures the difference between $V$ and $WH$, i.e.,
$\min_{(W,H)\in P}\quad f(W,H):=\frac{1}{2}\|WH-V\|_{F}^{2},$ (6.4)
where $P$ is a compact convex subset of $\mathbb{R}_{+}^{m\times r}\times\mathbb{R}_{+}^{r\times n}$. The objective function $f$ is weakly convex over $P$ due to Proposition 2.10. The gradient $\nabla f$ is not Lipschitz continuous, while $(f,\phi)$ is smooth adaptable [Mukkamala2019-mk] with $\phi(W,H)=\frac{1}{4}(\|W\|_{F}^{2}+\|H\|_{F}^{2})^{2}+\frac{1}{2}(\|W\|_{F}^{2}+\|H\|_{F}^{2})$.
[…]
Figure 6: NMF for $(m,n,r)=(100,5000,20)$.
Table 7: Average values of the primal gap, FW gap, and computational time (s) over 20 different instances of NMF for $(m,n,r)=(100,5000,20)$.
algorithm primal gap FW gap time (s)
BregFW 1.201735e-04 1.151457e-02 5.098349e+01
EucFW 1.281873e-04 7.625941e-04 4.359247e+01
[…]
Figure 7: NMF for MovieLens 100K Dataset.
Acknowledgement
This project has been partially supported by the Japan Society for the Promotion of Science (JSPS) through Invitational Fellowships for Research in Japan program (L24503), JSPS KAKENHI Grant Number JP23K19953 and JP25K21156; JSPS KAKENHI Grant Number JP23H03351 and JST ERATO Grant Number JPMJER1903. Research reported in this paper was also partially supported the Deutsche Forschungsgemeinschaft (DFG) through the DFG Cluster of Excellence MATH+ (grant number EXC-2046/1, project ID 390685689).
Appendix A Appendix
A.1 Results from Weak Convexity
Lemma A.1 (Primal gap bound from the quadratic growth).
Let $f$ be a $\rho$-weakly convex function that satisfies the local $\mu$-quadratic growth condition such that $\rho\leq\mu$. Let $x^{*}$ be the unique minimizer of $f$ over $P$ and let $\zeta>0$. Then, the following holds: for all $x\in[f\leq f^{*}+\zeta]\cap P$,
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{2},$
or equivalently,
$\left(\frac{\mu}{2}\right)^{1/2}\left(1-\frac{\rho}{\mu}\right)(f(x)-f^{*})^{1/2}\leq\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}.$
Proof.
By first applying weak convexity and then the local quadratic growth condition for $x^{*}\in\mathcal{X}^{*}$ with $f(x^{*})=f^{*}$ and for all $x\in[f\leq f^{*}+\zeta]\cap P$, it holds:
$f(x)-f^{*}$ $=f(x)-f(x^{*})$
$\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}$
$\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{\mu}\left(f(x)-f^{*}\right),$
which implies
$\left(1-\frac{\rho}{\mu}\right)\left(f(x)-f^{*}\right)\leq\langle\nabla f(x),x-x^{*}\rangle=\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\|x-x^{*}\|.$ (A.1)
Using the local quadratic growth condition again, we have
$\left(1-\frac{\rho}{\mu}\right)\left(f(x)-f^{*}\right)\leq\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\left(\frac{2}{\mu}(f(x)-f^{*})\right)^{1/2},$
which provides
$\left(\frac{\mu}{2}\right)^{1/2}\left(1-\frac{\rho}{\mu}\right)(f(x)-f^{*})^{1/2}\leq\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|},$
or equivalently
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{2}.$
Lemma A.2 (Upper bound on primal gap for weakly convex functions).
Suppose that Assumptions 3.1 and 5.2 hold. Let $P$ be a polytope with the pyramidal width $\delta>0$. Let $\mathcal{S}$ denote any set of vertices of $P$ with $x\in\operatorname{conv}\mathcal{S}$. Let $\psi$ be any vector, so that we define $v^{\textup{FW}}=\operatorname*{argmin}_{v\in P}\langle\psi,v\rangle$ and $v^{\textup{A}}=\operatorname*{argmax}_{v\in\mathcal{S}}\langle\psi,v\rangle$. If $\rho/\mu<1$, then it holds that, for all $x\in[f\leq f^{*}+\zeta]\cap P$,
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}\delta^{2}}\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}.$
Proof.
Using the weak convexity and the local quadratic growth condition of $f$, it holds that, for all $x\in[f\leq f^{*}+\zeta]\cap P$,
$f(x)-f(x^{*})$ $\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}$
$\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{\mu}(f(x)-f(x^{*})),$
which implies
$0\leq\left(1-\frac{\rho}{\mu}\right)(f(x)-f(x^{*}))\leq\langle\nabla f(x),x-x^{*}\rangle,$
where the first inequality follows from $01-\rho/\mu>$ 0. Using (2.10) with $\psi=\nabla f(x)$ and $y=x^{*}$ and the above inequality, we obtain
$\frac{\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}}{\delta^{2}}$ $\geq\frac{\langle\nabla f(x),x-x^{*}\rangle^{2}}{\|x-x^{*}\|^{2}}$
$\geq\left(1-\frac{\rho}{\mu}\right)^{2}\frac{(f(x)-f(x^{*}))^{2}}{\|x-x^{*}\|^{2}}$
$\geq\frac{\mu}{2}\left(1-\frac{\rho}{\mu}\right)^{2}\frac{(f(x)-f(x^{*}))^{2}}{f(x)-f(x^{*})}$
$=\frac{\mu}{2}\left(1-\frac{\rho}{\mu}\right)^{2}(f(x)-f(x^{*})),$
where the third inequality holds because of the local quadratic growth condition. ∎
A.2 Local Linear Convergence over Uniformly Convex Sets
In the convex optimization case, Canon and Cullum [Canon1968-sr] established an early lower bound on the convergence rate of the FW algorithm. However, in the special case of $P$ being strongly convex, Garber and Hazan [Garber2015-de] showed that one can improve upon that lower bound. Kerdreux et al. [Kerdreux2021-kc] establish it in the case of $P$ being uniformly convex. We will now carry over this result to establish local linear convergence for the case where $f$ is weakly convex and $L$-smad on $P$.
To this end, we recall the definition of uniformly convex sets.
Definition A.3 ($(\alpha,p)$-uniformly convex set [Braun2022-ew, Definition 2.18][Kerdreux2021-kc, Definition 1.1]).
[…]
$y+\gamma(x-y)+\gamma(1-\gamma)\cdot\alpha\|x-y\|^{p}z\in P.$
Moreover, $P$ is said to be strongly convex if $P$ is $(\alpha,2)$-uniformly convex.
We will use the scaling condition for uniformly convex sets to establish linear convergence.
Proposition A.4 (Scaling inequality [Braun2022-ew, Proposition 2.19][Kerdreux2021-kc, Lemma 2.1]).
Let $P$ be a full dimensional compact $(\alpha,p)$-uniformly convex set, $u$ any non-zero vector, and $v=\operatorname*{argmin}_{y\in P}\langle u,y\rangle$. Then for all $x\in P$
$\frac{\langle u,x-v\rangle}{\|x-v\|^{p}}\geq\alpha\|u\|.$
Now we establish local linear convergence for weakly convex optimization on uniformly convex sets.
Theorem A.5 (Local linear convergence over uniformly convex sets).
Suppose that Assumptions 3.1 and 5.2 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$ hold and that $P$ is $(\alpha,p)$-uniformly convex set. Let $\nabla f$ be bounded away from 0, i.e., $\|\nabla f(x)\|\geq c>0$ for all $x\in P$. Let $D_{\textup{Euc}}:=\sup_{x,y\in P}\|x-y\|$ be the diameter of $P$. Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L\|x_{t}-v_{t}\|^{2}},1\right\}$. Then, if $\rho<\mu$ and $\rho\leq L$, it holds that
$f(x_{t})-f^{*}\leq\begin{cases}\max\left\{\frac{1}{2}\left(1+\frac{\rho}{\mu}\right),1-\left(1-\frac{\rho}{\mu}\right)\frac{\alpha c}{2L}\right\}^{t-1}LD_{\textup{Euc}}^{2}&\text{if}\ p=2,\\ \left(\frac{1}{2}+\frac{\rho}{2\mu}\right)^{t-1}LD_{\textup{Euc}}^{2}&\text{if}\ 1\leq t\leq t_{0},p\geq 2,\\ \frac{L\left(\left(1-\rho/\mu\right)^{1-p/2}L/\alpha c\right)^{2/(p-2)}}{(1+(1-\rho/\mu)(1/2-1/p)(t-t_{0}))^{p/(p-2)}}=\mathcal{O}(1/t^{p/(p-2)})&\text{if}\ t\geq t_{0},p\geq 2,\end{cases}$
for all $t\geq 1$ where
$t_{0}:=\max\left\{\left\lfloor\log_{\frac{1}{2}\left(1+\frac{\rho}{\mu}\right)}\frac{L\left(\left(1-\rho/\mu\right)^{1-p/2}L/\alpha c\right)^{2/(p-2)}}{LD_{\textup{Euc}}^{2}}\right\rfloor+2,1\right\}.$
Proof.
Let $g_{t}:=\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle$ and $h_{t}:=f(x_{t})-f^{*}$. Lemma 3.3 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$ is followed by $f(x_{t})-f(x_{t+1})\geq\frac{g_{t}}{2}\gamma_{t}$. Using Proposition A.4, we have
$h_{t}-h_{t+1}$ $\geq\frac{g_{t}}{2}\min\left\{\frac{g_{t}}{L\|x_{t}-v_{t}\|^{2}},1\right\}$
$\geq\frac{g_{t}}{2}\min\left\{\frac{g_{t}^{1-2/p}\alpha^{2/p}\|\nabla f(x_{t})\|^{2/p}}{L},1\right\}$
$\geq\frac{1}{2}\left(h_{t}-\frac{\rho}{2}\|x-x^{*}\|^{2}\right)\min\left\{\left(h_{t}-\frac{\rho}{2}\|x-x^{*}\|^{2}\right)^{1-2/p}\frac{(\alpha c)^{2/p}}{L},1\right\}$
$\geq\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)\min\left\{\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c)^{2/p}}{L}h_{t}^{1-2/p},1\right\}\cdot h_{t},$
where the third inequality holds from Lemma 5.3 and $\|\nabla f(x)\|\geq c$, and the last inequality holds because of the local quadratic growth property of $f$. The initial bound $h_{1}\leq LD_{\textup{Euc}}^{2}=2LD^{2}$ holds from $\rho\leq L$ and (5.6). For $q=2$, we have $h_{t}-h_{t+1}\geq\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)\min\left\{\frac{\alpha c}{L},1\right\}\cdot h_{t}$, which implies
$h_{t+1}\leq\max\left\{\frac{1}{2}\left(1+\frac{\rho}{\mu}\right),1-\left(1-\frac{\rho}{\mu}\right)\frac{\alpha c}{2L}\right\}\cdot h_{t}.$
Thus, we have the claim. For $p>2$, we use Lemma 2.20 with $c_{0}=LD_{\textup{Euc}}^{2}$, $c_{1}=\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=c_{1}\cdot\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c)^{2/p}}{L}$, and $\theta_{0}=1-2/p$ and obtain the claim. ∎
Note that Assumption 3.1 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$ holds when $f$ is $L$-smooth over $P$. If $\rho\leq L$ does not hold, we can use the initial bound $h_{1}\leq\frac{\rho+L}{2}D_{\textup{Euc}}^{2}$ from (5.6). In that case, a local linear rate in Theorem A.5 is unchanged. Moreover, we have local linear convergence without assuming $\|\nabla f(x)\|>0$.
Theorem A.6 (Local linear convergence over uniformly convex sets without $\|\nabla f(x)\|>0$).
Suppose that Assumptions 3.1 and 5.2 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$ hold and that $P$ is $(\alpha,p)$-uniformly convex set. Let $D_{\textup{Euc}}:=\sup_{x,y\in P}\|x-y\|$ be the diameter of $P$. Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L\|x_{t}-v_{t}\|^{2}},1\right\}$. Then, if $\rho<\mu$ and $\rho\leq L$, it holds that
$f(x_{t})-f^{*}\leq\begin{cases}\left(\frac{1}{2}+\frac{\rho}{2\mu}\right)^{t-1}LD_{\textup{Euc}}^{2}&\text{if}\ 1\leq t\leq t_{0},p\geq 2,\\ \frac{\left(\left(1-\rho/\mu\right)^{2-p}Lc^{2}/\alpha^{2}\right)^{1/(p-1)}}{(1+(1/2)\cdot(1-\rho/\mu)(1-1/p)(t-t_{0}))^{p/(p-1)}}=\mathcal{O}(1/t^{p/(p-1)})&\text{if}\ t\geq t_{0},p\geq 2,\end{cases}$
for all $t\geq 1$ where $c=\frac{\sqrt{2\mu}}{\mu-\rho}$ and
$t_{0}:=\max\left\{\left\lfloor\log_{\frac{1}{2}\left(1+\frac{\rho}{\mu}\right)}\frac{L\left(\left(1-\rho/\mu\right)^{2-p}Lc^{2}/\alpha^{2}\right)^{1/(p-1)}}{LD_{\textup{Euc}}^{2}}\right\rfloor+2,1\right\}.$
Proof.
Using an argument similar to Theorem A.5, we obtain
$h_{t}-h_{t+1}$ $\geq\frac{g_{t}}{2}\min\left\{\frac{g_{t}}{L\|x_{t}-v_{t}\|^{2}},1\right\}$
$\geq\frac{g_{t}}{2}\min\left\{\frac{g_{t}^{1-2/p}\alpha^{2/p}\|\nabla f(x_{t})\|^{2/p}}{L},1\right\}$
$\geq\frac{1}{2}\left(h_{t}-\frac{\rho}{2}\|x-x^{*}\|^{2}\right)\min\left\{\left(h_{t}-\frac{\rho}{2}\|x-x^{*}\|^{2}\right)^{1-2/p}\frac{\alpha^{2/p}(h_{t}/c^{2})^{1/p}}{L},1\right\}$
$\geq\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)\min\left\{\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c^{-1})^{2/p}}{L}h_{t}^{1-1/p},1\right\}\cdot h_{t},$
where the third inequality holds from the PL inequality (2.8) with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$. Therefore, we use Lemma 2.20 with $c_{0}=LD_{\textup{Euc}}^{2}$, $c_{1}=\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=c_{1}\cdot\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c^{-1})^{2/p}}{L}$, and $\theta_{0}=1-1/p$ and obtain the claim. ∎

### v3

@@
-Accelerated Convergence of Frank–Wolfe Algorithms with Adaptive Bregman Step-Size Strategy
+Fast Frank–Wolfe Algorithms with Adaptive Bregman Step-Size for Weakly Convex Functions
 Abstract
@@
 Assuming that the objective function is weakly convex and satisfies the local quadratic growth condition, we provide both local sublinear and local linear convergence regarding the primal gap.
-We also propose a variant of the away-step FW algorithm using Bregman distances.
-We establish global accelerated (up to linear) convergence for convex optimization under the Hölder error bound condition and its local linear convergence for nonconvex optimization under the local quadratic growth condition over polytopes.
-Numerical experiments on nonnegative linear inverse problems, $\ell_{p}$ loss problems, phase retrieval, low-rank minimization, and nonnegative matrix factorization demonstrate that our proposed FW algorithms outperform existing methods.
+We also propose a variant of the away-step FW algorithm using Bregman distances over polytopes.
+We establish global faster (up to linear) convergence for convex optimization under the Hölder error bound condition and its local linear convergence for nonconvex optimization under the local quadratic growth condition.
+Numerical experiments demonstrate that our proposed FW algorithms outperform existing methods.
 1 Introduction
@@
 We are interested in both convex and nonconvex $f$ and assume that we have first-order oracle access to $f$, i.e., given $x\in\mathbb{R}^{n}$, we can compute $\nabla f(x)$.
-Usually, Problem (1.1) is solved by a variant of the projected gradient method (see, for example, [Nesterov2018-gj, Nocedal2006-sb]).
-However, even if $P$ is convex, it is not always possible to access the projection operator of $P$ or compute the projection efficiently.
-While interior point methods are a potential alternative approach to constrained optimization problems, their updates often require second-order information, which can be computationally expensive for large-scale problems.
-To alleviate the situation, Frank and Wolfe [Frank1956-cq] proposed the Frank–Wolfe (FW) algorithm (also known as the conditional gradient method), which is a projection-free first-order method.
+The Frank–Wolfe (FW) algorithm is a projection-free first-order method.
 Instead of requiring access to a projection oracle, the FW algorithm requires only access to a so-called linear minimization oracle (LMO), which for a given linear function $a\in\mathbb{R}^{n}$ computes $y\in\operatorname*{argmin}_{v\in P}\langle a,v\rangle$.
-Linear minimization oracles are often much cheaper than projection oracles, as shown in [CP2021] (see also summary in [Braun2022-ew, Table 1.1]), so that, in practice, FW algorithms are often faster than projected gradient methods when the projection operation is not trivial.
-Additionally, FW algorithms tend to be numerically quite robust and stable due to the affine-invariance of the algorithm and can be also used, e.g., to provide theoretical guarantees for the approximate Carathéodory problem [cp2019approxCara].
+LMOs are often much cheaper than projection oracles, as shown in [CP2021] (see also [Braun2022-ew, Table 1.1]), so that, in practice, FW algorithms are often faster than projected gradient methods even when the projection operation is not trivial.
+Additionally, FW algorithms tend to be numerically quite robust and stable due to their affine-invariance and can be also used, e.g., to provide theoretical guarantees for the approximate Carathéodory problem [cp2019approxCara].
+Overview of Existing FW Studies:
+The FW algorithm was originally proposed by Frank and Wolfe [Frank1956-cq] and was independently rediscovered and extended by Levitin and Polyak [Levitin1966-sm] as the conditional gradient method; we will use these terms interchangeably.
+Canon and Cullum [Canon1968-sr] established an initial lower bound for the convergence rate of the FW algorithm, which was later improved by GuéLat and Marcotte [GueLat1986-bg], which also provided the first analysis of the away-step FW algorithm by Wolfe [Wolfe1970-ax].
+Jaggi [Jaggi2013-oz] provided a more detailed convergence analysis of the FW algorithm, establishing a new lower bound that demonstrated a trade-off between sparsity and error.
+Concurrently, Lan [lan2013complexity] examined the complexity of linear programming-based first-order methods, establishing a similar lower bound.
+For a comprehensive review of FW algorithms, we refer the interested reader to the survey by Braun et al.
+[Braun2022-ew] and the brief introduction by Pokutta [Pokutta2024-qu].
+The classical FW algorithm (Algorithm 1) has been improved in two main directions to enhance both performance and convergence.
+One focuses on refining the step-size rule.
+The short-step strategy computes $\gamma_{t}$ using the Lipschitz constant $L$ of $\nabla f$.
+Pedregosa et al.
+[Pedregosa2020-ay] introduced an adaptive step-size strategy that dynamically estimates $L$.
+They proved that this strategy is at least as effective as the short-step strategy.
+It was enhanced by Pokutta [Pokutta2024-qu] to improve numerical stability.
+Another direction is to modify the classical FW algorithm to eliminate the zigzag behavior when approaching the optimal face containing the optimal solution $x^{*}$.
+This behavior inspired Wolfe [Wolfe1970-ax] to propose the away-step FW algorithm in 1970 that shortcuts the zigzagging by removing atoms that slow down the iterative sequence.
+Lacoste-Julien and Jaggi [Lacoste-Julien2015-gb] showed the linear convergence of the away-step FW algorithm.
+Their analysis***Similar analyses based on the pyramidal width have been developed for other advanced variants, such as the pairwise FW algorithm [Lacoste-Julien2015-gb], Wolfe’s algorithm [Lacoste-Julien2015-gb], the blended conditional gradient method [BPTW2018], and blended pairwise conditional gradient methods and similar variants [CP2020boost, TTP2021].
+Even Nesterov-style acceleration is possible under weak assumptions [CDP2019] building upon this analysis.
+introduced a geometric constant, the pyramidal width, that measures the conditioning of the polytope $P$, representing the feasible region.
+Our Research Idea:
+To derive the convergence rate of the above-mentioned FW algorithms, previous studies have typically required $L$-smoothness, i.e., that $\nabla f$ is Lipschitz continuous, and the convexity for the objective function $f$, although exceptions exist.Quite recently, Vyguzov and Stonyakin [Vyguzov2025-ha] proposed a FW algorithm with the Bregman distance, similar to ours, yet it is restricted to convex problems and provides only partial convergence guarantees.
+These assumptions often narrow the application of FW algorithms.
+Even if they do not satisfy $L$-smoothness, there are many functions that satisfy the $L$-smooth adaptable property ($L$-smad) with kernel generating distances $\phi$ (see Definition 2.1) by choosing $\phi$ well to match the function $f$.
+Since the $L$-smad property is consistent with $L$-smoothness when $\phi=\frac{1}{2}\|\cdot\|^{2}$, the class of $L$-smad functions includes the $L$-smooth function class.
+Other first-order algorithms, such as the Bregman proximal gradient algorithm, are often analyzed under the $L$-smad property (see e.g., [Bolte2018-zt, Hanzely2021-sz, Rebegoldi2018-pa, Takahashi2022-ml, Takahashi2024-wz, Yang2025-ls]).
+$L$-smad functions appear in many applications, such as nonnegative linear inverse problems [Bauschke1997-vk, Takahashi2024-wz], $\ell_{p}$ loss problems [Kyng2015-uo, Maddison2021-nt], phase retrieval [Bolte2018-zt, Takahashi2022-ml], nonnegative matrix factorization (NMF) [Mukkamala2019-mk, Takahashi2024-cu], and blind deconvolution [Takahashi2023-uh].
+Any $\mathcal{C}^{2}$ function is locally $L$-smooth over compact sets, but the resulting Lipschitz constant $L$ can be overly conservative, hindering practical performance.
+Functions outside $\mathcal{C}^{2}$ may not satisfy $L$-smoothness at all.
+Moreover, functions in the above-mentioned applications are weakly convex on compact sets.
+For nonconvex functions, although Lacoste-Julien [Lacoste-Julien2016-mj] established sublinear convergence, little else has been done, and simply tracing existing research does not achieve convergence rates better than sublinear convergence.
+Then the following question naturally arises:
+Is it possible to loosen the $L$-smoothness and convexity assumptions often made in FW algorithm studies while still achieving the theoretical guarantee of linear convergence rates?
+Contribution:
+The answer to the above question is yes.
+We develop FW algorithms for $L$-smad $f$ and give theoretical guarantees of linear convergence under weaker assumptions (Hölder error bound (HEB) condition with parameter $q\geq 1$ for convex $f$ and the special case with $q=2$ called quadratic growth for nonconvex $f$) than the strong convexity assumption that is often made when linear convergence is shown for FW algorithms.
+The convexity of $f$ is also weaken to weakly convex, i.e., $f+\frac{\rho}{2}\|\cdot\|^{2}$ is convex with some $\rho>0$ (see, e.g.[Davis2018-bb, Examples 3.1 and 3.2] and [Nurminskii1975-vp]).
+Moreover, it is easy to see that any $\mathcal{C}^{2}$ function defined on a compact set is weakly convex [Vial1983-oo, Proposition 4.11] and [Wu2007-dt]; since $\rho$ is only needed for the theoretical guarantee and not for the algorithm, there is no need for its estimation, and considering weakly-convex $f$ greatly expands the range of applications.
+Table 1 summarizes our contributions for the various cases that we consider.
+Relaxing the assumption from $L$-smoothness to the $L$-smad property makes the construction of a FW algorithm and its theoretical guarantee much more difficult.
+This is because we use the Bregman distance, which is an extension of the Euclidean distance, for the $L$-smad property.
+While (2.1) is simple with $\nu=1$ when using the Euclidean distance $\phi=\frac{1}{2}\|\cdot\|^{2}$, $\nu>0$ cannot be eliminated for the general Bregman distance; $1+\nu$ represents its scaling exponent.
+$\nu$ must be estimated to use in the algorithm since its exact value is unknown.
+Therefore, the extension to an $L$-smad $f$ is not straightforward.
+  • We establish sublinear convergence for the case of convex $L$-smad functions.
+Assuming that $f$ is convex, satisfies the HEB condition, and the minimizer $x^{*}\in\operatorname{int}P$, we provide the faster (up to linear) convergence.
+In this setting, our algorithm always converges linearly if the exponent of the HEB condition $q$ equals the scaling exponent of the Bregman distance $1+\nu$, i.e., $q=1+\nu$.As discussed in Sec.
+2, linear convergence with the Euclidean distance for convex optimization is shown in [garber2015faster] under the quadratic growth over strongly convex sets.
+This corresponds to the $q=1+\nu$ case with $\nu=1$.
+For $q>1+\nu$, if $t\leq t_{0}$ with some $t_{0}\in\mathbb{N}$, linear convergence holds; otherwise, $\mathcal{O}(\epsilon^{(1+\nu-q)/\nu q})$, which is faster than existing sublinear rates.
+A similar argument also applies to the nonconvex case ($q=2$) and the away-step FW algorithm.
+  • For nonconvex optimization, assuming that $f$ is weakly convex and satisfies the local quadratic growth condition, we establish local linear convergence.
+To our knowledge, this is the first proposal of a FW algorithm with linear convergence guarantees for a certain type of nonconvex optimization problem, albeit of a limited type.
+These results are also new for the standard setting: $L$-smooth $f$.
+Without weak convexity, we prove a global sublinear convergence to a stationary point of (1.1) for nonconvex $L$-smad $f$.
+  • We also propose a variant of the away-step FW algorithm with the Bregman distance and establish its linear convergence under the HEB condition for convex optimization and under the local quadratic growth condition for weakly convex optimization.
+Table 1: Convergence rates for Frank–Wolfe algorithms.
+Our convergence rates are generalizations of existing rates.
+Moreover, our convergence rate is better than $\mathcal{O}(\epsilon^{-1})$ in some cases.
+For nonconvex functions, convergence is measured using the Frank–Wolfe gap $\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\leq\epsilon$, instead of the primal gap $f(x_{t})-f^{*}\leq\epsilon$.
+Convergence rates for the weakly convex optimization locally hold.
+Assumptions Convergence rate
+FW $f$ convexity $f$ growth $x^{*}\in\operatorname{int}P$ $P$ polytope $L$-smooth $L$-smad ($\nu>0$)
+Alg.1 (any step-size) convex $\mathcal{O}(\epsilon^{-1})$ [Jaggi2013-oz] $\mathcal{O}(\epsilon^{-1/\nu})$ ([Vyguzov2025-ha], Thm.C.1)
+Alg.1 (short & adapt.) convex $q$-HEB $\mathcal{O}(\log\epsilon^{-1})$11footnotemark: 1[KDP2018jour] $\mathcal{O}(\log\epsilon^{-1})$ [^2] It holds if the exponent of the HEB condition $q$ equals the scaling exponent of the Bregman distance $1+\nu$, i.e., $q=1+\nu$ ($(\nu,q)=(1,2)$ for the away-step FW algorithm).
+For $q>1+\nu$, it also holds if $t\leq t_{0}$ with some $t_{0}\in\mathbb{N}$; otherwise $\mathcal{O}(\epsilon^{(1+\nu-q)/\nu q})$.(Thm.
+4.2)
+away-step FW convex $q$-HEB $\mathcal{O}(\log\epsilon^{-1})$11footnotemark: 1[KDP2018jour] $\mathcal{O}(\log\epsilon^{-1})$22footnotemark: 2(Thm. 4.4)
+Alg.1 (any step-size) weak 2-HEB $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(\epsilon^{-1/\nu})$ (Thm. D.3)
+Alg.1 (short & adapt.) weak 2-HEB $\mathcal{O}(\log\epsilon^{-1})$ $\mathcal{O}(\log\epsilon^{-1})$ [^4] It holds if $\nu=1$.
+For $\nu\in(0,1)$, it also holds if $t\leq t_{0}$ with some $t_{0}\in\mathbb{N}$; otherwise $\mathcal{O}(\epsilon^{(\nu-1)/2\nu})$.(Thm.
+5.2)
+away-step FW weak 2-HEB $\mathcal{O}(\log\epsilon^{-1})$ $\mathcal{O}(\log\epsilon^{-1})$44footnotemark: 4(Thm. 5.3)
+Alg.1 (any step-size) $\mathcal{O}(\epsilon^{-2})$ [Lacoste-Julien2016-mj] $\mathcal{O}(\epsilon^{-1-1/\nu})$ (Thm. D.1)
+2 Preliminaries
+[…]
+Let $C$ be a nonempty open convex subset of $\mathbb{R}^{n}$.
+Definition 2.1 (Kernel Generating Distance [Bolte2018-zt, Definition 2.1]).
+A function $\phi:\mathbb{R}^{n}\to(-\infty,+\infty]$ is called a kernel generating distance associated with $C$ if it satisfies the following conditions:
+  1. (i)
+    $\phi$ is proper, lower semicontinuous, and convex, with $\operatorname{dom}\phi\subset\operatorname{cl}C$ and $\operatorname{dom}\partial\phi=C$;
+  2. (ii)
+    $\phi$ is $\mathcal{C}^{1}$ on $\operatorname{int}\operatorname{dom}\phi\equiv C$.
+We denote the class of kernel generating distances associated with $C$ by $\mathcal{G}(C)$.
+Definition 2.2 (Bregman Distance [Bregman1967-rf]).
+Given $\phi\in\mathcal{G}(C)$, a Bregman distance $D_{\phi}:\operatorname{dom}\phi\times\operatorname{int}\operatorname{dom}\phi\to\mathbb{R}_{+}$ associated with $\phi$ is defined by $D_{\phi}(x,y)=\phi(x)-\phi(y)-\langle\nabla\phi(y),x-y\rangle\geq 0$.
+We use $D_{f}$ in Definition 2.2 for nonconvex $f$ instead of convex $\phi$ for brevity.
+In the remainder of the paper, if not stated otherwise, we will assume that as such for a given strictly convex $\phi$ and $C$, there exists $\nu>0$ such that the following holds for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and $\gamma\in[0,1]$ (see also Lemma A.3, which shows that a weaker one holds)):
+$D_{\phi}((1-\gamma)x+\gamma y,x)\leq\gamma^{1+\nu}D_{\phi}(y,x).$ (2.1)
+We will now recall the smooth adaptable property, which is a generalization of $L$-smoothness and was first introduced by [Bauschke2017-hg].
+The smooth adaptable property is also called relative smoothness [Lu2018-ii].
+Definition 2.3 ($L$-smooth Adaptable Property [Bolte2018-zt, Definition 2.2]).
+Let $\phi\in\mathcal{G}(C)$ and let $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ be proper and lower semicontinuous with $\operatorname{dom}\phi\subset\operatorname{dom}f$, which is $\mathcal{C}^{1}$ on $C\equiv\operatorname{int}\operatorname{dom}\phi$.
+The pair of functions $(f,\phi)$ is said to be $L$-smooth adaptable (for short: $L$-smad) on $C$ if there exists $L>0$ such that $L\phi-f$ and $L\phi+f$ are convex on $C$.
+Lemma 2.4 (Primal progress from the smooth adaptable property).
+Let the pair of functions $(f,\phi)$ be $L$-smad with a strictly convex function $\phi\in\mathcal{G}(C)$ and let $x_{+}=(1-\gamma)x+\gamma v$ with $x,v\in\operatorname{int}\operatorname{dom}\phi$ and $\gamma\in[0,1]$.
+Then it holds:
+$f(x)-f(x_{+})\geq\gamma\langle\nabla f(x),x-v\rangle-L\gamma^{1+\nu}D_{\phi}(v,x).$ (2.2)
+Often the point $v\in P\subset\operatorname{int}\operatorname{dom}\phi$ in Lemma 2.4 is chosen as a Frank–Wolfe vertex, i.e., $v\in\operatorname*{argmax}_{u\in P}\langle\nabla f(x),x-u\rangle$, but other choices, e.g., those arising from away-directions from the away-step FW algorithm, are also possible, as we will see in Section 3.2.
+In [garber2015faster] (see also [Garber2020-tr]) the linear convergence of the FW algorithm for convex optimization under the quadratic growth condition, which is a weaker assumption than assuming strong convexity, was established over strongly convex sets and later generalized to uniformly convex sets in [KAP2020] as well as to conditions weaker than quadratic growth in [KDP2018, KDP2018jour].
+Local variants of these notions, as necessary, e.g., for the nonconvex case, have been studied in [KAP2021] in the context of FW algorithms.
+For $f:\mathbb{R}^{n}\to[-\infty,+\infty]$, let $[f\leq\zeta]:=\{x\in\mathbb{R}^{n}\,|\,f(x)\leq\zeta\}$ be a $\zeta$-sublevel set of $f$ for some $\zeta\in\mathbb{R}$.
+Definition 2.5 (Hölder Error Bound [Braun2022-ew, Definition 3.25], [alex17sharp, Definition 1.2] and Quadratic Growth Conditions [garber2015faster, Garber2020-tr, Liao2024-vd]).
+Let $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ be a proper lower semicontinuous function and $P\subset\mathbb{R}^{n}$ be a compact convex set.
+Let $\mathcal{X}^{*}\neq\emptyset$ be the set of optimal solutions, i.e., $\mathcal{X}^{*}:=\operatorname*{argmin}_{x\in P}f(x)$, and let $f^{*}=\min_{x\in P}f(x)$ and $\zeta>0$.
+The function $f$ satisfies the Hölder error bound (HEB) condition on $P$ if there exists constants $q\geq 1$ and $\mu>0$ such that, for any $x\in P$,
+$\operatorname{dist}(x,\mathcal{X}^{*})^{q}\leq\frac{q}{\mu}(f(x)-f^{*}).$ (2.3)
+Particularly, when $q=2$, it is called the $\mu$-quadratic growth condition.
+Moreover, the function $f$ is said to be local $\mu$-quadratic growth (with $\zeta$) if there exists a constant $\mu>0$ such that, for any $x\in[f\leq f^{*}+\zeta]\cap P$, (2.3) holds with $q=2$.
+The HEB condition shows sharpness bounds on the primal gap (see [bolte17, alex17sharp]), which has also been extensively analyzed for FW algorithms [KDP2018, KDP2018jour].
+Convergence for nonconvex optimization under the quadratic growth or HEB condition was established in [Davis2024-rm, Davis2018-bb, Davis2024-og].
+We employ the HEB condition for convex optimization, while we use the local quadratic growth for nonconvex optimization.
+3 Proposed Bregman FW Algorithms
+Throughout this paper, we make the following assumptions.
+Assumption 3.1.
+  1. (i)
+    $\phi\in\mathcal{G}(C)$ with $\operatorname{cl}{C}=\operatorname{cl}{\operatorname{dom}\phi}$ is strictly convex on $C\equiv\operatorname{int}\operatorname{dom}\phi$;
+  2. (ii)
+    $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is proper and lower semicontinuous with $\operatorname{dom}\phi\subset\operatorname{dom}f$, which is $\mathcal{C}^{1}$ on $C$;
+  3. (iii)
+    The pair $(f,\phi)$ is $L$-smad on $P$;
+  4. (iv)
+    $P\subset\mathbb{R}^{n}$ is a nonempty compact convex set with $P\subset C$.
+Assumption 3.1(i)-(iii) are standard for Bregman-type algorithms [Bolte2018-zt, Takahashi2024-wz].
+The strict convexity of $\phi$ ensures the existence of $\nu$ by Lemma A.3 and is satisfied by many kernel generating distances (see, e.g., [Bauschke2017-hg, Bauschke1997-vk], and [Dhillon2008-zz, Table 2.1]).
+Assumption 3.1(iv) is standard for FW algorithms, and $P\subset C$ is natural ensuring that $D_{\phi}$ is well-defined on $P$.
 The original FW algorithm is given in Algorithm 1.
-Input: Initial point $x_{0}\in P$, objective function $f$, step-size strategy $\gamma_{t}\in[0,1]$
+Input: Initial point $x_{0}\in P$, objective function $f$, step-size $\gamma_{t}\in[0,1]$
 1
@@
 Algorithm 1 Frank–Wolfe algorithm
-Related Work
-In this section, we briefly review the work most closely related to this paper.
-The FW algorithm was originally proposed by Frank and Wolfe [Frank1956-cq] and was independently rediscovered and extended by Levitin and Polyak [Levitin1966-sm] as the conditional gradient method; we will use these terms interchangeably.
-Canon and Cullum [Canon1968-sr] established an initial lower bound for the convergence rate of the FW algorithm, which was later improved by GuéLat and Marcotte  [GueLat1986-bg], which also provided the first analysis of the away-step FW algorithm by Wolfe [Wolfe1970-ax].
-Jaggi [Jaggi2013-oz] provided a more detailed convergence analysis of the FW algorithm, establishing a new lower bound that demonstrated a trade-off between sparsity and error.
-Concurrently, Lan [lan2013complexity] examined the complexity of linear programming-based first-order methods, establishing a similar complexity lower bound.
-The classical FW algorithm is known to zigzag when approaching the optimal face containing the optimal solution $x^{*}$.
-This behavior inspired Wolfe [Wolfe1970-ax] to propose the away-step FW algorithm in 1970 that shortcuts the zigzagging by removing atoms that slow down the iterate sequence.
-Lacoste-Julien and Jaggi [Lacoste-Julien2015-gb] showed the linear convergence of the away-step FW algorithm.
-Their analysis introduced a geometric constant, the pyramidal width, that measures the conditioning of the polytope $P$, representing the feasible region.
-Similar analyses based on this constant have been developed for other advanced variants, such as the pairwise FW algorithm [Lacoste-Julien2015-gb], Wolfe’s algorithm [Lacoste-Julien2015-gb], the blended conditional gradient method [BPTW2018], and blended pairwise conditional gradient methods and similar variants [CP2020boost, TTP2021].
-Even Nesterov-style acceleration is possible under weak assumptions [CDP2019] building upon this analysis.
-Pedregosa et al.
-[Pedregosa2020-ay] introduced a powerful adaptive step-size strategy that dynamically approximates the local Lipschitz continuity of the function $f$.
-They proved that this strategy is at least as effective as the traditional short-step strategy, and it was later enhanced by Pokutta [Pokutta2024-qu] to improve numerical stability.
-For nonconvex functions, Lacoste-Julien [Lacoste-Julien2016-mj] established sublinear convergence, and Maskan et al.
-[Maskan2025-oa] incorporated the difference of convex functions (DC) into FW algorithms.
-For a comprehensive review of FW algorithms, we refer the interested reader to the survey by Braun et al.
-[Braun2022-ew] and the brief introduction by Pokutta [Pokutta2024-qu].
-Establishing the convergence rate of FW algorithms typically requires that $\nabla f$ is Lipschitz continuous, although exceptions exist, such as in the case of generalized self-concordant functions [CBP2021, SelfConcordantFW_2020].
-Some first-order methods using the Bregman distance do not require the global Lipschitz continuity of $\nabla f$.
-Bolte et al.
-[Bolte2018-zt] proposed the Bregman proximal gradient algorithm and established its global convergence under the smooth adaptable property (see Definition 2.6).
-Hanzely et al.
-[Hanzely2021-sz] introduced accelerated Bregman proximal gradient algorithms.
-Takahashi et al.
-[Takahashi2022-ml] developed a Bregman proximal gradient algorithm for the DC optimization and its accelerated version.
-Subproblems of Bregman proximal gradient algorithms are often not solvable in closed form.
-Takahashi and Takeda [Takahashi2024-wz] used approximations to Bregman distances to solve subproblems in closed form.
-Rebegoldi et al.
-[Rebegoldi2018-pa] proposed an inexact version of the Bregman proximal gradient algorithm.
-Contribution
-Our contributions can be roughly summarized as follows:
-Frank–Wolfe for Smooth Adaptable functions
-We study Frank–Wolfe algorithms for smooth adaptable functions (also called: relatively smooth) functions.
-Some functions whose gradients are not Lipschitz continuous satisfy the smooth adaptable property with kernel generating distances (see Definition 2.1).
-The class of smooth adaptable functions appears in many applications, such as nonnegative linear inverse problems [Bauschke1997-vk, Takahashi2024-wz], $\ell_{p}$ loss problems [Kyng2015-uo, Maddison2021-nt] phase retrieval [Bolte2018-zt, Takahashi2022-ml], nonnegative matrix factorization (NMF) [Mukkamala2019-mk, Takahashi2024-cu], and blind deconvolution [Takahashi2023-uh].
-While the gradient of any $\mathcal{C}^{2}$ function over compact sets is Lipschitz continuous, an $\ell_{p}$ loss function is not $\mathcal{C}^{2}$ when $21<p<$ 2, i.e., its gradient is not Lipschitz continuous over compact sets.
-Moreover, functions in these applications are weakly convex on compact sets.
-We establish sublinear to linear convergence rates for the respective variants, which are compatible and in line with existing results.
-Adaptive Bregman Step-Size Strategy
-We propose an adaptive Bregman step-size strategy and establish its convergence results for convex and nonconvex optimization.
-Table 1 summarizes our contributions for the various cases that we consider.
-While Vyguzov and Stonyakin [Vyguzov2024-tp] also proposed an adaptive FW algorithm with the Bregman distance similar to ours for convex optimization problems.
-They established a linear convergence rate assuming that the objective function is relatively strongly convex, the optimal solution belongs to $\operatorname{int}P$, and the angle condition holds; in contrast, we establish linear convergence under much weaker assumptions.
-Beyond that, we also establish sublinear convergence for convex optimization for the case of general relatively smooth functions.
-Assuming that $f$ is convex, satisfies the Hölder error bound condition, and the optimal solution $x^{*}\in\operatorname{int}P$, we also provide the accelerated (up to linear) convergence, however, in a different way than [Vyguzov2024-tp] with fewer assumptions.
-For nonconvex optimization, assuming that $f$ is weakly convex and satisfies the local quadratic growth condition, we establish local linear convergence for our step-size strategy.
-These results include a special case that $\nabla f$ is Lipschitz continuous, i.e., $L$-smoothness.
-Without weak convexity, we prove the global convergence rate of the FW algorithm to a stationary point of (1.1), similar to [Lacoste-Julien2016-mj].
-Away-Step FW Algorithm
-We propose a variant of the away-step FW algorithm with the Bregman distance and establish its linear convergence under the Hölder error bound condition for convex optimization and under the local quadratic growth condition for weakly convex optimization.
-Numerical Experiments
-In numerical experiments, we have applied the FW algorithm with the adaptive Bregman step-size strategy to nonnegative linear systems [Bauschke2017-hg], $\ell_{p}$ loss problems [Kyng2015-uo, Maddison2021-nt], phase retrieval [Patterson1934-sg, Patterson1944-nm], low-rank minimization [Dragomir2021-rv], and NMF [Gillis2020-dh].
-Bauschke et al.
-[Bauschke2017-hg] showed applications to linear inverse problems to solve nonnegative linear systems.
-Maddison et al.
-[Maddison2021-nt] considered $\ell_{p}$ loss problems, and Kyng et al.
-[Kyng2015-uo] showed its application as Lipschitz learning on graphs.
-Bolte et al.
-[Bolte2018-zt] and Takahashi et al.
-[Takahashi2022-ml] applied Bregman proximal algorithms to phase retrieval, Dragomir et al.
-[Dragomir2021-rv] applied them to low-rank minimization, and Mukkamala and Ochs [Mukkamala2019-mk] and Takahashi et al.
-[Takahashi2024-cu] applied them to NMF.
-Throughout these applications, we compare our proposed algorithms with existing FW algorithms and the mirror descent algorithm [Nemirovskij1983-sm] as well as away-step FW algorithms (see e.g.
-[Braun2022-ew, GueLat1986-bg, Lacoste-Julien2015-gb, Wolfe1970-ax]).
-We would also like to stress that while we formulate some of the results for the case that $x^{*}\in\operatorname{int}P$, the results can be extended to the case that $x^{*}\in\operatorname{ri}P$, i.e., the relative interior of $P$, which we did not do for the sake of clarity.
-Basically, the analysis is performed in the affine space spanned by the optimal face of $P$ in that case; the interested reader is referred to [Braun2022-ew] for details on how to extend the results.
-Table 1: Convergence rates for the Frank–Wolfe algorithm.
-For nonconvex functions, convergence is measured using the Frank-Wolfe gap, $\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\leq\epsilon$, instead of the primal gap, $f(x_{t})-f^{*}\leq\epsilon$.
-Weak convexity can be replaced by twice continuous differentiability (see Proposition 2.10).
-Convergence rates for the weakly convex optimization locally hold.
-AFW, HEB, and quad.
-denote the away-step Frank–Wolfe algorithm, the Hölder error bound condition with $q\geq 1$, and the quadratic growth condition, respectively.
-FW Assumptions Convergence rate
-$f$ convexity $f$ growth $x^{*}\in\operatorname{int}P$ $P$ polytope $L$-smooth $L$-smad ($\nu\in(0,1]$)
-any convex $\mathcal{O}(\epsilon^{-1})$ [Jaggi2013-oz] $\mathcal{O}(\epsilon^{-1/\nu})$ ([Vyguzov2024-tp], Thm. 4.2)
+In what follows, let $x^{*}\in\operatorname*{argmin}_{x\in P}f(x)$ and $f^{*}=f(x^{*})$.
+We recall a key property for analyzing convergence rates in convex optimization and derive its counterpart in weakly convex optimization.
+Lemma 3.2 (Primal gap, dual gap, and Frank–Wolfe gap for convex $f$ [Pokutta2024-qu, Lemma 4.1]).
+Suppose that Assumptions 3.1(ii), (iv), and that $f$ is convex. Then for all $x\in P$, it holds:
+$f(x)-f^{*}\leq\langle\nabla f(x),x-x^{*}\rangle\leq\max_{v\in P}\langle\nabla f(x),x-v\rangle.$ (3.1)
+Lemma 3.3 (Primal gap, dual gap, and Frank–Wolfe gap for weakly convex $f$).
+Suppose that Assumptions 3.1(ii), (iv), and that $f$ is $\rho$-weakly convex. Then for all $x\in P$, it holds:
+$f(x)-f^{*}\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}\leq\max_{v\in P}\langle\nabla f(x),x-v\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}.$ (3.2)
+3.1 Bregman Step-Size Strategy for $\gamma_{t}$
+Bregman Short Step-Size:
+Assume that $(f,\phi)$ is $L$-smad. Using the optimality condition of the right-hand side of (2.2) in terms of $\gamma$, we find
+$\gamma=\left(\frac{\langle\nabla f(x),x-v\rangle}{L(1+\nu)D_{\phi}(v,x)}\right)^{\frac{1}{\nu}}\quad\text{and}\quad f(x)-f(x_{+})\geq\frac{\nu}{1+\nu}\frac{\langle\nabla f(x),x-v\rangle^{1+1/\nu}}{(L(1+\nu)D_{\phi}(v,x))^{1/\nu}}.$
+Theoretically, when $x,v\in P$ and $\gamma\in[0,1]$, the Frank–Wolfe step $x_{+}=(1-\gamma)x+\gamma v\in P$ is well-defined.
+We update $\gamma=\min\left\{\left(\frac{\langle\nabla f(x),x-v\rangle}{L(1+\nu)D_{\phi}(v,x)}\right)^{1/\nu},\gamma_{\max}\right\}$ with some $\gamma_{\max}\in\mathbb{R}$ (usually set $\gamma_{\max}=1$).
+This step-size strategy provides the progress lemma.
+Lemma 3.4 (Progress lemma from the Bregman short step-size).
+Suppose that Assumption 3.1 holds.
+Define $x_{t+1}=(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$.
+Then if $x_{t+1}\in\operatorname{dom}f$, $f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle$ for $0\leq\gamma_{t}\leq\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu}$.
+Output : Estimates $\tilde{L}^{*}$ and $\tilde{\nu}^{*}$, step-size $\gamma$
 […]
-[…]
-any weak quad. [^3] We assume that $f$ is local quadratic growth. $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(\epsilon^{-1/\nu})$ (Thm. 5.5)
-[…]
-[…]
-any [^5] Its rate is established for the Frank–Wolfe gap. nonconvex $\mathcal{O}(\epsilon^{-2})$ [Lacoste-Julien2016-mj] $\mathcal{O}(\epsilon^{-1-1/\nu})$ (Thm. 5.1)
-Outline
-The structure of this paper is as follows.
-Section 2 introduces essential notations, including Bregman distances, the smooth adaptable property, weak convexity, growth conditions, and scaling inequalities, which are crucial for the convergence analyses that come later.
-In Section 3, we present the FW algorithm with an adaptive Bregman step-size strategy and the away-step FW algorithm utilizing the Bregman distance.
-Section 4 details the convergence results for convex optimization, demonstrating both sublinear and linear convergence rates.
-Section 5 addresses convergence results for the nonconvex case, showing that our algorithm achieves global convergence to a stationary point of (1.1).
-Under the assumption that $f$ is weakly convex and local quadratic growth, we establish local linear convergence to a minimizer of (1.1), indicating that the proposed algorithms converge to a minimizer when the initial point is sufficiently close.
-Section 6 presents applications to nonnegative linear inverse problems, phase retrieval, low-rank minimization, and nonnegative matrix factorization (NMF).
-2 Preliminaries
-[…]
-Let $\operatorname{int}{C}$, $\operatorname{ri}{C}$, and $\operatorname{cl}{C}$ be the interior, the relative interior, and the closure of a set $C\subset\mathbb{R}^{n}$, respectively.
-The set $\operatorname{conv}C$ denotes the convex hull of $C$, and the set $\operatorname{Vert}C$ denotes the set of vertices of $C$.
-We also define the distance from a point $x\in\mathbb{R}^{n}$ to $C$ by $\operatorname{dist}(x,C):=\inf_{y\in C}\|x-y\|$.
-For an extended-real-valued function $f:\mathbb{R}^{n}\to[-\infty,+\infty]$, the effective domain of $f$ is defined by $\operatorname{dom}f:=\{x\in\mathbb{R}^{n}\,|\,f(x)<+\infty\}$.
-The function $f$ is said to be proper if $f(x)>-\infty$ for all $x\in\mathbb{R}^{n}$ and $\operatorname{dom}f\neq\emptyset$.
-Let $[f\leq\zeta]:=\{x\in\mathbb{R}^{n}\,|\,f(x)\leq\zeta\}$ be a $\zeta$-sublevel set of $f$ for some $\zeta\in\mathbb{R}$.
-Let $\mathcal{C}^{k}$ be the class of $k$-times continuously differentiable functions for $k\geq 0$.
-The sign function $\operatorname{sgn}(x)$ is defined by $\operatorname{sgn}(x)=-1$ for $x<0$, $\operatorname{sgn}(x)=0$ for $x=0$, and $\operatorname{sgn}(x)=1$ for $x>0$.
-2.1 Bregman Distances
-Let $C$ be a nonempty open convex subset of $\mathbb{R}^{n}$.
-Definition 2.1 (Kernel Generating Distance [Bolte2018-zt, Definition 2.1]).
-A function $\phi:\mathbb{R}^{n}\to(-\infty,+\infty]$ is called a kernel generating distance associated with $C$ if it satisfies the following conditions:
-  1. (i)
-    $\phi$ is proper, lower semicontinuous, and convex, with $\operatorname{dom}\phi\subset\operatorname{cl}C$ and $\operatorname{dom}\partial\phi=C$.
-  2. (ii)
-    $\phi$ is $\mathcal{C}^{1}$ on $\operatorname{int}\operatorname{dom}\phi\equiv C$.
-We denote the class of kernel generating distances associated with $C$ by $\mathcal{G}(C)$.
-Definition 2.2 (Bregman Distance [Bregman1967-rf]).
-Given a kernel generating distance $\phi\in\mathcal{G}(C)$, a Bregman distance $D_{\phi}:\operatorname{dom}\phi\times\operatorname{int}\operatorname{dom}\phi\to\mathbb{R}_{+}$ associated with $\phi$ is defined by
-$D_{\phi}(x,y)=\phi(x)-\phi(y)-\langle\nabla\phi(y),x-y\rangle.$ (2.1)
+2      Choose $\beta<1$, $\eta\leq 1$, and $\tau>1$
+3      $M\leftarrow\eta\tilde{L}$, $\kappa\leftarrow 1$
+4      while $D_{f}(x_{+},x)\leq M\gamma^{1+\kappa}D_{\phi}(v,x)$ do
+5           $\gamma\leftarrow\min\left\{\left(\frac{\langle\nabla f(x),x-v\rangle}{M(1+\kappa)D_{\phi}(v,x)}\right)^{1/\kappa},\gamma_{\max}\right\}$
+6           $x_{+}\leftarrow(1-\gamma)x+\gamma v$
+7           $M\leftarrow\tau M$
+8           $\kappa\leftarrow\beta\kappa$
+9
+10           $\tilde{L}^{*}\leftarrow M$, $\tilde{\nu}^{*}\leftarrow\kappa$
+11           return $\tilde{L}^{*}$, $\tilde{\nu}^{*}$, $\gamma$
+12
+Algorithm 2 Adaptive Bregman step-size strategy
+If $\phi=\frac{1}{2}\|\cdot\|^{2}$, we have $\nu=1$ and $\gamma_{t}=\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L\|v_{t}-x_{t}\|^{2}}$, which is often called the (Euclidean) short step-size.
+Although the short step-size does not require line searches, it requires knowledge of the value of $L$.
+Adaptive Bregman Step-Size:
+The exact value or the tight upper bound of $L$ is often unknown, however, the algorithm’s performance heavily depends on it; an underestimation of $L$ might lead to non-convergence, and an overestimation of $L$ might lead to slow convergence.
+Moreover, a worst-case $L$ might be too conservative for regimes where the function is better behaved.
+Due to all these reasons, Pedregosa et al.
+[Pedregosa2020-ay] proposed an adaptive step-size strategy for FW algorithms in the case where $\phi=\frac{1}{2}\|\cdot\|^{2}$.
+We present our algorithm in Algorithm 2.
+This step-size strategy can be used as a drop-in replacement.
+For example, inserting $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},1)$ between lines 1 and 1 in Algorithm 1, we obtain that Algorithm 2 searches $L$ and $\nu$ satisfying (2.2).
+Remark 3.5 (Well-definedness and termination of Algorithm 2).
+By the $L$-smad property of $(f,\phi)$ and Lemma 2.4, we know that (2.2) holds for all $M\geq L$ and $\nu\geq\kappa>0$.
+Therefore, the loop in line 2 in Algorithm 2 is well-defined and guaranteed to terminate.
+It suffices to update $\kappa$ in line 2 only if (2.1) does not hold.
+Thus, $\kappa$ cannot become too small.
+3.2 Bregman Away-Step Frank–Wolfe Algorithm
+Input: $x_{0}\in\operatorname*{argmin}_{v\in P}\langle\nabla f(x),v\rangle$ for $x\in P$, $\beta<1$
+1 $\mathcal{S}_{0}\leftarrow\{x_{0}\}$, $\lambda_{x_{0},0}\leftarrow 1$
+2
+3for $t=0,\ldots$ do
+4      $v_{t}^{\textup{FW}}\leftarrow\operatorname*{argmin}_{v\in P}\langle\nabla f(x_{t}),v\rangle$, $v_{t}^{\textup{A}}\leftarrow\operatorname*{argmax}_{v\in\mathcal{S}_{t}}\langle\nabla f(x_{t}),v\rangle$
+5
+6     if $\langle\nabla f(x_{t}),x_{t}-v_{t}^{\textup{FW}}\rangle\geq\langle\nabla f(x_{t}),v_{t}^{\textup{A}}-x_{t}\rangle$ then
+7           $v_{t}\leftarrow v_{t}^{\textup{FW}}$, $d_{t}\leftarrow x_{t}-v_{t}^{\textup{FW}}$, $\gamma_{t,\max}\leftarrow 1$
+8
+9           else
+10                $v_{t}\leftarrow v_{t}^{\textup{A}}$, $d_{t}\leftarrow v_{t}^{\textup{A}}-x_{t}$, $\gamma_{t,\max}\leftarrow\lambda_{v_{t}^{\textup{A}},t}/(1-\lambda_{v_{t}^{\textup{A}},t})$
+11
+12                $\gamma_{t}\leftarrow\min\left\{\left(\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu},\gamma_{t,\max}\right\}$, $x_{t+1}\leftarrow x_{t}-\gamma_{t}d_{t}$
+13
+14               if $\langle\nabla f(x_{t}),x_{t}-v_{t}^{\textup{FW}}\rangle\geq\langle\nabla f(x_{t}),v_{t}^{\textup{A}}-x_{t}\rangle$ then
+15                     $\lambda_{v,t+1}\leftarrow(1-\gamma_{t})\lambda_{v,t}$ for all $v_{t}\in\mathcal{S}_{t}\setminus\{v_{t}^{\textup{FW}}\}$
+16                     $\lambda_{v_{t}^{\textup{FW}},t+1}\leftarrow\begin{cases}\gamma_{t}&\text{if}\ v_{t}^{\textup{FW}}\notin\mathcal{S}_{t}\\ (1-\gamma_{t})\lambda_{v_{t}^{\textup{FW}},t}+\gamma_{t}&\text{if}\ v_{t}^{\textup{FW}}\in\mathcal{S}_{t}\end{cases}$, $\mathcal{S}_{t+1}\leftarrow\begin{cases}\mathcal{S}_{t}\cup\{v_{t}^{\textup{FW}}\}&\text{if}\ \gamma_{t}<1\\ \{v_{t}^{\textup{FW}}\}&\text{if}\ \gamma_{t}=1\end{cases}$
+17
+18                     else
+19                          $\lambda_{v,t+1}\leftarrow(1+\gamma_{t})\lambda_{v,t}$ for all $v\in\mathcal{S}_{t}\setminus\{v_{t}^{\textup{A}}\}$, $\lambda_{v_{t}^{\textup{A}},t+1}\leftarrow(1+\gamma_{t})\lambda_{v_{t}^{\textup{A}},t}-\gamma_{t}$
+20                          $\mathcal{S}_{t+1}\leftarrow\begin{cases}\mathcal{S}_{t}\setminus\{v_{t}^{\textup{A}}\}&\text{if}\ \lambda_{v_{t}^{\textup{A}},t+1}=0\\ \mathcal{S}_{t}&\text{if}\ \lambda_{v_{t}^{\textup{A}},t+1}>0\end{cases}$
+21
+22
+Algorithm 3 Away-step Frank–Wolfe algorithm with the Bregman distance
+In the case where $P$ is a polytope, the classical FW algorithm might zigzag when approaching the optimal face and, in consequence, converge slowly.
+Away-step FW algorithm overcomes this drawback.
+Away steps allow the algorithm to move away from vertices in a convex combination of $x_{t}$, which can effectively shortcircuit the zigzagging.
+The convergence properties of this algorithm were unknown for a long time, and it was only quite recently that Lacoste-Julien and Jaggi [Lacoste-Julien2015-gb] established the linear convergence of the away-step FW algorithm.
+Inspired by [GueLat1986-bg, Lacoste-Julien2015-gb, Wolfe1970-ax], we propose a variant of the away-step FW algorithm utilizing Bregman distances as given in Algorithm 3.
+For the following discussion, we introduce some notions.
+In the same way [Braun2022-ew], we define an active set $\mathcal{S}\subset\operatorname{Vert}P$, where $\operatorname{Vert}P$ denotes the set of vertices of $P$, and the away vertex as $v_{t}^{\textup{A}}\in\operatorname*{argmax}_{v\in\mathcal{S}}\langle\nabla f(x_{t}),v\rangle$, where $x_{t}$ is a (strict) convex combination of elements in $\mathcal{S}$, i.e., $x_{t}=\sum_{v\in\mathcal{S}}\lambda_{v}x_{t}$ with $\lambda_{v}>0$ for all $v\in\mathcal{S}$ and $\sum_{v\in\mathcal{S}}\lambda_{v}=1$.
+If $\nu$ is unknown, we can add line search procedures to search $\nu$ until (2.1) holds.
+If $\nu$ and $L$ are unknown, we can use $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},\gamma_{t,\max})$ with $\gamma_{t}\leftarrow\min\left\{\left(\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{M(1+\kappa)D_{\phi}(v_{t},x_{t})}\right)^{1/\kappa},\gamma_{t,\max}\right\}$ instead of line 2 in Algorithm 2.
+When $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu\equiv 1$, Algorithm 3 corresponds to the Euclidean away-step FW algorithm.
+4 Linear Convergence for Convex Optimization
+In this section, we assume that $f$ is convex.
+Assumption 4.1.
+The objective function $f$ is convex.
+We will now establish faster convergence rates than $O(1/t)$ up to linear convergence depending on the choice of parameters.
+First, we establish the faster convergence of Algorithm 1 with the Bregman short step-size, i.e., $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu},1\right\}$, in the case where the optimal solution lies in the relative interior.
+Let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$.
+Theorem 4.2 (Linear convergence of FW algorithm with short step-size or adaptive step-size).
+Suppose that Assumptions 3.1 and 4.1 hold.
+Let $\nu>0$ and let $f$ satisfy the HEB condition (2.3) with $q\geq 1+\nu$ and $\mu>0$ and $D$ be the diameter of $P$.
+Assume that there exists a minimizer $x^{*}\in\operatorname{int}P$, i.e., there exists an $r>0$ with $B(x^{*},r)\subset P$.
+Consider the iterates of Algorithm 1 with short step-size $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu},1\right\}$ or the adaptive Bregman step-size strategy (Algorithm 2).
+Then, it holds that, for all $t\geq 1$,
+$f(x_{t})-f^{*}\leq\begin{cases}\max\left\{\frac{1}{1+\nu},1-\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{c^{1+1/\nu}D^{2/\nu}}\right\}^{t-1}LD^{2}&\text{if}\ q=1+\nu,\\ \frac{LD^{2}}{(1+\nu)^{t-1}}&\text{if}\ 1\leq t\leq t_{0},q>1+\nu,\\ \frac{(L(1+\nu)(c/r)^{1+\nu}D^{2})^{q/(q-1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}(t-t_{0})\right)^{\nu q/(q-1-\nu)}}=\mathcal{O}(1/t^{\nu q/(q-1-\nu)})&\text{if}\ t\geq t_{0},q>1+\nu,\end{cases}$
+where $c=(q/\mu)^{1/q}$ and $t_{0}:=\max\{\lfloor\log_{1/(1+\nu)}((L(1+\nu(c/r)^{1+\nu}D^{2})^{q/(q-1-\nu)}/LD^{2})\rfloor+2,1\}$.
+Next, we establish the linear convergence of Algorithm 3.
+Recall that the pyramidal width is the minimal $\delta>0$ satisfying Lemma B.9 (see [Braun2022-ew, Lemma 2.26] or [Lacoste-Julien2015-gb, Theorem 3] for an in-depth discussion).
+Under Assumption 4.3, it holds that $\delta\leq D$, and this will be important in establishing convergence rates of Algorithm 3.
+We make the following assumption:
+Assumption 4.3.
+The kernel generating distance $\phi$ is $\sigma$-strongly convex.
+Theorem 4.4 (Linear convergence of the away-step FW algorithm).
+Suppose that Assumptions 3.14.1, and 4.3 hold.
+Let $\nu>0$ and let $P$ be a polytope and $f$ satisfy the HEB condition with $q>1+\nu$ or $(\nu,q)=(1,2)$.
+The convergence rate of Algorithm 3 is linear: for all $t\geq 1$
+$f(x_{t})-f^{*}\leq\begin{cases}\left(1-\frac{\mu}{32L}\frac{\delta^{2}}{D^{2}}\right)^{\lceil(t-1)/2\rceil}LD^{2}&\text{if}\ (\nu,q)=(1,2),\\ \frac{1}{(1+\nu)^{\lceil(t-1)/2\rceil}}LD^{2}&\text{if}\ 1\leq t\leq t_{0},q>1+\nu,\\ \frac{(L(1+\nu)D^{2}c/(\delta/2)^{1+\nu})^{q/(q-1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}\lceil(t-t_{0})/2\rceil\right)^{\nu q/(q-1-\nu)}}=\mathcal{O}(1/t^{\nu q/(q-1-\nu)})&\text{if}\ t\geq t_{0},q>1+\nu,\end{cases}$
+where $c:=(q/\mu)^{1/q}$ $D$ and $\delta$ are the diameter and the pyramidal width of the polytope $P$, respectively, and $t_{0}:=\max\{\lfloor\log_{1/(1+\nu)}(L(1+\nu)D^{2}/((\mu/q)^{1/q}\delta/2)^{1+\nu})^{q/(q-1-\nu)}/LD^{2}\rfloor+2,1\}$.
+5 Local Linear Convergence for Nonconvex Optimization
+In this section, we consider a nonconvex objective function.
+We establish global sublinear, local sublinear, and local linear convergence (see, for global and local sublinear, Appendices D.1 and D.2).
+We show that the FW algorithm converges to a minimizer $x^{*}$ when an initial point is close enough to $x^{*}$.
+We need weak convexity and local quadratic growth for local convergence.
+Assumption 5.1.
+$f$ is $\rho$-weakly convex and local $\mu$-quadratic growth with $\zeta>0$.
+Theorem 5.2 (Local linear convergence of FW algorithm with short step-size or adaptive step-size).
+Suppose that Assumptions 3.1 and 5.1 hold.
+Let $\nu\in(0,1]$ and let $D$ be the diameter.
+Assume that there exists a minimizer $x^{*}\in\operatorname{int}P$, i.e., there exists an $r>0$ with $B(x^{*},r)\subset P$.
+Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu},1\right\}$ or the adaptive Bregman step-size strategy (Algorithm 2).
+Then, if $\rho/\mu<1$, it holds that, for all $t\geq 1$,
+$f(x_{t})-f^{*}\leq\begin{cases}\max\left\{\frac{1}{2}\left(1+\frac{\rho}{\mu}\right),1-\frac{r^{2}}{2Mc^{2}D^{2}}\right\}^{t-1}LD^{2}&\text{if}\ \nu=1,\\ \left(\frac{1}{1+\nu}\left(1+\frac{\nu\rho}{\mu}\right)\right)^{t-1}LD^{2}&\text{if}\ 1\leq t\leq t_{0},\nu\in(0,1),\\ \frac{(L(1+\nu)(1-\rho/\mu)^{\nu}c^{1+\nu}D^{2}/r^{1+\nu})^{2/(1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}\left(1-\frac{\rho}{\mu}\right)(t-t_{0})\right)^{2\nu/(1-\nu)}}=\mathcal{O}(1/t^{2\nu/(1-\nu)})&\text{if}\ t\geq t_{0},\nu\in(0,1),\end{cases}$
+where $c=\sqrt{2\mu}/(\mu-\rho)$, $M:=(L(1+\nu))^{1/\nu}$, and $t_{0}:=\max\{\lfloor\log_{\left(1+\nu\rho/\mu\right)/(1+\nu)}((L(1+\nu)(1-\rho/\mu)^{\nu}c^{1+\nu}D^{2}/r^{1+\nu})^{2/(1-\nu)}/LD^{2})\rfloor+2,1\}$.
+Because Algorithm 2 is well-defined (see also Remark 3.5), the convergence result of the FW algorithm with Algorithm 2 is the same as Theorem 5.2.
+Finally, we establish the local linear convergence of Algorithm 3. We also assume $\rho\leq L$, which is not restrictive (see also the discussion in Appendix D.2 and Example D.4).
+Theorem 5.3 (Local linear convergence by the away-step FW algorithm).
+Suppose that Assumptions 3.14.3, and 5.1 hold.
+Let $\nu\in(0,1]$ and let $P\subset\mathbb{R}^{n}$ be a polytope.
+The convergence rate of Algorithm 3 with $f$ is linear: if $\rho<\mu\leq L$, for all $t\geq 1$
+$f(x_{t})-f^{*}\leq\begin{cases}2\left(1-\frac{\omega}{4L}\frac{\delta^{2}}{D^{2}}\right)^{\lceil(t-1)/2\rceil}LD^{2}&\text{if}\ \nu=1,\\ 2\left(\frac{1}{1+\nu}\left(1+\frac{\nu\rho}{\mu}\right)\right)^{\lceil(t-1)/2\rceil}LD^{2}&\text{if}\ 1\leq t\leq t_{0},\nu\in(0,1),\\ \frac{(L(1+\nu)(1-\rho/\mu)^{\nu}D^{2}/(\sqrt{\omega}\delta)^{1+\nu})^{2/(1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}\left(1-\frac{\rho}{\mu}\right)\lceil(t-t_{0})/2\rceil\right)^{2\nu/(1-\nu)}}=\mathcal{O}(1/t^{2\nu/(1-\nu)})&\text{if}\ t\geq t_{0},\nu\in(0,1),\end{cases}$
+where $D$ and $\delta$ are the diameter and the pyramidal width of $P$, respectively, and $\omega:=(\mu-\rho)^{2}/8\mu$, and $t_{0}:=\max\{\lfloor\log_{(1+\nu\rho/\mu)/(1+\nu)}(L^{2}(1+\nu)^{2}D^{4}(1-\rho/\mu)^{2\nu}/(\omega\delta^{2})^{1+\nu})^{1/(1-\nu)}/2LD^{2}\rfloor+2,1\}$.
+Without loss of generality, we set $\sigma=1$.
+If $\rho\leq L$ does not hold, we can use $f(x_{1})-f^{*}\leq(\rho+L)D^{2}$, instead of $f(x_{1})-f^{*}\leq 2LD^{2}$ as the initial bound.
+Moreover, when $P$ is $\alpha$-strongly convex set and $\phi=\frac{1}{2}\|\cdot\|^{2}$, we also establish local linear convergence.
+Its proof can be found in Appendix D.5.
+6 Numerical Experiments
+In this section, we conducted numerical experiments to examine the performance of our proposed algorithms.
+All numerical experiments were performed in Julia 1.11 using the FrankWolfe.jl package [Besancon2022-lx] [^2] https://github.com/ZIB-IOL/FrankWolfe.jl on a MacBook Pro with an Apple M2 Max and 64GB LPDDR5 memory.
+We compare BregFW and BregAFW, which use the adaptive Bregman step-size strategy (Algorithm 2), with existing algorithms, EucFW, EucAFW, ShortFW, ShortAFW, OpenFW, OpenAFW, and MD (their notation is described in Appendix E).
+Note that we included OpenAFW only for comparison purposes; there is no established convergence theory for the away-step FW algorithm with the open loop.
+Indeed, there are no proper drop steps (operations corresponding to line 3 in Algorithm 3), and the favorable properties of the away-step FW algorithm are lost; see [Braun2022-ew].
+We use $\beta=0.9$, $\eta=0.9$, $\tau=2$, $\gamma_{\max}=1$, and the termination criterion $\max_{v\in P}\langle\nabla f(x_{t}),x_{t}-v\rangle\leq 10^{-7}$ throughout all experiments.
+Figure 1: Log plot of primal and FW gaps on $\ell_{p}$ loss for gas sensor data with $b_{\max}=130$.
+Figure 2: Log plot of primal and FW gaps on $\ell_{p}$ loss for gas sensor data with $b_{\max}=200$.
+Figure 3: Log plot of primal and FW gaps on phase retrieval for $(m,n)=(1000,10000)$ and $K=2000$.
+Figure 4: Log plot of primal and FW gaps on phase retrieval for $(m,n)=(2000,10000)$ and $K=2000$.
+6.1 $\ell_{p}$ Loss Problem
+We consider the $\ell_{p}$ loss problem [Kyng2015-uo, Maddison2021-nt] to find $x\in P$ such that $Ax\simeq b$, defined by a convex optimization problem $\min_{x\in P}\|Ax-b\|_{p}^{p}$, where $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, and $p>1$.
+Let $f(x)=\|Ax-b\|_{p}^{p}$.
+The gradient $\nabla f$ is not Lipschitz continuous on $\mathbb{R}^{n}$ when $p\neq 2$.
+Furthermore, when $p<2$, $f$ is not $\mathcal{C}^{2}$ but $\mathcal{C}^{1}$.
+Therefore, for $21<p<$ 2, $\nabla f$ is also not Lipschitz continuous over compact sets.
+Since $f$ is convex, the pair $(f,\phi)$ is 1-smad with $\phi:=f$.
+We consider real-world data using gas sensor data.§§§The dataset is freely available at https://archive.ics.uci.edu/dataset/270/gas+sensor+array+drift+dataset+at+different+concentrations The data size is $(m,n)=(13910,128)$.
+We use the $\ell_{2}$ constraint $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|^{2}\leq b_{\max}\}$.
+We compared BregFW with EucFW and OpenFW because ShortFW does not converge.
+The maximum iteration is 1000 except for OpenFW, and that of OpenFW is 5000.
+The initial point $x_{0}$ is an all-ones vector.
+We set $p=1.1$, i.e., $f$ is $\mathcal{C}^{1}$ not $\mathcal{C}^{2}$.
+Figures 4 and 4 show the primal and FW gaps per iteration and gaps per second for $b_{\max}=130$ and $b_{\max}=200$, respectively.
+Because $\nabla f$ is not Lipschitz continuous on $P$ from lacking $\mathcal{C}^{2}$, ShortFW and EucFW do not converge.
+The primal gap and the FW gap by BregFW are the smallest among these step-size strategies because only BregFW has a theoretical guarantee for this problem.
+6.2 Phase Retrieval
+We are interested in phase retrieval, which involves finding a signal $x\in\mathbb{R}^{n}$ such that $|\langle a_{i},x\rangle|^{2}\simeq b_{i}$, $i=1,\ldots,m$, where $a_{i}\in\mathbb{R}^{n}$ describes the model and $b\in\mathbb{R}^{m}$ is a vector of measurements.
+Phase retrieval arises in many fields of science and engineering, such as image processing [Candes2015-mk], X-ray crystallography [Patterson1934-sg, Patterson1944-nm], and optics [Shechtman2015-ic].
+To achieve the goal of phase retrieval, we focus on the nonconvex optimization problem $\min_{x\in P}f(x):=\frac{1}{4}\sum_{i=1}^{m}\left(|\langle a_{i},x\rangle|^{2}-b_{i}\right)^{2}$, where $P\subset\mathbb{R}^{n}$ is a compact convex set.
+Let $\phi$ be defined by $\phi(x)=\frac{1}{4}\|x\|^{4}+\frac{1}{2}\|x\|^{2}$, which is $\sigma$-strongly convex for $\sigma\leq 1$.
+For any $L$ satisfying $L\geq\sum_{i=1}^{m}\left(3\|a_{i}\|^{4}+\|a_{i}\|^{2}|b_{i}|\right)$, the pair $(f,\phi)$ is $L$-smad on $\mathbb{R}^{n}$ [Bolte2018-zt, Lemma 5.1].
+In addition, $\nabla^{2}f(x)+\rho I$ is positive semidefinite for $\rho\geq\sum_{i=1}^{m}\|a_{i}\|^{2}|b_{i}|$, i.e., $f$ is $\rho$-weakly convex.
+Therefore, $L\geq\rho/\sigma$ holds with $\sigma=1$.
+We use a $K$-sparse polytope as a constraint, i.e., $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|_{1}\leq K,\|x\|_{\infty}\leq 1\}$.
+We compared BregFW with EucFW, ShortFW, and OpenFW.
+The maximum iteration is 1000 except for OpenFW, and that of OpenFW is 5000.
+We generated $a_{i}$, $i=1,\ldots,m$ from an i.i.d.
+normal distribution and normalized it to have norm 1 (random seed 1234).
+We also generated $x^{*}$ from an i.i.d.
+uniform distribution in $[0,1]$ and normalized $x^{*}$ to have sum 1.
+The initial point $x_{0}$ was generated by computing an extreme point of $P$ that minimizes the linear approximation of $f$.
+Figures 4 and 4 show the primal and FW gaps per iteration and gaps per second for $(m,n)=(1000,10000)$ and $K=2000$ and for $(m,n)=(2000,10000)$ and $K=2000$, respectively.
+The primal gap and the FW gap by the adaptive Bregman step-size strategy are the smallest among these step-size strategies.
+In both cases, BregFW stopped before the 1000th iteration.
+7 Conclusion and Future Work
+Summary:
+We propose FW algorithms with an adaptive Bregman step-size strategy for a smooth adaptable function, which does not require Lipschitz continuity of its gradient.
+We have established sublinear to linear convergence rates depending on various assumptions for convex and nonconvex optimization.
+We also propose a variant of the away-step FW algorithm using Bregman distances over polytopes.
+For the away-step FW algorithm, we have also established global convergence for convex optimization under the HEB condition and its local linear convergence for nonconvex optimization under the local quadratic growth condition.
+Numerical experiments show that our algorithms outperform existing algorithms.
+Limitation and future work:
+Our step-size strategy for a smooth adaptable function requires the value of $\nu$, which can be estimated, because we use the Bregman distance.
+Theoretically, linear convergence of the FW algorithm always holds for convex optimization if the exponent of the HEB condition $q$ equals the scaling exponent of the Bregman distance $1+\nu$, i.e., $q=1+\nu$.
+For $q>1+\nu$, if $t\leq t_{0}$ with some $t_{0}\in\mathbb{N}$, linear convergence holds; otherwise, sublinear convergence.
+The condition $q=1+\nu$, which is derived from the growth of $f$ and $\phi$, is natural because $\phi$ has a similar structure to $f$.
+On the other hand, for nonconvex optimization, we assume the local quadratic growth condition, which is the special case of the HEB condition with $q=2$.
+Since this is derived from the weak convexity of $f$, relaxing the quadratic growth condition may not be possible.
+In that regard, one direction is to consider an extension using the difference of convex functions (DC) optimization.
+Recently, Maskan et al.
+[Maskan2025-oa] incorporated DC optimization into FW algorithms.
+While they established sublinear convergence, our algorithms may exploit DC optimization techniques.
+Acknowledgement
+This project has been partially supported by the Japan Society for the Promotion of Science (JSPS) through Invitational Fellowships for Research in Japan program (L24503), JSPS KAKENHI Grant Number JP23K19953 and JP25K21156; JSPS KAKENHI Grant Number JP23H03351 and JST ERATO Grant Number JPMJER1903.
+Research reported in this paper was also partially supported the Deutsche Forschungsgemeinschaft (DFG) through the DFG Cluster of Excellence MATH+ (grant number EXC-2046/1, project ID 390685689).
+Appendix A Results from Smooth Adaptable Property
+A.1 Properties of Bregman Distances
 The Bregman distance $D_{\phi}(x,y)$ measures the proximity between $x\in\operatorname{dom}\phi$ and $y\in\operatorname{int}\operatorname{dom}\phi$.
@@
 The Bregman distance is also called the Bregman divergence.
-Example 2.3.
+Example A.1.
 Well-known choices for $\phi$ and $D_{\phi}$ are listed below; for more examples, see, e.g., [Bauschke2017-hg, Example 1], [Bauschke1997-vk, Section 6], and [Dhillon2008-zz, Table 2.1].
@@
 We recall the triangle scaling property for Bregman distances from Hanzely et al. [Hanzely2021-sz, Section 2], where several properties of the triangle scaling property are also shown.
-Definition 2.4 (Triangle Scaling Property [Hanzely2021-sz, Definition 2]).
+Definition A.2 (Triangle Scaling Property [Hanzely2021-sz, Definition 2]).
 Given a kernel generating distance $\phi\in\mathcal{G}(C)$, the Bregman distance $D_{\phi}$ has the triangle scaling property if there exists a constant $\nu>0$ such that, for all $x,y,z\in\operatorname{int}\operatorname{dom}\phi$ and all $\gamma\in[0,1]$, it holds that
-$D_{\phi}((1-\gamma)x+\gamma y,(1-\gamma)x+\gamma z)\leq\gamma^{\nu}D_{\phi}(y,z).$ (2.2)
-Now, substituting $z\leftarrow x$ on the left-hand side of (2.2), we obtain
+$D_{\phi}((1-\gamma)x+\gamma y,(1-\gamma)x+\gamma z)\leq\gamma^{\nu}D_{\phi}(y,z).$ (A.1)
+Now, substituting $z\leftarrow x$ on the left-hand side of (A.1), we obtain
 $D_{\phi}((1-\gamma)x+\gamma y,x)$ $=\phi((1-\gamma)x+\gamma y)-\phi(x)-\gamma\langle\nabla\phi(x),y-x\rangle$
@@
 Therefore, there exists $\nu>0$ such that $D_{\phi}((1-\gamma)x+\gamma y,x)\leq\gamma^{\nu}D_{\phi}(y,x)$ holds for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and all $\gamma\in[0,1]$.
-We will now show that a stronger version can be obtained if $\phi$ is strictly convex, so that we can rephrase $\nu$ with $1+\nu$ where $\nu\in(0,1]$, i.e., the right-hand side is superlinear, which will be crucial for the convergence analysis later.
-Lemma 2.5.
-Given a kernel generating distance $\phi\in\mathcal{G}(C)$, if $\phi$ is strictly convex, then there exists $\nu\in(0,1]$ such that, for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and all $\gamma\in[0,1]$, it holds that
-$D_{\phi}((1-\gamma)x+\gamma y,x)\leq\gamma^{1+\nu}D_{\phi}(y,x).$ (2.3)
+We will now show that a weaker version of the triangle scaling property can be obtained if $\phi$ is strictly convex, so that we can rephrase $\nu$ with $1+\nu$ where $\nu>0$, i.e., the right-hand side is superlinear.
+Lemma A.3.
+Given a kernel generating distance $\phi\in\mathcal{G}(C)$, if $\phi$ is strictly convex, then for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and all $\gamma\in[0,1]$, there exists $\nu>0$ such that, it holds that
+$D_{\phi}((1-\gamma)x+\gamma y,x)\leq\gamma^{1+\nu}D_{\phi}(y,x).$ (A.2)
 Proof.
-Obviously, (2.3) holds if $y=x$ or $\gamma\in\{0,1\}$.
+Obviously, (A.2) holds if $y=x$ or $\gamma\in\{0,1\}$.
 In what follows, we assume $y\neq x$ and $10<\gamma<$ 1.
@@
 Because $\phi$ is quadratic when $D_{\phi}$ is symmetric [Bauschke2001-ry, Lemma 3.16], it always holds that $D_{\phi}((1-\gamma)x+\gamma y,x)=\gamma^{1+\nu}D_{\phi}(y,x)$ with $\nu=1$.
-In the remainder of the paper, if not stated otherwise, we will assume that $\phi$ is strictly convex and as such for a given $\phi$ and $C$ there exists a $\nu\in(0,1]$ such that $D_{\phi}((1-\gamma)x+\gamma y,x)\leq\gamma^{1+\nu}D_{\phi}(y,x)$ holds for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and all $\gamma\in[0,1]$.
-2.2 Smooth Adaptable Property
-We will now recall the smooth adaptable property, which is a generalization of $L$-smoothness and was first introduced by [Bauschke2017-hg].
-The smooth adaptable property is also called relative smoothness [Lu2018-ii].
-Definition 2.6 ($L$-smooth Adaptable Property [Bolte2018-zt, Definition 2.2]).
-Consider a pair of functions $(f,\phi)$ satisfying the following conditions:
-  1. (i)
-    $\phi\in\mathcal{G}(C)$,
-  2. (ii)
-    $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is proper and lower semicontinuous with $\operatorname{dom}\phi\subset\operatorname{dom}f$, which is $\mathcal{C}^{1}$ on $C\equiv\operatorname{int}\operatorname{dom}\phi$.
-The pair $(f,\phi)$ is said to be $L$-smooth adaptable (for short: $L$-smad) on $C$ if there exists $L>0$ such that $L\phi-f$ and $L\phi+f$ are convex on $C$.
+A.2 Lemmas of Smooth Adaptable Property
 The convexity of $L\phi-f$ and $L\phi+f$ plays a central role in developing and analyzing algorithms, and the smooth adaptable property implies the extended descent lemma.
-Lemma 2.7 (Extended Descent Lemma [Bolte2018-zt, Lemma 2.1]).
+Lemma A.4 (Extended Descent Lemma [Bolte2018-zt, Lemma 2.1]).
 The pair of functions $(f,\phi)$ is $L$-smad on $C$ if and only if for all $x,y\in\operatorname{int}\operatorname{dom}\phi$,
@@
 From this, it can be seen that the $L$-smad property for $(f,\phi)$ provides upper and lower approximations for $f$ majorized by $\phi$ with $L>0$.
-In addition, if $\phi(x)=\frac{1}{2}\|x\|^{2}$ on $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$, Lemma 2.7 corresponds to the classical descent lemma.
+In addition, if $\phi(x)=\frac{1}{2}\|x\|^{2}$ on $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$, Lemma A.4 corresponds to the classical descent lemma.
 While the $L$-smad property might seem unfamiliar at first, it is a natural generalization of $L$-smoothness, and examples of functions $f$ and $\phi$ satisfying the $L$-smad property are given, e.g., in [Bauschke2017-hg, Lemmas 7 and 8][Bolte2018-zt, Lemma 5.1], [Dragomir2021-rv, Propositions 2.1 and 2.3], [Mukkamala2019-mk, Proposition 2.1], [Takahashi2024-wz, Proposition 24], [Takahashi2023-uh, Theorem 1], and [Takahashi2024-cu, Theorem 15].
 The extended descent lemma immediately implies primal progress of FW algorithms under the $L$-smad property as a straightforward generalization of the $L$-smooth case.
-Lemma 2.8 (Primal progress from the smooth adaptable property).
-Let the pair of functions $(f,\phi)$ be $L$-smad with a strictly convex function $\phi$ and let $x_{+}=(1-\gamma)x+\gamma v$ with $x,v\in\operatorname{int}\operatorname{dom}\phi$ and $\gamma\in[0,1]$.
+Lemma A.5 (Primal progress from the smooth adaptable property).
+Let the pair of functions $(f,\phi)$ be $L$-smad with a strictly convex function $\phi\in\mathcal{G}(C)$ and let $x_{+}=(1-\gamma)x+\gamma v$ with $x,v\in\operatorname{int}\operatorname{dom}\phi$ and $\gamma\in[0,1]$.
 Then it holds:
@@
 Proof.
-Using Lemma 2.7 and substituting $x_{+}$ for $x$ and $x$ for $y$, we have
+Using Lemma A.4 and substituting $x_{+}$ for $x$ and $x$ for $y$, we have
 $f(x)-f(x_{+})$ $\geq\gamma\langle\nabla f(x),x-v\rangle-LD_{\phi}(x_{+},x).$
-It holds that $D_{\phi}(x_{+},x)\leq\gamma^{1+\nu}D_{\phi}(v,x)$ for some $\nu\in(0,1]$ by Lemma 2.5. Therefore, this provides the desired inequality. ∎
-Often the point $v\in P\subset\operatorname{int}\operatorname{dom}\phi$ in Lemma 2.8 is chosen as a Frank–Wolfe vertex, i.e.,
-$v\in\operatorname*{argmax}_{u\in P}\langle\nabla f(x),x-u\rangle,$
-but other choices, e.g., those arising from away-directions from the away-step FW method, are also possible, as we will see in Section 3.2.
-2.3 Weakly Convex Functions
-A function $f$ is called weakly convex if it becomes convex upon adding a quadratic perturbation.
-Definition 2.9 (Weakly Convex Function [Davis2018-bb, Nurminskii1975-vp]).
-A function $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is said to be $\rho$-weakly convex for some $\rho>0$ if the function $f+\frac{\rho}{2}\|\cdot\|^{2}$ is convex.
+It holds that $D_{\phi}(x_{+},x)\leq\gamma^{1+\nu}D_{\phi}(v,x)$ for some $\nu>0$ by Lemma A.3. Therefore, this provides the desired inequality. ∎
+We also have a progress lemma from the Bregman short step-size.
+Lemma A.6 (Progress lemma from the Bregman short step-size).
+Suppose that Assumption 3.1 holds. Define $x_{t+1}=(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$. Then if $x_{t+1}\in\operatorname{dom}f$,
+$f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\quad\text{for}\quad 0\leq\gamma_{t}\leq\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}}.$
+Proof.
+Substituting $x_{t+1}=(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$ with $x_{t},v_{t}\in P$ for $x_{+}$ in Lemma 2.4, we have
+$f(x_{t})-f(x_{t+1})$ $\geq\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle-L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t})$
+$\geq\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle-\frac{\gamma_{t}}{1+\nu}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle$
+$=\frac{\nu}{1+\nu}\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle,$
+where the second inequality holds because of an upper bound of $\gamma_{t}$. ∎
+Appendix B Properties from Sharpness
+B.1 Results from Weak Convexity
 Obviously, $f$ is $\rho$-weakly convex and differentiable if and only if, for all $x,y\in\operatorname{dom}f$,
 $f(y)-f(x)\geq\langle\nabla f(x),y-x\rangle-\frac{\rho}{2}\|y-x\|^{2}.$
-For examples of such $f$, see [Davis2018-bb, Examples 3.1 and 3.2].
-Moreover, it is easy to see that any $\mathcal{C}^{2}$ function defined on a compact set is weakly convex [Vial1983-oo, Proposition 4.11] and [Wu2007-dt], the rationale being that we can simply shift the function by the quadratic belonging to its smallest eigenvalue.
+Any $\mathcal{C}^{2}$ function defined on a compact set is weakly convex [Vial1983-oo, Proposition 4.11] and [Wu2007-dt].
 Since the proof of this fact is omitted therein, we provide the argument for the sake of completeness below.
-Proposition 2.10.
+Proposition B.1.
 Let $P\subset\mathbb{R}^{n}$ be a compact set and $f$ be a proper and $\mathcal{C}^{2}$ function on $P$. Then, $f$ is weakly convex on $P$.
 Proof.
-Let $g=\lambda_{\min}(\nabla^{2}f(\cdot))$.
+Let $g=\lambda_{\min}(\nabla^{2}f(\cdot))$, where $\lambda_{\min}(\cdot)$ is the smallest eigenvalue.
 Obviously, $g$ is continuous because $\lambda_{\min}(\cdot)$ and $\nabla^{2}f(\cdot)$ are continuous.
@@
 Minimizing a nonconvex $\mathcal{C}^{2}$-function over a compact set $P$ is equivalent to minimizing a weakly convex function over $P$.
-However, it is important to note that when $\rho$ becomes large, the assumptions of algorithms may not be satisfied (see also Theorems 5.5 and 5.9).
-2.4 Growth Conditions
-In [garber2015faster] (see also [Garber2020-tr]) the linear convergence of the FW algorithm for convex optimization under the quadratic growth condition, which is a weaker assumption than assuming strong convexity was established over strongly convex sets and later generalized to uniformly convex sets in [KAP2020] as well as to conditions weaker than quadratic growth in [KDP2018, KDP2018jour].
-Local variants of these notions, as necessary, e.g., for the nonconvex case, have been studied in [KAP2021] in the context of Frank–Wolfe methods.
-Definition 2.11 (Hölder Error Bound [Braun2022-ew, Definition 3.25], [alex17sharp, Definition 1.2] and Quadratic Growth Conditions [garber2015faster, Garber2020-tr, Liao2024-vd]).
-Let $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ be a proper lower semicontinuous function and $P\subset\mathbb{R}^{n}$ be a compact convex set.
-Let $\mathcal{X}^{*}\neq\emptyset$ be the set of optimal solutions, i.e., $\mathcal{X}^{*}:=\operatorname*{argmin}_{x\in P}f(x)$, and let $f^{*}=\min_{x\in P}f(x)$ and $\zeta>0$.
-The function $f$ satisfies the Hölder error bound condition on $P$ if there exists constants $q\geq 1$ and $\mu>0$ such that, for any $x\in P$,
-$\operatorname{dist}(x,\mathcal{X}^{*})^{q}\leq\frac{q}{\mu}(f(x)-f^{*}).$ (2.4)
-Particularly, when $q=2$, it is called the $\mu$-quadratic growth condition.
-Moreover, the function $f$ is said to be local $\mu$-quadratic growth (with $\zeta$) if there exists a constant $\mu>0$ such that, for any $x\in[f\leq f^{*}+\zeta]\cap P$, (2.4) holds with $q=2$.
-The Hölder error bound condition shows sharpness bounds on the primal gap (see, e.g., [bolte17, alex17sharp]), which has also been extensively analyzed for Frank–Wolfe methods in [KDP2018, KDP2018jour].
-The convergence analysis for nonconvex optimization under the quadratic growth or Hölder error bound condition was established in [Davis2024-rm, Davis2018-bb, Davis2024-og].
-We distinguish between the global condition and the local condition of Definition 2.11.
-We employ the Hölder error bound condition for convex optimization while we use the local quadratic growth for nonconvex optimization.
-Next, we show in the following lemma that the Hölder error bound condition immediately provides a bound on the primal optimality gap, which will be useful for establishing convergence rates of the proposed algorithms as well as the Łojasiewicz inequality [Bolte2007-xr, Lojasiewicz1963-sm, Lojasiewicz1965-ks], provided $f$ is convex.
-Lemma 2.12 (Primal gap bound from Hölder error bound).
-Let $f$ be a convex function and satisfy the Hölder error bound condition with $q\geq 1$ and $\mu>0$.
+However, it is important to note that when $\rho$ becomes large, the assumptions of algorithms may not be satisfied (see also Theorems D.3 and 5.3).
+We have the key lemma for convergence analysis from weak convexity.
+Lemma B.2 (Primal gap, dual gap, and Frank–Wolfe gap for weakly convex functions).
+Let $x^{*}\in P$ an optimal solution of (1.1) and let $f^{*}=f(x^{*})$. Suppose that Assumptions 3.1iiiv, and that $f$ is $\rho$-weakly convex. Then for all $x\in P$, it holds:
+$f(x)-f^{*}\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}\leq\max_{v\in P}\langle\nabla f(x),x-v\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}.$ (B.1)
+Proof.
+The first inequality follows from the weak convexity of $f$ and the second follows from the maximality of $\langle\nabla f(x),x-v\rangle$. ∎
+B.2 Primal Gap Bound
+We show in the following lemma that the HEB condition immediately provides a bound on the primal optimality gap, which will be useful for establishing convergence rates of the proposed algorithms as well as the Łojasiewicz inequality [Bolte2007-xr, Lojasiewicz1963-sm, Lojasiewicz1965-ks], provided $f$ is convex.
+Lemma B.3 (Primal gap bound from Hölder error bound).
+Let $f$ be a convex function and satisfy the HEB condition with $q\geq 1$ and $\mu>0$.
 Let $x^{*}$ be the unique minimizer of $f$ over $P$.
 Then, the following argument holds in general, for all $x\in P$:
-$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{\frac{q}{q-1}},$ (2.5)
+$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{\frac{q}{q-1}},$ (B.2)
 or equivalently,
@@
 Proof.
-By first applying convexity and then the Hölder error bound condition for any $x^{*}\in\mathcal{X}^{*}$ with $f^{*}=f(x^{*})$ it holds:
+By first applying convexity and then the HEB condition for any $x^{*}\in\mathcal{X}^{*}$ with $f^{*}=f(x^{*})$ it holds:
 $f(x)-f^{*}$ $\leq\langle\nabla f(x),x-x^{*}\rangle$
@@
 $f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{\frac{q}{q-1}}.$
-Remark 2.13.
-With the remark from above, we can immediately relate the Hölder error bound condition to the Łojasiewicz inequality [Bolte2007-xr, Lojasiewicz1963-sm, Lojasiewicz1965-ks]; the Łojasiewicz inequality is used to establish convergence analysis [Bolte2007-xr].
-To this end, using (2.5), we estimate $\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\leq\frac{\|\nabla f(x)\|\|x-x^{*}\|}{\|x-x^{*}\|}=\|\nabla f(x)\|$ to then obtain the weaker condition:
-$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\|\nabla f(x)\|^{\frac{q}{q-1}}=c^{\frac{1}{\theta}}\|\nabla f(x)\|^{\frac{1}{\theta}},$ (2.6)
+Remark B.4.
+With the remark from above, we can immediately relate the HEB condition to the Łojasiewicz inequality [Bolte2007-xr, Lojasiewicz1963-sm, Lojasiewicz1965-ks]; the Łojasiewicz inequality is used to establish convergence analysis [Bolte2007-xr].
+To this end, using (B.2), we estimate $\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\leq\frac{\|\nabla f(x)\|\|x-x^{*}\|}{\|x-x^{*}\|}=\|\nabla f(x)\|$ to then obtain the weaker condition:
+$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\|\nabla f(x)\|^{\frac{q}{q-1}}=c^{\frac{1}{\theta}}\|\nabla f(x)\|^{\frac{1}{\theta}},$ (B.3)
 where $c=\left(\frac{q}{\mu}\right)^{\frac{1}{q}}$ and $\theta=\frac{q-1}{q}$.
-Inequality (2.6) is called the $c$-Łojasiewicz inequality with $\theta\in[0,1)$.
+Inequality (B.3) is called the $c$-Łojasiewicz inequality with $\theta\in[0,1)$.
 If $\mathcal{X}^{*}\subseteq\operatorname{ri}P$, then the two conditions are equivalent.
-However, if the optimal solution(s) are on the boundary of $P$ as is not infrequently the case, then the two conditions are not equivalent as $\|\nabla f(x)\|$ might not vanish for $x\in\mathcal{X}^{*}$, whereas $\langle\nabla f(x),x-x^{*}\rangle$ does, i.e., the Hölder error bound condition is tighter than the one induced by the Łojasiewicz inequality.
-The next lemma shows that we can also obtain primal gap bounds in the weakly convex case together with (local) quadratic growth. Its proof can be found in Appendix A.1.
-Lemma 2.14 (Primal gap bound from the quadratic growth).
+However, if the optimal solution(s) are on the boundary of $P$ as is not infrequently the case, then the two conditions are not equivalent as $\|\nabla f(x)\|$ might not vanish for $x\in\mathcal{X}^{*}$, whereas $\langle\nabla f(x),x-x^{*}\rangle$ does, i.e., the HEB condition is tighter than the one induced by the Łojasiewicz inequality.
+The next lemma shows that we can also obtain primal gap bounds in the weakly convex case together with (local) quadratic growth.
+Lemma B.5 (Primal gap bound from the quadratic growth).
 Let $f$ be a $\rho$-weakly convex function that satisfies the local $\mu$-quadratic growth condition such that $\rho\leq\mu$.
@@
 Then, the following holds: for all $x\in[f\leq f^{*}+\zeta]\cap P$,
-$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{2},$ (2.7)
+$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{2},$ (B.4)
 or equivalently,
 $\left(\frac{\mu}{2}\right)^{1/2}\left(1-\frac{\rho}{\mu}\right)(f(x)-f^{*})^{1/2}\leq\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}.$
-Remark 2.15.
-In the same vein as the discussion in Remark 2.13, we can immediately relate the local $\mu$-quadratic growth condition with $\zeta>0$ to the local Polyak–Łojasiewicz (PL) inequality.
-Using (2.7), for all $x\in[f\leq f^{*}+\zeta]\cap P$, we have
-$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\|\nabla f(x)\|^{2},$ (2.8)
+Proof.
+By first applying weak convexity and then the local quadratic growth condition for $x^{*}\in\mathcal{X}^{*}$ with $f(x^{*})=f^{*}$ and for all $x\in[f\leq f^{*}+\zeta]\cap P$, it holds:
+$f(x)-f^{*}$ $=f(x)-f(x^{*})$
+$\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}$
+$\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{\mu}\left(f(x)-f^{*}\right),$
+which implies
+$\left(1-\frac{\rho}{\mu}\right)\left(f(x)-f^{*}\right)\leq\langle\nabla f(x),x-x^{*}\rangle=\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\|x-x^{*}\|.$ (B.5)
+Using the local quadratic growth condition again, we have
+$\left(1-\frac{\rho}{\mu}\right)\left(f(x)-f^{*}\right)\leq\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\left(\frac{2}{\mu}(f(x)-f^{*})\right)^{1/2},$
+which provides
+$\left(\frac{\mu}{2}\right)^{1/2}\left(1-\frac{\rho}{\mu}\right)(f(x)-f^{*})^{1/2}\leq\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|},$
+or equivalently
+$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{2}.$
+Remark B.6.
+In the same vein as the discussion in Remark B.4, we can immediately relate the local $\mu$-quadratic growth condition with $\zeta>0$ to the local Polyak–Łojasiewicz (PL) inequality.
+Using (B.4), for all $x\in[f\leq f^{*}+\zeta]\cap P$, we have
+$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\|\nabla f(x)\|^{2},$ (B.6)
 which is equivalent to the PL inequality [Polyak1963-dr], also called the gradient dominance property [Combettes2020-jj], with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$.
@@
 Strongly convex functions satisfy the PL inequality [Braun2022-ew, Lemma 2.13].
-Note that we will not have the primal gap bound under the Hölder error bound condition and weak convexity because an inequality like (A.1) does not hold from the Hölder error bound condition.
-2.5 Scaling Inequalities, Geometric Hölder Error Bounds, and Contractions
+Note that we will not have the primal gap bound under the HEB condition and weak convexity because an inequality like (B.5) does not hold from the HEB condition.
+B.3 Scaling Inequalities, Geometric Hölder Error Bounds, and Contractions
 In the following, we will now bring things together to derive tools that will be helpful in establishing convergence rates.
@@
 We restate it here for the sake of completeness.
-Proposition 2.16 (Scaling inequality for inner optima from convexity [Braun2022-ew, Proposition 2.16]).
+Proposition B.7 (Scaling inequality from convexity when $x^{*}\in\operatorname{int}P$ [Braun2022-ew, Proposition 2.16]).
 Let $P\subset\mathbb{R}^{n}$ be a nonempty compact convex set.
@@
 When $f$ is not convex, assuming that $f$ is weakly convex and local quadratic growth, we have the scaling inequality for a nonconvex objective function.
-Proposition 2.17 (Scaling inequality for inner optima from weak convexity).
+Proposition B.8 (Scaling inequality from weak convexity when $x^{*}\in\operatorname{int}P$).
 Let $P\subset\mathbb{R}^{n}$ be a nonempty compact convex set.
@@
 We consider $x^{*}-rz$, where $z$ is a point with $\|z\|=1$ and $\langle\nabla f(x),z\rangle=\|\nabla f(x)\|$. It holds that
-$\langle\nabla f(x),v\rangle\leq\langle\nabla f(x),x^{*}-rz\rangle=\langle\nabla f(x),x^{*}\rangle-r\|\nabla f(x)\|.$ (2.9)
+$\langle\nabla f(x),v\rangle\leq\langle\nabla f(x),x^{*}-rz\rangle=\langle\nabla f(x),x^{*}\rangle-r\|\nabla f(x)\|.$ (B.7)
 From weak convexity, for all $x\in[f\leq f^{*}+\zeta]\cap P$, we have
@@
 $\langle\nabla f(x),x-x^{*}\rangle\geq\left(1-\frac{\rho}{\mu}\right)(f(x)-f^{*})\geq 0.$
-By rearranging (2.9) and using $\langle\nabla f(x),x-x^{*}\rangle\geq 0$, we obtain
+By rearranging (B.7) and using $\langle\nabla f(x),x-x^{*}\rangle\geq 0$, we obtain
 $\langle\nabla f(x),x-v\rangle\geq\langle\nabla f(x),x-x^{*}\rangle+r\|\nabla f(x)\|\geq r\|\nabla f(x)\|.$
@@
 The authors in [Lacoste-Julien2015-gb] defined a geometric distance-like constant of a polytope, known as the pyramidal width, to analyze the convergence of the away-step Frank–Wolfe algorithm (and other variants that use the pyramidal width) over polytopes.
-It can be interpreted as the minimal $\delta>0$ satisfying the following scaling inequality (2.10), which plays a central role in establishing convergence rates for the away-step FW algorithm; see [Braun2022-ew] for an in-depth discussion.
-Lemma 2.18 (Scaling inequality via pyramidal width [Braun2022-ew, Theorem 2.26][Lacoste-Julien2015-gb, Theorem 3]).
+It can be interpreted as the minimal $\delta>0$ satisfying the following scaling inequality (B.8), which plays a central role in establishing convergence rates for the away-step FW algorithm; see [Braun2022-ew] for an in-depth discussion.
+Lemma B.9 (Scaling inequality via pyramidal width [Braun2022-ew, Theorem 2.26][Lacoste-Julien2015-gb, Theorem 3]).
 Let $P\subset\mathbb{R}^{n}$ be a polytope and let $\delta$ denote the pyramidal width of $P$.
-Let $x\in P$, and let $\mathcal{S}$ denote any set of vertices of $P$ with $x\in\operatorname{conv}\mathcal{S}$.
+Let $x\in P$, and let $\mathcal{S}$ denote any set of vertices of $P$ with $x\in\operatorname{conv}\mathcal{S}$, where $\operatorname{conv}\mathcal{S}$ denotes the convex hull of $\mathcal{S}$.
 Let $\psi$ be any vector, so that we define $v^{\textup{FW}}=\operatorname*{argmin}_{v\in P}\langle\psi,v\rangle$ and $v^{\textup{A}}=\operatorname*{argmax}_{v\in\mathcal{S}}\langle\psi,v\rangle$.
 Then for any $y\in P$
-$\langle\psi,v^{\textup{A}}-v^{\textup{FW}}\rangle\geq\delta\frac{\langle\psi,x-y\rangle}{\|x-y\|}.$ (2.10)
-Note that Lemma 2.18 does not require the convexity of $f$, and we will use it for nonconvex optimization.
-When $\phi$ is $\sigma$-strongly convex and $\langle\psi,x-y\rangle\geq 0$, Lemma 2.18 implies
+$\langle\psi,v^{\textup{A}}-v^{\textup{FW}}\rangle\geq\delta\frac{\langle\psi,x-y\rangle}{\|x-y\|}.$ (B.8)
+Note that Lemma B.9 does not require the convexity of $f$, and we will use it for nonconvex optimization. When $\phi$ is $\sigma$-strongly convex and $\langle\psi,x-y\rangle\geq 0$, Lemma B.9 implies
 $\langle\psi,v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}\geq\delta^{2}\frac{\langle\psi,x-y\rangle^{2}}{\|x-y\|^{2}}\geq\delta^{2}\sigma\frac{\langle\psi,x-y\rangle^{2}}{2D_{\phi}(x,y)}.$
@@
 Geometric Hölder Error Bound Condition
-We will now introduce the more compact notion of geometric Hölder error bound condition, which simply combines the pyramidal width and the Hölder error bound condition of the function $f$; see also [Braun2022-ew, Lemma 2.27] for details when $q=2$.
-Lemma 2.19 (Geometric Hölder Error Bound).
-Let $P$ be a polytope with pyramidal width $\delta>0$ and let $f$ be a convex function and satisfy the Hölder error bound condition with $v^{\textup{FW}}=\operatorname*{argmin}_{v\in P}\langle\nabla f(x),v\rangle$ and $v^{\textup{A}}=\operatorname*{argmax}_{v\in\mathcal{S}}\langle\nabla f(x),v\rangle$ with $\mathcal{S}\subseteq\operatorname{Vert}{P}$, so that $x\in\operatorname{conv}{\mathcal{S}}$, we have
+We will now introduce the more compact notion of geometric Hölder error bound condition, which simply combines the pyramidal width and the HEB condition of the function $f$; see also [Braun2022-ew, Lemma 2.27] for details when $q=2$.
+Lemma B.10 (Geometric Hölder Error Bound).
+Let $P$ be a polytope with pyramidal width $\delta>0$ and let $f$ be a convex function and satisfy the HEB condition with $v^{\textup{FW}}=\operatorname*{argmin}_{v\in P}\langle\nabla f(x),v\rangle$ and $v^{\textup{A}}=\operatorname*{argmax}_{v\in\mathcal{S}}\langle\nabla f(x),v\rangle$ with $\mathcal{S}\subseteq\operatorname{Vert}{P}$, so that $x\in\operatorname{conv}{\mathcal{S}}$, we have
 $f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle}{\delta}\right)^{\frac{q}{q-1}}.$
 Proof.
-Combining (2.10) with Lemma 2.12 for $\psi=\nabla f(x)$, we have
+Combining (B.8) with Lemma B.3 for $\psi=\nabla f(x)$, we have
 $f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{\frac{q}{q-1}}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle}{\delta}\right)^{\frac{q}{q-1}}.$
@@
 Besides scaling inequalities and the geometric Hölder error bound, we also utilize the following lemma for convex and nonconvex optimization, which allows us to turn a contraction into a convergence rate.
-Lemma 2.20 (From contractions to convergence rates [Braun2022-ew, Lemma 2.21]).
+Lemma B.11 (From contractions to convergence rates [Braun2022-ew, Lemma 2.21]).
 Let $\{h_{t}\}_{t}$ be a decreasing sequence of positive numbers and $c_{0}$, $c_{1}$, $c_{2}$, $\theta_{0}$ be positive numbers with $c_{1}<1$ such that $h_{1}\leq c_{0}$ and $h_{t}-h_{t+1}\geq h_{t}\min\{c_{1},c_{2}h_{t}^{\theta_{0}}\}$ for $t\geq 1$, then
@@
 In particular, we have $h_{t}\leq\epsilon$ if $t\geq t_{0}+\frac{1}{\theta_{0}c_{2}\epsilon^{\theta_{0}}}-\frac{1}{\theta_{0}c_{1}}$ and $\epsilon\leq(c_{1}/c_{2})^{1/\theta_{0}}$.
-3 Proposed Algorithms
-Throughout this paper, we make the following assumptions.
-Assumption 3.1.
-  1. (i)
-    $\phi\in\mathcal{G}(C)$ with $\operatorname{cl}{C}=\operatorname{cl}{\operatorname{dom}\phi}$ is strictly convex on $C\equiv\operatorname{int}\operatorname{dom}\phi$.
-  2. (ii)
-    $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is a proper and lower semicontinuous function with $\operatorname{dom}\phi\subset\operatorname{dom}f$, which is $\mathcal{C}^{1}$ on $C$.
-  3. (iii)
-    The pair $(f,\phi)$ is $L$-smad on $P$.
-  4. (iv)
-    $P\subset\mathbb{R}^{n}$ is a nonempty compact convex set with $P\subset C$.
-Assumption 3.1(i)-(iii) are standard for Bregman-type algorithms [Bolte2018-zt, Takahashi2024-wz].
-The strict convexity ensures the existence of $\nu$ by Lemma 2.5 and is satisfied by many kernel generating distances (see Example 2.3).
-Assumption 3.1(iv) is also standard for FW algorithms, and the condition $P\subset C$ is natural ensuring that the Bregman distance $D_{\phi}$ is well-defined on $P$.
-In what follows, let $x^{*}\in\operatorname*{argmin}_{x\in P}f(x)$ and let $f^{*}=f(x^{*})$. We recall a key property for analyzing the convergence rates of FW algorithms in convex optimization.
-Lemma 3.2 (Primal gap, dual gap, and Frank–Wolfe gap [Pokutta2024-qu, Lemma 4.1]).
-Let $x^{*}\in P$ be an optimal solution of (1.1) and let $f^{*}=f(x^{*})$. Suppose that Assumptions 3.1(ii), (iv) and that $f$ is convex. Then for all $x\in P$, it holds that
-$f(x)-f^{*}\leq\langle\nabla f(x),x-x^{*}\rangle\leq\max_{v\in P}\langle\nabla f(x),x-v\rangle.$ (3.1)
-3.1 Adaptive Bregman Step-Size Strategy
-Assume that $(f,\phi)$ is $L$-smad. Substituting $x_{t+1}=(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$ with $x_{t},v_{t}\in P$ for $x_{+}$ in Lemma 2.8 provides
-$f(x_{t})-f(x_{t+1})\geq\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle-L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t}).$ (3.2)
-Using the optimality condition of the right-hand side of (3.2) in terms of $\gamma_{t}$, we find
-$\gamma_{t}=\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}}\quad\text{and}\quad f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle^{1+1/\nu}}{(L(1+\nu)D_{\phi}(v_{t},x_{t}))^{1/\nu}}.$
-Theoretically, when $\gamma_{t}\in[0,1]$, the Frank–Wolfe step $x_{t+1}=(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}\in P$ is well-defined.
-We update $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},\gamma_{\max}\right\}$ with some $\gamma_{\max}\in\mathbb{R}$ (usually set $\gamma_{\max}=1$).
-This step-size strategy provides the progress lemma.
-Note that the following lemma does not require the convexity of $f$.
-Lemma 3.3 (Progress lemma from the Bregman short step-size).
-Suppose that Assumption 3.1 holds. Define $x_{t+1}=(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$. Then if $x_{t+1}\in\operatorname{dom}f$,
-$f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\quad\text{for}\quad 0\leq\gamma_{t}\leq\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}}.$
-Proof.
-Recalling (3.2), we have
-$f(x_{t})-f(x_{t+1})$ $\geq\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle-L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t})$
-$\geq\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle-\frac{\gamma_{t}}{1+\nu}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle$
-$=\frac{\nu}{1+\nu}\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle,$
-where the second inequality holds because of an upper bound of $\gamma_{t}$. ∎
-If $f$ is convex, using Lemma 3.3, we have
-$f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq\frac{\nu}{1+\nu}\gamma_{t}(f(x_{t})-f^{*}),$
-which implies that the primal progress is at least $\frac{\nu}{1+\nu}$ of the Frank–Wolfe gap and of the primal gap multiplied by $\gamma_{t}$. If $\phi=\frac{1}{2}\|\cdot\|^{2}$, we have $\nu=1$ and
-$\gamma_{t}=\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L\|v_{t}-x_{t}\|^{2}},$
-which is often called the (Euclidean) short step-size. Although the short step-size strategy does not require line searches, it requires knowledge of the value of $L$.
-The exact value or the tight upper bound of $L$ is often unknown, however, the algorithm’s performance heavily depends on it; an underestimation of $L$ might lead to non-convergence, and an overestimation of $L$ might lead to slow convergence.
-Moreover, a worst-case $L$ might be too conversative for regimes where the function is better behaved.
-Due to all these reasons, Pedregosa et al.
-[Pedregosa2020-ay] proposed an adaptive step-size strategy for FW algorithms in the case where $\phi=\frac{1}{2}\|\cdot\|^{2}$.
-Vyguzov and Stonyakin [Vyguzov2024-tp] proposed a variant using Bregman distances.
-We have concurrently and independently formulated an improved version similar (but different in details) to [Vyguzov2024-tp], which includes updates to $\nu$ and uses different coefficients for $\gamma_{t}$.
-We present our algorithm in Algorithm 2.
-Output : Updated estimates $\tilde{L}^{*}$ and $\tilde{\nu}^{*}$, step-size $\gamma$
-[…]
-2      Choose $\beta<1$, $\eta\leq 1$, and $\tau>1$
-3     $\kappa\leftarrow 1$
-4     $M\leftarrow\eta\tilde{L}$
-5     loop
-6           $\gamma\leftarrow\min\left\{\left(\frac{\langle\nabla f(x),x-v\rangle}{M(1+\kappa)D_{\phi}(v,x)}\right)^{\frac{1}{\kappa}},\gamma_{\max}\right\}$
-7          if $f((1-\gamma)x+\gamma v)-f(x)-\gamma\langle\nabla f(x),v-x\rangle\leq M\gamma^{1+\kappa}D_{\phi}(v,x)$ then
-8                $\tilde{L}^{*}\leftarrow M$
-9               $\tilde{\nu}^{*}\leftarrow\kappa$
-10               return $\tilde{L}^{*}$, $\tilde{\nu}^{*}$, $\gamma$
-11                $M\leftarrow\tau M$
-12               $\kappa\leftarrow\beta\kappa$
-13
-Algorithm 2 Adaptive Bregman step-size strategy
-This step-size strategy can be used as a drop-in replacement.
-For example, inserting $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},1)$ between lines 1 and 1 in Algorithm 1, we obtain that Algorithm 2 searches $L$ and $\nu$ until
-$f((1-\gamma)x+\gamma v)-f(x)-\gamma\langle\nabla f(x),v-x\rangle\leq M\gamma^{1+\kappa}D_{\phi}(v,x)$ (3.3)
-holds for the approximation values $M$ and $\kappa$ of $L$ and $\nu$, respectively.
-Remark 3.4 (Well-definedness and termination of Algorithm 2).
-By the $L$-smad property of $(f,\phi)$ and Lemma 2.8, we know that (3.3) holds for all $M\geq L$ and $\nu\geq\kappa>0$.
-Therefore, the loop in line 2 in Algorithm 2 is well-defined and guaranteed to terminate.
-Thus it suffices to update $\kappa$ in line 2 only if $D_{\phi}((1-\gamma)x+\gamma v,x)>\gamma^{1+\kappa}D_{\phi}(v,x)$ holds.
-Thus, $\kappa$ cannot become too small.
-3.2 The Away-Step Frank–Wolfe Algorithm
-In the case where $P$ is a polytope, the classical FW algorithm might zigzag when approaching the optimal face and, in consequence, converge slowly.
-To overcome this drawback, Wolfe [Wolfe1970-ax] introduced the away-step Frank–Wolfe algorithm and away steps, which allow the algorithm to move away from vertices in a convex combination of $x_{t}$, which can effectively shortcircuit the zigzagging.
-The convergence properties of this algorithm were unknown for a long time, and it was only quite recently that Lacoste-Julien and Jaggi [Lacoste-Julien2015-gb] established the linear convergence of the away-step Frank-Wolfe algorithm.
-Inspired by [GueLat1986-bg, Lacoste-Julien2015-gb, Wolfe1970-ax], we propose a variant of the away-step Frank–Wolfe algorithm utilizing Bregman distances as given in Algorithm 3.
-For the following discussion, we introduce the following notions.
-In the same way [Braun2022-ew], we define an active set $\mathcal{S}\subset\operatorname{Vert}P$ and the away vertex as $v_{t}^{\textup{A}}\in\operatorname*{argmax}_{v\in\mathcal{S}}\langle\nabla f(x_{t}),v\rangle$, where $x_{t}$ is a (strict) convex combination of elements in $\mathcal{S}$, i.e., $x_{t}=\sum_{v\in\mathcal{S}}\lambda_{v}x_{t}$ with $\lambda_{v}>0$ for all $v\in\mathcal{S}$ and $\sum_{v\in\mathcal{S}}\lambda_{v}=1$.
-Algorithm 3 updates $\nu$ until $D_{\phi}((1-\gamma)x_{t}+\gamma v_{t},x_{t})\leq\gamma^{1+\nu}D_{\phi}(v,x)$ holds in lines 3-3.
-If $\nu$ is known, one only updates $\gamma_{t}$.
-If $\nu$ and $L$ are unknown, we can use $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},\gamma_{t,\max})$ with $\gamma\leftarrow\min\left\{\left(\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{M(1+\kappa)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\kappa}},\gamma_{t,\max}\right\}$ instead of line 2 in Algorithm 2.
-When $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$, Algorithm 3 corresponds to the classical away-step FW algorithm.
-Input: Initial point $x_{0}\in\operatorname*{argmin}_{v\in P}\langle\nabla f(x),v\rangle$ for $x\in P$, $\beta<1$
-1 $\mathcal{S}_{0}\leftarrow\{x_{0}\}$, $\lambda_{x_{0},0}\leftarrow 1$
-2for $t=0,\ldots$ do
-3      $v_{t}^{\textup{FW}}\leftarrow\operatorname*{argmin}_{v\in P}\langle\nabla f(x_{t}),v\rangle$
-4     $v_{t}^{\textup{A}}\leftarrow\operatorname*{argmax}_{v\in\mathcal{S}_{t}}\langle\nabla f(x_{t}),v\rangle$
-5     if $\langle\nabla f(x_{t}),x_{t}-v_{t}^{\textup{FW}}\rangle\geq\langle\nabla f(x_{t}),v_{t}^{\textup{A}}-x_{t}\rangle$ then
-6           $v_{t}\leftarrow v_{t}^{\textup{FW}}$, $d_{t}\leftarrow x_{t}-v_{t}^{\textup{FW}}$, $\gamma_{t,\max}\leftarrow 1$
-7           else
-8                $v_{t}\leftarrow v_{t}^{\textup{A}}$, $d_{t}\leftarrow v_{t}^{\textup{A}}-x_{t}$, $\gamma_{t,\max}\leftarrow\frac{\lambda_{v_{t}^{\textup{A}},t}}{1-\lambda_{v_{t}^{\textup{A}},t}}$
-9                $\nu\leftarrow 1$
-10               while $D_{\phi}((1-\gamma)x_{t}+\gamma v_{t},x_{t})\leq\gamma^{1+\nu}D_{\phi}(v_{t},x_{t})$ do
-11                     $\gamma\leftarrow\min\left\{\left(\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},\gamma_{t,\max}\right\}$, $\nu\leftarrow\beta\nu$
-12
-13                    $\gamma_{t}\leftarrow\gamma$, $x_{t+1}\leftarrow x_{t}-\gamma_{t}d_{t}$
-14                    if $\langle\nabla f(x_{t}),x_{t}-v_{t}^{\textup{FW}}\rangle\geq\langle\nabla f(x_{t}),v_{t}^{\textup{A}}-x_{t}\rangle$ then
-15                          $\lambda_{v,t+1}\leftarrow(1-\gamma_{t})\lambda_{v,t}$ for all $v_{t}\in\mathcal{S}_{t}\setminus\{v_{t}^{\textup{FW}}\}$
-16                         $\lambda_{v_{t}^{\textup{FW}},t+1}\leftarrow\begin{cases}\gamma_{t}&\text{if}\ v_{t}^{\textup{FW}}\notin\mathcal{S}_{t}\\ (1-\gamma_{t})\lambda_{v_{t}^{\textup{FW}},t}+\gamma_{t}&\text{if}\ v_{t}^{\textup{FW}}\in\mathcal{S}_{t}\end{cases}$
-17                         $\mathcal{S}_{t+1}\leftarrow\begin{cases}\mathcal{S}_{t}\cup\{v_{t}^{\textup{FW}}\}&\text{if}\ \gamma_{t}<1\\ \{v_{t}^{\textup{FW}}\}&\text{if}\ \gamma_{t}=1\end{cases}$
-18                          else
-19                               $\lambda_{v,t+1}\leftarrow(1+\gamma_{t})\lambda_{v,t}$ for all $v_{t}\in\mathcal{S}_{t}\setminus\{v_{t}^{\textup{A}}\}$
-20                              $\lambda_{v_{t}^{\textup{A}},t+1}\leftarrow(1+\gamma_{t})\lambda_{v_{t}^{\textup{A}},t}-\gamma_{t}$
-21                              $\mathcal{S}_{t+1}\leftarrow\begin{cases}\mathcal{S}_{t}\setminus\{v_{t}^{\textup{A}}\}&\text{if}\ \lambda_{v_{t}^{\textup{A}},t+1}=0\\ \mathcal{S}_{t}&\text{if}\ \lambda_{v_{t}^{\textup{A}},t+1}>0\end{cases}$
-22
-Algorithm 3 Away-Step Frank–Wolfe algorithm with the Bregman distance
-4 Convergence Analysis for Convex Optimization
-In this section, we assume that $f$ is convex.
-Assumption 4.1.
-The objective function $f$ is convex.
-4.1 Sublinear Convergence
+Appendix C Proof of Convergence Analysis for Convex Optimization
+We would also like to stress that while we formulate some of the results for the case that $x^{*}\in\operatorname{int}P$, the results can be extended to the case that $x^{*}\in\operatorname{ri}P$, i.e., the relative interior of $P$, which we did not do for the sake of clarity.
+Basically, the analysis is performed in the affine space spanned by the optimal face of $P$ in that case; the interested reader is referred to [Braun2022-ew] for details on how to extend the results.
+C.1 Sublinear Convergence
 We establish a sublinear convergence rate of the FW algorithm under the smooth adaptable property.
-It is a similar result to [Vyguzov2024-tp, Theorem 1].
+It is a similar result to [Vyguzov2025-ha, Theorem 1].
 We conducted its proof following [Braun2022-ew, Jaggi2013-oz, Pokutta2024-qu].
 The FW algorithm uses the open loop step-size, i.e., $\gamma_{t}=\frac{2}{2+t}$ in the following theorem.
-Theorem 4.2 (Primal convergence of the Frank–Wolfe algorithm).
+Theorem C.1 (Primal convergence of the Frank–Wolfe algorithm).
 Suppose that Assumptions 3.1 and 4.1 hold.
-Let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$ characterized by the Bregman distance $D_{\phi}$.
+Let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$ characterized by the Bregman distance $D_{\phi}$ and let $\nu\in(0,1]$.
 Consider the iterates of Algorithm 1 with the open loop step-size, i.e., $\gamma_{t}=\frac{2}{2+t}$.
 Then, it holds that, for all $t\geq 1$,
-$f(x_{t})-f^{*}\leq\frac{2^{1+\nu}LD^{2}}{(t+2)^{\nu}},$ (4.1)
-and hence for any accuracy $\epsilon>0$ we have $f(x_{t})-f^{*}\leq\epsilon$ for all $t\geq\left(\frac{2^{1+\nu}LD^{2}}{\epsilon}\right)^{\frac{1}{\nu}}$.
+$f(x_{t})-f^{*}\leq\frac{2^{1+\nu}LD^{2}}{(t+2)^{\nu}},$ (C.1)
+and hence for any accuracy $\epsilon>0$ we have $f(x_{t})-f^{*}\leq\epsilon$ for all $t\geq(2^{1+\nu}LD^{2}/\epsilon)^{1/\nu}$.
 Proof.
@@
 where the last inequality holds due to $(3+t)^{\nu}(1+t)\leq(2+t)^{1+\nu}$ with $0<\nu\leq 1$. ∎
-If $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$ in (4.1), we have
-$f(x_{t})-f^{*}\leq\frac{4LD^{2}}{t+2},$
-which is the same as a sublinear convergence rate of the classical FW algorithm.
-Remark 4.3.
-While the convergence rate (4.1) is the same as [Vyguzov2024-tp], Vyguzov and Stonyakin assume that the triangle scaling property holds for $D_{\phi}$.
-In contrast, we require significantly weaker assumptions: it is enough to assume that $\phi$ is strictly convex due to Lemma 2.5 in order to establish Theorem 4.2.
-4.2 Accelerated Convergence
-We will now establish accelerated convergence rates better than $O(1/t)$ up to linear convergence depending on the choice of parameters.
-First, we establish the accelerated convergence of Algorithm 1 with the Bregman short step-size, i.e.,
-$\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},1\right\},$
-in the case where the optimal solution lies in the relative interior.
-Theorem 4.4 (Accelerated convergence for inner optima).
+If $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$ in (C.1), we have $f(x_{t})-f^{*}\leq 4LD^{2}/(t+2)$, which is the same as a sublinear convergence rate of the classical FW algorithm.
+C.2 Proof of Theorem 4.2
+Because Algorithm 2 is well-defined (see Remark 3.5), the convergence result of the FW algorithm using the adaptive step-size strategy (Algorithm 2) is essentially the same as the one that uses Bregman short steps (Theorems 4.2 and 5.2); up to small errors arising from the approximation whose precise analysis we skip for the sake of brevity.
+Theorem C.2 (Linear convergence FW algorithm with short step-size or adaptive step-size).
 Suppose that Assumptions 3.1 and 4.1 hold.
-Let $f$ satisfy the Hölder error bound condition with $q\geq 1+\nu$ and $\mu>0$ and $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$.
+Let $\nu>0$ and let $f$ satisfy the HEB condition with $q\geq 1+\nu$ and $\mu>0$ and $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$.
 Assume that there exists a minimizer $x^{*}\in\operatorname{int}P$, i.e., there exists an $r>0$ with $B(x^{*},r)\subset P$.
-Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},1\right\}$.
-Then, it holds that
+Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu},1\right\}$ or Algorithm 2.
+Then, it holds that, for all $t\geq 1$,
 $f(x_{t})-f^{*}\leq\begin{cases}h_{t}\leq\max\left\{\frac{1}{1+\nu},1-\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{c^{1+1/\nu}D^{2/\nu}}\right\}^{t-1}LD^{2}&\text{if}\ q=1+\nu,\\ \frac{LD^{2}}{(1+\nu)^{t-1}}&\text{if}\ 1\leq t\leq t_{0},q>1+\nu,\\ \frac{(L(1+\nu)(c/r)^{1+\nu}D^{2})^{q/(q-1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}(t-t_{0})\right)^{\nu q/(q-1-\nu)}}=\mathcal{O}(1/t^{\nu q/(q-1-\nu)})&\text{if}\ t\geq t_{0},q>1+\nu,\end{cases}$
-for all $t\geq 1$ where $c=(q/\mu)^{1/q}$ and
+where $c=(q/\mu)^{1/q}$ and
 $t_{0}:=\max\left\{\left\lfloor\log_{\frac{1}{1+\nu}}\left(\frac{(L(1+\nu(c/r)^{1+\nu}D^{2})^{q/(q-1-\nu)}}{LD^{2}}\right)\right\rfloor+2,1\right\}.$
@@
 $\geq\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}h_{t}^{\frac{(1+\nu)(q-1)}{\nu q}},$
-where the second inequality holds from $\langle f(x_{t}),x_{t}-v_{t}\rangle\geq r\|\nabla f(x_{t})\|$ in Proposition 2.16, and the last inequality holds because $f$ satisfies the $c$-Łojasiewicz inequality with $c=(q/\mu)^{1/q}$ (see Remark 2.13), and $M:=(L(1+\nu))^{1/\nu}$.
+where the second inequality holds from $\langle f(x_{t}),x_{t}-v_{t}\rangle\geq r\|\nabla f(x_{t})\|$ in Proposition B.7, and the last inequality holds because $f$ satisfies the $c$-Łojasiewicz inequality with $c=(q/\mu)^{1/q}$ (see Remark B.4), and $M:=(L(1+\nu))^{1/\nu}$.
 (ii) $\gamma_{t}=1$: Using (3.1), we have
@@
 $h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu},\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}\right\}\cdot h_{t}.$
-This inequality and the initial bound $f(x_{1})-f^{*}\leq LD^{2}$ due to Lemma 2.7 imply
+This inequality and the initial bound $f(x_{1})-f^{*}\leq LD^{2}$ due to Lemma A.4 imply
 $h_{t}\leq\max\left\{\frac{1}{1+\nu},1-\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}\right\}^{t-1}LD^{2}.$
@@
 $h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu},\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}h_{t}^{\frac{q-1-\nu}{\nu q}}\right\}\cdot h_{t}.$
-Using $f(x_{1})-f^{*}\leq LD^{2}$ and Lemma 2.20 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}$, $c_{2}=\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}$, and $\theta_{0}=\frac{q-1-\nu}{\nu q}>0$ from $q>1+\nu$, we have the claim.
+Using $f(x_{1})-f^{*}\leq LD^{2}$ and Lemma B.11 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}$, $c_{2}=\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}$, and $\theta_{0}=\frac{q-1-\nu}{\nu q}>0$ from $q>1+\nu$, we have the claim.
 ∎
-Because Algorithm 2 is well-defined (see Remark 3.4), the convergence result of the FW algorithm using the adaptive step-size strategy (Algorithm 2) is essentially the same as the one from Theorem 4.4 which uses Bregman short steps; up to small errors arising from the approximation whose precise analysis we skip for the sake of brevity.
-Note that Vyguzov and Stonyakin [Vyguzov2024-tp, Theorem 2] also established linear convergence, assuming that $f$ is relatively strongly convex, the optimal solution belongs to int $P$, the angle condition holds, and the triangle scaling property holds.
-However, these requirements are stronger than ours here, and they do not establish sublinear but accelerated rates for intermediate parameter regimes.
-Next, we establish the linear convergence of Algorithm 3.
-Recall that the pyramidal width is the minimal $\delta>0$ satisfying Lemma 2.18 (see [Braun2022-ew, Lemma 2.26] or [Lacoste-Julien2015-gb, Theorem 3] for an in-depth discussion).
-We make the following assumption:
-Assumption 4.5.
-The kernel generating distance $\phi$ is $\sigma$-strongly convex.
-Under Assumption 4.5, it holds that $\delta\leq D$, and this will be important in establishing convergence rates of Algorithm 3.
-Theorem 4.6 (Accelerated convergence of the away-step FW algorithm).
-Suppose that Assumptions 3.14.1, and 4.5 hold.
-Let $P\subset\mathbb{R}^{n}$ be a polytope and $f$ satisfy the Hölder error bound condition with $q>1+\nu$ or $(\nu,q)=(1,2)$.
+C.3 Proof of Theorem 4.4
+Theorem C.3 (Linear convergence of the away-step FW algorithm).
+Suppose that Assumptions 3.14.1, and 4.3 hold.
+Let $\nu>0$ and let $P\subset\mathbb{R}^{n}$ be a polytope and $f$ satisfy the HEB condition with $q>1+\nu$ or $(\nu,q)=(1,2)$.
 The convergence rate of Algorithm 3 is linear: for all $t\geq 1$
-$f(x_{t})-f^{*}\leq\begin{cases}\left(1-\frac{\mu}{32L}\frac{\delta^{2}}{D^{2}}\right)^{\lceil(t-1)/2\rceil}LD^{2}&\text{if}\ (\nu,q)=(1,2),\\ \frac{1}{(1+\nu)^{\lceil(t-1)/2\rceil}}LD^{2}&\text{if}\ 1\leq t\leq t_{0},q>1+\nu,\\ \frac{(L(1+\nu)D^{2}/((\mu/q)^{1/q}\delta/2)^{1+\nu})^{q/(q-1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}\lceil(t-t_{0})/2\rceil\right)^{\nu q/(q-1-\nu)}}=\mathcal{O}(1/t^{\nu q/(q-1-\nu)})&\text{if}\ t\geq t_{0},q>1+\nu,\end{cases}$
-where $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ and $\delta$ are the diameter and the pyramidal width of the polytope $P$, respectively, and
+$f(x_{t})-f^{*}\leq\begin{cases}\left(1-\frac{\mu}{32L}\frac{\delta^{2}}{D^{2}}\right)^{\lceil(t-1)/2\rceil}LD^{2}&\text{if}\ (\nu,q)=(1,2),\\ \frac{1}{(1+\nu)^{\lceil(t-1)/2\rceil}}LD^{2}&\text{if}\ 1\leq t\leq t_{0},q>1+\nu,\\ \frac{(L(1+\nu)D^{2}c/(\delta/2)^{1+\nu})^{q/(q-1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}\lceil(t-t_{0})/2\rceil\right)^{\nu q/(q-1-\nu)}}=\mathcal{O}(1/t^{\nu q/(q-1-\nu)})&\text{if}\ t\geq t_{0},q>1+\nu,\end{cases}$
+where $c:=(q/\mu)^{1/q}$ $D$ and $\delta$ are the diameter and the pyramidal width of the polytope $P$, respectively, and
 $t_{0}:=\max\left\{\left\lfloor\log_{\frac{1}{1+\nu}}\frac{(L(1+\nu)D^{2}/((\mu/q)^{1/q}\delta/2)^{1+\nu})^{q/(q-1-\nu)}}{LD^{2}}\right\rfloor+2,1\right\}.$
 Proof.
-By using the induced guarantee on the primal gap via Lemma 2.19, we have
-$h_{t}=f(x_{t})-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x_{t}),v_{t}^{\textup{A}}-v_{t}^{\textup{FW}}\rangle}{\delta}\right)^{\frac{q}{q-1}}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{2\langle\nabla f(x_{t}),d_{t}\rangle}{\delta}\right)^{\frac{q}{q-1}},$ (4.2)
+By using the induced guarantee on the primal gap via Lemma B.10, we have
+$h_{t}=f(x_{t})-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x_{t}),v_{t}^{\textup{A}}-v_{t}^{\textup{FW}}\rangle}{\delta}\right)^{\frac{q}{q-1}}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{2\langle\nabla f(x_{t}),d_{t}\rangle}{\delta}\right)^{\frac{q}{q-1}},$ (C.2)
 where the last inequality holds because $d_{t}$ is either $x_{t}-v_{t}^{\textup{FW}}$ or $v_{t}^{\textup{A}}-x_{t}$ with $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle/2$ in Lines 3 and 3 of Algorithm 3.
@@
 $=\min\left\{\frac{\nu h_{t}}{1+\nu}\gamma_{t,\max},\frac{\nu}{1+\nu}\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}h_{t}^{\frac{(1+\nu)(q-1)}{\nu q}}\right\}$
-where the first inequality holds due to Lemma 3.3 with $\gamma_{t,\max}=1$ (Frank–Wolfe steps) and $\gamma_{t,\max}=\frac{\lambda_{v_{t}^{\textup{A}},t}}{1-\lambda_{v_{t}^{\textup{A}},t}}$ (away steps), and the second inequality holds because of (4.2) and $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq h_{t}$.
+where the first inequality holds due to Lemma 3.4 with $\gamma_{t,\max}=1$ (Frank–Wolfe steps) and $\gamma_{t,\max}=\frac{\lambda_{v_{t}^{\textup{A}},t}}{1-\lambda_{v_{t}^{\textup{A}},t}}$ (away steps), and the second inequality holds because of (C.2) and $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq h_{t}$.
 In $(\nu,q)=(1,2)$, for Frank–Wolfe steps, $\gamma_{t,\max}=1\geq\mu\delta^{2}/LD^{2}\geq\mu\delta^{2}/32LD^{2}$.
@@
 In $q>1+\nu$, we have $h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu},\frac{\nu}{1+\nu}\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}h_{t}^{\frac{q-1-\nu}{\nu q}}\right\}\cdot h_{t}$ for at least half of the iterations (Frank–Wolfe steps) and $h_{t+1}\leq h_{t}$ for the rest (away steps).
-The initial bound $f(x_{1})-f^{*}\leq LD^{2}$ holds generally for the Frank–Wolfe algorithm (see the proof of Theorem 4.2 and [Braun2022-ew, Remark 2.4]).
-We use Lemma 2.20 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}$, $c_{2}=c_{1}\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}$, and $\theta_{0}=\frac{q-1-\nu}{\nu q}>0$ from $q>1+\nu$ and obtain the claim.
+The initial bound $f(x_{1})-f^{*}\leq LD^{2}$ holds generally for the Frank–Wolfe algorithm (see the proof of Theorem C.1 and [Braun2022-ew, Remark 2.4]).
+We use Lemma B.11 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}$, $c_{2}=c_{1}\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}$, and $\theta_{0}=\frac{q-1-\nu}{\nu q}>0$ from $q>1+\nu$ and obtain the claim.
 ∎
-Remark 4.7 (Compatibility of parameters).
+Remark C.4 (Compatibility of parameters).
 The condition $q>1+\nu$ is necessary in case that $(\nu,q)\neq(1,2)$. The reason is as follows. In the case $q=1+\nu$, for Frank–Wolfe steps, we have
-$\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}=\left(\frac{\mu}{L}\right)^{1/\nu}\left(\frac{\delta}{D}\right)^{\frac{1+\nu}{\nu}}\frac{1}{2^{\frac{1+\nu}{\nu}}(1+\nu)^{2/\nu}D^{(1-\nu)/\nu}},$ (4.3)
+$\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}=\left(\frac{\mu}{L}\right)^{1/\nu}\left(\frac{\delta}{D}\right)^{\frac{1+\nu}{\nu}}\frac{1}{2^{\frac{1+\nu}{\nu}}(1+\nu)^{2/\nu}D^{(1-\nu)/\nu}},$ (C.3)
 which might be greater than 1 because $\frac{1}{D^{(1-\nu)/\nu}}\geq 1$ if $D<1$ and $\nu$ are small enough.
-In order to make (4.3) smaller than 1, $\nu$ should be 1, i.e., $(\nu,q)=(1,2)$.
+In order to make (C.3) smaller than 1, $\nu$ should be 1, i.e., $(\nu,q)=(1,2)$.
 When $\nu=1$ and $q\neq 2$, it reduces to the $q>1+\nu$ case.
-5 Convergence Analysis for Nonconvex Optimization
-In this section, we consider a nonconvex objective function.
-We establish global sublinear convergence to a stationary point of (1.1), i.e., $x\in P$ such that $\max_{v\in P}\langle\nabla f(x),x-v\rangle=0$.
-We also obtain local sublinear and local linear convergence to a minimizer of (1.1).
-5.1 Global Convergence
+Appendix D Proof of Convergence Analysis for Nonconvex Optimization
+D.1 Global Convergence
 We show that Algorithm 1 with $\gamma_{t}=\gamma:=1/(T+1)^{\frac{1}{1+\nu}}$ globally converges to a stationary point, where $T\in\mathbb{N}$ is the number of iterations.
 Its proof is inspired by [Lacoste-Julien2016-mj, Theorem 1] and [Pokutta2024-qu, Theorem 4.7] and identical to those for the case when $\phi=\frac{1}{2}\|\cdot\|^{2}$.
-Theorem 5.1 (Global sublinear convergence for nonconvex optimization).
+Theorem D.1 (Global sublinear convergence for nonconvex optimization).
 Suppose that Assumption 3.1 holds.
-Let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$ characterized by $D_{\phi}$ and let $T\in\mathbb{N}$.
+Let $\nu>0$ and $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$ characterized by $D_{\phi}$ and let $T\in\mathbb{N}$.
 Then, the iterates of the FW algorithm with $\gamma_{t}=\gamma:=1/(T+1)^{\frac{1}{1+\nu}}$ satisfy
@@
 which is the same rate as [Lacoste-Julien2016-mj, Theorem 1].
-5.2 Local Convergence
-Next, we show that the FW algorithm converges to a minimizer $x^{*}\in\operatorname*{argmin}_{x\in P}f(x)$ when an initial point is close enough to $x^{*}$.
-We need the weak convexity of $f$ for its proof.
-Assumption 5.2.
-  1. (i)
-    $f$ is $\rho$-weakly convex.
-  2. (ii)
-    $f$ is local $\mu$-quadratic growth with $\zeta>0$.
-Under Assumption 5.2(i), the following primal gap lemma holds, whose proof is immediate.
-Lemma 5.3 (Primal gap, dual gap, and Frank–Wolfe gap for weakly convex functions).
-Suppose that Assumptions 3.1(ii)(iv), and 5.2(i). Let $x^{*}\in P$ an optimal solution of (1.1) and let $f^{*}=f(x^{*})$. For all $x\in P$, it holds:
-$f(x)-f^{*}$ $\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}$
-$\leq\max_{v\in P}\langle\nabla f(x),x-v\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}.$ (5.1)
-Proof.
-The first inequality follows from the weak convexity of $f$ and the second follows from the maximality of $\langle\nabla f(x),x-v\rangle$. ∎
-Note that Lemma 5.3 differs from Lemma 3.2 in the additional term $\frac{\rho}{2}\|x-x^{*}\|^{2}$. Moreover, we prove a lemma used in the proof of convergence analysis.
-Lemma 5.4.
+D.2 Local Sublinear Convergence
+We prove a lemma used in the proof of convergence analysis.
+Lemma D.2.
 For any $\nu\in(0,1]$ and $t\geq 0$, it holds that
-$h(t):=\frac{(t+3)^{\nu}(t+2-\nu)}{(t+2)^{1+\nu}}\leq 1.$ (5.2)
+$h(t):=\frac{(t+3)^{\nu}(t+2-\nu)}{(t+2)^{1+\nu}}\leq 1.$ (D.1)
 Proof.
@@
 ∎
-Now we show sublinear convergence with $\gamma_{t}=\frac{2}{2+t}$. The proof is a modified version of Theorem 4.2.
-Theorem 5.5 (Local sublinear convergence).
-Suppose that Assumptions 3.14.5, and 5.2 hold.
-Let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$ characterized by the Bregman distance $D_{\phi}$.
+Now we show sublinear convergence with $\gamma_{t}=\frac{2}{2+t}$. The proof is a modified version of Theorem C.1.
+Theorem D.3 (Local sublinear convergence).
+Suppose that Assumptions 3.14.3, and 5.1 hold.
+Let $\nu\in(0,1]$ and let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$ characterized by the Bregman distance $D_{\phi}$.
 Consider the iterates of Algorithm 1 with the open loop step-size, i.e., $\gamma_{t}=\frac{2}{2+t}$.
 Then, if $\rho/\sigma\leq L$ and $\frac{3\rho}{\mu}\leq 2-\nu$ hold, it holds that, for all $t\geq 1$,
-$f(x_{t})-f^{*}\leq\frac{2^{1+\nu}\mu LD^{2}}{\rho(t+2)^{\nu}},$ (5.3)
-and hence for any accuracy $\epsilon>0$ we have $f(x_{t})-f^{*}\leq\epsilon$ for all $t\geq\left(\frac{2^{1+\nu}\mu LD^{2}}{\rho\epsilon}\right)^{\frac{1}{\nu}}$.
+$f(x_{t})-f^{*}\leq\frac{2^{1+\nu}\mu LD^{2}}{\rho(t+2)^{\nu}},$ (D.2)
+and hence for any accuracy $\epsilon>0$ we have $f(x_{t})-f^{*}\leq\epsilon$ for all $t\geq(2^{1+\nu}\mu LD^{2}/\rho\epsilon)^{1/\nu}$.
 Proof.
@@
 where the last inequality holds because of (3.2) in Lemma 3.3. Subtracting $f^{*}$ and rearranging provides
-$f(x_{t+1})-f^{*}$ $\leq(1-\gamma_{t})(f(x_{t})-f^{*})+\frac{\rho\gamma_{t}}{2}\|x_{t}-x^{*}\|^{2}+L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t})$ (5.4)
-$\leq\left(1-\left(1-\frac{\rho}{\mu}\right)\gamma_{t}\right)(f(x_{t})-f^{*})+L\gamma_{t}^{1+\nu}D^{2},$ (5.5)
-where the last inequality holds due to the quadratic growth condition. For $t=0$, using (5.4), we have
-$f(x_{1})-f^{*}\leq\frac{\rho}{2}\|x_{0}-x^{*}\|^{2}+LD_{\phi}(v_{0},x_{0})\leq\left(\frac{\rho}{\sigma}+L\right)D^{2}\leq 2LD^{2}\leq\frac{2\mu}{\rho}LD^{2},$ (5.6)
+$f(x_{t+1})-f^{*}$ $\leq(1-\gamma_{t})(f(x_{t})-f^{*})+\frac{\rho\gamma_{t}}{2}\|x_{t}-x^{*}\|^{2}+L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t})$ (D.3)
+$\leq\left(1-\left(1-\frac{\rho}{\mu}\right)\gamma_{t}\right)(f(x_{t})-f^{*})+L\gamma_{t}^{1+\nu}D^{2},$ (D.4)
+where the last inequality holds due to the quadratic growth condition. For $t=0$, using (D.3), we have
+$f(x_{1})-f^{*}\leq\frac{\rho}{2}\|x_{0}-x^{*}\|^{2}+LD_{\phi}(v_{0},x_{0})\leq\left(\frac{\rho}{\sigma}+L\right)D^{2}\leq 2LD^{2}\leq\frac{2\mu}{\rho}LD^{2},$ (D.5)
 where the second inequality holds because $\phi$ is $\sigma$-strongly convex and the last inequality holds because of $3<6/(2-\nu)\leq 2\mu/\rho$.
 Now we consider $t\geq 1$.
-Using (5.5) and $\gamma_{t}=\frac{2}{2+t}$, we have
+Using (D.4) and $\gamma_{t}=\frac{2}{2+t}$, we have
 $f(x_{t+1})-f^{*}$ $\leq\left(t+\frac{2\rho}{\mu}\right)\frac{2^{1+\nu}\mu LD^{2}}{\rho(t+2)^{1+\nu}}+\frac{2^{1+\nu}LD^{2}}{(t+2)^{1+\nu}}$
@@
 $\leq\frac{2^{1+\nu}\mu LD^{2}}{\rho(t+3)^{\nu}},$
-where the last inequality holds because of $(t+3)^{\nu}(t+3\rho/\mu)\leq(t+3)^{\nu}(t+2-\nu)\leq(t+2)^{1+\nu}$ from (5.2). ∎
+where the last inequality holds because of $(t+3)^{\nu}(t+3\rho/\mu)\leq(t+3)^{\nu}(t+2-\nu)\leq(t+2)^{1+\nu}$ from (D.1). ∎
 We can apply the quadratic growth condition again as before and obtain
@@
 $\operatorname{dist}(x_{t},\mathcal{X}^{*})\leq\frac{2^{1+\nu/2}D\sqrt{L}}{\sqrt{\rho}(t+2)^{\nu/2}}.$
-Note that Theorem 5.5 does not require knowledge of the number of iterations $T$ ahead of time compared to Theorem 5.1.
+Note that Theorem D.3 does not require knowledge of the number of iterations $T$ ahead of time compared to Theorem D.1.
 We stress, nonetheless, that the latter can also be adjusted using a different step-size strategy to obtain an any-time guarantee; see [Braun2022-ew] for details for the standard Euclidean case, which can be generalized to our setup.
-Moreover, we can apply Theorem 5.5 to the classical FW algorithm, i.e., $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$.
-Using (5.3), $D_{\textup{Euc}}:=\sup_{x,y\in P}\|x-y\|^{2}$, and $D_{\textup{Euc}}=\sqrt{2}D$, we obtain
+Moreover, we can apply Theorem D.3 to the Euclidean FW algorithm, i.e., $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$.
+Using (D.2), $D_{\textup{Euc}}:=\sup_{x,y\in P}\|x-y\|^{2}$, and $D_{\textup{Euc}}=\sqrt{2}D$, we obtain
 $f(x_{t})-f^{*}\leq\frac{2\mu LD_{\textup{Euc}}^{2}}{\rho(t+2)},$
-Theorem 5.5 requires $\rho/\sigma\leq L$ and $\frac{3\rho}{\mu}\leq 2-\nu$. These assumptions are easy to satisfy.
-Example 5.6 (Example 3.1 in the arXiv version of [Liao2024-vd]).
+Theorem D.3 requires $\rho/\sigma\leq L$ and $\frac{3\rho}{\mu}\leq 2-\nu$. These assumptions are easy to satisfy.
+Example D.4 (Example 3.1 in the arXiv version of [Liao2024-vd]).
 Let us consider
@@
 A global optimal solution of $f$ is $x^{*}=-1$ and its value is $f(x^{*})=0$.
-Moreover, $f$ has the quadratic growth property with $0<\mu\leq 6$.
+Moreover, $f$ satisfies the quadratic growth condition with $0<\mu\leq 6$.
 Because $f$ is a quadratic function, we have $L\geq 6$.
 It holds that $2=\rho\leq L$ and $1=\frac{3\rho}{\mu}\leq 1<2-\nu$ (set $\mu=6$).
-Therefore, the assumption of Theorem 5.5 holds.
+Therefore, the assumption of Theorem D.3 holds.
 In order to verify $\frac{3\rho}{\mu}\leq 2-\nu$, it is easier to examine the sufficient condition $\frac{3\rho}{\mu}\leq 1$ instead because the exact value of $\nu$ is difficult to estimate.
@@
 Therefore, it suffices to verify $\rho/\sigma\leq\rho\leq L$, which often holds (for example, see an example of phase retrieval in Section 6.2).
-Next we establish local accelerated convergence with the short step step-size, i.e., $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},1\right\}$.
-Theorem 5.7 (Local accelerated convergence for inner optima).
-Suppose that Assumptions 3.1 and 5.2 hold.
-Let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter.
+Next we establish local linear convergence with the short step step-size, i.e., $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},1\right\}$.
+D.3 Proof of Theorem 5.2
+Theorem D.5 (Local linear convergence of FW algorithm with short step-size or adaptive step-size).
+Suppose that Assumptions 3.1 and 5.1 hold.
+Let $\nu\in(0,1]$ and let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter.
 Assume that there exists a minimizer $x^{*}\in\operatorname{int}P$, i.e., there exists an $r>0$ with $B(x^{*},r)\subset P$.
-Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},1\right\}$.
+Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},1\right\}$ or Algorithm 2.
 Then, if $\rho/\mu<1$, it holds that, for all $t\geq 1$,
@@
 $\geq\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}h_{t}^{\frac{1+\nu}{2\nu}},$
-where the third inequality holds due to Proposition 2.17 and the definition of $D$ and the last inequality holds due to the local PL inequality from Remark 2.15 with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$ and $M:=(L(1+\nu))^{1/\nu}$.
+where the third inequality holds due to Proposition B.8 and the definition of $D$ and the last inequality holds due to the local PL inequality from Remark B.6 with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$ and $M:=(L(1+\nu))^{1/\nu}$.
 (ii) In the case where $\gamma_{t}=1$, i.e., $\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}}\geq 1$, this implies
-$\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq L(1+\nu)D_{\phi}(v_{t},x_{t}),$ (5.7)
-because of $11/\nu>$ 1. Using Lemma 2.8 and (5.7), we have
+$\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq L(1+\nu)D_{\phi}(v_{t},x_{t}),$ (D.6)
+because of $11/\nu>$ 1. Using Lemma 2.4 and (D.6), we have
 $h_{t+1}-h_{t}$ $\leq LD_{\phi}(v_{t},x_{t})-\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle$
@@
 When $\nu=1$, we have $h_{t}-h_{t+1}\geq\min\left\{\frac{1}{2}\left(1-\frac{\rho}{\mu}\right),\frac{r^{2}}{2Mc^{2}D^{2}}\right\}\cdot h_{t}$.
-This inequality and the initial bound $f(x_{1})-f^{*}\leq LD^{2}$ due to Lemma 2.7 imply
+This inequality and the initial bound $f(x_{1})-f^{*}\leq LD^{2}$ due to Lemma A.4 imply
 $h_{t}\leq\max\left\{\frac{1}{2}\left(1+\frac{\rho}{\mu}\right),1-\frac{r^{2}}{2Mc^{2}D^{2}}\right\}^{t-1}LD^{2}.$
@@
 $h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right),\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}h_{t}^{\frac{1-\nu}{2\nu}}\right\}\cdot h_{t}.$
-Using $f(x_{1})-f^{*}\leq LD^{2}$ and Lemma 2.20 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}$, and $\theta_{0}=\frac{1-\nu}{2\nu}>0$, we have the claim.
+Using $f(x_{1})-f^{*}\leq LD^{2}$ and Lemma B.11 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}$, and $\theta_{0}=\frac{1-\nu}{2\nu}>0$, we have the claim.
 ∎
@@
 $f(x_{t})-f^{*}\leq\max\left\{\frac{1}{2}\left(1+\frac{\rho}{\mu}\right),1-\frac{r^{2}}{2Lc^{2}D_{\textup{Euc}}^{2}}\right\}^{t-1}\frac{LD_{\textup{Euc}}^{2}}{2}.$
-Because Algorithm 2 is well-defined (see also Remark 3.4), the convergence result of the FW algorithm with Algorithm 2 is the same as Theorem 5.7.
-Finally, we establish the local linear convergence of Algorithm 3.
-In the same way as Theorem 4.6, the pyramidal width is the minimal $\delta>0$ satisfying Lemma 2.18 (see also [Braun2022-ew, Lemma 2.26] or [Lacoste-Julien2015-gb, Theorem 3]).
+D.4 Proof of Theorem 5.3
+In the same way as Theorem 4.4, the pyramidal width is the minimal $\delta>0$ satisfying Lemma B.9 (see also [Braun2022-ew, Lemma 2.26] or [Lacoste-Julien2015-gb, Theorem 3]).
 An upper bound exists on the primal gap for weakly convex functions.
-Its proof can be found in Appendix A.1.
-Lemma 5.8 (Upper bound on primal gap for weakly convex functions).
-Suppose that Assumptions 3.1 and 5.2 hold.
+Lemma D.6 (Upper bound on primal gap for weakly convex functions).
+Suppose that Assumptions 3.1 and 5.1 hold.
 Let $P$ be a polytope with the pyramidal width $\delta>0$.
@@
 If $\rho/\mu<1$, then it holds that, for all $x\in[f\leq f^{*}+\zeta]\cap P$,
-$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}\delta^{2}}\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}.$ (5.8)
-We are ready to prove the local linear convergence of Algorithm 3 for nonconvex optimization.
-We also assume $\rho\leq L$, which is not restrictive (see also the discussion of Theorem 5.5 and Example 5.6).
-Theorem 5.9 (Local linear convergence by the away-step FW algorithm).
-Suppose that Assumptions 3.14.5, and 5.2 hold. Let $P\subset\mathbb{R}^{n}$ be a polytope. The convergence rate of Algorithm 3 with $f$ is linear: if $\rho<\mu\leq L$, for all $t\geq 1$
-$f(x_{t})-f^{*}\leq\begin{cases}2\left(1-\frac{\omega}{4L}\frac{\delta^{2}}{D^{2}}\right)^{\lceil(t-1)/2\rceil}LD^{2}&\text{if}\ \nu=1,\\ 2\left(\frac{1}{1+\nu}\left(1+\frac{\nu\rho}{\mu}\right)\right)^{\lceil(t-1)/2\rceil}LD^{2}&\text{if}\ 1\leq t\leq t_{0},\nu\in(0,1),\\ \frac{(L^{2}(1+\nu)^{2}D^{4}(1-\rho/\mu)^{2\nu}/(\omega\delta^{2})^{1+\nu})^{1/(1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}\left(1-\frac{\rho}{\mu}\right)\lceil(t-t_{0})/2\rceil\right)^{2\nu/(1-\nu)}}=\mathcal{O}(1/t^{2\nu/(1-\nu)})&\text{if}\ t\geq t_{0},\nu\in(0,1),\end{cases}$
+$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}\delta^{2}}\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}.$ (D.7)
+Proof.
+Using the weak convexity and the local quadratic growth condition of $f$, it holds that, for all $x\in[f\leq f^{*}+\zeta]\cap P$,
+$f(x)-f(x^{*})$ $\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}$
+$\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{\mu}(f(x)-f(x^{*})),$
+which implies
+$0\leq\left(1-\frac{\rho}{\mu}\right)(f(x)-f(x^{*}))\leq\langle\nabla f(x),x-x^{*}\rangle,$
+where the first inequality follows from $01-\rho/\mu>$ 0. Using (B.8) with $\psi=\nabla f(x)$ and $y=x^{*}$ and the above inequality, we obtain
+$\frac{\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}}{\delta^{2}}$ $\geq\frac{\langle\nabla f(x),x-x^{*}\rangle^{2}}{\|x-x^{*}\|^{2}}$
+$\geq\left(1-\frac{\rho}{\mu}\right)^{2}\frac{(f(x)-f(x^{*}))^{2}}{\|x-x^{*}\|^{2}}$
+$\geq\frac{\mu}{2}\left(1-\frac{\rho}{\mu}\right)^{2}\frac{(f(x)-f(x^{*}))^{2}}{f(x)-f(x^{*})}$
+$=\frac{\mu}{2}\left(1-\frac{\rho}{\mu}\right)^{2}(f(x)-f(x^{*})),$
+where the third inequality holds because of the local quadratic growth condition. ∎
+Theorem D.7 (Local linear convergence by the away-step FW algorithm).
+Suppose that Assumptions 3.14.3, and 5.1 hold.
+Let $\nu\in(0,1]$ and let $P\subset\mathbb{R}^{n}$ be a polytope.
+The convergence rate of Algorithm 3 with $f$ is linear: if $\rho<\mu\leq L$, for all $t\geq 1$
+$f(x_{t})-f^{*}\leq\begin{cases}2\left(1-\frac{\omega}{4L}\frac{\delta^{2}}{D^{2}}\right)^{\lceil(t-1)/2\rceil}LD^{2}&\text{if}\ \nu=1,\\ 2\left(\frac{1}{1+\nu}\left(1+\frac{\nu\rho}{\mu}\right)\right)^{\lceil(t-1)/2\rceil}LD^{2}&\text{if}\ 1\leq t\leq t_{0},\nu\in(0,1),\\ \frac{(L(1+\nu)(1-\rho/\mu)^{\nu}D^{2}/(\sqrt{\omega}\delta)^{1+\nu})^{2/(1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}\left(1-\frac{\rho}{\mu}\right)\lceil(t-t_{0})/2\rceil\right)^{2\nu/(1-\nu)}}=\mathcal{O}(1/t^{2\nu/(1-\nu)})&\text{if}\ t\geq t_{0},\nu\in(0,1),\end{cases}$
 where $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ and $\delta$ are the diameter and the pyramidal width of the polytope $P$, respectively, and $\omega:=\frac{(\mu-\rho)^{2}}{8\mu}$, and
@@
 Proof.
-By letting $h_{t}=f(x_{t})-f^{*}$ and using Lemma 5.8, we have
-$h_{t}\leq\frac{2\mu}{(\mu-\rho)^{2}\delta^{2}}\langle\nabla f(x_{t}),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}\leq\frac{8\mu\langle\nabla f(x_{t}),d_{t}\rangle^{2}}{(\mu-\rho)^{2}\delta^{2}}=\frac{\langle\nabla f(x_{t}),d_{t}\rangle^{2}}{\omega\delta^{2}},$ (5.9)
+By letting $h_{t}=f(x_{t})-f^{*}$ and using Lemma D.6, we have
+$h_{t}\leq\frac{2\mu}{(\mu-\rho)^{2}\delta^{2}}\langle\nabla f(x_{t}),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}\leq\frac{8\mu\langle\nabla f(x_{t}),d_{t}\rangle^{2}}{(\mu-\rho)^{2}\delta^{2}}=\frac{\langle\nabla f(x_{t}),d_{t}\rangle^{2}}{\omega\delta^{2}},$ (D.8)
 where the second inequality holds because $d_{t}$ is either $x_{t}-v_{t}^{\textup{FW}}$ or $v_{t}^{\textup{A}}-x_{t}$ with $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle/2$ in Lines 3 and 3 of Algorithm 3.
@@
 $=\min\left\{\frac{\nu\gamma_{t,\max}}{1+\nu}\left(1-\frac{\rho}{\mu}\right)h_{t},\frac{\nu}{1+\nu}\frac{(\omega\delta^{2})^{\frac{1+\nu}{2\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}h_{t}^{\frac{1+\nu}{2\nu}}\right\},$
-where the second inequality holds because of $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq h_{t}-\frac{\rho}{2}\|x_{t}-v_{t}\|^{2}\geq\left(1-\frac{\rho}{\mu}\right)h_{t}$ (from Lemma 5.3 and the quadratic growth condition of $f$) and (5.9).
+where the second inequality holds because of $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq h_{t}-\frac{\rho}{2}\|x_{t}-v_{t}\|^{2}\geq\left(1-\frac{\rho}{\mu}\right)h_{t}$ (from Lemma 3.3 and the quadratic growth condition of $f$) and (D.8).
 In $\nu=1$, for Frank–Wolfe steps, we have $\gamma_{t,\max}\left(1-\frac{\rho}{\mu}\right)=\left(1-\frac{\rho}{\mu}\right)\geq\frac{1}{8}\left(1-\frac{\rho}{\mu}\right)^{2}\frac{\delta^{2}}{D^{2}}=\frac{\omega\delta^{2}}{\mu D^{2}}\geq\frac{\omega\delta^{2}}{2LD^{2}}$ by $0<\rho<\mu\leq L$ and $\delta\leq D$.
@@
 Therefore, at most half of iterations until $t$ iterations are in away steps, i.e., in all other steps, we have $\gamma_{t}=\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{2LD_{\phi}(v_{t},x_{t})}<\gamma_{t,\max}$ and then $h_{t+1}\leq\left(1-\frac{\omega\delta^{2}}{4LD^{2}}\right)h_{t}$.
-Because we have $h_{t+1}\leq\left(1-\frac{\omega\delta^{2}}{4LD^{2}}\right)h_{t}$ for at least half of the iterations and $h_{t+1}\leq h_{t}$ for the rest, using $h_{1}\leq 2LD^{2}$ from (5.6) and $\rho<L$, we obtain
+Because we have $h_{t+1}\leq\left(1-\frac{\omega\delta^{2}}{4LD^{2}}\right)h_{t}$ for at least half of the iterations and $h_{t+1}\leq h_{t}$ for the rest, using $h_{1}\leq 2LD^{2}$ from (D.5) and $\rho<L$, we obtain
 $f(x_{t})-f^{*}\leq 2\left(1-\frac{\omega}{4L}\frac{\delta^{2}}{D^{2}}\right)^{\lceil(t-1)/2\rceil}LD^{2}.$
 In $\nu\in(0,1)$, we have $h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right),\frac{\nu}{1+\nu}\frac{(\omega\delta^{2})^{\frac{1+\nu}{2\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}h_{t}^{\frac{1-\nu}{2\nu}}\right\}\cdot h_{t}$ for at least half of the iterations and $h_{t+1}\leq h_{t}$ for the rest.
-The initial bound $f(x_{1})-f^{*}\leq(\rho+L)D^{2}\leq 2LD^{2}$ holds from (5.6) and $\rho<L$.
-We use Lemma 2.20 with $c_{0}=2LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=\frac{\nu}{1+\nu}\frac{(\omega\delta^{2})^{\frac{1+\nu}{2\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}$, and $\theta_{0}=\frac{1-\nu}{2\nu}$ and obtain the claim.
+The initial bound $f(x_{1})-f^{*}\leq(\rho+L)D^{2}\leq 2LD^{2}$ holds from (D.5) and $\rho<L$.
+We use Lemma B.11 with $c_{0}=2LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=\frac{\nu}{1+\nu}\frac{(\omega\delta^{2})^{\frac{1+\nu}{2\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}$, and $\theta_{0}=\frac{1-\nu}{2\nu}$ and obtain the claim.
 ∎
@@
 $f(x_{t})-f(x^{*})\leq\left(1-\frac{\omega}{2L}\frac{\delta^{2}}{D_{\textup{Euc}}^{2}}\right)^{\lceil(t-1)/2\rceil}LD_{\textup{Euc}}^{2}.$
-Without loss of generality, we set $\sigma=1$. If $\rho\leq L$ does not hold, we can use $h_{1}\leq(\rho+L)D^{2}$ as the initial bound.
-Moreover, when $P$ is $\alpha$-strongly convex set and $\phi=\frac{1}{2}\|\cdot\|^{2}$, we also establish local linear convergence. Its proof can be found in Appendix A.2.
-6 Numerical Experiments
-In this section, we conducted numerical experiments to examine the performance of our proposed algorithms.
-All numerical experiments were performed in Julia 1.11 using the FrankWolfe.jl package [Besancon2022-lx] [^2] https://github.com/ZIB-IOL/FrankWolfe.jl on a MacBook Pro with an Apple M2 Max and 64GB LPDDR5 memory.
-We compare our algorithm with the following algorithms and use the following notation:
-  • BregFW: the FW algorithm with the adaptive Bregman step-size strategy (Algorithm 2, our proposed)
-  • BregAFW: the away-step FW algorithm with the adaptive Bregman step-size strategy (Algorithm 3 using Algorithm 2, i.e., $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},\gamma_{t,\max})$, our proposed update)
-  • EucFW: the FW algorithm with the adaptive (Euclidean) step-size strategy [Pedregosa2020-ay]
-  • EucAFW: the away-step FW algorithm with the adaptive (Euclidean) step-size strategy [Pedregosa2020-ay]
-  • ShortFW: the FW algorithm with the (Euclidean) short step
-  • ShortAFW: the away-step FW algorithm with the (Euclidean) short step
-  • OpenFW: the FW algorithm with the open loop with $\gamma_{t}=\frac{2}{2+t}$
-  • OpenAFW: the away-step FW algorithm with the open loop with $\gamma_{t}=\frac{2}{2+t}$
-  • MD: the mirror descent [Nemirovskij1983-sm]
-Note that we included OpenAFW only for comparison purposes; there is no established convergence theory for AFW with open-loop strategies.
-In particular, there are no proper drop steps, and the favorable properties of the away-step Frank-Wolfe algorithm are lost; see [Braun2022-ew].
-We use $\beta=0.9$, $\eta=0.9$, $\tau=2$, and $\gamma_{\max}=1$ throughout all numerical experiments.
-6.1 Nonnegative Linear Inverse Problems
-Given a nonnegative matrix $A\in\mathbb{R}_{+}^{m\times n}$ and a nonnegative vector $b\in\mathbb{R}_{+}^{m}$, the goal of nonnegative (Poisson) linear inverse problems is to recover a signal $x\in\mathbb{R}_{+}^{n}$ such that $Ax\simeq b$.
-This class of problems has been studied in image deblurring [Bertero2009-jq] and positron emission tomography [Vardi1985-de] as well as optimization [Bauschke2017-hg, Takahashi2024-wz].
-Since the dimension of $x$ is often larger than the number of observations $m$, the system is indeterminate.
-From this point of view, we consider the constraint $\Delta_{n}:=\{x\in\mathbb{R}_{+}^{n}\,|\,\sum_{j=1}^{n}x_{j}\leq 1\}$.
-Recovering $x$ can be formulated as a minimization problem:
-$\min_{x\in\Delta_{n}}\quad f(x):=d(Ax,b),$ (6.1)
-where $d(x,y):=\sum_{i=1}^{m}\left(x_{i}\log\frac{x_{i}}{y_{i}}+y_{i}-x_{i}\right)$ is the Kullback–Leibler divergence (see also Example 2.3).
-Problem (6.1) is convex, while $\nabla f$ is not Lipschitz continuous on $\mathbb{R}_{+}^{n}$.
-The pair $(f,\phi)$ is $L$-smad on $\mathbb{R}_{+}^{n}$ with $\phi(x)=\sum_{j=1}^{n}x_{j}\log x_{j}$ and $L\geq\max_{1\leq j\leq n}\sum_{i=1}^{m}a_{ij}$ from [Bauschke2017-hg, Lemma 8].
-We compared BregFW with EucFW, ShortFW, OpenFW, and the mirror descent algorithm (MD[Nemirovskij1983-sm].
-The subproblem of MD can be solved in closed-form for $\{x\in\mathbb{R}_{+}^{n}\,|\,\sum_{j=1}^{n}x_{j}=1\}$ by [Beck2017-qc, Example 3.71], and this can be readily extended to $\Delta_{n}$.
-We used $1000$ as maximum iteration limit and we generated $\tilde{A}$ from an i.i.d.
-normal distribution and set $a_{ij}=|\tilde{a}_{ij}|/\sum_{i=1}^{m}|\tilde{a}_{ij}|$.
-We also generated $\tilde{x}$ from an i.i.d.
-uniform distribution in $[0,1]$ and set the ground truth $x^{*}=0.8\tilde{x}/\sum_{j=1}^{n}\tilde{x}_{j}$ so that $x^{*}\in\operatorname{int}\Delta_{n}$.
-All components of the initial point $x_{0}$ were $1/n$.
-For $(m,n)=(100,1000)$, Figure 1 shows the primal gap $f(x_{t})-f^{*}$ and the FW gap $\max_{v\in P}\langle\nabla f(x_{t}),x_{t}-v\rangle$ per iteration (left) and the primal gap per second (right).
-Table 2 shows average numbers of the primal gap, the FW gap, and computational time over 20 different instances for $(m,n)=(100,1000)$.
-Here, BregFW outperformed other algorithms, both in iterations and time.
-Figure 1: Nonnegative linear inverse problem for $(m,n)=(100,1000)$.
-Table 2: Average numbers of the primal gap, the FW gap, and computational time (s) over 20 different instances of nonnegative linear inverse problems for $(m,n)=(100,1000)$.
-algorithm primal gap FW gap time (s)
-BregFW 6.963691e-08 1.145520e-05 1.457719e-01
-EucFW 3.028696e-07 2.922331e-05 1.077826e-01
-ShortFW 4.747044e-05 4.613479e-04 6.251837e-02
-OpenFW 4.957628e-07 8.538245e-05 6.503594e-02
-MD 2.249368e-06 1.721359e-04 1.888499e-01
-6.2 $\ell_{p}$oss Problem
-We consider the $\ell_{p}$ loss problem [Kyng2015-uo, Maddison2021-nt] to find $x\in P$ such that $Ax\simeq b$, defined by
-$\min_{x\in P}\quad\|Ax-b\|_{p}^{p},$ (6.2)
-where $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, and $p>1$. Defining $f(x)=\|Ax-b\|_{p}^{p}$, we have
-$\nabla f(x)$ $=p\sum_{i=1}^{m}|\langle a_{i},x\rangle-b_{i}|^{p-1}\operatorname{sgn}(\langle a_{i},x\rangle-b_{i}),$
-$\nabla^{2}f(x)$ $=p(p-1)\sum_{i=1}^{m}|\langle a_{i},x\rangle-b_{i}|^{p-2}a_{i}a_{i}^{\mathsf{T}},$
-where $a_{i}$ is the $i$th row vector of $A$.
-Problem (6.2) is convex, but $\nabla f$ is not Lipschitz continuous on $\mathbb{R}^{n}$ when $p\neq 2$.
-Furthermore, when $p<2$, $f$ is not $\mathcal{C}^{2}$ but $\mathcal{C}^{1}$.
-Therefore, for $21<p<$ 2, $\nabla f$ is also not Lipschitz continuous over compact sets.
-Since $f$ is convex, the pair $(f,\phi)$ is 1-smad with $\phi=f$.
-We use an $\ell_{2}$ norm ball as a constraint, i.e., $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|_{2}\leq 1\}$.
-We compared BregFW with EucFW, ShortFW, and OpenFW.
-We generated $A$ from an i.i.d.
-normal distribution and normalized it so that $\|a_{i}\|=1$.
-We also generated $\tilde{x}$ from an i.i.d.
-normal distribution and set $x^{*}=0.8\tilde{x}/\|\tilde{x}\|$ to ensure $x^{*}\in\operatorname{int}P$.
-The initial point $x_{0}$ was generated by computing an extreme point of $P$ that minimizes the linear approximation of $f$.
-For $(n,m)=(100,100)$ and $p=1.1$, Figure 2 shows the primal gap $f(x_{t})-f^{*}$ and the FW gap $\max_{v\in P}\langle\nabla f(x_{t}),x_{t}-v\rangle$ per iteration and those gaps per second up to 1000 iterations.
-Table 3 also shows the average performance over 20 different instances.
-BregFW outperformed the other algorithms.
-Since $\nabla f$ is not Lipschitz continuous, ShortFW did not reduce the primal and FW gap.
-Figure 2: The $\ell_{p}$ loss problem for $(m,n)=(1000,100)$.
-Table 3: Average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of $\ell_{p}$ loss problems for $(m,n)=(1000,100)$.
-algorithm primal gap FW gap time (s)
-BregFW 1.056988e-13 3.764084e-01 5.600876e-01
-EucFW 6.341301e-10 8.576574e-01 3.138045e-01
-ShortFW 2.465054e+01 4.754413e+01 9.198800e-02
-OpenFW 1.698968e-02 4.007941e+00 8.713342e-02
-6.3 Phase Retrieval
-We are interested in phase retrieval, which involves finding a signal $x\in\mathbb{R}^{n}$ such that
-$|\langle a_{i},x\rangle|^{2}\simeq b_{i},\quad i=1,\ldots,m,$
-where $a_{i}\in\mathbb{R}^{n}$ describes the model and $b\in\mathbb{R}^{m}$ is a vector of measurements.
-Phase retrieval has a long history.
-Patterson studied phase retrieval in X-ray crystallography in 1934 [Patterson1934-sg] and 1944 [Patterson1944-nm].
-Phase retrieval arises in many fields of science and engineering, such as image processing [Candes2015-mk], X-ray crystallography [Patterson1934-sg, Patterson1944-nm], and optics [Shechtman2015-ic].
-Bolte et al.
-[Bolte2018-zt] applied the Bregman proximal gradient algorithm to phase retrieval.
-In applications, $a_{i}$ and $x$ are often complex vectors. Now, we consider real vectors. To achieve the goal of phase retrieval, we focus on the following nonconvex optimization problem:
-$\min_{x\in P}\quad f(x):=\frac{1}{4}\sum_{i=1}^{m}\left(|\langle a_{i},x\rangle|^{2}-b_{i}\right)^{2},$
-where $P\subset\mathbb{R}^{n}$ is a compact convex set. We introduce some properties of $f$. Let $\phi$ be defined by
-$\phi(x)=\frac{1}{4}\|x\|^{4}+\frac{1}{2}\|x\|^{2},$
-which is $\sigma$-strongly convex for $\sigma\leq 1$. For any $L$ satisfying
-$L\geq\sum_{i=1}^{m}\left(3\|a_{i}\|^{4}+\|a_{i}\|^{2}|b_{i}|\right),$ (6.3)
-the pair $(f,\phi)$ is $L$-smad on $\mathbb{R}^{n}$ [Bolte2018-zt, Lemma 5.1].
-The right-hand side of (6.3) is relatively large in applications, while Takahashi et al.
-derived a smaller lower bound of $L$ [Takahashi2023-uh, Propositions 5 and 6].
-In addition, because we have
-$\nabla^{2}f(x)=\sum_{i=1}^{m}\left((3|\langle a_{i},x\rangle|^{2}-b_{i})a_{i}a_{i}^{\mathsf{T}}\right),$
-the matrix $\nabla^{2}f(x)+\rho I$ is positive semidefinite for $\rho\geq\sum_{i=1}^{m}\|a_{i}\|^{2}|b_{i}|$, i.e., $f$ is $\rho$-weakly convex.
-For example, $L=\sum_{i=1}^{m}\left(3\|a_{i}\|^{4}+\|a_{i}\|^{2}|b_{i}|\right)\geq\sum_{i=1}^{m}\|a_{i}\|^{2}|b_{i}|=\rho/\sigma$ holds with $\sigma=1$.
-We use a $K$-sparse polytope as a constraint, i.e., $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|_{1}\leq K,\|x\|_{\infty}\leq 1\}$.
-We compared BregFW with EucFW, ShortFW, and OpenFW.
-The maximum iteration is 1000.
-We generated $a_{i}$, $i=1,\ldots,m$ from an i.i.d.
-normal distribution and normalized it to have norm 1.
-We also generated $x^{*}$ from an i.i.d.
-uniform distribution in $[0,1]$ and normalized $x^{*}$ to have sum 1.
-The initial point $x_{0}$ was generated by computing an extreme point of $P$ that minimizes the linear approximation of $f$.
-For $(m,n)=(100,2000)$ and $K=200$, Figure 3 shows the primal and FW gaps per iteration and gaps per second.
-Because $\nabla f$ is not Lipschitz continuous, ShortFW does not converge, and EucFW performs slowly.
-The primal gap and the FW gap by the Bregman adaptive step-size strategy are the smallest among these step-size strategies.
-Table 4 shows the average performance over 20 different instances of $(m,n)=(100,2000)$.
-We compared BregAFW with EucAFW, ShortAFW, and OpenAFW.
-In only this setting, we generated $x^{*}$ from an i.i.d.
-uniform distribution in $[0,1]$ and did not normalize it; that is, $x^{*}$ might be in the face of $P$.
-Figure 4 shows another setting’s results for $(m,n)=(200,200)$ and $K=110$.
-Table 5 shows the average performance over 20 different instances for $(m,n)=(200,200)$ and $K=110$.
-The primal gap by BregAFW is the smallest among these algorithms, while ShortAFW has the smallest value of the FW gap.
-Figure 3: Phase retrieval for $(m,n)=(100,2000)$ and $K=200$.
-Table 4: Average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of phase retrieval for $(m,n)=(100,2000)$ and $K=200$.
-algorithm primal gap FW gap time
-BregFW 3.307714e-09 4.372800e-06 9.532406e-01
-EucFW 1.419221e-04 1.271594e+01 8.291155e-01
-ShortFW 5.488362e+04 5.389255e+03 1.822046e-01
-OpenFW 4.710460e-08 2.747374e+00 1.945798e-01
-Figure 4: Phase retrieval for $(m,n)=(200,200)$ and $K=110$ via away-step FW algorithms.
-Table 5: Average numbers of the primal gap, FW gap, and computational time (sec) over 20 different instances of phase retrieval for $(m,n)=(200,200)$ and $K=110$ via away-step FW algorithms.
-algorithm primal gap FW gap time (s)
-BregAFW 4.343027e+00 9.757773e-01 2.333602e-01
-EucAFW 4.396803e+00 1.221226e+00 3.942057e-01
-ShortAFW 4.409245e+00 2.597756e-02 7.353968e-02
-OpenAFW 4.409352e+00 7.199201e-02 5.899826e-02
-6.4 Low-Rank Minimization
-Given a symmetric matrix $M\in\mathbb{R}^{n\times n}$, our goal is to find $X\in\mathbb{R}^{n\times r}$ such that $M\simeq XX^{\mathsf{T}}$. This is accomplished by minimizing the function
-$\min_{X\in P}\quad f(X):=\frac{1}{2}\|XX^{\mathsf{T}}-M\|_{F}^{2},$
-where $P\subset\mathbb{R}^{n\times r}$.
-We assume that $r\leq n$.
-This problem is known as low-rank minimization [Dragomir2021-rv].
-In this paper, we define $P=\{X\in\mathbb{R}^{n\times r}\,|\,\|X\|_{*}\leq\xi\}$, where $\|\cdot\|_{*}$ denotes the nuclear norm and $r\in\mathbb{R}_{+}$ for the low-rank assumption.
-We define
-$\phi(X)=\frac{1}{4}\|X\|_{F}^{4}+\frac{1}{2}\|X\|_{F}^{2}.$
-There exists a constant $L$ such that the pair $(f,\phi)$ is $L$-smad on $\mathbb{R}^{n}$ [Dragomir2021-rv].
-Additionally, $f(X)$ is weakly convex on any compact set due to Proposition 2.10, which follows from the twice continuous differentiability of $f$.
-We also compared BregFW with EucFW, ShortFW, and OpenFW.
-The parameter settings are the same as those in the previous subsection.
-We generated $X^{*}$ from an i.i.d.
-uniform distribution in $[0,1]$, normalized each column of $X^{*}$, and set $M=X^{*}(X^{*})^{\mathsf{T}}$.
-The initial point $X_{0}$ was generated from an i.i.d.
-uniform distribution in $[0,1]$.
-We set $\xi=10\lambda_{\max}(M)$ for $P$.
-Figure 5 shows the primal and FW gaps per iteration and gaps per second for $(n,r)=(1000,20)$ up to the 1000th iteration.
-Table 6 presents the average performance over 20 different instances for $(n,r)=(1000,20)$.
-BregFW performs slightly better than EucFW.
-OpenFW also performed as fast as BregFW and EucFW, but its performance was unstable.
-ShortFW did not converge due to the lack of Lipschitz continuity of $\nabla f$.
-Figure 5: Low-rank minimization for $(n,r)=(1000,20)$.
-Table 6: Average values of the primal gap, FW gap, and computational time (s) over 20 different instances of low-rank minimization for $(n,r)=(1000,20)$.
-algorithm primal gap FW gap time (s)
-BregFW 5.651947e-01 1.472666e+01 1.260218e+01
-EucFW 5.776476e-01 1.723729e+01 1.228013e+01
-ShortFW 2.710108e+08 2.169351e+09 4.137911e+00
-OpenFW 1.978179e+00 2.121183e+03 4.176979e+00
-6.5 Nonnegative Matrix Factorization
-Given a nonnegative matrix $V\in\mathbb{R}_{+}^{m\times n}$, nonnegative matrix factorization (NMF) aims to find nonnegative matrices $W\in\mathbb{R}_{+}^{m\times r}$ and $H\in\mathbb{R}_{+}^{r\times n}$ such that $V\simeq WH$.
-NMF can be formulated as a minimization problem of the loss function that measures the difference between $V$ and $WH$, i.e.,
-$\min_{(W,H)\in P}\quad f(W,H):=\frac{1}{2}\|WH-V\|_{F}^{2},$ (6.4)
-where $P$ is a compact convex subset of $\mathbb{R}_{+}^{m\times r}\times\mathbb{R}_{+}^{r\times n}$.
-The objective function $f$ is weakly convex over $P$ due to Proposition 2.10.
-The gradient $\nabla f$ is not Lipschitz continuous, while $(f,\phi)$ is smooth adaptable [Mukkamala2019-mk] with $\phi(W,H)=\frac{1}{4}(\|W\|_{F}^{2}+\|H\|_{F}^{2})^{2}+\frac{1}{2}(\|W\|_{F}^{2}+\|H\|_{F}^{2})$.
-[…]
-Figure 6: NMF for $(m,n,r)=(100,5000,20)$.
-Table 7: Average values of the primal gap, FW gap, and computational time (s) over 20 different instances of NMF for $(m,n,r)=(100,5000,20)$.
-algorithm primal gap FW gap time (s)
-BregFW 1.201735e-04 1.151457e-02 5.098349e+01
-EucFW 1.281873e-04 7.625941e-04 4.359247e+01
-[…]
-Figure 7: NMF for MovieLens 100K Dataset.
-Acknowledgement
-This project has been partially supported by the Japan Society for the Promotion of Science (JSPS) through Invitational Fellowships for Research in Japan program (L24503), JSPS KAKENHI Grant Number JP23K19953 and JP25K21156; JSPS KAKENHI Grant Number JP23H03351 and JST ERATO Grant Number JPMJER1903.
-Research reported in this paper was also partially supported the Deutsche Forschungsgemeinschaft (DFG) through the DFG Cluster of Excellence MATH+ (grant number EXC-2046/1, project ID 390685689).
-Appendix A Appendix
-A.1 Results from Weak Convexity
-Lemma A.1 (Primal gap bound from the quadratic growth).
-Let $f$ be a $\rho$-weakly convex function that satisfies the local $\mu$-quadratic growth condition such that $\rho\leq\mu$.
-Let $x^{*}$ be the unique minimizer of $f$ over $P$ and let $\zeta>0$.
-Then, the following holds: for all $x\in[f\leq f^{*}+\zeta]\cap P$,
-$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{2},$
-or equivalently,
-$\left(\frac{\mu}{2}\right)^{1/2}\left(1-\frac{\rho}{\mu}\right)(f(x)-f^{*})^{1/2}\leq\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}.$
-Proof.
-By first applying weak convexity and then the local quadratic growth condition for $x^{*}\in\mathcal{X}^{*}$ with $f(x^{*})=f^{*}$ and for all $x\in[f\leq f^{*}+\zeta]\cap P$, it holds:
-$f(x)-f^{*}$ $=f(x)-f(x^{*})$
-$\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}$
-$\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{\mu}\left(f(x)-f^{*}\right),$
-which implies
-$\left(1-\frac{\rho}{\mu}\right)\left(f(x)-f^{*}\right)\leq\langle\nabla f(x),x-x^{*}\rangle=\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\|x-x^{*}\|.$ (A.1)
-Using the local quadratic growth condition again, we have
-$\left(1-\frac{\rho}{\mu}\right)\left(f(x)-f^{*}\right)\leq\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\left(\frac{2}{\mu}(f(x)-f^{*})\right)^{1/2},$
-which provides
-$\left(\frac{\mu}{2}\right)^{1/2}\left(1-\frac{\rho}{\mu}\right)(f(x)-f^{*})^{1/2}\leq\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|},$
-or equivalently
-$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{2}.$
-Lemma A.2 (Upper bound on primal gap for weakly convex functions).
-Suppose that Assumptions 3.1 and 5.2 hold.
-Let $P$ be a polytope with the pyramidal width $\delta>0$.
-Let $\mathcal{S}$ denote any set of vertices of $P$ with $x\in\operatorname{conv}\mathcal{S}$.
-Let $\psi$ be any vector, so that we define $v^{\textup{FW}}=\operatorname*{argmin}_{v\in P}\langle\psi,v\rangle$ and $v^{\textup{A}}=\operatorname*{argmax}_{v\in\mathcal{S}}\langle\psi,v\rangle$.
-If $\rho/\mu<1$, then it holds that, for all $x\in[f\leq f^{*}+\zeta]\cap P$,
-$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}\delta^{2}}\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}.$
-Proof.
-Using the weak convexity and the local quadratic growth condition of $f$, it holds that, for all $x\in[f\leq f^{*}+\zeta]\cap P$,
-$f(x)-f(x^{*})$ $\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}$
-$\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{\mu}(f(x)-f(x^{*})),$
-which implies
-$0\leq\left(1-\frac{\rho}{\mu}\right)(f(x)-f(x^{*}))\leq\langle\nabla f(x),x-x^{*}\rangle,$
-where the first inequality follows from $01-\rho/\mu>$ 0. Using (2.10) with $\psi=\nabla f(x)$ and $y=x^{*}$ and the above inequality, we obtain
-$\frac{\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}}{\delta^{2}}$ $\geq\frac{\langle\nabla f(x),x-x^{*}\rangle^{2}}{\|x-x^{*}\|^{2}}$
-$\geq\left(1-\frac{\rho}{\mu}\right)^{2}\frac{(f(x)-f(x^{*}))^{2}}{\|x-x^{*}\|^{2}}$
-$\geq\frac{\mu}{2}\left(1-\frac{\rho}{\mu}\right)^{2}\frac{(f(x)-f(x^{*}))^{2}}{f(x)-f(x^{*})}$
-$=\frac{\mu}{2}\left(1-\frac{\rho}{\mu}\right)^{2}(f(x)-f(x^{*})),$
-where the third inequality holds because of the local quadratic growth condition. ∎
-A.2 Local Linear Convergence over Uniformly Convex Sets
+D.5 Local Linear Convergence over Uniformly Convex Sets
 In the convex optimization case, Canon and Cullum [Canon1968-sr] established an early lower bound on the convergence rate of the FW algorithm.
@@
 To this end, we recall the definition of uniformly convex sets.
-Definition A.3 ($(\alpha,p)$-uniformly convex set [Braun2022-ew, Definition 2.18][Kerdreux2021-kc, Definition 1.1]).
+Definition D.8 ($(\alpha,p)$-uniformly convex set [Braun2022-ew, Definition 2.18][Kerdreux2021-kc, Definition 1.1]).
 […]
@@
 We will use the scaling condition for uniformly convex sets to establish linear convergence.
-Proposition A.4 (Scaling inequality [Braun2022-ew, Proposition 2.19][Kerdreux2021-kc, Lemma 2.1]).
+Proposition D.9 (Scaling inequality [Braun2022-ew, Proposition 2.19][Kerdreux2021-kc, Lemma 2.1]).
 Let $P$ be a full dimensional compact $(\alpha,p)$-uniformly convex set, $u$ any non-zero vector, and $v=\operatorname*{argmin}_{y\in P}\langle u,y\rangle$. Then for all $x\in P$
@@
 Now we establish local linear convergence for weakly convex optimization on uniformly convex sets.
-Theorem A.5 (Local linear convergence over uniformly convex sets).
-Suppose that Assumptions 3.1 and 5.2 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$ hold and that $P$ is $(\alpha,p)$-uniformly convex set.
+Theorem D.10 (Local linear convergence over uniformly convex sets).
+Suppose that Assumptions 3.1 and 5.1 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$ hold and that $P$ is $(\alpha,p)$-uniformly convex set.
 Let $\nabla f$ be bounded away from 0, i.e., $\|\nabla f(x)\|\geq c>0$ for all $x\in P$.
@@
 Lemma 3.4 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$ is followed by $f(x_{t})-f(x_{t+1})\geq\frac{g_{t}}{2}\gamma_{t}$.
-Using Proposition A.4, we have
+Using Proposition D.9, we have
 $h_{t}-h_{t+1}$ $\geq\frac{g_{t}}{2}\min\left\{\frac{g_{t}}{L\|x_{t}-v_{t}\|^{2}},1\right\}$
@@
 where the third inequality holds from Lemma 3.3 and $\|\nabla f(x)\|\geq c$, and the last inequality holds because of the local quadratic growth property of $f$.
-The initial bound $h_{1}\leq LD_{\textup{Euc}}^{2}=2LD^{2}$ holds from $\rho\leq L$ and (5.6).
+The initial bound $h_{1}\leq LD_{\textup{Euc}}^{2}=2LD^{2}$ holds from $\rho\leq L$ and (D.5).
 For $q=2$, we have $h_{t}-h_{t+1}\geq\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)\min\left\{\frac{\alpha c}{L},1\right\}\cdot h_{t}$, which implies
@@
 Thus, we have the claim.
-For $p>2$, we use Lemma 2.20 with $c_{0}=LD_{\textup{Euc}}^{2}$, $c_{1}=\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=c_{1}\cdot\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c)^{2/p}}{L}$, and $\theta_{0}=1-2/p$ and obtain the claim.
+For $p>2$, we use Lemma B.11 with $c_{0}=LD_{\textup{Euc}}^{2}$, $c_{1}=\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=c_{1}\cdot\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c)^{2/p}}{L}$, and $\theta_{0}=1-2/p$ and obtain the claim.
 ∎
 Note that Assumption 3.1 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$ holds when $f$ is $L$-smooth over $P$.
-If $\rho\leq L$ does not hold, we can use the initial bound $h_{1}\leq\frac{\rho+L}{2}D_{\textup{Euc}}^{2}$ from (5.6).
-In that case, a local linear rate in Theorem A.5 is unchanged.
+If $\rho\leq L$ does not hold, we can use the initial bound $h_{1}\leq\frac{\rho+L}{2}D_{\textup{Euc}}^{2}$ from (D.5).
+In that case, a local linear rate in Theorem D.10 is unchanged.
 Moreover, we have local linear convergence without assuming $\|\nabla f(x)\|>0$.
-Theorem A.6 (Local linear convergence over uniformly convex sets without $\|\nabla f(x)\|>0$).
-Suppose that Assumptions 3.1 and 5.2 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$ hold and that $P$ is $(\alpha,p)$-uniformly convex set.
+Theorem D.11 (Local linear convergence over uniformly convex sets without $\|\nabla f(x)\|>0$).
+Suppose that Assumptions 3.1 and 5.1 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$ hold and that $P$ is $(\alpha,p)$-uniformly convex set.
 Let $D_{\textup{Euc}}:=\sup_{x,y\in P}\|x-y\|$ be the diameter of $P$.
@@
 Proof.
-Using an argument similar to Theorem A.5, we obtain
+Using an argument similar to Theorem D.10, we obtain
 $h_{t}-h_{t+1}$ $\geq\frac{g_{t}}{2}\min\left\{\frac{g_{t}}{L\|x_{t}-v_{t}\|^{2}},1\right\}$
@@
 $\geq\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)\min\left\{\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c^{-1})^{2/p}}{L}h_{t}^{1-1/p},1\right\}\cdot h_{t},$
-where the third inequality holds from the PL inequality (2.8) with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$.
-Therefore, we use Lemma 2.20 with $c_{0}=LD_{\textup{Euc}}^{2}$, $c_{1}=\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=c_{1}\cdot\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c^{-1})^{2/p}}{L}$, and $\theta_{0}=1-1/p$ and obtain the claim.
+where the third inequality holds from the PL inequality (B.6) with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$.
+Therefore, we use Lemma B.11 with $c_{0}=LD_{\textup{Euc}}^{2}$, $c_{1}=\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=c_{1}\cdot\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c^{-1})^{2/p}}{L}$, and $\theta_{0}=1-1/p$ and obtain the claim.
 ∎
+Appendix E Experiments
+We use the following notation:
+  • BregFW: the FW algorithm with the adaptive Bregman step-size strategy (Algorithm 2, our proposed)
+  • BregAFW: the away-step FW algorithm with the adaptive Bregman step-size strategy (Algorithm 3 using Algorithm 2, i.e., $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},\gamma_{t,\max})$, our proposed update)
+  • EucFW: the FW algorithm with the adaptive (Euclidean) step-size strategy [Pedregosa2020-ay]
+  • EucAFW: the away-step FW algorithm with the adaptive (Euclidean) step-size strategy [Pedregosa2020-ay]
+  • ShortFW: the FW algorithm with the (Euclidean) short step
+  • ShortAFW: the away-step FW algorithm with the (Euclidean) short step
+  • OpenFW: the FW algorithm with the open loop with $\gamma_{t}=\frac{2}{2+t}$
+  • OpenAFW: the away-step FW algorithm with the open loop with $\gamma_{t}=\frac{2}{2+t}$
+  • MD: the mirror descent [Nemirovskij1983-sm]
+E.1 Nonnegative Linear Inverse Problems
+Given a nonnegative matrix $A\in\mathbb{R}_{+}^{m\times n}$ and a nonnegative vector $b\in\mathbb{R}_{+}^{m}$, the goal of nonnegative (Poisson) linear inverse problems is to recover a signal $x\in\mathbb{R}_{+}^{n}$ such that $Ax\simeq b$.
+This class of problems has been studied in image deblurring [Bertero2009-jq] and positron emission tomography [Vardi1985-de] as well as optimization [Bauschke2017-hg, Takahashi2024-wz].
+Since the dimension of $x$ is often larger than the number of observations $m$, the system is indeterminate.
+From this point of view, we consider the constraint $\Delta_{n}:=\{x\in\mathbb{R}_{+}^{n}\,|\,\sum_{j=1}^{n}x_{j}\leq 1\}$.
+Recovering $x$ can be formulated as a minimization problem:
+$\min_{x\in\Delta_{n}}\quad f(x):=d(Ax,b),$ (E.1)
+where $d(x,y):=\sum_{i=1}^{m}\left(x_{i}\log\frac{x_{i}}{y_{i}}+y_{i}-x_{i}\right)$ is the Kullback–Leibler divergence.
+Problem (E.1) is convex, while $\nabla f$ is not Lipschitz continuous on $\mathbb{R}_{+}^{n}$.
+The pair $(f,\phi)$ is $L$-smad on $\mathbb{R}_{+}^{n}$ with $\phi(x)=\sum_{j=1}^{n}x_{j}\log x_{j}$ and $L\geq\max_{1\leq j\leq n}\sum_{i=1}^{m}a_{ij}$ from [Bauschke2017-hg, Lemma 8].
+We compared BregFW with EucFW, ShortFW, OpenFW, and the mirror descent algorithm (MD[Nemirovskij1983-sm].
+The subproblem of MD can be solved in closed-form for $\{x\in\mathbb{R}_{+}^{n}\,|\,\sum_{j=1}^{n}x_{j}=1\}$ by [Beck2017-qc, Example 3.71], and this can be readily extended to $\Delta_{n}$.
+We used $1000$ as maximum iteration limit and we generated $\tilde{A}$ from an i.i.d.
+normal distribution and set $a_{ij}=|\tilde{a}_{ij}|/\sum_{i=1}^{m}|\tilde{a}_{ij}|$.
+We also generated $\tilde{x}$ from an i.i.d.
+uniform distribution in $[0,1]$ and set the ground truth $x^{*}=0.8\tilde{x}/\sum_{j=1}^{n}\tilde{x}_{j}$ so that $x^{*}\in\operatorname{int}\Delta_{n}$ (random seed 1234).
+All components of the initial point $x_{0}$ were $1/n$.
+For $(m,n)=(100,1000)$, Figure 5 shows the primal gap $f(x_{t})-f^{*}$ and the FW gap $\max_{v\in P}\langle\nabla f(x_{t}),x_{t}-v\rangle$ per iteration (left) and the primal gap per second (right).
+Table 2 shows average numbers of the primal gap, the FW gap, and computational time over 20 different instances for $(m,n)=(100,1000)$.
+Here, BregFW outperformed other algorithms, both in iterations and time.
+Figure 5: Nonnegative linear inverse problem for $(m,n)=(100,1000)$.
+Table 2: Average numbers of the primal gap, the FW gap, and computational time (s) over 20 different instances of nonnegative linear inverse problems for $(m,n)=(100,1000)$.
+The results are reported as mean $\pm$ standard deviation.
+algorithm primal gap FW gap time (s)
+BregFW 6.9637e-08 $\pm$ 2.0e-08 1.1455e-05 $\pm$ 1.4e-06 1.4577e-01 $\pm$ 4.1e-02
+EucFW 3.0287e-07 $\pm$ 9.4e-08 2.9223e-05 $\pm$ 4.1e-06 1.0778e-01 $\pm$ 5.8e-03
+ShortFW 4.7470e-05 $\pm$ 3.5e-06 4.6135e-04 $\pm$ 7.2e-05 6.2518e-02 $\pm$ 3.6e-03
+OpenFW 4.9576e-07 $\pm$ 6.3e-08 8.5382e-05 $\pm$ 1.3e-05 6.5036e-02 $\pm$ 6.3e-03
+MD 2.2494e-06 $\pm$ 3.1e-07 1.7214e-04 $\pm$ 1.9e-05 1.8885e-01 $\pm$ 6.4e-03
+E.2 $\ell_{p}$ Loss Problem
+Figure 6: The $\ell_{p}$ loss problem for $(m,n)=(1000,100)$.
+Table 3: Average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of $\ell_{p}$ loss problems for $(m,n)=(1000,100)$.
+The results are reported as mean $\pm$ standard deviation.
+algorithm primal gap FW gap time (s)
+BregFW 1.0589e-13 $\pm$ 8.6e-15 3.7686e-01 $\pm$ 8.6e-03 5.6037e-01 $\pm$ 1.8e-02
+EucFW 6.2980e-10 $\pm$ 9.9e-11 8.5940e-01 $\pm$ 1.0e-01 3.1421e-01 $\pm$ 7.4e-03
+ShortFW 2.4697e+01 $\pm$ 1.2e+00 4.7621e+01 $\pm$ 8.6e-01 9.2352e-02 $\pm$ 3.7e-03
+OpenFW 1.6961e-02 $\pm$ 9.9e-04 4.$0$$\pm$ 1.0e-01 8.7292e-02 $\pm$ 3.5e-03
+We use an $\ell_{2}$ norm ball as a constraint, i.e., $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|_{2}\leq 1\}$.
+We compared BregFW with EucFW, ShortFW, and OpenFW.
+We generated $A$ from an i.i.d.
+normal distribution and normalized it so that $\|a_{i}\|=1$.
+We also generated $\tilde{x}$ from an i.i.d.
+normal distribution and set $x^{*}=0.8\tilde{x}/\|\tilde{x}\|$ to ensure $x^{*}\in\operatorname{int}P$ (random seed 1234).
+The initial point $x_{0}$ was generated by computing an extreme point of $P$ that minimizes the linear approximation of $f$.
+For $(n,m)=(100,100)$ and $p=1.1$, Figure 6 shows the primal gap $f(x_{t})-f^{*}$ and the FW gap $\max_{v\in P}\langle\nabla f(x_{t}),x_{t}-v\rangle$ per iteration and those gaps per second up to 1000 iterations.
+Table 3 also shows the average performance over 20 different instances.
+BregFW outperformed the other algorithms.
+Since $\nabla f$ is not Lipschitz continuous, ShortFW did not reduce the primal and FW gap.
+E.3 Phase Retrieval
+We show the average performance for the setting of Section 6.2. Tables 4 and 5 show the average performance over 20 different instances of $(m,n)=(1000,10000)$ and $(m,n)=(2000,10000)$, respectively.
+We also compared BregAFW with EucAFW, ShortAFW, and OpenAFW.
+In only this setting, we generated $x^{*}$ from an i.i.d.
+uniform distribution in $[0,1]$ and did not normalize it (random seed 1234); that is, $x^{*}$ might be in the face of $P$.
+Figure 7 shows another setting’s results for $(m,n)=(200,200)$ and $K=110$.
+Table 6 shows the average performance over 20 different instances for $(m,n)=(200,200)$ and $K=110$.
+The primal gap by BregAFW is the smallest among these algorithms, while ShortAFW has the smallest value of the FW gap.
+Figure 7: Phase retrieval for $(m,n)=(200,200)$ and $K=110$ via AFW.
+Table 4: Average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of phase retrieval for $(m,n)=(1000,10000)$ and $K=2000$.
+The results are reported as mean $\pm$ standard deviation.
+algorithm primal gap FW gap time
+BregFW $-1$$\pm$ 1.1e-12 2.1518e-02 $\pm$ 5.9e-02 4.6990e+01 $\pm$ 1.5e+01
+EucFW 1.7170e+00 $\pm$ 3.9e+00 2.4612e+02 $\pm$ 4.5e+02 4.4986e+01 $\pm$ 3.2e+01
+ShortFW 3.2739e+05 $\pm$ 1.9e+04 4.2232e+04 $\pm$ 5.0e+02 1.0942e+01 $\pm$ 3.2e-01
+OpenFW 5.2496e-10 $\pm$ 3.9e-11 4.3802e+00 $\pm$ 5.5e-02 6.0353e+01 $\pm$ 1.6e+00
+Table 5: Average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of phase retrieval for $(m,n)=(2000,10000)$ and $K=2000$.
+The results are reported as mean $\pm$ standard deviation.
+algorithm primal gap FW gap time
+BregFW 6.6743e-12 $\pm$ 5.5e-13 9.5656e-08 $\pm$ 7.9e-09 7.2592e+01 $\pm$ 8.8e+00
+EucFW 3.2499e+00 $\pm$ 7.1e+00 3.1304e+02 $\pm$ 4.8e+02 1.4348e+02 $\pm$ 4.3e+01
+ShortFW 6.0937e+04 $\pm$ 2.3e+03 2.5461e+04 $\pm$ 2.3e+02 1.2648e+01 $\pm$ 2.4e-01
+OpenFW 9.3513e-11 $\pm$ 4.9e-12 2.6628e+00 $\pm$ 1.7e-02 7.1916e+01 $\pm$ 1.3e+00
+Table 6: Average numbers of the primal gap, FW gap, and computational time (sec) over 20 different instances of phase retrieval for $(m,n)=(200,200)$ and $K=110$ via away-step FW algorithms.
+The results are reported as mean $\pm$ standard deviation.
+algorithm primal gap FW gap time (s)
+BregAFW 4.3430e+00 $\pm$ 1.7e+00 9.7578e-01 $\pm$ 9.1e-01 2.3336e-01 $\pm$ 6.4e-02
+EucAFW 4.3968e+00 $\pm$ 1.7e+00 1.2212e+00 $\pm$ 2.3e+00 3.9421e-01 $\pm$ 1.6e-01
+ShortAFW 4.4092e+00 $\pm$ 1.7e+00 2.5978e-02 $\pm$ 7.3e-03 7.3540e-02 $\pm$ 3.5e-02
+OpenAFW 4.4094e+00 $\pm$ 1.7e+00 7.1992e-02 $\pm$ 4.9e-02 5.8998e-02 $\pm$ 2.4e-03
+E.4 Low-Rank Minimization
+Given a symmetric matrix $M\in\mathbb{R}^{n\times n}$, our goal is to find $X\in\mathbb{R}^{n\times r}$ such that $M\simeq XX^{\mathsf{T}}$. This is accomplished by minimizing the function
+$\min_{X\in P}\quad f(X):=\frac{1}{2}\|XX^{\mathsf{T}}-M\|_{F}^{2},$
+where $P\subset\mathbb{R}^{n\times r}$.
+We assume that $r\leq n$.
+This problem is known as low-rank minimization [Dragomir2021-rv].
+In this paper, we define $P=\{X\in\mathbb{R}^{n\times r}\,|\,\|X\|_{*}\leq\xi\}$, where $\|\cdot\|_{*}$ denotes the nuclear norm and $r\in\mathbb{R}_{+}$ for the low-rank assumption.
+We define
+$\phi(X)=\frac{1}{4}\|X\|_{F}^{4}+\frac{1}{2}\|X\|_{F}^{2}.$
+There exists a constant $L$ such that the pair $(f,\phi)$ is $L$-smad on $\mathbb{R}^{n}$ [Dragomir2021-rv].
+Additionally, $f(X)$ is weakly convex on any compact set due to Proposition B.1, which follows from the twice continuous differentiability of $f$.
+We also compared BregFW with EucFW, ShortFW, and OpenFW.
+The parameter settings are the same as those in the previous subsection.
+We generated $X^{*}$ from an i.i.d.
+uniform distribution in $[0,1]$, normalized each column of $X^{*}$, and set $M=X^{*}(X^{*})^{\mathsf{T}}$ (random seed 42).
+The initial point $X_{0}$ was generated from an i.i.d.
+uniform distribution in $[0,1]$.
+We set $\xi=10\lambda_{\max}(M)$ for $P$.
+Figure 8 shows the primal and FW gaps per iteration and gaps per second for $(n,r)=(1000,20)$ up to the 1000th iteration.
+Table 7 presents the average performance over 20 different instances for $(n,r)=(1000,20)$.
+BregFW performs slightly better than EucFW.
+OpenFW also performed as fast as BregFW and EucFW, but its performance was unstable.
+ShortFW did not converge due to the lack of Lipschitz continuity of $\nabla f$.
+Figure 8: Low-rank minimization for $(n,r)=(1000,20)$.
+Table 7: average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of low-rank minimization for $(n,r)=(1000,20)$.
+The results are reported as mean $\pm$ standard deviation.
+algorithm primal gap FW gap time (s)
+BregFW 5.6519e-01 $\pm$ 4.2e-02 1.4727e+01 $\pm$ 1.0e+00 1.2602e+01 $\pm$ 4.5e-01
+EucFW 5.7765e-01 $\pm$ 4.6e-02 1.7237e+01 $\pm$ 2.0e+00 1.2280e+01 $\pm$ 3.8e-01
+ShortFW 2.7101e+08 $\pm$ 1.9e+06 2.1694e+09 $\pm$ 1.5e+07 4.1379e+00 $\pm$ 7.9e-02
+OpenFW 1.9782e+00 $\pm$ 3.6e-01 2.1212e+03 $\pm$ 3.1e+02 4.1770e+00 $\pm$ 1.7e-01
+E.5 Nonnegative Matrix Factorization
+Given a nonnegative matrix $V\in\mathbb{R}_{+}^{m\times n}$, nonnegative matrix factorization (NMF) aims to find nonnegative matrices $W\in\mathbb{R}_{+}^{m\times r}$ and $H\in\mathbb{R}_{+}^{r\times n}$ such that $V\simeq WH$.
+NMF can be formulated as a minimization problem of the loss function that measures the difference between $V$ and $WH$, i.e., $\min_{(W,H)\in P}f(W,H):=\frac{1}{2}\|WH-V\|_{F}^{2}$, where $P$ is a compact convex subset of $\mathbb{R}_{+}^{m\times r}\times\mathbb{R}_{+}^{r\times n}$.
+The objective function $f$ is weakly convex over $P$ due to Proposition B.1.
+The gradient $\nabla f$ is not Lipschitz continuous, while $(f,\phi)$ is smooth adaptable [Mukkamala2019-mk] with $\phi(W,H)=\frac{1}{4}(\|W\|_{F}^{2}+\|H\|_{F}^{2})^{2}+\frac{1}{2}(\|W\|_{F}^{2}+\|H\|_{F}^{2})$.
+[…]
+[…]
+Figure 9: Log-log plot of the primal and the FW gaps on NMF for $(m,n,r)=(100,5000,20)$.
+Figure 10: Log-log plot of the primal and the FW gaps on NMF for MovieLens 100K Dataset.
+Table 8: average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of NMF for $(m,n,r)=(100,5000,20)$.
+The results are reported as mean $\pm$ standard deviation.
+algorithm primal gap FW gap time (s)
+BregFW 1.2017e-04 $\pm$ 5.7e-06 1.1515e-02 $\pm$ 1.4e-02 5.0983e+01 $\pm$ 1.1e+00
+EucFW 1.2819e-04 $\pm$ 4.6e-06 7.6259e-04 $\pm$ 8.0e-04 4.3592e+01 $\pm$ 1.0e+00

### v4

@@
 Abstract
-We propose a Frank–Wolfe (FW) algorithm with an adaptive Bregman step-size strategy for smooth adaptable (also called: relatively smooth) (weakly-) convex functions.
+We propose Frank–Wolfe (FW) algorithms with an adaptive Bregman step-size strategy for smooth adaptable (also called: relatively smooth) (weakly-) convex functions.
 This means that the gradient of the objective function is not necessarily Lipschitz continuous, and we only require the smooth adaptable property.
-Compared to existing FW algorithms, our assumptions are less restrictive.
-We establish convergence guarantees in various settings, such as sublinear to linear convergence rates, depending on the assumptions for convex and nonconvex objective functions.
-Assuming that the objective function is weakly convex and satisfies the local quadratic growth condition, we provide both local sublinear and local linear convergence regarding the primal gap.
+Compared with existing FW algorithms, our assumptions are less restrictive.
+We establish convergence guarantees in various settings, including convergence rates ranging from sublinear to linear, depending on the assumptions for convex and nonconvex objective functions.
+Assuming that the objective function is weakly convex and satisfies the local quadratic growth condition, we provide both local sublinear and local linear convergence with respect to the primal gap.
 We also propose a variant of the away-step FW algorithm using Bregman distances over polytopes.
-We establish global faster (up to linear) convergence for convex optimization under the Hölder error bound condition and its local linear convergence for nonconvex optimization under the local quadratic growth condition.
+We establish faster global convergence (up to a linear rate) for convex optimization under the Hölder error bound condition and local linear convergence for nonconvex optimization under the local quadratic growth condition.
 Numerical experiments demonstrate that our proposed FW algorithms outperform existing methods.
@@
 Instead of requiring access to a projection oracle, the FW algorithm requires only access to a so-called linear minimization oracle (LMO), which for a given linear function $a\in\mathbb{R}^{n}$ computes $y\in\operatorname*{arg\,min}_{v\in P}\langle a,v\rangle$.
-LMOs are often much cheaper than projection oracles, as shown in [CP2021] (see also [Braun2022-ew, Table 1.1]), so that, in practice, FW algorithms are often faster than projected gradient methods even when the projection operation is not trivial.
-Additionally, FW algorithms tend to be numerically quite robust and stable due to their affine-invariance and can be also used, e.g., to provide theoretical guarantees for the approximate Carathéodory problem [cp2019approxCara].
+LMOs are often much cheaper than projection oracles, as shown in Combettes and Pokutta (2021) (see also (Braun et al., 2025, Table 1.1)).
+Consequently, in practice, FW algorithms are often faster than projected gradient methods even when the projection operation is nontrivial.
+Additionally, FW algorithms tend to be numerically quite robust and stable due to their affine-invariance and can also be used, e.g., to provide theoretical guarantees for the approximate Carathéodory problem by Combettes and Pokutta (2023).
 Overview of Existing FW Studies:
-The FW algorithm was originally proposed by Frank and Wolfe [Frank1956-cq] and was independently rediscovered and extended by Levitin and Polyak [Levitin1966-sm] as the conditional gradient method; we will use these terms interchangeably.
-Canon and Cullum [Canon1968-sr] established an initial lower bound for the convergence rate of the FW algorithm, which was later improved by GuéLat and Marcotte [GueLat1986-bg], which also provided the first analysis of the away-step FW algorithm by Wolfe [Wolfe1970-ax].
-Jaggi [Jaggi2013-oz] provided a more detailed convergence analysis of the FW algorithm, establishing a new lower bound that demonstrated a trade-off between sparsity and error.
-Concurrently, Lan [lan2013complexity] examined the complexity of linear programming-based first-order methods, establishing a similar lower bound.
+The FW algorithm was originally proposed by Frank and Wolfe (1956) and was independently rediscovered and extended by Levitin and Polyak (1966) as the conditional gradient method; we will use these terms interchangeably.
+Canon and Cullum (1968) established an initial lower bound for the rate of the FW algorithm.
+GuéLat and Marcotte (1986) improved this bound and also provided the first analysis of the away-step FW algorithm by Wolfe (1970).
+Jaggi (2013) provided a more detailed convergence analysis of the FW algorithm, establishing a new lower bound that demonstrated a trade-off between sparsity and error.
+Concurrently, Lan (2013) examined the complexity of linear programming-based first-order methods, establishing a similar lower bound.
 For a comprehensive review of FW algorithms, we refer the interested reader to the survey by Braun et al.
-[Braun2022-ew] and the brief introduction by Pokutta [Pokutta2024-qu].
+(2025) and the brief introduction by Pokutta (2024).
 The classical FW algorithm (Algorithm 1) has been improved in two main directions to enhance both performance and convergence.
@@
 Pedregosa et al.
-[Pedregosa2020-ay] introduced an adaptive step-size strategy that dynamically estimates $L$.
+(2020) introduced an adaptive step-size strategy that dynamically estimates $L$.
 They proved that this strategy is at least as effective as the short-step strategy.
-It was enhanced by Pokutta [Pokutta2024-qu] to improve numerical stability.
+It was enhanced by Pokutta (2024) to improve numerical stability.
 Another direction is to modify the classical FW algorithm to eliminate the zigzag behavior when approaching the optimal face containing the optimal solution $x^{*}$.
-This behavior inspired Wolfe [Wolfe1970-ax] to propose the away-step FW algorithm in 1970 that shortcuts the zigzagging by removing atoms that slow down the iterative sequence.
-Lacoste-Julien and Jaggi [Lacoste-Julien2015-gb] showed the linear convergence of the away-step FW algorithm.
-Their analysis***Similar analyses based on the pyramidal width have been developed for other advanced variants, such as the pairwise FW algorithm [Lacoste-Julien2015-gb], Wolfe’s algorithm [Lacoste-Julien2015-gb], the blended conditional gradient method [BPTW2018], and blended pairwise conditional gradient methods and similar variants [CP2020boost, TTP2021].
-Even Nesterov-style acceleration is possible under weak assumptions [CDP2019] building upon this analysis.
+This behavior inspired Wolfe (1970) to propose the away-step FW algorithm in 1970 that shortcuts the zigzagging by removing atoms that slow down the iterative sequence.
+Lacoste-Julien and Jaggi (2015) showed the linear convergence of the away-step FW algorithm.
+Their analysis [^1] Similar analyses based on the pyramidal width have been developed for other advanced variants, such as the pairwise FW algorithm (Lacoste-Julien and Jaggi, 2015), Wolfe’s algorithm (Lacoste-Julien and Jaggi, 2015), the blended conditional gradient method (Braun et al., 2019), and blended pairwise conditional gradient methods and similar variants (Combettes and Pokutta, 2020; Tsuji et al., 2022).
+Even Nesterov-style acceleration is possible under weak assumptions (Diakonikolas et al., 2020) building upon this analysis.
 introduced a geometric constant, the pyramidal width, that measures the conditioning of the polytope $P$, representing the feasible region.
 Our Research Idea:
-To derive the convergence rate of the above-mentioned FW algorithms, previous studies have typically required $L$-smoothness, i.e., that $\nabla f$ is Lipschitz continuous, and the convexity for the objective function $f$, although exceptions exist.Quite recently, Vyguzov and Stonyakin [Vyguzov2025-ha] proposed a FW algorithm with the Bregman distance, similar to ours, yet it is restricted to convex problems and provides only partial convergence guarantees.
+To derive the convergence rate of the above-mentioned FW algorithms, previous studies have typically required $L$-smoothness, i.e., that $\nabla f$ is Lipschitz continuous, and convexity for the objective function $f$, although exceptions exist.
+[^2] Quite recently, Vyguzov and Stonyakin (2025) proposed a FW algorithm with the Bregman distance, similar to ours, yet it is restricted to convex problems and provides only partial convergence guarantees.
 These assumptions often narrow the application of FW algorithms.
@@
 Since the $L$-smad property is consistent with $L$-smoothness when $\phi=\frac{1}{2}\|\cdot\|^{2}$, the class of $L$-smad functions includes the $L$-smooth function class.
-Other first-order algorithms, such as the Bregman proximal gradient algorithm, are often analyzed under the $L$-smad property (see e.g., [Bolte2018-zt, Hanzely2021-sz, Rebegoldi2018-pa, Takahashi2022-ml, Takahashi2024-wz, Yang2025-ls]).
-$L$-smad functions appear in many applications, such as nonnegative linear inverse problems [Bauschke1997-vk, Takahashi2024-wz], $\ell_{p}$ loss problems [Kyng2015-uo, Maddison2021-nt], phase retrieval [Bolte2018-zt, Takahashi2022-ml], nonnegative matrix factorization (NMF) [Mukkamala2019-mk, Takahashi2024-cu], and blind deconvolution [Takahashi2023-uh].
+Other first-order algorithms, such as the Bregman proximal gradient algorithm, are often analyzed under the $L$-smad property (see Appendix A and Bolte et al.
+(2018); Hanzely et al.
+(2021); Rebegoldi et al.
+(2018); Takahashi et al.
+(2022); Takahashi and Takeda (2025); Yang and Toh (2025)).
+$L$-smad functions appear in many applications, such as nonnegative linear inverse problems (Bauschke and Borwein, 1997; Takahashi and Takeda, 2025), $\ell_{p}$ loss problems (Kyng et al., 2015; Maddison et al., 2021), phase retrieval (Bolte et al., 2018; Takahashi et al., 2022), nonnegative matrix factorization (NMF) (Mukkamala and Ochs, 2019; Takahashi et al., 2026), and blind deconvolution (Takahashi et al., 2023).
 Any $\mathcal{C}^{2}$ function is locally $L$-smooth over compact sets, but the resulting Lipschitz constant $L$ can be overly conservative, hindering practical performance.
@@
 Moreover, functions in the above-mentioned applications are weakly convex on compact sets.
-For nonconvex functions, although Lacoste-Julien [Lacoste-Julien2016-mj] established sublinear convergence, little else has been done, and simply tracing existing research does not achieve convergence rates better than sublinear convergence.
+For nonconvex functions, although Lacoste-Julien (2016) established sublinear convergence, little else has been done, and simply tracing existing research does not achieve convergence rates better than sublinear convergence.
 Then the following question naturally arises:
-Is it possible to loosen the $L$-smoothness and convexity assumptions often made in FW algorithm studies while still achieving the theoretical guarantee of linear convergence rates?
+Is it possible to relax the $L$-smoothness and convexity assumptions often made in FW algorithm studies while still achieving the theoretical guarantee of linear convergence rates?
 Contribution:
@@
 We develop FW algorithms for $L$-smad $f$ and give theoretical guarantees of linear convergence under weaker assumptions (Hölder error bound (HEB) condition with parameter $q\geq 1$ for convex $f$ and the special case with $q=2$ called quadratic growth for nonconvex $f$) than the strong convexity assumption that is often made when linear convergence is shown for FW algorithms.
-The convexity of $f$ is also weaken to weakly convex, i.e., $f+\frac{\rho}{2}\|\cdot\|^{2}$ is convex with some $\rho>0$ (see, e.g.[Davis2018-bb, Examples 3.1 and 3.2] and [Nurminskii1975-vp]).
-Moreover, it is easy to see that any $\mathcal{C}^{2}$ function defined on a compact set is weakly convex [Vial1983-oo, Proposition 4.11] and [Wu2007-dt]; since $\rho$ is only needed for the theoretical guarantee and not for the algorithm, there is no need for its estimation, and considering weakly-convex $f$ greatly expands the range of applications.
+The convexity assumption on $f$ is also weakened to weakly convex, i.e., $f+\frac{\rho}{2}\|\cdot\|^{2}$ is convex with some $\rho>0$ (see, e.g.(Davis et al., 2018, Examples 3.1 and 3.2) and Nurminskii (1975)).
+Any $\mathcal{C}^{2}$ function on a compact set is weakly convex (Vial, 1983; Wu, 2007) (see Proposition C.1); since $\rho$ is only for the theoretical guarantee and not for the algorithm, there is no need for its estimation, and considering weakly convex $f$ greatly broadens applicability.
 Table 1 summarizes our contributions for the various cases that we consider.
-Relaxing the assumption from $L$-smoothness to the $L$-smad property makes the construction of a FW algorithm and its theoretical guarantee much more difficult.
-This is because we use the Bregman distance, which is an extension of the Euclidean distance, for the $L$-smad property.
-While (2.1) is simple with $\nu=1$ when using the Euclidean distance $\phi=\frac{1}{2}\|\cdot\|^{2}$, $\nu>0$ cannot be eliminated for the general Bregman distance; $1+\nu$ represents its scaling exponent.
-$\nu$ must be estimated to use in the algorithm since its exact value is unknown.
-Therefore, the extension to an $L$-smad $f$ is not straightforward.
+Relaxing the assumption from $L$-smoothness to the $L$-smad property extends the FW algorithm to a broader range of problems, while making the construction of the FW algorithm and its theoretical guarantee much more difficult.
+This is because we use the Bregman distance $D_{\phi}$, which is an extension of the Euclidean distance, for the $L$-smad property.
+While Equation 2.1 shown later is simple with $\nu=1$ when using the Euclidean distance, $\nu>0$ cannot be eliminated for the general Bregman distance; $1+\nu$ represents its scaling exponent.
+$\nu$ needs to be estimated during the algorithm since its exact value is unknown, and thus, the extension to an $L$-smad $f$ is not straightforward.
+Table 1: Our convergence rates are generalizations of existing rates.
+Moreover, our convergence rate is better than $\mathcal{O}(\epsilon^{-1})$ in some cases.
+For nonconvex functions, convergence is measured using the Frank–Wolfe gap $\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\leq\epsilon$, instead of the primal gap $f(x_{t})-f^{*}\leq\epsilon$.
+Convergence rates for weakly convex optimization hold locally.
+Assumptions333ff conv. means the convexity of $f$, IP means $x^{*}\in\operatorname{int}P$, and poly. means that $P$ is a polytope. Convergence rate
+FW $f$ conv. $f$ growth IP poly. $L$-smooth $L$-smad ($\nu>0$)
+Alg.1 (any step-size) conv. $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(\epsilon^{-1/\nu})$ (Thm.D.2)
+[…]
+away-step FW conv.
+$q$-HEB $\mathcal{O}(\log\epsilon^{-1})$, $\mathcal{O}(\epsilon^{\frac{2-q}{q}})$44footnotemark: 4 $\mathcal{O}(\log\epsilon^{-1})$, $\mathcal{O}(\epsilon^{\frac{1+\nu-q}{\nu q}})$ (Thm.
+4.4)55footnotemark: 5
+Alg.1 (any step-size) weak 2-HEB $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(\epsilon^{-1/\nu})$ (Thm. E.3)
+[…]
+away-step FW weak 2-HEB $\mathcal{O}(\log\epsilon^{-1})$ $\mathcal{O}(\log\epsilon^{-1})$, $\mathcal{O}(\epsilon^{\frac{\nu-1}{2\nu}})$ (Thm. 5.4)66footnotemark: 6
+Alg.1 (any step-size) $\mathcal{O}(\epsilon^{-2})$ $\mathcal{O}(\epsilon^{-1-1/\nu})$ (Thm. E.1)
   • We establish sublinear convergence for the case of convex $L$-smad functions.
-Assuming that $f$ is convex, satisfies the HEB condition, and the minimizer $x^{*}\in\operatorname{int}P$, we provide the faster (up to linear) convergence.
-In this setting, our algorithm always converges linearly if the exponent of the HEB condition $q$ equals the scaling exponent of the Bregman distance $1+\nu$, i.e., $q=1+\nu$.As discussed in Sec.
-2, linear convergence with the Euclidean distance for convex optimization is shown in [garber2015faster] under the quadratic growth over strongly convex sets.
+Assuming that $f$ is convex and satisfies the HEB condition, and that the minimizer $x^{*}\in\operatorname{int}P$, we provide faster convergence (up to a linear rate).
+In this setting, our algorithm always converges linearly if the exponent of the HEB condition $q$ equals the scaling exponent of the Bregman distance $1+\nu$, i.e., $q=1+\nu$.
+[^7] As discussed in Sec.
+2, linear convergence with the Euclidean distance is shown in Garber and Hazan (2015) under the quadratic growth over strongly convex sets.
 This corresponds to the $q=1+\nu$ case with $\nu=1$.
@@
   • For nonconvex optimization, assuming that $f$ is weakly convex and satisfies the local quadratic growth condition, we establish local linear convergence.
-To our knowledge, this is the first proposal of a FW algorithm with linear convergence guarantees for a certain type of nonconvex optimization problem, albeit of a limited type.
+To the best of our knowledge, this is the first FW algorithm with linear convergence guarantees for a certain class of nonconvex optimization problems, albeit for a rather restricted subclass.
 These results are also new for the standard setting: $L$-smooth $f$.
-Without weak convexity, we prove a global sublinear convergence to a stationary point of (1.1) for nonconvex $L$-smad $f$.
+Without weak convexity, we prove global sublinear convergence to a stationary point of Equation 1.1 for nonconvex $L$-smad $f$.
   • We also propose a variant of the away-step FW algorithm with the Bregman distance and establish its linear convergence under the HEB condition for convex optimization and under the local quadratic growth condition for weakly convex optimization.
-Table 1: Convergence rates for Frank–Wolfe algorithms.
-Our convergence rates are generalizations of existing rates.
-Moreover, our convergence rate is better than $\mathcal{O}(\epsilon^{-1})$ in some cases.
-For nonconvex functions, convergence is measured using the Frank–Wolfe gap $\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\leq\epsilon$, instead of the primal gap $f(x_{t})-f^{*}\leq\epsilon$.
-Convergence rates for the weakly convex optimization locally hold.
-Assumptions Convergence rate
-FW $f$ convexity $f$ growth $x^{*}\in\operatorname{int}P$ $P$ polytope $L$-smooth $L$-smad ($\nu>0$)
-Alg.1 (any step-size) convex $\mathcal{O}(\epsilon^{-1})$ [Jaggi2013-oz] $\mathcal{O}(\epsilon^{-1/\nu})$ ([Vyguzov2025-ha], Thm.C.1)
-Alg.1 (short & adapt.) convex $q$-HEB $\mathcal{O}(\log\epsilon^{-1})$11footnotemark: 1[KDP2018jour] $\mathcal{O}(\log\epsilon^{-1})$ [^2] It holds if the exponent of the HEB condition $q$ equals the scaling exponent of the Bregman distance $1+\nu$, i.e., $q=1+\nu$ ($(\nu,q)=(1,2)$ for the away-step FW algorithm).
-For $q>1+\nu$, it also holds if $t\leq t_{0}$ with some $t_{0}\in\mathbb{N}$; otherwise $\mathcal{O}(\epsilon^{(1+\nu-q)/\nu q})$.(Thm.
-4.2)
-away-step FW convex $q$-HEB $\mathcal{O}(\log\epsilon^{-1})$11footnotemark: 1[KDP2018jour] $\mathcal{O}(\log\epsilon^{-1})$22footnotemark: 2(Thm. 4.4)
-Alg.1 (any step-size) weak 2-HEB $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(\epsilon^{-1/\nu})$ (Thm. D.3)
-Alg.1 (short & adapt.) weak 2-HEB $\mathcal{O}(\log\epsilon^{-1})$ $\mathcal{O}(\log\epsilon^{-1})$ [^4] It holds if $\nu=1$.
-For $\nu\in(0,1)$, it also holds if $t\leq t_{0}$ with some $t_{0}\in\mathbb{N}$; otherwise $\mathcal{O}(\epsilon^{(\nu-1)/2\nu})$.(Thm.
-5.2)
-away-step FW weak 2-HEB $\mathcal{O}(\log\epsilon^{-1})$ $\mathcal{O}(\log\epsilon^{-1})$44footnotemark: 4(Thm. 5.3)
-Alg.1 (any step-size) $\mathcal{O}(\epsilon^{-2})$ [Lacoste-Julien2016-mj] $\mathcal{O}(\epsilon^{-1-1/\nu})$ (Thm. D.1)
 2 Preliminaries
-[…]
 Let $C$ be a nonempty open convex subset of $\mathbb{R}^{n}$.
-Definition 2.1 (Kernel Generating Distance [Bolte2018-zt, Definition 2.1]).
-A function $\phi:\mathbb{R}^{n}\to(-\infty,+\infty]$ is called a kernel generating distance associated with $C$ if it satisfies the following conditions:
-  1. (i)
-    $\phi$ is proper, lower semicontinuous, and convex, with $\operatorname{dom}\phi\subset\operatorname{cl}C$ and $\operatorname{dom}\partial\phi=C$;
-  2. (ii)
-    $\phi$ is $\mathcal{C}^{1}$ on $\operatorname{int}\operatorname{dom}\phi\equiv C$.
+Let $\operatorname{dist}(x,C):=\inf_{y\in C}\|x-y\|$ and let $\mathcal{C}^{k}$ be the class of $k$-times continuously differentiable functions for $k\geq 0$.
+Definition 2.1 (Kernel Generating Distance (Bolte et al., 2018)).
+A function $\phi:\mathbb{R}^{n}\to(-\infty,+\infty]$ is called a kernel generating distance associated with $C$ if it satisfies the following conditions: (i) $\phi$ is proper, lower semicontinuous, and convex, with $\operatorname{dom}\phi\subset\operatorname{cl}C$ and $\operatorname{dom}\partial\phi=C$; (ii) $\phi$ is $\mathcal{C}^{1}$ on $\operatorname{int}\operatorname{dom}\phi\equiv C$.
 We denote the class of kernel generating distances associated with $C$ by $\mathcal{G}(C)$.
-Definition 2.2 (Bregman Distance [Bregman1967-rf]).
-Given $\phi\in\mathcal{G}(C)$, a Bregman distance $D_{\phi}:\operatorname{dom}\phi\times\operatorname{int}\operatorname{dom}\phi\to\mathbb{R}_{+}$ associated with $\phi$ is defined by $D_{\phi}(x,y)=\phi(x)-\phi(y)-\langle\nabla\phi(y),x-y\rangle\geq 0$.
+Definition 2.2 (Bregman Distance (Bregman, 1967)).
+Given $\phi\in\mathcal{G}(C)$, a Bregman distance $D_{\phi}:\operatorname{dom}\phi\times\operatorname{int}\operatorname{dom}\phi\to\mathbb{R}_{+}$ associated with $\phi$ is defined by $D_{\phi}(x,y):=\phi(x)-\phi(y)-\langle\nabla\phi(y),x-y\rangle$.
+The Bregman distance $D_{\phi}(x,y)$ measures the proximity between $x\in\operatorname{dom}\phi$ and $y\in\operatorname{int}\operatorname{dom}\phi$.
+Since $\phi$ is convex, $D_{\phi}(x,y)\geq 0$ holds.
+Moreover, when $\phi$ is strictly convex, $D_{\phi}(x,y)=0$ holds if and only if $x=y$.
+However, the Bregman distance is not always symmetric and does not have to satisfy the triangle inequality.
+See Example B.1.
 We use $D_{f}$ in Definition 2.2 for nonconvex $f$ instead of convex $\phi$ for brevity.
-In the remainder of the paper, if not stated otherwise, we will assume that as such for a given strictly convex $\phi$ and $C$, there exists $\nu>0$ such that the following holds for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and $\gamma\in[0,1]$ (see also Lemma A.3, which shows that a weaker one holds)):
+In the remainder of the paper, we assume that, for a given strictly convex $\phi$ and $C$, there exists $\nu>0$ such that the following holds for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and $\gamma\in[0,1]$ (see Lemma B.3):
 $D_{\phi}((1-\gamma)x+\gamma y,x)\leq\gamma^{1+\nu}D_{\phi}(y,x).$ (2.1)
-We will now recall the smooth adaptable property, which is a generalization of $L$-smoothness and was first introduced by [Bauschke2017-hg].
-The smooth adaptable property is also called relative smoothness [Lu2018-ii].
-Definition 2.3 ($L$-smooth Adaptable Property [Bolte2018-zt, Definition 2.2]).
+We recall the smooth adaptable property, which is a generalization of $L$-smoothness and was introduced by Bauschke et al.
+(2017); Bolte et al.
+(2018); Lu et al.
+(2018).
+Its characterizations can be found in Appendix B.2, Example B.5, Remark B.6, and (Bauschke et al., 2017, Proposition 1).
+Definition 2.3 ($L$-smooth Adaptable Property).
 Let $\phi\in\mathcal{G}(C)$ and let $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ be proper and lower semicontinuous with $\operatorname{dom}\phi\subset\operatorname{dom}f$, which is $\mathcal{C}^{1}$ on $C\equiv\operatorname{int}\operatorname{dom}\phi$.
 The pair of functions $(f,\phi)$ is said to be $L$-smooth adaptable (for short: $L$-smad) on $C$ if there exists $L>0$ such that $L\phi-f$ and $L\phi+f$ are convex on $C$.
-Lemma 2.4 (Primal progress from the smooth adaptable property).
-Let the pair of functions $(f,\phi)$ be $L$-smad with a strictly convex function $\phi\in\mathcal{G}(C)$ and let $x_{+}=(1-\gamma)x+\gamma v$ with $x,v\in\operatorname{int}\operatorname{dom}\phi$ and $\gamma\in[0,1]$.
-Then it holds:
-$f(x)-f(x_{+})\geq\gamma\langle\nabla f(x),x-v\rangle-L\gamma^{1+\nu}D_{\phi}(v,x).$ (2.2)
-Often the point $v\in P\subset\operatorname{int}\operatorname{dom}\phi$ in Lemma 2.4 is chosen as a Frank–Wolfe vertex, i.e., $v\in\operatorname*{argmax}_{u\in P}\langle\nabla f(x),x-u\rangle$, but other choices, e.g., those arising from away-directions from the away-step FW algorithm, are also possible, as we will see in Section 3.2.
-In [garber2015faster] (see also [Garber2020-tr]) the linear convergence of the FW algorithm for convex optimization under the quadratic growth condition, which is a weaker assumption than assuming strong convexity, was established over strongly convex sets and later generalized to uniformly convex sets in [KAP2020] as well as to conditions weaker than quadratic growth in [KDP2018, KDP2018jour].
-Local variants of these notions, as necessary, e.g., for the nonconvex case, have been studied in [KAP2021] in the context of FW algorithms.
-For $f:\mathbb{R}^{n}\to[-\infty,+\infty]$, let $[f\leq\zeta]:=\{x\in\mathbb{R}^{n}\,|\,f(x)\leq\zeta\}$ be a $\zeta$-sublevel set of $f$ for some $\zeta\in\mathbb{R}$.
-Definition 2.5 (Hölder Error Bound [Braun2022-ew, Definition 3.25], [alex17sharp, Definition 1.2] and Quadratic Growth Conditions [garber2015faster, Garber2020-tr, Liao2024-vd]).
+The $L$-smad property is equivalent to the extended descent lemma (Bolte et al., 2018, Lemma 2.1), which implies that the $L$-smad property for $(f,\phi)$ provides upper and lower approximations for $f$ majorized by $\phi$ with $L>0$.
+For example, $\log x$ is not $L$-smooth on $x>0$ and $\frac{1}{4}x^{4}$ is not $L$-smooth on $\mathbb{R}$.
+It is the Bregman distance, rather than the Euclidean distance, that allows us to bound the first-order approximation of these functions.
+See Example B.5 for further examples of $L$-smad pairs.
+When $(f,\phi)$ is $L$-smad with a strictly convex function $\phi\in\mathcal{G}(C)$, we have Lemma B.7, i.e.,
+$f(x)-f(x_{+})\geq\gamma\langle\nabla f(x),x-v\rangle-L\gamma^{1+\nu}D_{\phi}(v,x),$ (2.2)
+where $x_{+}=(1-\gamma)x+\gamma v$ with $x,v\in\operatorname{int}\operatorname{dom}\phi$ and $\gamma\in[0,1]$.
+Often $v\in P\subset\operatorname{int}\operatorname{dom}\phi$ in Lemma B.7 is chosen as a Frank–Wolfe vertex, i.e., $v\in\operatorname*{arg\,max}_{u\in P}\langle\nabla f(x),x-u\rangle$, but other choices, e.g., those arising from away-directions, are also possible, as we will see in Section 3.2.
+In Garber and Hazan (2015) (see also Garber (2020)) the linear convergence of the FW algorithm for convex optimization under the quadratic growth condition, which is a weaker assumption than assuming strong convexity, was established over strongly convex sets and later generalized to uniformly convex sets in Kerdreux et al.
+(2021b) as well as to conditions weaker than quadratic growth in Kerdreux et al.
+(2019; 2022).
+Local variants of these notions, as necessary, e.g., for the nonconvex case, have been studied in Kerdreux et al.
+(2021a) in the context of FW algorithms.
+For $f:\mathbb{R}^{n}\to[-\infty,+\infty]$, let $f\leq\zeta]:=\{x\in\mathbb{R}^{n}\,|\,f(x)\leq\zeta\}$ be a $\zeta$-sublevel set of $f$ for some $\zeta\in\mathbb{R}$.
+Definition 2.4 (Hölder Error Bound (Braun et al., 2025; Roulet and d’Aspremont, 2017) and Quadratic Growth Conditions (Garber and Hazan, 2015; Garber, 2020; Liao et al., 2024)).
 Let $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ be a proper lower semicontinuous function and $P\subset\mathbb{R}^{n}$ be a compact convex set.
 Let $\mathcal{X}^{*}\neq\emptyset$ be the set of optimal solutions, i.e., $\mathcal{X}^{*}:=\operatorname*{arg\,min}_{x\in P}f(x)$, and let $f^{*}=\min_{x\in P}f(x)$ and $\zeta>0$.
-The function $f$ satisfies the Hölder error bound (HEB) condition on $P$ if there exists constants $q\geq 1$ and $\mu>0$ such that, for any $x\in P$,
-$\operatorname{dist}(x,\mathcal{X}^{*})^{q}\leq\frac{q}{\mu}(f(x)-f^{*}).$ (2.3)
-Particularly, when $q=2$, it is called the $\mu$-quadratic growth condition.
-Moreover, the function $f$ is said to be local $\mu$-quadratic growth (with $\zeta$) if there exists a constant $\mu>0$ such that, for any $x\in[f\leq f^{*}+\zeta]\cap P$, (2.3) holds with $q=2$.
-The HEB condition shows sharpness bounds on the primal gap (see [bolte17, alex17sharp]), which has also been extensively analyzed for FW algorithms [KDP2018, KDP2018jour].
-Convergence for nonconvex optimization under the quadratic growth or HEB condition was established in [Davis2024-rm, Davis2018-bb, Davis2024-og].
-We employ the HEB condition for convex optimization, while we use the local quadratic growth for nonconvex optimization.
+The function $f$ satisfies the $q$-Hölder error bound (HEB) condition on $P$ if there exist constants $q\geq 1$ and $\mu>0$ such that $\operatorname{dist}(x,\mathcal{X}^{*})^{q}\leq\frac{q}{\mu}(f(x)-f^{*})$ holds for any $x\in P$.
+In particular, when $q=2$, it is called the $\mu$-quadratic growth condition.
+Moreover, the function $f$ is said to satisfy local $\mu$-quadratic growth (with $\zeta$) if there exists a constant $\mu>0$ such that $\operatorname{dist}(x,\mathcal{X}^{*})^{2}\leq\frac{2}{\mu}(f(x)-f^{*})$ holds for any $x\in[f\leq f^{*}+\zeta]\cap P$.
+For example, $f(x)=\log(1+x^{2})$ is nonconvex but satisfies local quadratic growth.
+This function is used for image restoration (Boţ et al., 2016; Stella et al., 2017).
+See, for more examples, Examples E.4 and E.5.
+The HEB condition shows sharpness bounds on the primal gap (see Bolte et al.
+(2017); Roulet and d’Aspremont (2017)), which has been extensively analyzed for FW algorithms (Kerdreux et al., 2019; 2022).
+Convergence of (sub)gradient algorithms for nonconvex optimization under the quadratic growth or HEB condition was established in Davis et al.
+(2018; 2024); Davis and Jiang (2024).
+We assume the HEB condition for convex optimization and the local quadratic growth for nonconvex optimization.
 3 Proposed Bregman FW Algorithms
@@
 Assumption 3.1.
-  1. (i)
-    $\phi\in\mathcal{G}(C)$ with $\operatorname{cl}{C}=\operatorname{cl}{\operatorname{dom}\phi}$ is strictly convex on $C\equiv\operatorname{int}\operatorname{dom}\phi$;
-  2. (ii)
-    $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is proper and lower semicontinuous with $\operatorname{dom}\phi\subset\operatorname{dom}f$, which is $\mathcal{C}^{1}$ on $C$;
-  3. (iii)
-    The pair $(f,\phi)$ is $L$-smad on $P$;
-  4. (iv)
-    $P\subset\mathbb{R}^{n}$ is a nonempty compact convex set with $P\subset C$.
-Assumption 3.1(i)-(iii) are standard for Bregman-type algorithms [Bolte2018-zt, Takahashi2024-wz].
-The strict convexity of $\phi$ ensures the existence of $\nu$ by Lemma A.3 and is satisfied by many kernel generating distances (see, e.g., [Bauschke2017-hg, Bauschke1997-vk], and [Dhillon2008-zz, Table 2.1]).
-Assumption 3.1(iv) is standard for FW algorithms, and $P\subset C$ is natural ensuring that $D_{\phi}$ is well-defined on $P$.
-The original FW algorithm is given in Algorithm 1.
+(i) $\phi\in\mathcal{G}(C)$with $\operatorname{cl}{C}=\operatorname{cl}{\operatorname{dom}\phi}$ is strictly convex on $C\equiv\operatorname{int}\operatorname{dom}\phi$ and has $\nu>0$ satisfying Equation 2.1; (ii) $f:\mathbb{R}^{n}\to(-\infty,+\infty]$is proper and lower semicontinuous with $\operatorname{dom}\phi\subset\operatorname{dom}f$, which is $\mathcal{C}^{1}$ on $C$; (iii) The pair $(f,\phi)$ is $L$-smad on $P$; (iv) $P\subset\mathbb{R}^{n}$is a nonempty compact convex set with $P\subset C$.
 Input: Initial point $x_{0}\in P$, objective function $f$, step-size $\gamma_{t}\in[0,1]$
@@
 2for $t=0,\ldots$ do
-3      $v_{t}\leftarrow\operatorname*{argmin}_{v\in P}\langle\nabla f(x_{t}),v\rangle$
-4     $x_{t+1}\leftarrow(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$
+[…]
+[…]
 Algorithm 1 Frank–Wolfe algorithm
+Assumption 3.1(i)-(iii) are standard for Bregman-type algorithms (Bolte et al., 2018; Takahashi and Takeda, 2025).
+The strict convexity of $\phi$ ensures the existence of $\nu$ by Lemma B.3 and is satisfied by many kernel generating distances (see, e.g., Bauschke et al.
+(2017); Bauschke and Borwein (1997), and (Dhillon and Tropp, 2008, Table 2.1)).
+Assumption 3.1(iv) is standard for FW algorithms, and $P\subset C$ naturally ensures that $D_{\phi}$ is well-defined on $P$.
+The original FW algorithm is given in Algorithm 1.
 In what follows, let $x^{*}\in\operatorname*{arg\,min}_{x\in P}f(x)$ and $f^{*}=f(x^{*})$.
-We recall a key property for analyzing convergence rates in convex optimization and derive its counterpart in weakly convex optimization.
-Lemma 3.2 (Primal gap, dual gap, and Frank–Wolfe gap for convex $f$ [Pokutta2024-qu, Lemma 4.1]).
-Suppose that Assumptions 3.1(ii), (iv), and that $f$ is convex. Then for all $x\in P$, it holds:
-$f(x)-f^{*}\leq\langle\nabla f(x),x-x^{*}\rangle\leq\max_{v\in P}\langle\nabla f(x),x-v\rangle.$ (3.1)
-Lemma 3.3 (Primal gap, dual gap, and Frank–Wolfe gap for weakly convex $f$).
-Suppose that Assumptions 3.1(ii), (iv), and that $f$ is $\rho$-weakly convex. Then for all $x\in P$, it holds:
-$f(x)-f^{*}\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}\leq\max_{v\in P}\langle\nabla f(x),x-v\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}.$ (3.2)
 3.1 Bregman Step-Size Strategy for $\gamma_{t}$
 Bregman Short Step-Size:
-Assume that $(f,\phi)$ is $L$-smad. Using the optimality condition of the right-hand side of (2.2) in terms of $\gamma$, we find
+Assume that $(f,\phi)$ is $L$-smad. Maximizing the right-hand side of Equation 2.2 in terms of $\gamma$, we find
 $\gamma=\left(\frac{\langle\nabla f(x),x-v\rangle}{L(1+\nu)D_{\phi}(v,x)}\right)^{\frac{1}{\nu}}\quad\text{and}\quad f(x)-f(x_{+})\geq\frac{\nu}{1+\nu}\frac{\langle\nabla f(x),x-v\rangle^{1+1/\nu}}{(L(1+\nu)D_{\phi}(v,x))^{1/\nu}}.$
@@
 We update $\gamma=\min\left\{\left(\frac{\langle\nabla f(x),x-v\rangle}{L(1+\nu)D_{\phi}(v,x)}\right)^{1/\nu},\gamma_{\max}\right\}$ with some $\gamma_{\max}\in\mathbb{R}$ (usually set $\gamma_{\max}=1$).
-This step-size strategy provides the progress lemma.
-Lemma 3.4 (Progress lemma from the Bregman short step-size).
-Suppose that Assumption 3.1 holds.
-Define $x_{t+1}=(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$.
-Then if $x_{t+1}\in\operatorname{dom}f$, $f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle$ for $0\leq\gamma_{t}\leq\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu}$.
+This step-size strategy provides $f(x)-f(x_{+})\geq\frac{\nu}{1+\nu}\gamma\langle\nabla f(x),x-v\rangle$ for $0\leq\gamma\leq\left(\frac{\langle\nabla f(x),x-v\rangle}{L(1+\nu)D_{\phi}(v,x)}\right)^{1/\nu}$.
+Its proof can be found in Lemma B.8.
 Output : Estimates $\tilde{L}^{*}$ and $\tilde{\nu}^{*}$, step-size $\gamma$
+1 Procedure $\textnormal{{step\_size}}(f,\phi,x,v,\tilde{L},\gamma_{\max})$
+2  Choose $\beta\in(0,1)$, $\eta\in(0,1]$, and $\tau>1$
 […]
-2      Choose $\beta<1$, $\eta\leq 1$, and $\tau>1$
-3      $M\leftarrow\eta\tilde{L}$, $\kappa\leftarrow 1$
-4      while $D_{f}(x_{+},x)\leq M\gamma^{1+\kappa}D_{\phi}(v,x)$ do
+4loop
 5   $\gamma\leftarrow\min\left\{\left(\frac{\langle\nabla f(x),x-v\rangle}{M(1+\kappa)D_{\phi}(v,x)}\right)^{1/\kappa},\gamma_{\max}\right\}$
 6   $x_{+}\leftarrow(1-\gamma)x+\gamma v$
-7           $M\leftarrow\tau M$
-8           $\kappa\leftarrow\beta\kappa$
-9
+7   if $D_{f}(x_{+},x)\leq M\gamma^{1+\kappa}D_{\phi}(v,x)$ then
 8    $\tilde{L}^{*}\leftarrow M$, $\tilde{\nu}^{*}\leftarrow\kappa$
 9    return $\tilde{L}^{*}$, $\tilde{\nu}^{*}$, $\gamma$
+10
+11    $M\leftarrow\tau M$, $\kappa\leftarrow\beta\kappa$
 12
+13
 Algorithm 2 Adaptive Bregman step-size strategy
@@
 Adaptive Bregman Step-Size:
-The exact value or the tight upper bound of $L$ is often unknown, however, the algorithm’s performance heavily depends on it; an underestimation of $L$ might lead to non-convergence, and an overestimation of $L$ might lead to slow convergence.
+The exact value or the tight upper bound of $L$ is often unknown; however, the algorithm’s performance heavily depends on it; an underestimation of $L$ might lead to non-convergence, and an overestimation of $L$ might lead to slow convergence.
 Moreover, a worst-case $L$ might be too conservative for regimes where the function is better behaved.
-Due to all these reasons, Pedregosa et al.
-[Pedregosa2020-ay] proposed an adaptive step-size strategy for FW algorithms in the case where $\phi=\frac{1}{2}\|\cdot\|^{2}$.
+For these reasons, Pedregosa et al.
+(2020) proposed an adaptive step-size strategy for FW algorithms in the case where $\phi=\frac{1}{2}\|\cdot\|^{2}$.
 We present our algorithm in Algorithm 2.
 This step-size strategy can be used as a drop-in replacement.
-For example, inserting $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},1)$ between lines 1 and 1 in Algorithm 1, we obtain that Algorithm 2 searches $L$ and $\nu$ satisfying (2.2).
+For example, inserting $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},1)$ between lines 1 and 1 in Algorithm 1, we obtain that Algorithm 2 searches $L$ and $\nu$ satisfying Equation 2.2.
+For $\phi=\frac{1}{2}\|\cdot\|^{2}$, Algorithm 2 corresponds to a (Euclidean) adaptive step-size strategy (Pedregosa et al., 2020).
 Remark 3.2 (Well-definedness and termination of Algorithm 2).
-By the $L$-smad property of $(f,\phi)$ and Lemma 2.4, we know that (2.2) holds for all $M\geq L$ and $\nu\geq\kappa>0$.
-Therefore, the loop in line 2 in Algorithm 2 is well-defined and guaranteed to terminate.
-It suffices to update $\kappa$ in line 2 only if (2.1) does not hold.
+By the $L$-smad property of $(f,\phi)$ and Lemma B.7, we know that Equation 2.2 holds for all $M\geq L$ and $\nu\geq\kappa>0$.
+Therefore, the condition in line 2 in Algorithm 2 is well-defined and guaranteed to terminate.
+It suffices to update $\kappa$ in line 2 only if Equation 2.1 does not hold.
 Thus, $\kappa$ cannot become too small.
+We have a bound on the total number of evaluations of Algorithm 2. Its proof can be found in Appendix B.3.
+Theorem 3.3.
+Let $L_{-1}$ be the initial $L$-smad estimate and $n_{t}$ be the total number of evaluations of Equation 2.2 up to iteration $t$.
+Then we have $n_{t}\leq\max\{(1-\log\eta/\log\tau)(t+1)+\max\{\log(\tau L/L_{-1}),0\}/\log\tau,(1+\log\nu/\log\beta)(t+1)\}$.
+Pedregosa et al.
+(2020) showed that, asymptotically, no more than 16% of the iterations require more than one evaluation of the line search procedures when $\eta=0.9$ and $\tau=2$.
+This follows from the bound $(1-\log\eta/\log\tau)\leq 1.16$.
 3.2 Bregman Away-Step Frank–Wolfe Algorithm
+In the case where $P$ is a polytope, the classical FW algorithm might zigzag when approaching the optimal face and, in consequence, converge slowly.
+The away-step FW algorithm overcomes this drawback.
+Away steps allow the algorithm to move away from vertices in a convex combination of $x_{t}$, which can effectively short-circuit the zigzagging.
+The convergence properties of this algorithm were unknown for a long time, and it was only quite recently that Lacoste-Julien and Jaggi (2015) established the linear convergence of the away-step FW algorithm.
+Inspired by GuéLat and Marcotte (1986); Lacoste-Julien and Jaggi (2015); Wolfe (1970), we propose a variant of the away-step FW algorithm utilizing Bregman distances as given in Algorithm 3.
+The main difference between Algorithm 3 and the existing away-step FW algorithm lies in the update in line 3.
+For the following discussion, we introduce some notions.
+In the same way as Braun et al.
+(2025), we define an active set $\mathcal{S}\subset\operatorname{Vert}P$, where $\operatorname{Vert}P$ denotes the set of vertices of $P$, and the away vertex as $v_{t}^{\textup{A}}\in\operatorname*{arg\,max}_{v\in\mathcal{S}}\langle\nabla f(x_{t}),v\rangle$, where $x_{t}$ is a (strict) convex combination of elements in $\mathcal{S}$, i.e., $x_{t}=\sum_{v\in\mathcal{S}}\lambda_{v}v$ with $\lambda_{v}>0$ for all $v\in\mathcal{S}$ and $\sum_{v\in\mathcal{S}}\lambda_{v}=1$.
+If $\nu$ is unknown, we can add line search procedures to search $\nu$ until Equation 2.1 holds.
+If $\nu$ and $L$ are unknown, we can use $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},\gamma_{t,\max})$ with $\gamma_{t}\leftarrow\min\left\{\left(\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{M(1+\kappa)D_{\phi}(v_{t},x_{t})}\right)^{1/\kappa},\gamma_{t,\max}\right\}$ instead of line 2 in Algorithm 2.
+When $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu\equiv 1$, Algorithm 3 corresponds to the Euclidean away-step FW algorithm.
 Input: $x_{0}\in\operatorname*{arg\,min}_{v\in P}\langle\nabla f(x),v\rangle$ for $x\in P$, $\beta<1$
@@
 3for $t=0,\ldots$ do
-4      $v_{t}^{\textup{FW}}\leftarrow\operatorname*{argmin}_{v\in P}\langle\nabla f(x_{t}),v\rangle$, $v_{t}^{\textup{A}}\leftarrow\operatorname*{argmax}_{v\in\mathcal{S}_{t}}\langle\nabla f(x_{t}),v\rangle$
+[…]
 5
-6     if $\langle\nabla f(x_{t}),x_{t}-v_{t}^{\textup{FW}}\rangle\geq\langle\nabla f(x_{t}),v_{t}^{\textup{A}}-x_{t}\rangle$ then
+6if $\langle\nabla f(x_{t}),x_{t}-v_{t}^{\textup{FW}}\rangle\geq\langle\nabla f(x_{t}),v_{t}^{\textup{A}}-x_{t}\rangle$ then
 7   $v_{t}\leftarrow v_{t}^{\textup{FW}}$, $d_{t}\leftarrow x_{t}-v_{t}^{\textup{FW}}$, $\gamma_{t,\max}\leftarrow 1$
@@
 Algorithm 3 Away-step Frank–Wolfe algorithm with the Bregman distance
-In the case where $P$ is a polytope, the classical FW algorithm might zigzag when approaching the optimal face and, in consequence, converge slowly.
-Away-step FW algorithm overcomes this drawback.
-Away steps allow the algorithm to move away from vertices in a convex combination of $x_{t}$, which can effectively shortcircuit the zigzagging.
-The convergence properties of this algorithm were unknown for a long time, and it was only quite recently that Lacoste-Julien and Jaggi [Lacoste-Julien2015-gb] established the linear convergence of the away-step FW algorithm.
-Inspired by [GueLat1986-bg, Lacoste-Julien2015-gb, Wolfe1970-ax], we propose a variant of the away-step FW algorithm utilizing Bregman distances as given in Algorithm 3.
-For the following discussion, we introduce some notions.
-In the same way [Braun2022-ew], we define an active set $\mathcal{S}\subset\operatorname{Vert}P$, where $\operatorname{Vert}P$ denotes the set of vertices of $P$, and the away vertex as $v_{t}^{\textup{A}}\in\operatorname*{argmax}_{v\in\mathcal{S}}\langle\nabla f(x_{t}),v\rangle$, where $x_{t}$ is a (strict) convex combination of elements in $\mathcal{S}$, i.e., $x_{t}=\sum_{v\in\mathcal{S}}\lambda_{v}x_{t}$ with $\lambda_{v}>0$ for all $v\in\mathcal{S}$ and $\sum_{v\in\mathcal{S}}\lambda_{v}=1$.
-If $\nu$ is unknown, we can add line search procedures to search $\nu$ until (2.1) holds.
-If $\nu$ and $L$ are unknown, we can use $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},\gamma_{t,\max})$ with $\gamma_{t}\leftarrow\min\left\{\left(\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{M(1+\kappa)D_{\phi}(v_{t},x_{t})}\right)^{1/\kappa},\gamma_{t,\max}\right\}$ instead of line 2 in Algorithm 2.
-When $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu\equiv 1$, Algorithm 3 corresponds to the Euclidean away-step FW algorithm.
 4 Linear Convergence for Convex Optimization
@@
 The objective function $f$ is convex.
-We will now establish faster convergence rates than $O(1/t)$ up to linear convergence depending on the choice of parameters.
-First, we establish the faster convergence of Algorithm 1 with the Bregman short step-size, i.e., $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu},1\right\}$, in the case where the optimal solution lies in the relative interior.
 Let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$.
+We will now establish faster convergence rates than $O(1/t)$ up to linear convergence depending on $\nu$ in Equation B.2 and $q$ in Definition 2.4.
+First, we establish the faster convergence of Algorithm 1 with the Bregman short step-size, i.e., $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu},1\right\}$, or Algorithm 2 in the case where the optimal solution lies in the relative interior.
+Its proof can be found in Appendix D.2.
 Theorem 4.2 (Linear convergence of FW algorithm with short step-size or adaptive step-size).
 Suppose that Assumptions 3.1 and 4.1 hold.
-Let $\nu>0$ and let $f$ satisfy the HEB condition (2.3) with $q\geq 1+\nu$ and $\mu>0$ and $D$ be the diameter of $P$.
+Let $\nu>0$ and let $f$ satisfy the HEB condition with $q\geq 1+\nu$ and $\mu>0$ and $D$ be the diameter of $P$.
 Assume that there exists a minimizer $x^{*}\in\operatorname{int}P$, i.e., there exists an $r>0$ with $B(x^{*},r)\subset P$.
@@
 Next, we establish the linear convergence of Algorithm 3.
-Recall that the pyramidal width is the minimal $\delta>0$ satisfying Lemma B.9 (see [Braun2022-ew, Lemma 2.26] or [Lacoste-Julien2015-gb, Theorem 3] for an in-depth discussion).
+Recall that the pyramidal width is the minimal $\delta>0$ satisfying Lemma C.9 (see (Braun et al., 2025, Lemma 2.26) or (Lacoste-Julien and Jaggi, 2015, Theorem 3) for an in-depth discussion).
 Under Assumption 4.3, it holds that $\delta\leq D$, and this will be important in establishing convergence rates of Algorithm 3.
@@
 The kernel generating distance $\phi$ is $\sigma$-strongly convex.
+We have the linear convergence of Algorithm 3. Its proof can be found in Appendix D.3.
 Theorem 4.4 (Linear convergence of the away-step FW algorithm).
@@
 where $c:=(q/\mu)^{1/q}$ $D$ and $\delta$ are the diameter and the pyramidal width of the polytope $P$, respectively, and $t_{0}:=\max\{\lfloor\log_{1/(1+\nu)}(L(1+\nu)D^{2}/((\mu/q)^{1/q}\delta/2)^{1+\nu})^{q/(q-1-\nu)}/LD^{2}\rfloor+2,1\}$.
+When $\phi=\frac{1}{2}\|\cdot\|^{2}$, i.e., $\nu=1$, Theorems 4.2 and 4.4 correspond to existing results (Braun et al., 2025; Kerdreux et al., 2022).
 5 Local Linear Convergence for Nonconvex Optimization
 In this section, we consider a nonconvex objective function.
-We establish global sublinear, local sublinear, and local linear convergence (see, for global and local sublinear, Appendices D.1 and D.2).
+We establish global sublinear, local sublinear, and local linear convergence (see Theorem E.1 and Theorem E.3 for global and local sublinear convergence, respectively).
 We show that the FW algorithm converges to a minimizer $x^{*}$ when an initial point is close enough to $x^{*}$.
@@
 Assumption 5.1.
-$f$ is $\rho$-weakly convex and local $\mu$-quadratic growth with $\zeta>0$.
+$f$ is $\rho$-weakly convex and has local $\mu$-quadratic growth with $\zeta>0$.
+Assumption 5.1 is not too restrictive because the (local) strong convexity or the (local) Polyak–Łojasiewicz (PL) inequality implies the (local) quadratic growth condition for weakly convex functions (Liao et al., 2024, Theorem 3.1).
+For more examples satisfying Assumption 5.1 and $\rho\leq\mu$, see Examples E.4 and E.5.
+Remark 5.2.
+Linear convergence rates for nonconvex optimization problems are, to the best of our knowledge, the first such result for the Frank–Wolfe algorithm.
+Their proofs are technically challenging.
+Proposition C.8 is new and key for establishing a linear rate.
+In the nonconvex case, one cannot derive primal-gap inequalities such as Lemma D.1.
+We overcome this by restricting our function class to weakly convex functions and proving Lemma C.2.
+There remains an obstacle that could be resolved by the assumption of the quadratic growth condition.
+Assumption 5.1 enables us to derive Lemma C.3 and Proposition C.8 (as well as Remark C.4), which are crucial for establishing linear convergence.
+We have local linear convergence, and its proof can be found in Appendix E.3.
 Theorem 5.3 (Local linear convergence of FW algorithm with short step-size or adaptive step-size).
@@
 where $c=\sqrt{2\mu}/(\mu-\rho)$, $M:=(L(1+\nu))^{1/\nu}$, and $t_{0}:=\max\{\lfloor\log_{\left(1+\nu\rho/\mu\right)/(1+\nu)}((L(1+\nu)(1-\rho/\mu)^{\nu}c^{1+\nu}D^{2}/r^{1+\nu})^{2/(1-\nu)}/LD^{2})\rfloor+2,1\}$.
-Because Algorithm 2 is well-defined (see also Remark 3.5), the convergence result of the FW algorithm with Algorithm 2 is the same as Theorem 5.2.
-Finally, we establish the local linear convergence of Algorithm 3. We also assume $\rho\leq L$, which is not restrictive (see also the discussion in Appendix D.2 and Example D.4).
+Finally, we establish the local linear convergence of Algorithm 3. Its proof can be found in Appendix E.4. We assume $\rho\leq L$, which is not restrictive (see Appendix E.2 and Example E.4).
 Theorem 5.4 (Local linear convergence by the away-step FW algorithm).
@@
 where $D$ and $\delta$ are the diameter and the pyramidal width of $P$, respectively, and $\omega:=(\mu-\rho)^{2}/8\mu$, and $t_{0}:=\max\{\lfloor\log_{(1+\nu\rho/\mu)/(1+\nu)}(L^{2}(1+\nu)^{2}D^{4}(1-\rho/\mu)^{2\nu}/(\omega\delta^{2})^{1+\nu})^{1/(1-\nu)}/2LD^{2}\rfloor+2,1\}$.
-Without loss of generality, we set $\sigma=1$.
-If $\rho\leq L$ does not hold, we can use $f(x_{1})-f^{*}\leq(\rho+L)D^{2}$, instead of $f(x_{1})-f^{*}\leq 2LD^{2}$ as the initial bound.
-Moreover, when $P$ is $\alpha$-strongly convex set and $\phi=\frac{1}{2}\|\cdot\|^{2}$, we also establish local linear convergence.
-Its proof can be found in Appendix D.5.
+Without loss of generality, $\sigma=1$.
+If $\rho>L$, we can use $(\rho+L)D^{2}$, instead of $2LD^{2}$ as the initial bound of $f(x_{1})-f^{*}$.
+When $\phi=\frac{1}{2}\|\cdot\|^{2}$, i.e., $\nu=1$, Theorems 5.3 and 5.4 show linear convergence rates for Euclidean cases.
+Even if we use the Euclidean distance, our convergence results are novel in that linear convergence is achieved for the first time for a certain type of nonconvex problem.
+Moreover, when $P$ is $(\alpha,p)$-uniformly convex set (see Definition E.7) and $\phi=\frac{1}{2}\|\cdot\|^{2}$, we also establish local linear convergence (Theorems E.9 and E.10) in Appendix E.5.
 6 Numerical Experiments
-In this section, we conducted numerical experiments to examine the performance of our proposed algorithms.
-All numerical experiments were performed in Julia 1.11 using the FrankWolfe.jl package [Besancon2022-lx] [^2] https://github.com/ZIB-IOL/FrankWolfe.jl on a MacBook Pro with an Apple M2 Max and 64GB LPDDR5 memory.
-We compare BregFW and BregAFW, which use the adaptive Bregman step-size strategy (Algorithm 2), with existing algorithms, EucFW, EucAFW, ShortFW, ShortAFW, OpenFW, OpenAFW, and MD (their notation is described in Appendix E).
+In this section, we conduct numerical experiments to examine the performance of our algorithms.
+All numerical experiments were performed in Julia 1.11 using the FrankWolfe.jl package (Besançon et al., 2022) [^8] https://github.com/ZIB-IOL/FrankWolfe.jl on a MacBook Pro with an Apple M2 Max and 64GB LPDDR5 memory.
+We compare BregFW (Algorithm 1) and BregAFW (Algorithm 3, whose results are shown in Figures 7) using the adaptive Bregman step-size strategy (Algorithm 2), with existing algorithms, EucFW, EucAFW, ShortFW, ShortAFW, OpenFW, OpenAFW, MD, and ProjGD (their notation is described in Appendix F).
+In this section, we consider $\ell_{p}$ loss problems and phase retrieval, and we report additional experiments on nonnegative linear inverse problems, low-rank minimization, and NMF in Appendix F.
 Note that we included OpenAFW only for comparison purposes; there is no established convergence theory for the away-step FW algorithm with the open loop.
-Indeed, there are no proper drop steps (operations corresponding to line 3 in Algorithm 3), and the favorable properties of the away-step FW algorithm are lost; see [Braun2022-ew].
+Indeed, there are no proper drop steps (operations corresponding to line 3 in Algorithm 3), and the favorable properties of the away-step FW algorithm are lost; see Braun et al.
+(2025).
 We use $\beta=0.9$, $\eta=0.9$, $\tau=2$, $\gamma_{\max}=1$, and the termination criterion $\max_{v\in P}\langle\nabla f(x_{t}),x_{t}-v\rangle\leq 10^{-7}$ throughout all experiments.
@@
 Figure 2: Log plot of primal and FW gaps on $\ell_{p}$ loss for gas sensor data with $b_{\max}=200$.
-Figure 3: Log plot of primal and FW gaps on phase retrieval for $(m,n)=(1000,10000)$ and $K=2000$.
-Figure 4: Log plot of primal and FW gaps on phase retrieval for $(m,n)=(2000,10000)$ and $K=2000$.
+Figure 3: Log plot of primal and FW gaps on phase retrieval for $(m,n)=(1000,10000)$.
+Figure 4: Log plot of primal and FW gaps on phase retrieval for $(m,n)=(2000,10000)$.
 6.1 $\ell_{p}$ Loss Problem
-We consider the $\ell_{p}$ loss problem [Kyng2015-uo, Maddison2021-nt] to find $x\in P$ such that $Ax\simeq b$, defined by a convex optimization problem $\min_{x\in P}\|Ax-b\|_{p}^{p}$, where $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, and $p>1$.
+We consider the $\ell_{p}$ loss problem (Kyng et al., 2015; Maddison et al., 2021) to find $x\in P$ such that $Ax\simeq b$, defined by a convex optimization problem $\min_{x\in P}\|Ax-b\|_{p}^{p}$, where $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, and $p>1$.
 Let $f(x)=\|Ax-b\|_{p}^{p}$.
@@
 Furthermore, when $p<2$, $f$ is not $\mathcal{C}^{2}$ but $\mathcal{C}^{1}$.
-Therefore, for $21<p<$ 2, $\nabla f$ is also not Lipschitz continuous over compact sets.
+Therefore, for $1<p<2$, $\nabla f$ is also not Lipschitz continuous over compact sets.
 Since $f$ is convex, the pair $(f,\phi)$ is 1-smad with $\phi:=f$.
-We consider real-world data using gas sensor data.§§§The dataset is freely available at https://archive.ics.uci.edu/dataset/270/gas+sensor+array+drift+dataset+at+different+concentrations The data size is $(m,n)=(13910,128)$.
+We consider real-world gas sensor data.
+[^9] The dataset is freely available at https://archive.ics.uci.edu/dataset/270/gas+sensor+array+drift+dataset+at+different+concentrations The data size is $(m,n)=(13910,128)$.
 We use the $\ell_{2}$ constraint $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|^{2}\leq b_{\max}\}$.
-We compared BregFW with EucFW and OpenFW because ShortFW does not converge.
-The maximum iteration is 1000 except for OpenFW, and that of OpenFW is 5000.
+We compare BregFW with EucFW, OpenFW, and ProjGD because ShortFW does not converge.
+The maximum number of iterations is 1000 for all methods except OpenFW, for which it is 5000.
 The initial point $x_{0}$ is an all-ones vector.
@@
 Figures 4 and 4 show the primal and FW gaps per iteration and gaps per second for $b_{\max}=130$ and $b_{\max}=200$, respectively.
-Because $\nabla f$ is not Lipschitz continuous on $P$ from lacking $\mathcal{C}^{2}$, ShortFW and EucFW do not converge.
-The primal gap and the FW gap by BregFW are the smallest among these step-size strategies because only BregFW has a theoretical guarantee for this problem.
+Because $\nabla f$ is not Lipschitz continuous, ShortFW and EucFW do not converge.
+The primal gap and the FW gap achieved by BregFW are the smallest among these methods because only BregFW has a theoretical guarantee for this problem.
 6.2 Phase Retrieval
-We are interested in phase retrieval, which involves finding a signal $x\in\mathbb{R}^{n}$ such that $|\langle a_{i},x\rangle|^{2}\simeq b_{i}$, $i=1,\ldots,m$, where $a_{i}\in\mathbb{R}^{n}$ describes the model and $b\in\mathbb{R}^{m}$ is a vector of measurements.
-Phase retrieval arises in many fields of science and engineering, such as image processing [Candes2015-mk], X-ray crystallography [Patterson1934-sg, Patterson1944-nm], and optics [Shechtman2015-ic].
-To achieve the goal of phase retrieval, we focus on the nonconvex optimization problem $\min_{x\in P}f(x):=\frac{1}{4}\sum_{i=1}^{m}\left(|\langle a_{i},x\rangle|^{2}-b_{i}\right)^{2}$, where $P\subset\mathbb{R}^{n}$ is a compact convex set.
-Let $\phi$ be defined by $\phi(x)=\frac{1}{4}\|x\|^{4}+\frac{1}{2}\|x\|^{2}$, which is $\sigma$-strongly convex for $\sigma\leq 1$.
-For any $L$ satisfying $L\geq\sum_{i=1}^{m}\left(3\|a_{i}\|^{4}+\|a_{i}\|^{2}|b_{i}|\right)$, the pair $(f,\phi)$ is $L$-smad on $\mathbb{R}^{n}$ [Bolte2018-zt, Lemma 5.1].
-In addition, $\nabla^{2}f(x)+\rho I$ is positive semidefinite for $\rho\geq\sum_{i=1}^{m}\|a_{i}\|^{2}|b_{i}|$, i.e., $f$ is $\rho$-weakly convex.
-Therefore, $L\geq\rho/\sigma$ holds with $\sigma=1$.
+We are interested in phase retrieval, which involves finding a signal $x\in P\subset\mathbb{R}^{n}$ such that $|\langle a_{i},x\rangle|^{2}\simeq b_{i}$, $i=1,\ldots,m$, where $a_{i}\in\mathbb{R}^{n}$ describes the model and $b\in\mathbb{R}^{m}$ is a vector of measurements.
+Phase retrieval arises in many fields of science and engineering, such as image processing (Candés et al., 2015), X-ray crystallography (Patterson, 1934; 1944), and optics (Shechtman et al., 2015).
+To achieve the goal of phase retrieval, we focus on the nonconvex optimization problem $\min_{x\in P}f(x):=\frac{1}{4}\sum_{i}(|\langle a_{i},x\rangle|^{2}-b_{i})^{2}$.
+Let $\phi$ be defined by $\sigma$-strongly convex $\phi(x)=\frac{1}{4}\|x\|^{4}+\frac{1}{2}\|x\|^{2}$ for $\sigma\leq 1$.
+For any $L$ satisfying $L\geq\sum_{i}(3\|a_{i}\|^{4}+\|a_{i}\|^{2}|b_{i}|)$, the pair $(f,\phi)$ is $L$-smad on $\mathbb{R}^{n}$ (Bolte et al., 2018, Lemma 5.1).
+In addition, $\nabla^{2}f(x)+\rho I$ is positive semidefinite for $\rho\geq\sum_{i}\|a_{i}\|^{2}|b_{i}|$, i.e., $f$ is $\rho$-weakly convex.
+Thus, $L\geq\rho/\sigma$ holds with $\sigma=1$.
 We use a $K$-sparse polytope as a constraint, i.e., $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|_{1}\leq K,\|x\|_{\infty}\leq 1\}$.
-We compared BregFW with EucFW, ShortFW, and OpenFW.
-The maximum iteration is 1000 except for OpenFW, and that of OpenFW is 5000.
+We compare BregFW with EucFW, ShortFW, and OpenFW.
+The maximum number of iterations is 1000 for all methods except OpenFW, for which it is 5000.
 We generated $a_{i}$, $i=1,\ldots,m$ from an i.i.d.
-normal distribution and normalized it to have norm 1 (random seed 1234).
+normal distribution and normalized them to have norm 1 (random seed 1234).
 We also generated $x^{*}$ from an i.i.d.
@@
 Summary:
-We propose FW algorithms with an adaptive Bregman step-size strategy for a smooth adaptable function, which does not require Lipschitz continuity of its gradient.
-We have established sublinear to linear convergence rates depending on various assumptions for convex and nonconvex optimization.
-We also propose a variant of the away-step FW algorithm using Bregman distances over polytopes.
-For the away-step FW algorithm, we have also established global convergence for convex optimization under the HEB condition and its local linear convergence for nonconvex optimization under the local quadratic growth condition.
+We propose FW algorithms with an adaptive Bregman step-size strategy for smooth adaptable functions, which do not require Lipschitz continuous gradients.
+We have established convergence rates ranging from sublinear to linear, depending on various assumptions for convex and nonconvex optimization.
+We also propose a variant of the away-step FW algorithm using Bregman distances over polytopes and have established global convergence for convex optimization under the HEB condition and local linear convergence for nonconvex optimization under the local quadratic growth condition.
 Numerical experiments show that our algorithms outperform existing algorithms.
-Limitation and future work:
-Our step-size strategy for a smooth adaptable function requires the value of $\nu$, which can be estimated, because we use the Bregman distance.
+Limitations and future work:
+Our step-size strategy for $L$-smad functions requires the value of $\nu$, which must be estimated because it depends on the choice of the Bregman distance.
 Theoretically, linear convergence of the FW algorithm always holds for convex optimization if the exponent of the HEB condition $q$ equals the scaling exponent of the Bregman distance $1+\nu$, i.e., $q=1+\nu$.
-For $q>1+\nu$, if $t\leq t_{0}$ with some $t_{0}\in\mathbb{N}$, linear convergence holds; otherwise, sublinear convergence.
-The condition $q=1+\nu$, which is derived from the growth of $f$ and $\phi$, is natural because $\phi$ has a similar structure to $f$.
-On the other hand, for nonconvex optimization, we assume the local quadratic growth condition, which is the special case of the HEB condition with $q=2$.
+For $q>1+\nu$, if $t\leq t_{0}$, linear convergence holds; otherwise, sublinear convergence.
+The condition $q=1+\nu$, derived from the growth of $f$ and $\phi$, is natural because $\phi$ has a similar structure to $f$.
+On the other hand, for nonconvex optimization, we assume the local quadratic growth condition, i.e., the $2$-HEB condition.
 Since this is derived from the weak convexity of $f$, relaxing the quadratic growth condition may not be possible.
-In that regard, one direction is to consider an extension using the difference of convex functions (DC) optimization.
+One direction is to consider an extension using the difference of convex functions (DC) optimization.
 Recently, Maskan et al.
-[Maskan2025-oa] incorporated DC optimization into FW algorithms.
-While they established sublinear convergence, our algorithms may exploit DC optimization techniques.
-Acknowledgement
+(2025) incorporated DC optimization into FW algorithms and established sublinear convergence.
+Our algorithms could potentially benefit from DC optimization techniques.
+Acknowledgments
 This project has been partially supported by the Japan Society for the Promotion of Science (JSPS) through Invitational Fellowships for Research in Japan program (L24503), JSPS KAKENHI Grant Number JP23K19953 and JP25K21156; JSPS KAKENHI Grant Number JP23H03351 and JST ERATO Grant Number JPMJER1903.
-Research reported in this paper was also partially supported the Deutsche Forschungsgemeinschaft (DFG) through the DFG Cluster of Excellence MATH+ (grant number EXC-2046/1, project ID 390685689).
-Appendix A Results from Smooth Adaptable Property
-A.1 Properties of Bregman Distances
+Research reported in this paper was also partially supported by the Deutsche Forschungsgemeinschaft (DFG) through the DFG Cluster of Excellence MATH+ (grant number EXC-2046/1, project ID 390685689).
+LLM Usage
+We used large language models, ChatGPT-5/5.1/5.2, solely to polish the grammar and wording of the author-written text (e.g., to improve clarity and tone).
+The model did not generate novel ideas, analyses, code, or results.
+All outputs were reviewed and edited by the authors, who take full responsibility for the content.
+Ethics Statement
+This study uses only publicly available datasets that contain no personally identifiable information.
+We did not collect any new human-subject data; therefore, IRB approval was not required.
+We also reviewed the datasets for potentially harmful content.
+Reproducibility Statement
+We include all evaluation code, configuration files, and scripts in the supplementary material.
+We provide the exact preprocessing steps for all datasets used.
+We fix and disclose random seeds for all experiments.
+Results are reported as mean $\pm$ standard deviation over multiple runs.
+Appendix A Related Work
+Overview of Existing Bregman-Based Algorithm Studies:
+The Bregman gradient method was first introduced as the mirror descent by Nemirovskij and Yudin (1983).
+Accelerated mirror descent methods were proposed by Krichene et al.
+(2015); Allen-Zhu and Orecchia (2017).
+Some first-order methods based on the Bregman distance do not require the global Lipschitz continuity of $\nabla f$.
+Bolte et al.
+(2018) proposed the Bregman proximal gradient algorithm (BPG) for nonconvex optimization problems and established its global convergence under the smooth adaptable property (see Definition 2.3).
+Hanzely et al.
+(2021) introduced the accelerated BPGs.
+Takahashi et al.
+(2022) developed the BPG for the DC optimization and its accelerated version.
+The subproblem of BPG is often not solvable in closed form.
+Takahashi and Takeda (2025); Fujiki et al.
+(2025) used approximations to Bregman distances to solve subproblems in closed form.
+Takahashi et al.
+(2026) developed the majorization-minimization BPG and its accelerated version, and applied them to NMF.
+Rebegoldi et al.
+(2018) proposed an inexact version of the Bregman proximal gradient algorithm.
+Appendix B Results from Smooth Adaptable Property
+B.1 Properties of Bregman Distances
 The Bregman distance $D_{\phi}(x,y)$ measures the proximity between $x\in\operatorname{dom}\phi$ and $y\in\operatorname{int}\operatorname{dom}\phi$.
@@
 The Bregman distance is also called the Bregman divergence.
-Example A.1.
-Well-known choices for $\phi$ and $D_{\phi}$ are listed below; for more examples, see, e.g., [Bauschke2017-hg, Example 1], [Bauschke1997-vk, Section 6], and [Dhillon2008-zz, Table 2.1].
+Example B.1.
+Well-known choices for $\phi$ and $D_{\phi}$ are listed below; for more examples, see, e.g., (Bauschke et al., 2017, Example 1), (Bauschke and Borwein, 1997, Section 6), and (Dhillon and Tropp, 2008, Table 2.1).
   1. (i)
@@
   2. (ii)
-    The Boltzmann–Shannon entropy $\phi(x)=\sum_{i=1}^{n}x_{i}\log x_{i}$ with $00\log 0=$ 0 and $\operatorname{dom}\phi=\mathbb{R}_{+}^{n}$.
-Then, $D_{\phi}(x,y)=\sum_{i=1}^{n}(x_{i}\log\frac{x_{i}}{y_{i}}-x_{i}+y_{i})$ is called the Kullback–Leibler divergence [Kullback1951-yv].
+    The Boltzmann–Shannon entropy $\phi(x)=\sum_{i=1}^{n}x_{i}\log x_{i}$ with $0\log 0=0$ and $\operatorname{dom}\phi=\mathbb{R}_{+}^{n}$.
+Then, $D_{\phi}(x,y)=\sum_{i=1}^{n}(x_{i}\log\frac{x_{i}}{y_{i}}-x_{i}+y_{i})$ is called the Kullback–Leibler (KL) divergence (Kullback and Leibler, 1951).
   3. (iii)
     The Burg entropy $\phi(x)=-\sum_{i=1}^{n}\log x_{i}$ and $\operatorname{dom}\phi=\mathbb{R}_{++}^{n}$.
-Then, $D_{\phi}(x,y)=\sum_{i=1}^{n}(\frac{x_{i}}{y_{i}}-\log\frac{x_{i}}{y_{i}}-1)$ is called the Itakura–Saito divergence [Itakura1968-en].
+Then, $D_{\phi}(x,y)=\sum_{i=1}^{n}(\frac{x_{i}}{y_{i}}-\log\frac{x_{i}}{y_{i}}-1)$ is called the Itakura–Saito divergence (Itakura and Saito, 1968).
   4. (iv)
     Let $\phi(x)=\frac{1}{4}\|x\|^{4}+\frac{1}{2}\|x\|^{2}$ and $\operatorname{dom}\phi=\mathbb{R}^{n}$.
-The Bregman distance $D_{\phi}$ is used in phase retrieval [Bolte2018-zt], low-rank minimization [Dragomir2021-rv], NMF [Mukkamala2019-mk], and blind deconvolution [Takahashi2023-uh].
-We recall the triangle scaling property for Bregman distances from Hanzely et al. [Hanzely2021-sz, Section 2], where several properties of the triangle scaling property are also shown.
-Definition A.2 (Triangle Scaling Property [Hanzely2021-sz, Definition 2]).
+The Bregman distance $D_{\phi}$ is used in phase retrieval (Bolte et al., 2018), low-rank minimization (Dragomir et al., 2021), NMF (Mukkamala and Ochs, 2019), and blind deconvolution (Takahashi et al., 2023).
+We recall the triangle scaling property for Bregman distances from (Hanzely et al., 2021, Section 2), where several properties of the triangle scaling property are also shown.
+Definition B.2 (Triangle Scaling Property (Hanzely et al., 2021, Definition 2)).
 Given a kernel generating distance $\phi\in\mathcal{G}(C)$, the Bregman distance $D_{\phi}$ has the triangle scaling property if there exists a constant $\nu>0$ such that, for all $x,y,z\in\operatorname{int}\operatorname{dom}\phi$ and all $\gamma\in[0,1]$, it holds that
-$D_{\phi}((1-\gamma)x+\gamma y,(1-\gamma)x+\gamma z)\leq\gamma^{\nu}D_{\phi}(y,z).$ (A.1)
-Now, substituting $z\leftarrow x$ on the left-hand side of (A.1), we obtain
+$D_{\phi}((1-\gamma)x+\gamma y,(1-\gamma)x+\gamma z)\leq\gamma^{\nu}D_{\phi}(y,z).$ (B.1)
+Now, substituting $z\leftarrow x$ on the left-hand side of Equation B.1, we obtain
 $D_{\phi}((1-\gamma)x+\gamma y,x)$ $=\phi((1-\gamma)x+\gamma y)-\phi(x)-\gamma\langle\nabla\phi(x),y-x\rangle$
@@
 Therefore, there exists $\nu>0$ such that $D_{\phi}((1-\gamma)x+\gamma y,x)\leq\gamma^{\nu}D_{\phi}(y,x)$ holds for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and all $\gamma\in[0,1]$.
-We will now show that a weaker version of the triangle scaling property can be obtained if $\phi$ is strictly convex, so that we can rephrase $\nu$ with $1+\nu$ where $\nu>0$, i.e., the right-hand side is superlinear.
-Lemma A.3.
-Given a kernel generating distance $\phi\in\mathcal{G}(C)$, if $\phi$ is strictly convex, then for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and all $\gamma\in[0,1]$, there exists $\nu>0$ such that, it holds that
-$D_{\phi}((1-\gamma)x+\gamma y,x)\leq\gamma^{1+\nu}D_{\phi}(y,x).$ (A.2)
+We will now show that a stronger version can be obtained if $\phi$ is strictly convex, so that we can rephrase $\nu$ with $1+\nu$ where $\nu>0$, i.e., the right-hand side is superlinear, which will be crucial for the convergence analysis later.
+Lemma B.3.
+Given a kernel generating distance $\phi\in\mathcal{G}(C)$, if $\phi$ is strictly convex, then there exists $\nu>0$ such that, for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and all $\gamma\in[0,1]$, it holds that
+$D_{\phi}((1-\gamma)x+\gamma y,x)\leq\gamma^{1+\nu}D_{\phi}(y,x).$ (B.2)
 Proof.
-Obviously, (A.2) holds if $y=x$ or $\gamma\in\{0,1\}$.
-In what follows, we assume $y\neq x$ and $10<\gamma<$ 1.
-Let $g(\nu):=\gamma^{1+\nu}D_{\phi}(y,x)-D_{\phi}((1-\gamma)x+\gamma y,x)$ for all $x,y\in\operatorname{int}\operatorname{dom}\phi$, $y\neq x$ and all $\gamma\in(0,1)$.
+Equation B.2 holds if $y=x$ or $\gamma\in\{0,1\}$.
+In what follows, we assume $y\neq x$ and $0<\gamma<1$.
+Let $g(\nu):=\gamma^{1+\nu}D_{\phi}(y,x)-D_{\phi}((1-\gamma)x+\gamma y,x)$ for all $\in\operatorname{int}\operatorname{dom}\phi, yxy\neq x$ and all $\gamma\in(0,1)$.
 Using $D_{\phi}(y,x)>0$ due to the strict convexity of $\phi$, we have, for any $\nu\geq 0$,
@@
 ∎
-Because $\phi$ is quadratic when $D_{\phi}$ is symmetric [Bauschke2001-ry, Lemma 3.16], it always holds that $D_{\phi}((1-\gamma)x+\gamma y,x)=\gamma^{1+\nu}D_{\phi}(y,x)$ with $\nu=1$.
-A.2 Lemmas of Smooth Adaptable Property
+Because $\phi$ is quadratic when $D_{\phi}$ is symmetric (Bauschke and Borwein, 2001, Lemma 3.16), it always holds that $D_{\phi}((1-\gamma)x+\gamma y,x)=\gamma^{1+\nu}D_{\phi}(y,x)$ with $\nu=1$.
+B.2 Properties of Smooth Adaptable Property
 The convexity of $L\phi-f$ and $L\phi+f$ plays a central role in developing and analyzing algorithms, and the smooth adaptable property implies the extended descent lemma.
-Lemma A.4 (Extended Descent Lemma [Bolte2018-zt, Lemma 2.1]).
+Lemma B.4 (Extended Descent Lemma (Bolte et al., 2018, Lemma 2.1)).
 The pair of functions $(f,\phi)$ is $L$-smad on $C$ if and only if for all $x,y\in\operatorname{int}\operatorname{dom}\phi$,
@@
 From this, it can be seen that the $L$-smad property for $(f,\phi)$ provides upper and lower approximations for $f$ majorized by $\phi$ with $L>0$.
-In addition, if $\phi(x)=\frac{1}{2}\|x\|^{2}$ on $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$, Lemma A.4 corresponds to the classical descent lemma.
-While the $L$-smad property might seem unfamiliar at first, it is a natural generalization of $L$-smoothness, and examples of functions $f$ and $\phi$ satisfying the $L$-smad property are given, e.g., in [Bauschke2017-hg, Lemmas 7 and 8][Bolte2018-zt, Lemma 5.1], [Dragomir2021-rv, Propositions 2.1 and 2.3], [Mukkamala2019-mk, Proposition 2.1], [Takahashi2024-wz, Proposition 24], [Takahashi2023-uh, Theorem 1], and [Takahashi2024-cu, Theorem 15].
+In addition, if $\phi(x)=\frac{1}{2}\|x\|^{2}$ on $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$, Lemma B.4 corresponds to the classical descent lemma.
+While the $L$-smad property might seem unfamiliar at first, it is a natural generalization of $L$-smoothness, and examples of functions $f$ and $\phi$ satisfying the $L$-smad property are given, e.g., in (Bauschke et al., 2017, Lemmas 7 and 8)(Bolte et al., 2018, Lemma 5.1), (Dragomir et al., 2021, Propositions 2.1 and 2.3), (Mukkamala and Ochs, 2019, Proposition 2.1), (Takahashi and Takeda, 2025, Proposition 24), (Takahashi et al., 2023, Theorem 1), and (Takahashi et al., 2026, Theorem 15).
+Example B.5.
+Other examples can be found in Sections 6.16.2, Appendices F.1F.4, and F.5.
+  1. (i)
+    Blind deconvolution: Let $f:\mathbb{R}^{d_{1}}\times\mathbb{R}^{d_{2}}\to\mathbb{R}$ be $f(h,x)=\frac{1}{4}\|Bh\odot Ax-y\|^{2}$, where $\odot$ denotes the Hadamard product (also known as the element-wise product).
+The function $f$ can be decomposed into a DC function $f(h,x)=f_{1}(h,x)-f_{2}(h,x)$ with $f_{1}(h,x)=\frac{1}{4}\|Bh\|_{4}^{4}+\frac{1}{4}\|Ax\|_{4}^{4}+\frac{1}{2}\left(\|Bh\odot Ax\|^{2}+\|y\odot Bh\|^{2}+\|Ax\|^{2}+\|y\|^{2}\right)$ and $f_{2}(h,x)=\frac{1}{4}\|Bh\|_{4}^{4}+\frac{1}{4}\|Ax\|_{4}^{4}+\frac{1}{2}\|y\odot Bh+Ax\|^{2}$, where $B\in\mathbb{R}^{m\times d_{1}}$, $A\in\mathbb{R}^{m\times d_{2}}$, and $y\in\mathbb{R}^{m}$.
+Let $\phi(h,x)=\frac{1}{4}(\|h\|^{2}+\|x\|^{2})^{2}+\frac{1}{2}(\|h\|^{2}+\|x\|^{2})$ For any $L$ satisfying
+    $L\geq\sum_{j=1}^{m}\left(3\|b_{j}\|^{4}+3\|a_{j}\|^{4}+\|b_{j}\|\|a_{j}\|+|y_{j}|\|b_{j}\|^{2}+\|a_{j}\|^{2}\right),$
+    where $b_{j}$ and $a_{j}$ are the $j$th row vectors of $B$ and $A$, respectively, $(f_{1},\phi)$ is $L$-smad (Takahashi et al., 2023, Theorem 1).
+Note that Bregman proximal DC algorithms require the smooth adaptable property of $(f_{1},\phi)$, instead of $(f,\phi)$.
+  2. (ii)
+    KL-NMF: Let $f:\mathbb{R}_{+}^{m\times r}\times\mathbb{R}_{+}^{r\times n}\to\mathbb{R}$ be $f(W,H)=\sum_{i=1}^{m}\sum_{j=1}^{n}(X_{ij}\log\frac{X_{ij}}{(WH)_{ij}}-X_{ij}+(WH)_{ij})$, where $X\in\mathbb{R}_{+}^{m\times n}$.
+$f$ is bounded by $\hat{f}(W,H)=\sum_{i,j}(X_{ij}\log X_{ij}-X_{ij}\sum_{l=1}^{r}\alpha_{ilj}\log\frac{W_{il}H_{lj}}{\alpha_{ilj}}-X_{ij}+(WH)_{ij})$ with $\alpha_{ilj}\in[0,1]$, $i=1,\ldots,m$, $l=1,\ldots,r$, $j=1,\ldots,n$.
+Let $\phi(W,H)=\sum_{i,l}(-\log W_{il}+\frac{1}{2}W_{il}^{2})+\sum_{l,j}(-\log H_{lj}+\frac{1}{2}H_{lj}^{2})$.
+Then, for any $L_{k}>0$ satisfying
+    $L\geq\max\left\{\max_{i,l}\left\{\sum_{j=1}^{n}\alpha_{ilj}X_{ij}\right\},\max_{l,j}\left\{\sum_{i=1}^{m}\alpha_{ilj}X_{ij}\right\},m,n\right\},$
+    the pair $(\hat{f},\phi)$ is $L$-smad on $\mathbb{R}_{++}^{m\times r}\times\mathbb{R}_{++}^{r\times n}$ (Takahashi et al., 2026, Theorem 4.1).
+Note that the majorization-minimization BPGs (Takahashi et al., 2026) require the smooth adaptable property of $(\hat{f},\phi)$, instead of $(f,\phi)$.
+Remark B.6.
+In general, it is difficult to choose the best $\phi$ such that $(f,\phi)$ is $L$-smad.
+On the other hand, it is known how to choose a suitable $\phi$ empirically.
+An easy way to find $\phi$ is to use elementary functions similar to $f$.
+See also the above examples, Section 6, and Appendix F.
 The extended descent lemma immediately implies primal progress of FW algorithms under the $L$-smad property as a straightforward generalization of the $L$-smooth case.
-Lemma A.5 (Primal progress from the smooth adaptable property).
+Lemma B.7 (Primal progress from the smooth adaptable property).
 Let the pair of functions $(f,\phi)$ be $L$-smad with a strictly convex function $\phi\in\mathcal{G}(C)$ and let $x_{+}=(1-\gamma)x+\gamma v$ with $x,v\in\operatorname{int}\operatorname{dom}\phi$ and $\gamma\in[0,1]$.
@@
 Proof.
-Using Lemma A.4 and substituting $x_{+}$ for $x$ and $x$ for $y$, we have
+Using Lemma B.4 and substituting $x_{+}$ for $x$ and $x$ for $y$, we have
 $f(x)-f(x_{+})$ $\geq\gamma\langle\nabla f(x),x-v\rangle-LD_{\phi}(x_{+},x).$
-It holds that $D_{\phi}(x_{+},x)\leq\gamma^{1+\nu}D_{\phi}(v,x)$ for some $\nu>0$ by Lemma A.3. Therefore, this provides the desired inequality. ∎
+It holds that $D_{\phi}(x_{+},x)\leq\gamma^{1+\nu}D_{\phi}(v,x)$ for some $\nu>0$ by Lemma B.3. Therefore, this provides the desired inequality. ∎
 We also have a progress lemma from the Bregman short step-size.
-Lemma A.6 (Progress lemma from the Bregman short step-size).
+Lemma B.8 (Progress lemma from the Bregman short step-size).
 Suppose that Assumption 3.1 holds. Define $x_{t+1}=(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$. Then if $x_{t+1}\in\operatorname{dom}f$,
@@
 Proof.
-Substituting $x_{t+1}=(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$ with $x_{t},v_{t}\in P$ for $x_{+}$ in Lemma 2.4, we have
+Substituting $x_{t+1}=(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$ with $x_{t},v_{t}\in P$ for $x_{+}$ in Lemma B.7, we have
 $f(x_{t})-f(x_{t+1})$ $\geq\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle-L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t})$
@@
 where the second inequality holds because of an upper bound of $\gamma_{t}$. ∎
-Appendix B Properties from Sharpness
-B.1 Results from Weak Convexity
-Obviously, $f$ is $\rho$-weakly convex and differentiable if and only if, for all $x,y\in\operatorname{dom}f$,
+B.3 Proof of Theorem 3.3
+Proof of Theorem 3.3.
+By following the same argument as (Pedregosa et al., 2020, Theorem 1), we have a bound on estimating $L$ as $(1-\log\eta/\log\tau)(t+1)+\max\{\log(\tau L/L_{-1}),0\}/\log\tau$.
+Let $n_{i}$ be the number of evaluations needed to estimate $\nu$.
+We have $\nu=\beta^{n_{i}-1}$, which provides $n_{i}=1+\log\nu/\log\beta$.
+Therefore, we obtain $n_{t}\leq\max\{(1-\log\eta/\log\tau)(t+1)+\max\{\log(\tau L/L_{-1}),0\}/\log\tau,(1+\log\nu/\log\beta)(t+1)\}$
+Appendix C Properties from Sharpness
+C.1 Results from Weak Convexity
+A function $f$ is $\rho$-weakly convex and differentiable if and only if, for all $x,y\in\operatorname{dom}f$,
 $f(y)-f(x)\geq\langle\nabla f(x),y-x\rangle-\frac{\rho}{2}\|y-x\|^{2}.$
-Any $\mathcal{C}^{2}$ function defined on a compact set is weakly convex [Vial1983-oo, Proposition 4.11] and [Wu2007-dt].
+Any $\mathcal{C}^{2}$ function defined on a compact set is weakly convex (Vial, 1983, Proposition 4.11) and Wu (2007).
 Since the proof of this fact is omitted therein, we provide the argument for the sake of completeness below.
-Proposition B.1.
+Proposition C.1.
 Let $P\subset\mathbb{R}^{n}$ be a compact set and $f$ be a proper and $\mathcal{C}^{2}$ function on $P$. Then, $f$ is weakly convex on $P$.
@@
 Let $g=\lambda_{\min}(\nabla^{2}f(\cdot))$, where $\lambda_{\min}(\cdot)$ is the smallest eigenvalue.
-Obviously, $g$ is continuous because $\lambda_{\min}(\cdot)$ and $\nabla^{2}f(\cdot)$ are continuous.
+The function $g$ is continuous because $\lambda_{\min}(\cdot)$ and $\nabla^{2}f(\cdot)$ are continuous.
 Therefore, there exists the minimum value of $g$ on $P$ due to the continuity of $g$ and the compactness of $P$.
@@
 Minimizing a nonconvex $\mathcal{C}^{2}$-function over a compact set $P$ is equivalent to minimizing a weakly convex function over $P$.
-However, it is important to note that when $\rho$ becomes large, the assumptions of algorithms may not be satisfied (see also Theorems D.3 and 5.3).
+However, it is important to note that when $\rho$ becomes large, the assumptions of algorithms may not be satisfied (see also Theorems E.3 and 5.4).
 We have the key lemma for convergence analysis from weak convexity.
-Lemma B.2 (Primal gap, dual gap, and Frank–Wolfe gap for weakly convex functions).
-Let $x^{*}\in P$ an optimal solution of (1.1) and let $f^{*}=f(x^{*})$. Suppose that Assumptions 3.1iiiv, and that $f$ is $\rho$-weakly convex. Then for all $x\in P$, it holds:
-$f(x)-f^{*}\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}\leq\max_{v\in P}\langle\nabla f(x),x-v\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}.$ (B.1)
+Lemma C.2 (Primal gap, dual gap, and Frank–Wolfe gap for weakly convex $f$).
+Suppose that Assumptions 3.1(ii), (iv), and that $f$ is $\rho$-weakly convex. Then for all $x\in P$, it holds:
+$f(x)-f^{*}\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}\leq\max_{v\in P}\langle\nabla f(x),x-v\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}.$ (C.1)
 Proof.
 The first inequality follows from the weak convexity of $f$ and the second follows from the maximality of $\langle\nabla f(x),x-v\rangle$. ∎
-B.2 Primal Gap Bound
-We show in the following lemma that the HEB condition immediately provides a bound on the primal optimality gap, which will be useful for establishing convergence rates of the proposed algorithms as well as the Łojasiewicz inequality [Bolte2007-xr, Lojasiewicz1963-sm, Lojasiewicz1965-ks], provided $f$ is convex.
-Lemma B.3 (Primal gap bound from Hölder error bound).
+C.2 Primal Gap Bound
+We show in the following lemma that the HEB condition immediately provides a bound on the primal optimality gap, which will be useful for establishing convergence rates of the proposed algorithms as well as the Łojasiewicz inequality (Bolte et al., 2007; Łojasiewicz, 1963; 1965), provided $f$ is convex.
+Lemma C.3 (Primal gap bound from Hölder error bound).
 Let $f$ be a convex function and satisfy the HEB condition with $q\geq 1$ and $\mu>0$.
@@
 Then, the following argument holds in general, for all $x\in P$:
-$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{\frac{q}{q-1}},$ (B.2)
+$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{\frac{q}{q-1}},$ (C.2)
 or equivalently,
@@
 $f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{\frac{q}{q-1}}.$
-Remark B.4.
-With the remark from above, we can immediately relate the HEB condition to the Łojasiewicz inequality [Bolte2007-xr, Lojasiewicz1963-sm, Lojasiewicz1965-ks]; the Łojasiewicz inequality is used to establish convergence analysis [Bolte2007-xr].
-To this end, using (B.2), we estimate $\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\leq\frac{\|\nabla f(x)\|\|x-x^{*}\|}{\|x-x^{*}\|}=\|\nabla f(x)\|$ to then obtain the weaker condition:
-$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\|\nabla f(x)\|^{\frac{q}{q-1}}=c^{\frac{1}{\theta}}\|\nabla f(x)\|^{\frac{1}{\theta}},$ (B.3)
+Remark C.4.
+With the remark from above, we can immediately relate the HEB condition to the Łojasiewicz inequality (Bolte et al., 2007; Łojasiewicz, 1963; 1965); the Łojasiewicz inequality is used to establish convergence analysis (Bolte et al., 2007).
+To this end, using Equation C.2, we estimate $\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\leq\frac{\|\nabla f(x)\|\|x-x^{*}\|}{\|x-x^{*}\|}=\|\nabla f(x)\|$ to then obtain the weaker condition:
+$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\|\nabla f(x)\|^{\frac{q}{q-1}}=c^{\frac{1}{\theta}}\|\nabla f(x)\|^{\frac{1}{\theta}},$ (C.3)
 where $c=\left(\frac{q}{\mu}\right)^{\frac{1}{q}}$ and $\theta=\frac{q-1}{q}$.
-Inequality (B.3) is called the $c$-Łojasiewicz inequality with $\theta\in[0,1)$.
+Inequality Equation C.3 is called the cc-Łojasiewicz inequality with $\theta\in[0,1)$.
 If $\mathcal{X}^{*}\subseteq\operatorname{ri}P$, then the two conditions are equivalent.
@@
 The next lemma shows that we can also obtain primal gap bounds in the weakly convex case together with (local) quadratic growth.
-Lemma B.5 (Primal gap bound from the quadratic growth).
-Let $f$ be a $\rho$-weakly convex function that satisfies the local $\mu$-quadratic growth condition such that $\rho\leq\mu$.
+Lemma C.5 (Primal gap bound from the quadratic growth).
+Let $f$ be a $\rho$-weakly convex function that satisfies the local $\mu$-quadratic growth condition such that $\rho<\mu$.
 Let $x^{*}$ be the unique minimizer of $f$ over $P$ and let $\zeta>0$.
 Then, the following holds: for all $x\in[f\leq f^{*}+\zeta]\cap P$,
-$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{2},$ (B.4)
+$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{2},$ (C.4)
 or equivalently,
@@
 which implies
-$\left(1-\frac{\rho}{\mu}\right)\left(f(x)-f^{*}\right)\leq\langle\nabla f(x),x-x^{*}\rangle=\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\|x-x^{*}\|.$ (B.5)
+$\left(1-\frac{\rho}{\mu}\right)\left(f(x)-f^{*}\right)\leq\langle\nabla f(x),x-x^{*}\rangle=\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\|x-x^{*}\|.$ (C.5)
 Using the local quadratic growth condition again, we have
@@
 $f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{2}.$
-Remark B.6.
-In the same vein as the discussion in Remark B.4, we can immediately relate the local $\mu$-quadratic growth condition with $\zeta>0$ to the local Polyak–Łojasiewicz (PL) inequality.
-Using (B.4), for all $x\in[f\leq f^{*}+\zeta]\cap P$, we have
-$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\|\nabla f(x)\|^{2},$ (B.6)
-which is equivalent to the PL inequality [Polyak1963-dr], also called the gradient dominance property [Combettes2020-jj], with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$.
+Remark C.6.
+In the same vein as the discussion in Remark C.4, we can immediately relate the local $\mu$-quadratic growth condition with $\zeta>0$ to the local Polyak–Łojasiewicz (PL) inequality.
+Using Equation C.4 in Lemma C.5, for all $x\in[f\leq f^{*}+\zeta]\cap P$, we have
+$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\|\nabla f(x)\|^{2},$ (C.6)
+which is equivalent to the PL inequality (Polyak, 1963), also called the gradient dominance property (Combettes and Pokutta, 2020), with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$.
 The PL inequality is also equivalent to the Łojasiewicz inequality with $\theta=\frac{1}{2}$.
-Strongly convex functions satisfy the PL inequality [Braun2022-ew, Lemma 2.13].
-Note that we will not have the primal gap bound under the HEB condition and weak convexity because an inequality like (B.5) does not hold from the HEB condition.
-B.3 Scaling Inequalities, Geometric Hölder Error Bounds, and Contractions
+Strongly convex functions satisfy the PL inequality (Braun et al., 2025, Lemma 2.13).
+Note that we will not have the primal gap bound under the HEB condition and weak convexity because an inequality like Equation C.5 does not follow from the HEB condition.
+C.3 Scaling Inequalities, Geometric Hölder Error Bounds, and Contractions
 In the following, we will now bring things together to derive tools that will be helpful in establishing convergence rates.
@@
 The first scaling inequality is useful for analyzing the case where the optimal solution lies in the relative interior of the feasible region.
-While its formulation in [Braun2022-ew, Proposition 2.16] required $L$-smoothness of $f$, it is actually not used in the proof, and the results hold more broadly.
+While its formulation in (Braun et al., 2025, Proposition 2.16) required $L$-smoothness of $f$, it is actually not used in the proof, and the results hold more broadly.
 We restate it here for the sake of completeness.
-Proposition B.7 (Scaling inequality from convexity when $x^{*}\in\operatorname{int}P$ [Braun2022-ew, Proposition 2.16]).
+Proposition C.7 (Scaling inequality from convexity when $x^{*}\in\operatorname{int}P$ (Braun et al., 2025, Proposition 2.16)).
 Let $P\subset\mathbb{R}^{n}$ be a nonempty compact convex set.
@@
 When $f$ is not convex, assuming that $f$ is weakly convex and local quadratic growth, we have the scaling inequality for a nonconvex objective function.
-Proposition B.8 (Scaling inequality from weak convexity when $x^{*}\in\operatorname{int}P$).
+Proposition C.8 (Scaling inequality from weak convexity when $x^{*}\in\operatorname{int}P$).
 Let $P\subset\mathbb{R}^{n}$ be a nonempty compact convex set.
@@
 We consider $x^{*}-rz$, where $z$ is a point with $\|z\|=1$ and $\langle\nabla f(x),z\rangle=\|\nabla f(x)\|$. It holds that
-$\langle\nabla f(x),v\rangle\leq\langle\nabla f(x),x^{*}-rz\rangle=\langle\nabla f(x),x^{*}\rangle-r\|\nabla f(x)\|.$ (B.7)
+$\langle\nabla f(x),v\rangle\leq\langle\nabla f(x),x^{*}-rz\rangle=\langle\nabla f(x),x^{*}\rangle-r\|\nabla f(x)\|.$ (C.7)
 From weak convexity, for all $x\in[f\leq f^{*}+\zeta]\cap P$, we have
@@
 $\langle\nabla f(x),x-x^{*}\rangle\geq\left(1-\frac{\rho}{\mu}\right)(f(x)-f^{*})\geq 0.$
-By rearranging (B.7) and using $\langle\nabla f(x),x-x^{*}\rangle\geq 0$, we obtain
+By rearranging Equation C.7 and using $\langle\nabla f(x),x-x^{*}\rangle\geq 0$, we obtain
 $\langle\nabla f(x),x-v\rangle\geq\langle\nabla f(x),x-x^{*}\rangle+r\|\nabla f(x)\|\geq r\|\nabla f(x)\|.$
@@
 where the first inequality holds because of the Cauchy–Schwarz inequality. ∎
-The authors in [Lacoste-Julien2015-gb] defined a geometric distance-like constant of a polytope, known as the pyramidal width, to analyze the convergence of the away-step Frank–Wolfe algorithm (and other variants that use the pyramidal width) over polytopes.
-It can be interpreted as the minimal $\delta>0$ satisfying the following scaling inequality (B.8), which plays a central role in establishing convergence rates for the away-step FW algorithm; see [Braun2022-ew] for an in-depth discussion.
-Lemma B.9 (Scaling inequality via pyramidal width [Braun2022-ew, Theorem 2.26][Lacoste-Julien2015-gb, Theorem 3]).
+Lacoste-Julien and Jaggi (2015) defined a geometric distance-like constant of a polytope, known as the pyramidal width, to analyze the convergence of the away-step Frank–Wolfe algorithm (and other variants that use the pyramidal width) over polytopes.
+It can be interpreted as the minimal $\delta>0$ satisfying the following scaling inequality Equation C.8, which plays a central role in establishing convergence rates for the away-step FW algorithm; see Braun et al.
+(2025) for an in-depth discussion.
+Lemma C.9 (Scaling inequality via pyramidal width (Braun et al., 2025, Theorem 2.26)(Lacoste-Julien and Jaggi, 2015, Theorem 3)).
 Let $P\subset\mathbb{R}^{n}$ be a polytope and let $\delta$ denote the pyramidal width of $P$.
@@
 Then for any $y\in P$
-$\langle\psi,v^{\textup{A}}-v^{\textup{FW}}\rangle\geq\delta\frac{\langle\psi,x-y\rangle}{\|x-y\|}.$ (B.8)
-Note that Lemma B.9 does not require the convexity of $f$, and we will use it for nonconvex optimization. When $\phi$ is $\sigma$-strongly convex and $\langle\psi,x-y\rangle\geq 0$, Lemma B.9 implies
+$\langle\psi,v^{\textup{A}}-v^{\textup{FW}}\rangle\geq\delta\frac{\langle\psi,x-y\rangle}{\|x-y\|}.$ (C.8)
+Note that Lemma C.9 does not require the convexity of $f$, and we will use it for nonconvex optimization. When $\phi$ is $\sigma$-strongly convex and $\langle\psi,x-y\rangle\geq 0$, Lemma C.9 implies
 $\langle\psi,v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}\geq\delta^{2}\frac{\langle\psi,x-y\rangle^{2}}{\|x-y\|^{2}}\geq\delta^{2}\sigma\frac{\langle\psi,x-y\rangle^{2}}{2D_{\phi}(x,y)}.$
@@
 Geometric Hölder Error Bound Condition
-We will now introduce the more compact notion of geometric Hölder error bound condition, which simply combines the pyramidal width and the HEB condition of the function $f$; see also [Braun2022-ew, Lemma 2.27] for details when $q=2$.
-Lemma B.10 (Geometric Hölder Error Bound).
+We will now introduce the more compact notion of geometric Hölder error bound condition, which simply combines the pyramidal width and the HEB condition of the function $f$; see also (Braun et al., 2025, Lemma 2.27) for details when $q=2$.
+Lemma C.10 (Geometric Hölder error bound).
 Let $P$ be a polytope with pyramidal width $\delta>0$ and let $f$ be a convex function and satisfy the HEB condition with $v^{\textup{FW}}=\operatorname*{arg\,min}_{v\in P}\langle\nabla f(x),v\rangle$ and $v^{\textup{A}}=\operatorname*{arg\,max}_{v\in\mathcal{S}}\langle\nabla f(x),v\rangle$ with $\mathcal{S}\subseteq\operatorname{Vert}{P}$, so that $x\in\operatorname{conv}{\mathcal{S}}$, we have
@@
 Proof.
-Combining (B.8) with Lemma B.3 for $\psi=\nabla f(x)$, we have
+Combining Equation C.8 with Lemma C.3 for $\psi=\nabla f(x)$, we have
 $f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{\frac{q}{q-1}}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle}{\delta}\right)^{\frac{q}{q-1}}.$
@@
 Besides scaling inequalities and the geometric Hölder error bound, we also utilize the following lemma for convex and nonconvex optimization, which allows us to turn a contraction into a convergence rate.
-Lemma B.11 (From contractions to convergence rates [Braun2022-ew, Lemma 2.21]).
+Lemma C.11 (From contractions to convergence rates (Braun et al., 2025, Lemma 2.21)).
 Let $\{h_{t}\}_{t}$ be a decreasing sequence of positive numbers and $c_{0}$, $c_{1}$, $c_{2}$, $\theta_{0}$ be positive numbers with $c_{1}<1$ such that $h_{1}\leq c_{0}$ and $h_{t}-h_{t+1}\geq h_{t}\min\{c_{1},c_{2}h_{t}^{\theta_{0}}\}$ for $t\geq 1$, then
@@
 In particular, we have $h_{t}\leq\epsilon$ if $t\geq t_{0}+\frac{1}{\theta_{0}c_{2}\epsilon^{\theta_{0}}}-\frac{1}{\theta_{0}c_{1}}$ and $\epsilon\leq(c_{1}/c_{2})^{1/\theta_{0}}$.
-Appendix C Proof of Convergence Analysis for Convex Optimization
+Appendix D Proof of Convergence Analysis for Convex Optimization
 We would also like to stress that while we formulate some of the results for the case that $x^{*}\in\operatorname{int}P$, the results can be extended to the case that $x^{*}\in\operatorname{ri}P$, i.e., the relative interior of $P$, which we did not do for the sake of clarity.
-Basically, the analysis is performed in the affine space spanned by the optimal face of $P$ in that case; the interested reader is referred to [Braun2022-ew] for details on how to extend the results.
-C.1 Sublinear Convergence
+Basically, the analysis is performed in the affine space spanned by the optimal face of $P$ in that case; the interested reader is referred to Braun et al.
+(2025) for details on how to extend the results.
+D.1 Sublinear Convergence
+We recall a key property for analyzing convergence rates in convex optimization.
+Lemma D.1 (Primal gap, dual gap, and Frank–Wolfe gap for convex $f$ (Pokutta, 2024, Lemma 4.1)).
+Suppose that Assumptions 3.1(ii), (iv), and that $f$ is convex. Then for all $x\in P$, it holds:
+$f(x)-f^{*}\leq\langle\nabla f(x),x-x^{*}\rangle\leq\max_{v\in P}\langle\nabla f(x),x-v\rangle.$ (D.1)
 We establish a sublinear convergence rate of the FW algorithm under the smooth adaptable property.
-It is a similar result to [Vyguzov2025-ha, Theorem 1].
-We conducted its proof following [Braun2022-ew, Jaggi2013-oz, Pokutta2024-qu].
+It is a similar result to (Vyguzov and Stonyakin, 2025, Theorem 1).
+We conducted its proof following Braun et al.
+(2025); Jaggi (2013); Pokutta (2024).
 The FW algorithm uses the open loop step-size, i.e., $\gamma_{t}=\frac{2}{2+t}$ in the following theorem.
-Theorem C.1 (Primal convergence of the Frank–Wolfe algorithm).
+Theorem D.2 (Primal convergence of the Frank–Wolfe algorithm).
 Suppose that Assumptions 3.1 and 4.1 hold.
@@
 Then, it holds that, for all $t\geq 1$,
-$f(x_{t})-f^{*}\leq\frac{2^{1+\nu}LD^{2}}{(t+2)^{\nu}},$ (C.1)
+$f(x_{t})-f^{*}\leq\frac{2^{1+\nu}LD^{2}}{(t+2)^{\nu}},$ (D.2)
 and hence for any accuracy $\epsilon>0$ we have $f(x_{t})-f^{*}\leq\epsilon$ for all $t\geq(2^{1+\nu}LD^{2}/\epsilon)^{1/\nu}$.
@@
 $\geq\gamma_{t}(f(x_{t})-f^{*})-L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t}),$
-where the first inequality holds because of Lemma 2.4 and the last inequality holds because of Lemma 3.2.
+where the first inequality holds because of Lemma B.7 and the last inequality holds because of Lemma D.1.
 Subtracting $f^{*}$ on both sides, using $D_{\phi}(v_{t},x_{t})\leq D^{2}$, and rearranging leads to
@@
 where the last inequality holds due to $(3+t)^{\nu}(1+t)\leq(2+t)^{1+\nu}$ with $0<\nu\leq 1$. ∎
-If $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$ in (C.1), we have $f(x_{t})-f^{*}\leq 4LD^{2}/(t+2)$, which is the same as a sublinear convergence rate of the classical FW algorithm.
-C.2 Proof of Theorem 4.2
-Because Algorithm 2 is well-defined (see Remark 3.5), the convergence result of the FW algorithm using the adaptive step-size strategy (Algorithm 2) is essentially the same as the one that uses Bregman short steps (Theorems 4.2 and 5.2); up to small errors arising from the approximation whose precise analysis we skip for the sake of brevity.
-Theorem C.2 (Linear convergence FW algorithm with short step-size or adaptive step-size).
-Suppose that Assumptions 3.1 and 4.1 hold.
-Let $\nu>0$ and let $f$ satisfy the HEB condition with $q\geq 1+\nu$ and $\mu>0$ and $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$.
-Assume that there exists a minimizer $x^{*}\in\operatorname{int}P$, i.e., there exists an $r>0$ with $B(x^{*},r)\subset P$.
-Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu},1\right\}$ or Algorithm 2.
-Then, it holds that, for all $t\geq 1$,
-$f(x_{t})-f^{*}\leq\begin{cases}h_{t}\leq\max\left\{\frac{1}{1+\nu},1-\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{c^{1+1/\nu}D^{2/\nu}}\right\}^{t-1}LD^{2}&\text{if}\ q=1+\nu,\\ \frac{LD^{2}}{(1+\nu)^{t-1}}&\text{if}\ 1\leq t\leq t_{0},q>1+\nu,\\ \frac{(L(1+\nu)(c/r)^{1+\nu}D^{2})^{q/(q-1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}(t-t_{0})\right)^{\nu q/(q-1-\nu)}}=\mathcal{O}(1/t^{\nu q/(q-1-\nu)})&\text{if}\ t\geq t_{0},q>1+\nu,\end{cases}$
-where $c=(q/\mu)^{1/q}$ and
-$t_{0}:=\max\left\{\left\lfloor\log_{\frac{1}{1+\nu}}\left(\frac{(L(1+\nu(c/r)^{1+\nu}D^{2})^{q/(q-1-\nu)}}{LD^{2}}\right)\right\rfloor+2,1\right\}.$
-Proof.
-Let $h_{t}:=f(x_{t})-f^{*}$. Using Lemma 3.4, we have
+If $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$ in Equation D.2, we have $f(x_{t})-f^{*}\leq 4LD^{2}/(t+2)$, which is the same as a sublinear convergence rate of the classical FW algorithm.
+Remark D.3.
+While the convergence rate Equation D.2 is the same as Vyguzov and Stonyakin (2025), Vyguzov and Stonyakin assume that the triangle scaling property holds for $D_{\phi}$.
+In contrast, we require significantly weaker assumptions: it is enough to assume that $\phi$ is strictly convex due to Lemma B.3 in order to establish Theorem D.2.
+D.2 Proof of Theorem 4.2
+Because Algorithm 2 is well-defined (see Remark 3.2), the convergence result of the FW algorithm using the adaptive step-size strategy (Algorithm 2) is essentially the same as the one that uses Bregman short steps (Theorems 4.2 and 5.3); up to small errors arising from the approximation whose precise analysis we skip for the sake of brevity.
+Proof of Theorem 4.2.
+Let $h_{t}:=f(x_{t})-f^{*}$. Using Lemma B.8, we have
 $h_{t}-h_{t+1}=f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\gamma_{t},$
@@
 $\geq\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}h_{t}^{\frac{(1+\nu)(q-1)}{\nu q}},$
-where the second inequality holds from $\langle f(x_{t}),x_{t}-v_{t}\rangle\geq r\|\nabla f(x_{t})\|$ in Proposition B.7, and the last inequality holds because $f$ satisfies the $c$-Łojasiewicz inequality with $c=(q/\mu)^{1/q}$ (see Remark B.4), and $M:=(L(1+\nu))^{1/\nu}$.
-(ii) $\gamma_{t}=1$: Using (3.1), we have
+where the second inequality holds from $\langle f(x_{t}),x_{t}-v_{t}\rangle\geq r\|\nabla f(x_{t})\|$ in Proposition C.7, and the last inequality holds because $f$ satisfies the cc-Łojasiewicz inequality with $c=(q/\mu)^{1/q}$ (see Remark C.4), and $M:=(L(1+\nu))^{1/\nu}$.
+(ii) $\gamma_{t}=1$: Using Equation D.1, we have
 $h_{t}-h_{t+1}=f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq\frac{\nu}{1+\nu}(f(x_{t})-f^{*})=\frac{\nu}{1+\nu}h_{t},$
@@
 $h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu},\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}\right\}\cdot h_{t}.$
-This inequality and the initial bound $f(x_{1})-f^{*}\leq LD^{2}$ due to Lemma A.4 imply
+This inequality and the initial bound $f(x_{1})-f^{*}\leq LD^{2}$ due to Lemma B.4 imply
 $h_{t}\leq\max\left\{\frac{1}{1+\nu},1-\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}\right\}^{t-1}LD^{2}.$
@@
 $h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu},\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}h_{t}^{\frac{q-1-\nu}{\nu q}}\right\}\cdot h_{t}.$
-Using $f(x_{1})-f^{*}\leq LD^{2}$ and Lemma B.11 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}$, $c_{2}=\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}$, and $\theta_{0}=\frac{q-1-\nu}{\nu q}>0$ from $q>1+\nu$, we have the claim.
+Using $f(x_{1})-f^{*}\leq LD^{2}$ and Lemma C.11 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}$, $c_{2}=\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}$, and $\theta_{0}=\frac{q-1-\nu}{\nu q}>0$ from $q>1+\nu$, we have the claim.
 ∎
-C.3 Proof of Theorem 4.4
-Theorem C.3 (Linear convergence of the away-step FW algorithm).
-Suppose that Assumptions 3.14.1, and 4.3 hold.
-Let $\nu>0$ and let $P\subset\mathbb{R}^{n}$ be a polytope and $f$ satisfy the HEB condition with $q>1+\nu$ or $(\nu,q)=(1,2)$.
-The convergence rate of Algorithm 3 is linear: for all $t\geq 1$
-$f(x_{t})-f^{*}\leq\begin{cases}\left(1-\frac{\mu}{32L}\frac{\delta^{2}}{D^{2}}\right)^{\lceil(t-1)/2\rceil}LD^{2}&\text{if}\ (\nu,q)=(1,2),\\ \frac{1}{(1+\nu)^{\lceil(t-1)/2\rceil}}LD^{2}&\text{if}\ 1\leq t\leq t_{0},q>1+\nu,\\ \frac{(L(1+\nu)D^{2}c/(\delta/2)^{1+\nu})^{q/(q-1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}\lceil(t-t_{0})/2\rceil\right)^{\nu q/(q-1-\nu)}}=\mathcal{O}(1/t^{\nu q/(q-1-\nu)})&\text{if}\ t\geq t_{0},q>1+\nu,\end{cases}$
-where $c:=(q/\mu)^{1/q}$ $D$ and $\delta$ are the diameter and the pyramidal width of the polytope $P$, respectively, and
-$t_{0}:=\max\left\{\left\lfloor\log_{\frac{1}{1+\nu}}\frac{(L(1+\nu)D^{2}/((\mu/q)^{1/q}\delta/2)^{1+\nu})^{q/(q-1-\nu)}}{LD^{2}}\right\rfloor+2,1\right\}.$
-Proof.
-By using the induced guarantee on the primal gap via Lemma B.10, we have
-$h_{t}=f(x_{t})-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x_{t}),v_{t}^{\textup{A}}-v_{t}^{\textup{FW}}\rangle}{\delta}\right)^{\frac{q}{q-1}}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{2\langle\nabla f(x_{t}),d_{t}\rangle}{\delta}\right)^{\frac{q}{q-1}},$ (C.2)
+D.3 Proof of Theorem 4.4
+Proof of Theorem 4.4.
+By using the induced guarantee on the primal gap via Lemma C.10, we have
+$h_{t}=f(x_{t})-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x_{t}),v_{t}^{\textup{A}}-v_{t}^{\textup{FW}}\rangle}{\delta}\right)^{\frac{q}{q-1}}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{2\langle\nabla f(x_{t}),d_{t}\rangle}{\delta}\right)^{\frac{q}{q-1}},$ (D.3)
 where the last inequality holds because $d_{t}$ is either $x_{t}-v_{t}^{\textup{FW}}$ or $v_{t}^{\textup{A}}-x_{t}$ with $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle/2$ in Lines 3 and 3 of Algorithm 3.
@@
 $=\min\left\{\frac{\nu h_{t}}{1+\nu}\gamma_{t,\max},\frac{\nu}{1+\nu}\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}h_{t}^{\frac{(1+\nu)(q-1)}{\nu q}}\right\}$
-where the first inequality holds due to Lemma 3.4 with $\gamma_{t,\max}=1$ (Frank–Wolfe steps) and $\gamma_{t,\max}=\frac{\lambda_{v_{t}^{\textup{A}},t}}{1-\lambda_{v_{t}^{\textup{A}},t}}$ (away steps), and the second inequality holds because of (C.2) and $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq h_{t}$.
+where the first inequality holds due to Lemma B.8 with $\gamma_{t,\max}=1$ (Frank–Wolfe steps) and $\gamma_{t,\max}=\frac{\lambda_{v_{t}^{\textup{A}},t}}{1-\lambda_{v_{t}^{\textup{A}},t}}$ (away steps), and the second inequality holds because of Equation D.3 and $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq h_{t}$.
 In $(\nu,q)=(1,2)$, for Frank–Wolfe steps, $\gamma_{t,\max}=1\geq\mu\delta^{2}/LD^{2}\geq\mu\delta^{2}/32LD^{2}$.
@@
 In $q>1+\nu$, we have $h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu},\frac{\nu}{1+\nu}\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}h_{t}^{\frac{q-1-\nu}{\nu q}}\right\}\cdot h_{t}$ for at least half of the iterations (Frank–Wolfe steps) and $h_{t+1}\leq h_{t}$ for the rest (away steps).
-The initial bound $f(x_{1})-f^{*}\leq LD^{2}$ holds generally for the Frank–Wolfe algorithm (see the proof of Theorem C.1 and [Braun2022-ew, Remark 2.4]).
-We use Lemma B.11 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}$, $c_{2}=c_{1}\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}$, and $\theta_{0}=\frac{q-1-\nu}{\nu q}>0$ from $q>1+\nu$ and obtain the claim.
+The initial bound $f(x_{1})-f^{*}\leq LD^{2}$ holds generally for the Frank–Wolfe algorithm (see the proof of Theorem D.2 and (Braun et al., 2025, Remark 2.4)).
+We use Lemma C.11 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}$, $c_{2}=c_{1}\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}$, and $\theta_{0}=\frac{q-1-\nu}{\nu q}>0$ from $q>1+\nu$ and obtain the claim.
 ∎
-Remark C.4 (Compatibility of parameters).
+Remark D.4 (Compatibility of parameters).
 The condition $q>1+\nu$ is necessary in case that $(\nu,q)\neq(1,2)$. The reason is as follows. In the case $q=1+\nu$, for Frank–Wolfe steps, we have
-$\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}=\left(\frac{\mu}{L}\right)^{1/\nu}\left(\frac{\delta}{D}\right)^{\frac{1+\nu}{\nu}}\frac{1}{2^{\frac{1+\nu}{\nu}}(1+\nu)^{2/\nu}D^{(1-\nu)/\nu}},$ (C.3)
+$\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}=\left(\frac{\mu}{L}\right)^{1/\nu}\left(\frac{\delta}{D}\right)^{\frac{1+\nu}{\nu}}\frac{1}{2^{\frac{1+\nu}{\nu}}(1+\nu)^{2/\nu}D^{(1-\nu)/\nu}},$ (D.4)
 which might be greater than 1 because $\frac{1}{D^{(1-\nu)/\nu}}\geq 1$ if $D<1$ and $\nu$ are small enough.
-In order to make (C.3) smaller than 1, $\nu$ should be 1, i.e., $(\nu,q)=(1,2)$.
+In order to make Equation D.4 smaller than 1, $\nu$ should be 1, i.e., $(\nu,q)=(1,2)$.
 When $\nu=1$ and $q\neq 2$, it reduces to the $q>1+\nu$ case.
-Appendix D Proof of Convergence Analysis for Nonconvex Optimization
-D.1 Global Convergence
+Appendix E Proof of Convergence Analysis for Nonconvex Optimization
+E.1 Global Convergence
 We show that Algorithm 1 with $\gamma_{t}=\gamma:=1/(T+1)^{\frac{1}{1+\nu}}$ globally converges to a stationary point, where $T\in\mathbb{N}$ is the number of iterations.
-Its proof is inspired by [Lacoste-Julien2016-mj, Theorem 1] and [Pokutta2024-qu, Theorem 4.7] and identical to those for the case when $\phi=\frac{1}{2}\|\cdot\|^{2}$.
-Theorem D.1 (Global sublinear convergence for nonconvex optimization).
+Its proof is inspired by (Lacoste-Julien, 2016, Theorem 1) and (Pokutta, 2024, Theorem 4.7) and identical to those for the case when $\phi=\frac{1}{2}\|\cdot\|^{2}$.
+Theorem E.1 (Global sublinear convergence for nonconvex optimization).
 Suppose that Assumption 3.1 holds.
@@
 Proof.
-Substituting $x_{t+1}$ for $x_{+}$ and $x_{t}$ for $x$ in Lemma 2.4, we have the following inequality:
+Substituting $x_{t+1}$ for $x_{+}$ and $x_{t}$ for $x$ in Lemma B.7, we have the following inequality:
 $f(x_{t})-f(x_{t+1})$ $\geq\gamma\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle-L\gamma^{1+\nu}D_{\phi}(v_{t},x_{t}).$
@@
 $\min_{0\leq t\leq T}\max_{v_{t}\in P}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\leq\frac{2\max\{h_{0},LD^{2}\}}{\sqrt{T+1}},$
-which is the same rate as [Lacoste-Julien2016-mj, Theorem 1].
-D.2 Local Sublinear Convergence
+which is the same rate as (Lacoste-Julien, 2016, Theorem 1).
+E.2 Local Sublinear Convergence
 We prove a lemma used in the proof of convergence analysis.
-Lemma D.2.
+Lemma E.2.
 For any $\nu\in(0,1]$ and $t\geq 0$, it holds that
-$h(t):=\frac{(t+3)^{\nu}(t+2-\nu)}{(t+2)^{1+\nu}}\leq 1.$ (D.1)
+$h(t):=\frac{(t+3)^{\nu}(t+2-\nu)}{(t+2)^{1+\nu}}\leq 1.$ (E.1)
 Proof.
@@
 ∎
-Now we show sublinear convergence with $\gamma_{t}=\frac{2}{2+t}$. The proof is a modified version of Theorem C.1.
-Theorem D.3 (Local sublinear convergence).
+Now we show sublinear convergence with $\gamma_{t}=\frac{2}{2+t}$. The proof is a modified version of Theorem D.2.
+Theorem E.3 (Local sublinear convergence).
 Suppose that Assumptions 3.14.3, and 5.1 hold.
@@
 Then, if $\rho/\sigma\leq L$ and $\frac{3\rho}{\mu}\leq 2-\nu$ hold, it holds that, for all $t\geq 1$,
-$f(x_{t})-f^{*}\leq\frac{2^{1+\nu}\mu LD^{2}}{\rho(t+2)^{\nu}},$ (D.2)
+$f(x_{t})-f^{*}\leq\frac{2^{1+\nu}\mu LD^{2}}{\rho(t+2)^{\nu}},$ (E.2)
 and hence for any accuracy $\epsilon>0$ we have $f(x_{t})-f^{*}\leq\epsilon$ for all $t\geq(2^{1+\nu}\mu LD^{2}/\rho\epsilon)^{1/\nu}$.
 Proof.
-Using Lemma 2.4, we have
+Using Lemma B.7, we have
 $f(x_{t})-f(x_{t+1})$ $\geq\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle-L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t})$
 $\geq\gamma_{t}\left(f(x_{t})-f^{*}-\frac{\rho}{2}\|x_{t}-x^{*}\|^{2}\right)-L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t}),$
-where the last inequality holds because of (3.2) in Lemma 3.3. Subtracting $f^{*}$ and rearranging provides
-$f(x_{t+1})-f^{*}$ $\leq(1-\gamma_{t})(f(x_{t})-f^{*})+\frac{\rho\gamma_{t}}{2}\|x_{t}-x^{*}\|^{2}+L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t})$ (D.3)
-$\leq\left(1-\left(1-\frac{\rho}{\mu}\right)\gamma_{t}\right)(f(x_{t})-f^{*})+L\gamma_{t}^{1+\nu}D^{2},$ (D.4)
-where the last inequality holds due to the quadratic growth condition. For $t=0$, using (D.3), we have
-$f(x_{1})-f^{*}\leq\frac{\rho}{2}\|x_{0}-x^{*}\|^{2}+LD_{\phi}(v_{0},x_{0})\leq\left(\frac{\rho}{\sigma}+L\right)D^{2}\leq 2LD^{2}\leq\frac{2\mu}{\rho}LD^{2},$ (D.5)
+where the last inequality holds because of Equation C.1 in Lemma C.2. Subtracting $f^{*}$ and rearranging provides
+$f(x_{t+1})-f^{*}$ $\leq(1-\gamma_{t})(f(x_{t})-f^{*})+\frac{\rho\gamma_{t}}{2}\|x_{t}-x^{*}\|^{2}+L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t})$ (E.3)
+$\leq\left(1-\left(1-\frac{\rho}{\mu}\right)\gamma_{t}\right)(f(x_{t})-f^{*})+L\gamma_{t}^{1+\nu}D^{2},$ (E.4)
+where the last inequality holds due to the quadratic growth condition. For $t=0$, using Equation E.3, we have
+$f(x_{1})-f^{*}\leq\frac{\rho}{2}\|x_{0}-x^{*}\|^{2}+LD_{\phi}(v_{0},x_{0})\leq\left(\frac{\rho}{\sigma}+L\right)D^{2}\leq 2LD^{2}\leq\frac{2\mu}{\rho}LD^{2},$ (E.5)
 where the second inequality holds because $\phi$ is $\sigma$-strongly convex and the last inequality holds because of $3<6/(2-\nu)\leq 2\mu/\rho$.
 Now we consider $t\geq 1$.
-Using (D.4) and $\gamma_{t}=\frac{2}{2+t}$, we have
+Using Equation E.4 and $\gamma_{t}=\frac{2}{2+t}$, we have
 $f(x_{t+1})-f^{*}$ $\leq\left(t+\frac{2\rho}{\mu}\right)\frac{2^{1+\nu}\mu LD^{2}}{\rho(t+2)^{1+\nu}}+\frac{2^{1+\nu}LD^{2}}{(t+2)^{1+\nu}}$
@@
 $\leq\frac{2^{1+\nu}\mu LD^{2}}{\rho(t+3)^{\nu}},$
-where the last inequality holds because of $(t+3)^{\nu}(t+3\rho/\mu)\leq(t+3)^{\nu}(t+2-\nu)\leq(t+2)^{1+\nu}$ from (D.1). ∎
+where the last inequality holds because of $(t+3)^{\nu}(t+3\rho/\mu)\leq(t+3)^{\nu}(t+2-\nu)\leq(t+2)^{1+\nu}$ from Equation E.1. ∎
 We can apply the quadratic growth condition again as before and obtain
@@
 $\operatorname{dist}(x_{t},\mathcal{X}^{*})\leq\frac{2^{1+\nu/2}D\sqrt{L}}{\sqrt{\rho}(t+2)^{\nu/2}}.$
-Note that Theorem D.3 does not require knowledge of the number of iterations $T$ ahead of time compared to Theorem D.1.
-We stress, nonetheless, that the latter can also be adjusted using a different step-size strategy to obtain an any-time guarantee; see [Braun2022-ew] for details for the standard Euclidean case, which can be generalized to our setup.
-Moreover, we can apply Theorem D.3 to the Euclidean FW algorithm, i.e., $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$.
-Using (D.2), $D_{\textup{Euc}}:=\sup_{x,y\in P}\|x-y\|^{2}$, and $D_{\textup{Euc}}=\sqrt{2}D$, we obtain
+Note that Theorem E.3 does not require knowledge of the number of iterations $T$ ahead of time compared to Theorem E.1.
+We stress, nonetheless, that the latter can also be adjusted using a different step-size strategy to obtain an any-time guarantee; see Braun et al.
+(2025) for details for the standard Euclidean case, which can be generalized to our setup.
+Moreover, we can apply Theorem E.3 to the Euclidean FW algorithm, i.e., $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$.
+Using Equation E.2, $D_{\textup{Euc}}:=\sup_{x,y\in P}\|x-y\|^{2}$, and $D_{\textup{Euc}}=\sqrt{2}D$, we obtain
 $f(x_{t})-f^{*}\leq\frac{2\mu LD_{\textup{Euc}}^{2}}{\rho(t+2)},$
-Theorem D.3 requires $\rho/\sigma\leq L$ and $\frac{3\rho}{\mu}\leq 2-\nu$. These assumptions are easy to satisfy.
-Example D.4 (Example 3.1 in the arXiv version of [Liao2024-vd]).
+Theorem E.3 requires $\rho/\sigma\leq L$ and $\frac{3\rho}{\mu}\leq 2-\nu$. These assumptions are easy to satisfy.
+Example E.4 (Example 3.1 in the arXiv version of Liao et al. (2024)).
 Let us consider
@@
 It holds that $2=\rho\leq L$ and $1=\frac{3\rho}{\mu}\leq 1<2-\nu$ (set $\mu=6$).
-Therefore, the assumption of Theorem D.3 holds.
+Therefore, the assumption of Theorem E.3 holds.
 In order to verify $\frac{3\rho}{\mu}\leq 2-\nu$, it is easier to examine the sufficient condition $\frac{3\rho}{\mu}\leq 1$ instead because the exact value of $\nu$ is difficult to estimate.
@@
 Therefore, it suffices to verify $\rho/\sigma\leq\rho\leq L$, which often holds (for example, see an example of phase retrieval in Section 6.2).
-Next we establish local linear convergence with the short step step-size, i.e., $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},1\right\}$.
-D.3 Proof of Theorem 5.2
-Theorem D.5 (Local linear convergence of FW algorithm with short step-size or adaptive step-size).
-Suppose that Assumptions 3.1 and 5.1 hold.
-Let $\nu\in(0,1]$ and let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter.
-Assume that there exists a minimizer $x^{*}\in\operatorname{int}P$, i.e., there exists an $r>0$ with $B(x^{*},r)\subset P$.
-Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},1\right\}$ or Algorithm 2.
-Then, if $\rho/\mu<1$, it holds that, for all $t\geq 1$,
-$f(x_{t})-f^{*}\leq\begin{cases}\max\left\{\frac{1}{2}\left(1+\frac{\rho}{\mu}\right),1-\frac{r^{2}}{2Mc^{2}D^{2}}\right\}^{t-1}LD^{2}&\text{if}\ \nu=1,\\ \left(\frac{1}{1+\nu}\left(1+\frac{\nu\rho}{\mu}\right)\right)^{t-1}LD^{2}&\text{if}\ 1\leq t\leq t_{0},\nu\in(0,1),\\ \frac{(L(1+\nu)(1-\rho/\mu)^{\nu}c^{1+\nu}D^{2}/r^{1+\nu})^{2/(1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}\left(1-\frac{\rho}{\mu}\right)(t-t_{0})\right)^{2\nu/(1-\nu)}}=\mathcal{O}(1/t^{2\nu/(1-\nu)})&\text{if}\ t\geq t_{0},\nu\in(0,1),\end{cases}$
-where $c=\frac{\sqrt{2\mu}}{\mu-\rho}$, $M:=(L(1+\nu))^{1/\nu}$, and
-$t_{0}:=\max\left\{\left\lfloor\log_{\frac{1}{1+\nu}\left(1+\frac{\nu\rho}{\mu}\right)}\left(\frac{(L(1+\nu)(1-\rho/\mu)^{\nu}c^{1+\nu}D^{2}/r^{1+\nu})^{2/(1-\nu)}}{LD^{2}}\right)\right\rfloor+2,1\right\}.$
-Proof.
-Let $h_{t}:=f(x_{t})-f^{*}$. Using Lemma 3.4, we have
+E.3 Proof of Theorem 5.3
+Theorems 5.35.4, and E.9 require the weak convexity and the local quadratic growth condition of $f$ with $\rho\leq\mu$. We show several examples as follows.
+Example E.5.
+We show some examples of $\rho$-weakly convex functions satisfying the (local) quadratic growth condition with $\rho<\mu$.
+  1. (i)
+    Let $(x)=\log(1+x^{2}).
+ff$ attains the minimum $f^{*}=0$ at $x^{*}=0$.
+The function $f$ is $1/4$-weakly convex because $f^{\prime\prime}(x)=-\frac{2(1-x^{2})}{(1+x^{2})^{2}}$ and its minimum is 1/4-1/4.
+Moreover, $(x-x^{*})^{2}=x^{2}\leq\frac{2}{\mu}\log(1+x^{2})=\frac{2}{\mu}(f(x)-f^{*})$ holds for any $x\in[f\leq\zeta]$.
+Letting $\mu=1/3$, we have $6\log(1+x^{2})-x^{2}\geq 0$ for any $x\in[f\leq 16]$.
+Note that $\zeta=16$ is not an exact value.
+In this case, we have $1/4=\rho<\mu=1/3$.
+Through simple calculations and visualization, we can obtain a larger value of $\mu$ than $1/3$.
+This function is used for image restoration (Boţ et al., 2016; Stella et al., 2017).
+  2. (ii)
+    Let $(x)=\frac{1}{2}x^{2}+2(1-e^{-x^{2}}).
+ff$ attains the minimum $f^{*}=0$ at $x^{*}=0$.
+The function $f$ is $(8e^{-3/2}-1)$-weakly convex because $f^{\prime\prime}(x)=1+4e^{-x^{2}}(1-2x^{2})$ and its minimum is $1-8e^{-3/2}\simeq-0.785$.
+Moreover, $(x-x^{*})=x^{2}\leq\frac{2}{\mu}(f(x)-f^{*})$ holds for any $x\in[f\leq\zeta]$.
+Letting $\mu=1$, we have $\frac{2}{\mu}(f(x)-f^{*})-x^{2}=4(1-e^{-x^{2}})\geq 0$ for any $x\in\mathbb{R}$.
+The function $f$ satisfies the global quadratic growth condition with $8e^{-3/2}-1=\rho<\mu=1$.
+  3. (iii)
+    Let $f$ be the function defined in Example E.4. We set $\mu=\mu_{0}$ for some $\mu_{0}\in(2,6]$ because $\mu$ can be from $(0,6]$. It holds that $2=\rho<\mu$.
+Because Algorithm 2 is well-defined (see also Remark 3.2), the convergence result of the FW algorithm with Algorithm 2 is the same as Theorem 5.3.
+Proof of Theorem 5.3.
+Let $h_{t}:=f(x_{t})-f^{*}$. Using Lemma B.8, we have
 $h_{t}-h_{t+1}=f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\gamma_{t},$
@@
 $\geq\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}h_{t}^{\frac{1+\nu}{2\nu}},$
-where the third inequality holds due to Proposition B.8 and the definition of $D$ and the last inequality holds due to the local PL inequality from Remark B.6 with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$ and $M:=(L(1+\nu))^{1/\nu}$.
+where the third inequality holds due to Proposition C.8 and the definition of $D$ and the last inequality holds due to the local PL inequality from Remark C.6 with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$ and $M:=(L(1+\nu))^{1/\nu}$.
 (ii) In the case where $\gamma_{t}=1$, i.e., $\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}}\geq 1$, this implies
-$\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq L(1+\nu)D_{\phi}(v_{t},x_{t}),$ (D.6)
-because of $11/\nu>$ 1. Using Lemma 2.4 and (D.6), we have
+$\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq L(1+\nu)D_{\phi}(v_{t},x_{t}),$ (E.6)
+because of $1/\nu>1$. Using Lemma B.7 and Equation E.6, we have
 $h_{t+1}-h_{t}$ $\leq LD_{\phi}(v_{t},x_{t})-\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle$
@@
 $\leq-\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right)h_{t},$
-where the third inequality holds because of (3.2) in Lemma 3.3, and the last inequality holds because of the local quadratic growth condition of $f$ with $\mu>0$. Therefore, we have
+where the third inequality holds because of Equation C.1 in Lemma C.2, and the last inequality holds because of the local quadratic growth condition of $f$ with $\mu>0$. Therefore, we have
 $h_{t+1}\leq\frac{1}{1+\nu}\left(1+\frac{\rho\nu}{\mu}\right)h_{t}.$
@@
 When $\nu=1$, we have $h_{t}-h_{t+1}\geq\min\left\{\frac{1}{2}\left(1-\frac{\rho}{\mu}\right),\frac{r^{2}}{2Mc^{2}D^{2}}\right\}\cdot h_{t}$.
-This inequality and the initial bound $f(x_{1})-f^{*}\leq LD^{2}$ due to Lemma A.4 imply
+This inequality and the initial bound $f(x_{1})-f^{*}\leq LD^{2}$ due to Lemma B.4 imply
 $h_{t}\leq\max\left\{\frac{1}{2}\left(1+\frac{\rho}{\mu}\right),1-\frac{r^{2}}{2Mc^{2}D^{2}}\right\}^{t-1}LD^{2}.$
@@
 $h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right),\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}h_{t}^{\frac{1-\nu}{2\nu}}\right\}\cdot h_{t}.$
-Using $f(x_{1})-f^{*}\leq LD^{2}$ and Lemma B.11 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}$, and $\theta_{0}=\frac{1-\nu}{2\nu}>0$, we have the claim.
+Using $f(x_{1})-f^{*}\leq LD^{2}$ and Lemma C.11 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}$, and $\theta_{0}=\frac{1-\nu}{2\nu}>0$, we have the claim.
 ∎
@@
 $f(x_{t})-f^{*}\leq\max\left\{\frac{1}{2}\left(1+\frac{\rho}{\mu}\right),1-\frac{r^{2}}{2Lc^{2}D_{\textup{Euc}}^{2}}\right\}^{t-1}\frac{LD_{\textup{Euc}}^{2}}{2}.$
-D.4 Proof of Theorem 5.3
-In the same way as Theorem 4.4, the pyramidal width is the minimal $\delta>0$ satisfying Lemma B.9 (see also [Braun2022-ew, Lemma 2.26] or [Lacoste-Julien2015-gb, Theorem 3]).
+E.4 Proof of Theorem 5.4
+In the same way as Theorem 4.4, the pyramidal width is the minimal $\delta>0$ satisfying Lemma C.9 (see also (Braun et al., 2025, Lemma 2.26) or (Lacoste-Julien and Jaggi, 2015, Theorem 3)).
 An upper bound exists on the primal gap for weakly convex functions.
-Lemma D.6 (Upper bound on primal gap for weakly convex functions).
+Lemma E.6 (Upper bound on primal gap for weakly convex functions).
 Suppose that Assumptions 3.1 and 5.1 hold.
@@
 If $\rho/\mu<1$, then it holds that, for all $x\in[f\leq f^{*}+\zeta]\cap P$,
-$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}\delta^{2}}\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}.$ (D.7)
+$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}\delta^{2}}\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}.$
 Proof.
@@
 $0\leq\left(1-\frac{\rho}{\mu}\right)(f(x)-f(x^{*}))\leq\langle\nabla f(x),x-x^{*}\rangle,$
-where the first inequality follows from $01-\rho/\mu>$ 0. Using (B.8) with $\psi=\nabla f(x)$ and $y=x^{*}$ and the above inequality, we obtain
+where the first inequality follows from $1-\rho/\mu>0$. Using Equation C.8 with $\psi=\nabla f(x)$ and $y=x^{*}$ and the above inequality, we obtain
 $\frac{\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}}{\delta^{2}}$ $\geq\frac{\langle\nabla f(x),x-x^{*}\rangle^{2}}{\|x-x^{*}\|^{2}}$
@@
 where the third inequality holds because of the local quadratic growth condition. ∎
-Theorem D.7 (Local linear convergence by the away-step FW algorithm).
-Suppose that Assumptions 3.14.3, and 5.1 hold.
-Let $\nu\in(0,1]$ and let $P\subset\mathbb{R}^{n}$ be a polytope.
-The convergence rate of Algorithm 3 with $f$ is linear: if $\rho<\mu\leq L$, for all $t\geq 1$
-$f(x_{t})-f^{*}\leq\begin{cases}2\left(1-\frac{\omega}{4L}\frac{\delta^{2}}{D^{2}}\right)^{\lceil(t-1)/2\rceil}LD^{2}&\text{if}\ \nu=1,\\ 2\left(\frac{1}{1+\nu}\left(1+\frac{\nu\rho}{\mu}\right)\right)^{\lceil(t-1)/2\rceil}LD^{2}&\text{if}\ 1\leq t\leq t_{0},\nu\in(0,1),\\ \frac{(L(1+\nu)(1-\rho/\mu)^{\nu}D^{2}/(\sqrt{\omega}\delta)^{1+\nu})^{2/(1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}\left(1-\frac{\rho}{\mu}\right)\lceil(t-t_{0})/2\rceil\right)^{2\nu/(1-\nu)}}=\mathcal{O}(1/t^{2\nu/(1-\nu)})&\text{if}\ t\geq t_{0},\nu\in(0,1),\end{cases}$
-where $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ and $\delta$ are the diameter and the pyramidal width of the polytope $P$, respectively, and $\omega:=\frac{(\mu-\rho)^{2}}{8\mu}$, and
-$t_{0}:=\max\left\{\left\lfloor\log_{\frac{1}{1+\nu}\left(1+\frac{\nu\rho}{\mu}\right)}\frac{(L^{2}(1+\nu)^{2}D^{4}(1-\rho/\mu)^{2\nu}/(\omega\delta^{2})^{1+\nu})^{1/(1-\nu)}}{2LD^{2}}\right\rfloor+2,1\right\}.$
-Proof.
-By letting $h_{t}=f(x_{t})-f^{*}$ and using Lemma D.6, we have
-$h_{t}\leq\frac{2\mu}{(\mu-\rho)^{2}\delta^{2}}\langle\nabla f(x_{t}),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}\leq\frac{8\mu\langle\nabla f(x_{t}),d_{t}\rangle^{2}}{(\mu-\rho)^{2}\delta^{2}}=\frac{\langle\nabla f(x_{t}),d_{t}\rangle^{2}}{\omega\delta^{2}},$ (D.8)
+Proof of Theorem 5.4.
+By letting $h_{t}=f(x_{t})-f^{*}$ and using Lemma E.6, we have
+$h_{t}\leq\frac{2\mu}{(\mu-\rho)^{2}\delta^{2}}\langle\nabla f(x_{t}),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}\leq\frac{8\mu\langle\nabla f(x_{t}),d_{t}\rangle^{2}}{(\mu-\rho)^{2}\delta^{2}}=\frac{\langle\nabla f(x_{t}),d_{t}\rangle^{2}}{\omega\delta^{2}},$ (E.7)
 where the second inequality holds because $d_{t}$ is either $x_{t}-v_{t}^{\textup{FW}}$ or $v_{t}^{\textup{A}}-x_{t}$ with $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle/2$ in Lines 3 and 3 of Algorithm 3.
-From Lemma 3.4 with $\gamma_{t,\max}=1$ (Frank–Wolfe steps) and $\gamma_{t,\max}=\frac{\lambda_{v_{t}^{\textup{A}},t}}{1-\lambda_{v_{t}^{\textup{A}},t}}$ (away steps), we obtain
+From Lemma B.8 with $\gamma_{t,\max}=1$ (Frank–Wolfe steps) and $\gamma_{t,\max}=\frac{\lambda_{v_{t}^{\textup{A}},t}}{1-\lambda_{v_{t}^{\textup{A}},t}}$ (away steps), we obtain
 $h_{t}-h_{t+1}$ $\geq\frac{\nu}{1+\nu}\langle\nabla f(x_{t}),d_{t}\rangle\min\left\{\gamma_{t,\max},\left(\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}}\right\}$
@@
 $=\min\left\{\frac{\nu\gamma_{t,\max}}{1+\nu}\left(1-\frac{\rho}{\mu}\right)h_{t},\frac{\nu}{1+\nu}\frac{(\omega\delta^{2})^{\frac{1+\nu}{2\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}h_{t}^{\frac{1+\nu}{2\nu}}\right\},$
-where the second inequality holds because of $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq h_{t}-\frac{\rho}{2}\|x_{t}-v_{t}\|^{2}\geq\left(1-\frac{\rho}{\mu}\right)h_{t}$ (from Lemma 3.3 and the quadratic growth condition of $f$) and (D.8).
+where the second inequality holds because of $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq h_{t}-\frac{\rho}{2}\|x_{t}-v_{t}\|^{2}\geq\left(1-\frac{\rho}{\mu}\right)h_{t}$ (from Lemma C.2 and the quadratic growth condition of $f$) and Equation E.7.
 In $\nu=1$, for Frank–Wolfe steps, we have $\gamma_{t,\max}\left(1-\frac{\rho}{\mu}\right)=\left(1-\frac{\rho}{\mu}\right)\geq\frac{1}{8}\left(1-\frac{\rho}{\mu}\right)^{2}\frac{\delta^{2}}{D^{2}}=\frac{\omega\delta^{2}}{\mu D^{2}}\geq\frac{\omega\delta^{2}}{2LD^{2}}$ by $0<\rho<\mu\leq L$ and $\delta\leq D$.
@@
 Therefore, at most half of iterations until $t$ iterations are in away steps, i.e., in all other steps, we have $\gamma_{t}=\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{2LD_{\phi}(v_{t},x_{t})}<\gamma_{t,\max}$ and then $h_{t+1}\leq\left(1-\frac{\omega\delta^{2}}{4LD^{2}}\right)h_{t}$.
-Because we have $h_{t+1}\leq\left(1-\frac{\omega\delta^{2}}{4LD^{2}}\right)h_{t}$ for at least half of the iterations and $h_{t+1}\leq h_{t}$ for the rest, using $h_{1}\leq 2LD^{2}$ from (D.5) and $\rho<L$, we obtain
+Because we have $h_{t+1}\leq\left(1-\frac{\omega\delta^{2}}{4LD^{2}}\right)h_{t}$ for at least half of the iterations and $h_{t+1}\leq h_{t}$ for the rest, using $h_{1}\leq 2LD^{2}$ from Equation E.5 and $\rho<L$, we obtain
 $f(x_{t})-f^{*}\leq 2\left(1-\frac{\omega}{4L}\frac{\delta^{2}}{D^{2}}\right)^{\lceil(t-1)/2\rceil}LD^{2}.$
 In $\nu\in(0,1)$, we have $h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right),\frac{\nu}{1+\nu}\frac{(\omega\delta^{2})^{\frac{1+\nu}{2\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}h_{t}^{\frac{1-\nu}{2\nu}}\right\}\cdot h_{t}$ for at least half of the iterations and $h_{t+1}\leq h_{t}$ for the rest.
-The initial bound $f(x_{1})-f^{*}\leq(\rho+L)D^{2}\leq 2LD^{2}$ holds from (D.5) and $\rho<L$.
-We use Lemma B.11 with $c_{0}=2LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=\frac{\nu}{1+\nu}\frac{(\omega\delta^{2})^{\frac{1+\nu}{2\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}$, and $\theta_{0}=\frac{1-\nu}{2\nu}$ and obtain the claim.
+The initial bound $f(x_{1})-f^{*}\leq(\rho+L)D^{2}\leq 2LD^{2}$ holds from Equation E.5 and $\rho<L$.
+We use Lemma C.11 with $c_{0}=2LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=\frac{\nu}{1+\nu}\frac{(\omega\delta^{2})^{\frac{1+\nu}{2\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}$, and $\theta_{0}=\frac{1-\nu}{2\nu}$ and obtain the claim.
 ∎
@@
 $f(x_{t})-f(x^{*})\leq\left(1-\frac{\omega}{2L}\frac{\delta^{2}}{D_{\textup{Euc}}^{2}}\right)^{\lceil(t-1)/2\rceil}LD_{\textup{Euc}}^{2}.$
-D.5 Local Linear Convergence over Uniformly Convex Sets
-In the convex optimization case, Canon and Cullum [Canon1968-sr] established an early lower bound on the convergence rate of the FW algorithm.
-However, in the special case of $P$ being strongly convex, Garber and Hazan [Garber2015-de] showed that one can improve upon that lower bound.
+E.5 Local Linear Convergence over Uniformly Convex Sets
+In the convex optimization case, Canon and Cullum (1968) established an early lower bound on the convergence rate of the FW algorithm.
+However, in the special case of $P$ being strongly convex, Garber and Hazan (2015) showed that one can improve upon that lower bound.
 Kerdreux et al.
-[Kerdreux2021-kc] establish it in the case of $P$ being uniformly convex.
+(2021b) establish it in the case of $P$ being uniformly convex.
 We will now carry over this result to establish local linear convergence for the case where $f$ is weakly convex and $L$-smad on $P$.
 To this end, we recall the definition of uniformly convex sets.
-Definition D.8 ($(\alpha,p)$-uniformly convex set [Braun2022-ew, Definition 2.18][Kerdreux2021-kc, Definition 1.1]).
-[…]
+Definition E.7 ($(\alpha,p)$-uniformly convex set (Braun et al., 2025, Definition 2.18)(Kerdreux et al., 2021b, Definition 1.1)).
+Let $\alpha$ and $p$ be positive numbers.
+The set $P\subset\mathbb{R}^{n}$ is $(\alpha,p)$-uniformly convex with respect to the norm $\|\cdot\|$ if for any $x,y\in P$, any $\gamma\in[0,1]$, and any $z\in\mathbb{R}^{n}$ with $\|z\|\leq 1$ the following holds:
 $y+\gamma(x-y)+\gamma(1-\gamma)\cdot\alpha\|x-y\|^{p}z\in P.$
@@
 We will use the scaling condition for uniformly convex sets to establish linear convergence.
-Proposition D.9 (Scaling inequality [Braun2022-ew, Proposition 2.19][Kerdreux2021-kc, Lemma 2.1]).
+Proposition E.8 (Scaling inequality (Braun et al., 2025, Proposition 2.19)(Kerdreux et al., 2021b, Lemma 2.1)).
 Let $P$ be a full dimensional compact $(\alpha,p)$-uniformly convex set, $u$ any non-zero vector, and $v=\operatorname*{arg\,min}_{y\in P}\langle u,y\rangle$. Then for all $x\in P$
@@
 Now we establish local linear convergence for weakly convex optimization on uniformly convex sets.
-Theorem D.10 (Local linear convergence over uniformly convex sets).
+Theorem E.9 (Local linear convergence over uniformly convex sets).
 Suppose that Assumptions 3.1 and 5.1 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$ hold and that $P$ is $(\alpha,p)$-uniformly convex set.
@@
 Let $g_{t}:=\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle$ and $h_{t}:=f(x_{t})-f^{*}$.
-Lemma 3.4 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$ is followed by $f(x_{t})-f(x_{t+1})\geq\frac{g_{t}}{2}\gamma_{t}$.
-Using Proposition D.9, we have
+Lemma B.8 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$ is followed by $f(x_{t})-f(x_{t+1})\geq\frac{g_{t}}{2}\gamma_{t}$.
+Using Proposition E.8, we have
 $h_{t}-h_{t+1}$ $\geq\frac{g_{t}}{2}\min\left\{\frac{g_{t}}{L\|x_{t}-v_{t}\|^{2}},1\right\}$
@@
 $\geq\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)\min\left\{\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c)^{2/p}}{L}h_{t}^{1-2/p},1\right\}\cdot h_{t},$
-where the third inequality holds from Lemma 3.3 and $\|\nabla f(x)\|\geq c$, and the last inequality holds because of the local quadratic growth property of $f$.
-The initial bound $h_{1}\leq LD_{\textup{Euc}}^{2}=2LD^{2}$ holds from $\rho\leq L$ and (D.5).
+where the third inequality holds from Lemma C.2 and $\|\nabla f(x)\|\geq c$, and the last inequality holds because of the local quadratic growth property of $f$.
+The initial bound $h_{1}\leq LD_{\textup{Euc}}^{2}=2LD^{2}$ holds from $\rho\leq L$ and Equation E.5.
 For $q=2$, we have $h_{t}-h_{t+1}\geq\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)\min\left\{\frac{\alpha c}{L},1\right\}\cdot h_{t}$, which implies
@@
 Thus, we have the claim.
-For $p>2$, we use Lemma B.11 with $c_{0}=LD_{\textup{Euc}}^{2}$, $c_{1}=\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=c_{1}\cdot\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c)^{2/p}}{L}$, and $\theta_{0}=1-2/p$ and obtain the claim.
+For $p>2$, we use Lemma C.11 with $c_{0}=LD_{\textup{Euc}}^{2}$, $c_{1}=\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=c_{1}\cdot\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c)^{2/p}}{L}$, and $\theta_{0}=1-2/p$ and obtain the claim.
 ∎
 Note that Assumption 3.1 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$ holds when $f$ is $L$-smooth over $P$.
-If $\rho\leq L$ does not hold, we can use the initial bound $h_{1}\leq\frac{\rho+L}{2}D_{\textup{Euc}}^{2}$ from (D.5).
-In that case, a local linear rate in Theorem D.10 is unchanged.
+If $\rho\leq L$ does not hold, we can use the initial bound $h_{1}\leq\frac{\rho+L}{2}D_{\textup{Euc}}^{2}$ from Equation E.5.
+In that case, a local linear rate in Theorem E.9 is unchanged.
 Moreover, we have local linear convergence without assuming $\|\nabla f(x)\|>0$.
-Theorem D.11 (Local linear convergence over uniformly convex sets without $\|\nabla f(x)\|>0$).
+Theorem E.10 (Local linear convergence over uniformly convex sets without $\|\nabla f(x)\|>0$).
 Suppose that Assumptions 3.1 and 5.1 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$ hold and that $P$ is $(\alpha,p)$-uniformly convex set.
@@
 Proof.
-Using an argument similar to Theorem D.10, we obtain
+Using an argument similar to Theorem E.9, we obtain
 $h_{t}-h_{t+1}$ $\geq\frac{g_{t}}{2}\min\left\{\frac{g_{t}}{L\|x_{t}-v_{t}\|^{2}},1\right\}$
@@
 $\geq\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)\min\left\{\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c^{-1})^{2/p}}{L}h_{t}^{1-1/p},1\right\}\cdot h_{t},$
-where the third inequality holds from the PL inequality (B.6) with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$.
-Therefore, we use Lemma B.11 with $c_{0}=LD_{\textup{Euc}}^{2}$, $c_{1}=\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=c_{1}\cdot\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c^{-1})^{2/p}}{L}$, and $\theta_{0}=1-1/p$ and obtain the claim.
+where the third inequality holds from the PL inequality Equation C.6 with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$.
+Therefore, we use Lemma C.11 with $c_{0}=LD_{\textup{Euc}}^{2}$, $c_{1}=\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=c_{1}\cdot\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c^{-1})^{2/p}}{L}$, and $\theta_{0}=1-1/p$ and obtain the claim.
 ∎
-Appendix E Experiments
+Appendix F Experiments
 We use the following notation:
@@
   • BregAFW: the away-step FW algorithm with the adaptive Bregman step-size strategy (Algorithm 3 using Algorithm 2, i.e., $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},\gamma_{t,\max})$, our proposed update)
-  • EucFW: the FW algorithm with the adaptive (Euclidean) step-size strategy [Pedregosa2020-ay]
-  • EucAFW: the away-step FW algorithm with the adaptive (Euclidean) step-size strategy [Pedregosa2020-ay]
+  • EucFW: the FW algorithm with the adaptive (Euclidean) step-size strategy (Pedregosa et al., 2020)
+  • EucAFW: the away-step FW algorithm with the adaptive (Euclidean) step-size strategy (Pedregosa et al., 2020)
   • ShortFW: the FW algorithm with the (Euclidean) short step
@@
   • OpenAFW: the away-step FW algorithm with the open loop with $\gamma_{t}=\frac{2}{2+t}$
-  • MD: the mirror descent [Nemirovskij1983-sm]
-E.1 Nonnegative Linear Inverse Problems
+  • MD: the mirror descent (Nemirovskij and Yudin, 1983)
+  • ProjGD: the projected gradient descent algorithm (see, e.g., Beck (2017))
+F.1 Nonnegative Linear Inverse Problems
 Given a nonnegative matrix $A\in\mathbb{R}_{+}^{m\times n}$ and a nonnegative vector $b\in\mathbb{R}_{+}^{m}$, the goal of nonnegative (Poisson) linear inverse problems is to recover a signal $x\in\mathbb{R}_{+}^{n}$ such that $Ax\simeq b$.
-This class of problems has been studied in image deblurring [Bertero2009-jq] and positron emission tomography [Vardi1985-de] as well as optimization [Bauschke2017-hg, Takahashi2024-wz].
+This class of problems has been studied in image deblurring (Bertero et al., 2009) and positron emission tomography (Vardi et al., 1985) as well as optimization (Bauschke et al., 2017; Takahashi and Takeda, 2025).
 Since the dimension of $x$ is often larger than the number of observations $m$, the system is indeterminate.
@@
 Recovering $x$ can be formulated as a minimization problem:
-$\min_{x\in\Delta_{n}}\quad f(x):=d(Ax,b),$ (E.1)
-where $d(x,y):=\sum_{i=1}^{m}\left(x_{i}\log\frac{x_{i}}{y_{i}}+y_{i}-x_{i}\right)$ is the Kullback–Leibler divergence.
-Problem (E.1) is convex, while $\nabla f$ is not Lipschitz continuous on $\mathbb{R}_{+}^{n}$.
-The pair $(f,\phi)$ is $L$-smad on $\mathbb{R}_{+}^{n}$ with $\phi(x)=\sum_{j=1}^{n}x_{j}\log x_{j}$ and $L\geq\max_{1\leq j\leq n}\sum_{i=1}^{m}a_{ij}$ from [Bauschke2017-hg, Lemma 8].
-We compared BregFW with EucFW, ShortFW, OpenFW, and the mirror descent algorithm (MD[Nemirovskij1983-sm].
-The subproblem of MD can be solved in closed-form for $\{x\in\mathbb{R}_{+}^{n}\,|\,\sum_{j=1}^{n}x_{j}=1\}$ by [Beck2017-qc, Example 3.71], and this can be readily extended to $\Delta_{n}$.
+$\min_{x\in\Delta_{n}}\quad f(x):=d(Ax,b),$ (F.1)
+where $d(x,y):=\sum_{i=1}^{m}\left(x_{i}\log\frac{x_{i}}{y_{i}}+y_{i}-x_{i}\right)$ is the KL divergence.
+Problem Equation F.1 is convex, while $\nabla f$ is not Lipschitz continuous on $\mathbb{R}_{+}^{n}$.
+The pair $(f,\phi)$ is $L$-smad on $\mathbb{R}_{+}^{n}$ with $\phi(x)=\sum_{j=1}^{n}x_{j}\log x_{j}$ and $L\geq\max_{1\leq j\leq n}\sum_{i=1}^{m}a_{ij}$ from (Bauschke et al., 2017, Lemma 8).
+We compared BregFW with EucFW, ShortFW, OpenFW, and the mirror descent algorithm (MD(Nemirovskij and Yudin, 1983).
+The subproblem of MD can be solved in closed-form for $\{x\in\mathbb{R}_{+}^{n}\,|\,\sum_{j=1}^{n}x_{j}=1\}$ by (Beck, 2017, Example 3.71), and this can be readily extended to $\Delta_{n}$.
 We used $1000$ as maximum iteration limit and we generated $\tilde{A}$ from an i.i.d.
@@
 For $(m,n)=(100,1000)$, Figure 5 shows the primal gap $f(x_{t})-f^{*}$ and the FW gap $\max_{v\in P}\langle\nabla f(x_{t}),x_{t}-v\rangle$ per iteration (left) and the primal gap per second (right).
-Table 2 shows average numbers of the primal gap, the FW gap, and computational time over 20 different instances for $(m,n)=(100,1000)$.
+Table 2 shows the average values of the primal gap, the FW gap, and computational time over 20 different instances for $(m,n)=(100,1000)$.
 Here, BregFW outperformed other algorithms, both in iterations and time.
-Figure 5: Nonnegative linear inverse problem for $(m,n)=(100,1000)$.
-Table 2: Average numbers of the primal gap, the FW gap, and computational time (s) over 20 different instances of nonnegative linear inverse problems for $(m,n)=(100,1000)$.
+Figure 5: Log-log plot of primal and FW gaps on nonnegative linear inverse problem for $(m,n)=(100,1000)$.
+Table 2: Average values of the primal gap, the FW gap, and computational time (s) over 20 different instances of nonnegative linear inverse problems for $(m,n)=(100,1000)$.
 The results are reported as mean $\pm$ standard deviation.
@@
 MD 2.2494e-06 $\pm$ 3.1e-07 1.7214e-04 $\pm$ 1.9e-05 1.8885e-01 $\pm$ 6.4e-03
-E.2 $\ell_{p}$ Loss Problem
-Figure 6: The $\ell_{p}$ loss problem for $(m,n)=(1000,100)$.
-Table 3: Average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of $\ell_{p}$ loss problems for $(m,n)=(1000,100)$.
+F.2 $\ell_{p}$ Loss Problem
+Figure 6: Log-log plot of primal and FW gaps on the $\ell_{p}$ loss problem for $(m,n)=(1000,100)$.
+Table 3: Average values of the primal gap, FW gap, and computational time (s) over 20 different instances of $\ell_{p}$ loss problems for $(m,n)=(1000,100)$.
 The results are reported as mean $\pm$ standard deviation.
@@
 EucFW 6.2980e-10 $\pm$ 9.9e-11 8.5940e-01 $\pm$ 1.0e-01 3.1421e-01 $\pm$ 7.4e-03
-ShortFW 2.4697e+01 $\pm$ 1.2e+00 4.7621e+01 $\pm$ 8.6e-01 9.2352e-02 $\pm$ 3.7e-03
-OpenFW 1.6961e-02 $\pm$ 9.9e-04 4.$0$$\pm$ 1.0e-01 8.7292e-02 $\pm$ 3.5e-03
-We use an $\ell_{2}$ norm ball as a constraint, i.e., $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|_{2}\leq 1\}$.
+ShortFW 2.4697e+01 $\pm$ 1.2e+$0$ 4.7621e+01 $\pm$ 8.6e-01 9.2352e-02 $\pm$ 3.7e-03
+OpenFW 1.6961e-02 $\pm$ 9.9e-04 4.0077e+$0$ $\pm$ 1.0e-01 8.7292e-02 $\pm$ 3.5e-03
+We use an $\ell_{2}$ norm ball as a constraint, i.e., $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|\leq 1\}$.
 We compared BregFW with EucFW, ShortFW, and OpenFW.
@@
 BregFW outperformed the other algorithms.
-Since $\nabla f$ is not Lipschitz continuous, ShortFW did not reduce the primal and FW gap.
-E.3 Phase Retrieval
+Since $\nabla f$ is not Lipschitz continuous, ShortFW did not reduce the primal and FW gaps.
+F.3 Phase Retrieval
 We show the average performance for the setting of Section 6.2. Tables 4 and 5 show the average performance over 20 different instances of $(m,n)=(1000,10000)$ and $(m,n)=(2000,10000)$, respectively.
@@
 The primal gap by BregAFW is the smallest among these algorithms, while ShortAFW has the smallest value of the FW gap.
-Figure 7: Phase retrieval for $(m,n)=(200,200)$ and $K=110$ via AFW.
-Table 4: Average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of phase retrieval for $(m,n)=(1000,10000)$ and $K=2000$.
+Figure 7: Log-log plot of primal and FW gaps on phase retrieval for $(m,n)=(200,200)$ and $K=110$ via AFW.
+Table 4: Average values of the primal gap, FW gap, and computational time (s) over 20 different instances of phase retrieval for $(m,n)=(1000,10000)$ and $K=2000$.
 The results are reported as mean $\pm$ standard deviation.
 algorithm primal gap FW gap time
-BregFW $-1$$\pm$ 1.1e-12 2.1518e-02 $\pm$ 5.9e-02 4.6990e+01 $\pm$ 1.5e+01
-EucFW 1.7170e+00 $\pm$ 3.9e+00 2.4612e+02 $\pm$ 4.5e+02 4.4986e+01 $\pm$ 3.2e+01
+BregFW 1.3107e-11 $\pm$ 1.1e-12 2.1518e-02 $\pm$ 5.9e-02 4.6990e+01 $\pm$ 1.5e+01
+EucFW 1.7170e+$0$ $\pm$ 3.9e+$0$ 2.4612e+02 $\pm$ 4.5e+02 4.4986e+01 $\pm$ 3.2e+01
 ShortFW 3.2739e+05 $\pm$ 1.9e+04 4.2232e+04 $\pm$ 5.0e+02 1.0942e+01 $\pm$ 3.2e-01
-OpenFW 5.2496e-10 $\pm$ 3.9e-11 4.3802e+00 $\pm$ 5.5e-02 6.0353e+01 $\pm$ 1.6e+00
-Table 5: Average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of phase retrieval for $(m,n)=(2000,10000)$ and $K=2000$.
+OpenFW 5.2496e-10 $\pm$ 3.9e-11 4.3802e+$0$ $\pm$ 5.5e-02 6.0353e+01 $\pm$ 1.6e+$0$
+Table 5: Average values of the primal gap, FW gap, and computational time (s) over 20 different instances of phase retrieval for $(m,n)=(2000,10000)$ and $K=2000$.
 The results are reported as mean $\pm$ standard deviation.
 algorithm primal gap FW gap time
-BregFW 6.6743e-12 $\pm$ 5.5e-13 9.5656e-08 $\pm$ 7.9e-09 7.2592e+01 $\pm$ 8.8e+00
-EucFW 3.2499e+00 $\pm$ 7.1e+00 3.1304e+02 $\pm$ 4.8e+02 1.4348e+02 $\pm$ 4.3e+01
+BregFW 6.6743e-12 $\pm$ 5.5e-13 9.5656e-08 $\pm$ 7.9e-09 7.2592e+01 $\pm$ 8.8e+$0$
+EucFW 3.2499e+$0$ $\pm$ 7.1e+$0$ 3.1304e+02 $\pm$ 4.8e+02 1.4348e+02 $\pm$ 4.3e+01
 ShortFW 6.0937e+04 $\pm$ 2.3e+03 2.5461e+04 $\pm$ 2.3e+02 1.2648e+01 $\pm$ 2.4e-01
-OpenFW 9.3513e-11 $\pm$ 4.9e-12 2.6628e+00 $\pm$ 1.7e-02 7.1916e+01 $\pm$ 1.3e+00
-Table 6: Average numbers of the primal gap, FW gap, and computational time (sec) over 20 different instances of phase retrieval for $(m,n)=(200,200)$ and $K=110$ via away-step FW algorithms.
+OpenFW 9.3513e-11 $\pm$ 4.9e-12 2.6628e+$0$ $\pm$ 1.7e-02 7.1916e+01 $\pm$ 1.3e+$0$
+Table 6: Average values of the primal gap, FW gap, and computational time (s) over 20 different instances of phase retrieval for $(m,n)=(200,200)$ and $K=110$ via away-step FW algorithms.
 The results are reported as mean $\pm$ standard deviation.
 algorithm primal gap FW gap time (s)
-BregAFW 4.3430e+00 $\pm$ 1.7e+00 9.7578e-01 $\pm$ 9.1e-01 2.3336e-01 $\pm$ 6.4e-02
-EucAFW 4.3968e+00 $\pm$ 1.7e+00 1.2212e+00 $\pm$ 2.3e+00 3.9421e-01 $\pm$ 1.6e-01
-ShortAFW 4.4092e+00 $\pm$ 1.7e+00 2.5978e-02 $\pm$ 7.3e-03 7.3540e-02 $\pm$ 3.5e-02
-OpenAFW 4.4094e+00 $\pm$ 1.7e+00 7.1992e-02 $\pm$ 4.9e-02 5.8998e-02 $\pm$ 2.4e-03
-E.4 Low-Rank Minimization
+BregAFW 4.3430e+$0$ $\pm$ 1.7e+$0$ 9.7578e-01 $\pm$ 9.1e-01 2.3336e-01 $\pm$ 6.4e-02
+EucAFW 4.3968e+$0$ $\pm$ 1.7e+$0$ 1.2212e+$0$ $\pm$ 2.3e+$0$ 3.9421e-01 $\pm$ 1.6e-01
+ShortAFW 4.4092e+$0$ $\pm$ 1.7e+$0$ 2.5978e-02 $\pm$ 7.3e-03 7.3540e-02 $\pm$ 3.5e-02
+OpenAFW 4.4094e+$0$ $\pm$ 1.7e+$0$ 7.1992e-02 $\pm$ 4.9e-02 5.8998e-02 $\pm$ 2.4e-03
+F.4 Low-Rank Minimization
 Given a symmetric matrix $M\in\mathbb{R}^{n\times n}$, our goal is to find $X\in\mathbb{R}^{n\times r}$ such that $M\simeq XX^{\mathsf{T}}$. This is accomplished by minimizing the function
@@
 We assume that $r\leq n$.
-This problem is known as low-rank minimization [Dragomir2021-rv].
-In this paper, we define $P=\{X\in\mathbb{R}^{n\times r}\,|\,\|X\|_{*}\leq\xi\}$, where $\|\cdot\|_{*}$ denotes the nuclear norm and $r\in\mathbb{R}_{+}$ for the low-rank assumption.
+This problem is known as low-rank minimization (Dragomir et al., 2021).
+In this paper, we define $P=\{X\in\mathbb{R}^{n\times r}\,|\,\|X\|_{*}\leq\xi\}$, where $\|\cdot\|_{*}$ denotes the nuclear norm and $\xi\in\mathbb{R}_{+}$ for the low-rank assumption.
 We define
 $\phi(X)=\frac{1}{4}\|X\|_{F}^{4}+\frac{1}{2}\|X\|_{F}^{2}.$
-There exists a constant $L$ such that the pair $(f,\phi)$ is $L$-smad on $\mathbb{R}^{n}$ [Dragomir2021-rv].
-Additionally, $f(X)$ is weakly convex on any compact set due to Proposition B.1, which follows from the twice continuous differentiability of $f$.
+There exists a constant $L$ such that the pair $(f,\phi)$ is $L$-smad on $\mathbb{R}^{n}$ (Dragomir et al., 2021).
+Additionally, $f(X)$ is weakly convex on any compact set due to Proposition C.1, which follows from the twice continuous differentiability of $f$.
 We also compared BregFW with EucFW, ShortFW, and OpenFW.
@@
 Table 7 presents the average performance over 20 different instances for $(n,r)=(1000,20)$.
-BregFW performs slightly better than EucFW.
+BregFW performed slightly better than EucFW.
 OpenFW also performed as fast as BregFW and EucFW, but its performance was unstable.
 ShortFW did not converge due to the lack of Lipschitz continuity of $\nabla f$.
-Figure 8: Low-rank minimization for $(n,r)=(1000,20)$.
-Table 7: average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of low-rank minimization for $(n,r)=(1000,20)$.
+Figure 8: Log-log plot of primal and FW gaps on low-rank minimization for $(n,r)=(1000,20)$.
+Table 7: Average values of the primal gap, FW gap, and computational time (s) over 20 different instances of low-rank minimization for $(n,r)=(1000,20)$.
 The results are reported as mean $\pm$ standard deviation.
 algorithm primal gap FW gap time (s)
-BregFW 5.6519e-01 $\pm$ 4.2e-02 1.4727e+01 $\pm$ 1.0e+00 1.2602e+01 $\pm$ 4.5e-01
-EucFW 5.7765e-01 $\pm$ 4.6e-02 1.7237e+01 $\pm$ 2.0e+00 1.2280e+01 $\pm$ 3.8e-01
-ShortFW 2.7101e+08 $\pm$ 1.9e+06 2.1694e+09 $\pm$ 1.5e+07 4.1379e+00 $\pm$ 7.9e-02
-OpenFW 1.9782e+00 $\pm$ 3.6e-01 2.1212e+03 $\pm$ 3.1e+02 4.1770e+00 $\pm$ 1.7e-01
-E.5 Nonnegative Matrix Factorization
+BregFW 5.6519e-01 $\pm$ 4.2e-02 1.4727e+01 $\pm$ 1.0e+$0$ 1.2602e+01 $\pm$ 4.5e-01
+EucFW 5.7765e-01 $\pm$ 4.6e-02 1.7237e+01 $\pm$ 2.0e+$0$ 1.2280e+01 $\pm$ 3.8e-01
+ShortFW 2.7101e+08 $\pm$ 1.9e+06 2.1694e+09 $\pm$ 1.5e+07 4.1379e+$0$ $\pm$ 7.9e-02
+OpenFW 1.9782e+$0$ $\pm$ 3.6e-01 2.1212e+03 $\pm$ 3.1e+02 4.1770e+$0$ $\pm$ 1.7e-01
+F.5 Nonnegative Matrix Factorization
 Given a nonnegative matrix $V\in\mathbb{R}_{+}^{m\times n}$, nonnegative matrix factorization (NMF) aims to find nonnegative matrices $W\in\mathbb{R}_{+}^{m\times r}$ and $H\in\mathbb{R}_{+}^{r\times n}$ such that $V\simeq WH$.
 NMF can be formulated as a minimization problem of the loss function that measures the difference between $V$ and $WH$, i.e., $\min_{(W,H)\in P}f(W,H):=\frac{1}{2}\|WH-V\|_{F}^{2}$, where $P$ is a compact convex subset of $\mathbb{R}_{+}^{m\times r}\times\mathbb{R}_{+}^{r\times n}$.
-The objective function $f$ is weakly convex over $P$ due to Proposition B.1.
-The gradient $\nabla f$ is not Lipschitz continuous, while $(f,\phi)$ is smooth adaptable [Mukkamala2019-mk] with $\phi(W,H)=\frac{1}{4}(\|W\|_{F}^{2}+\|H\|_{F}^{2})^{2}+\frac{1}{2}(\|W\|_{F}^{2}+\|H\|_{F}^{2})$.
-[…]
-[…]
+The objective function $f$ is weakly convex over $P$ due to Proposition C.1.
+The gradient $\nabla f$ is not Lipschitz continuous, while $(f,\phi)$ is smooth adaptable (Mukkamala and Ochs, 2019) with $\phi(W,H)=\frac{1}{4}(\|W\|_{F}^{2}+\|H\|_{F}^{2})^{2}+\frac{1}{2}(\|W\|_{F}^{2}+\|H\|_{F}^{2})$.
+We used a box constraint $P=\{(W,H)\in\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}\,|\,0\leq W_{lj}\leq 3,0\leq H_{lj}\leq 1\}$.
+We compared BregFW with EucFW because ShortFW and OpenFW stopped at the 2nd iteration.
+We generated $W^{*}$ from an i.i.d.
+uniform distribution in $[0,1]$ and normalized each column of $W^{*}$ (random seed 42).
+We also generated $H^{*}$ from an i.i.d.
+Dirichlet distribution.
+The initial point $(W_{0},H_{0})$ was generated from an i.i.d.
+uniform distribution in $[0,1]$.
+Figure 10 shows the primal and FW gaps for $(m,n,r)=(100,5000,20)$ up to the 5000th iteration.
+Table 8 shows the average performance over 20 different instances for $(m,n,r)=(100,5000,20)$ up to the 5000th iteration.
+BregFW is slightly better than EucFW in terms of the primal gap, while the FW gap for EucFW is smaller than that for BregFW.
+Furthermore, we consider real-world data using the MovieLens 100K Dataset.
+We set $P=\{(W,H)\in\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}\,|\,0\leq W_{lj}\leq 5.0,\|H\|_{*}\leq\xi\}$.
+For $\phi(W,H)=\frac{3}{4}(\|W\|_{F}^{2}+\|H\|_{F}^{2})^{2}+\frac{\|V\|_{F}}{2}(\|W\|_{F}^{2}+\|H\|_{F}^{2})$ by (Mukkamala and Ochs, 2019, Proposition 2.1), $(f,\phi)$ is also $L$-smooth adaptable.
+Figure 10 shows the primal and FW gaps up to the 1000th iteration with $\xi=10\sqrt{\lambda_{\max}(VV^{\mathsf{T}})}$, where $\lambda_{\max}(\cdot)$ is the largest eigenvalue.
+The primal gap of BregFW is 6.041269e+05, and that of EucFW is 6.041272e+05.
+In this setting, the primal and FW gaps of BregFW are slightly better than those of EucFW.
 Figure 9: Log-log plot of the primal and the FW gaps on NMF for $(m,n,r)=(100,5000,20)$.
 Figure 10: Log-log plot of the primal and the FW gaps on NMF for MovieLens 100K Dataset.
-Table 8: average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of NMF for $(m,n,r)=(100,5000,20)$.
-The results are reported as mean $\pm$ standard deviation.
+Table 8: Average values of the primal gap, FW gap, and computational time (s) over 20 different instances of NMF for $(m,n,r)=(100,5000,20)$. The results are reported as mean $\pm$ standard deviation.
 algorithm primal gap FW gap time (s)
-BregFW 1.2017e-04 $\pm$ 5.7e-06 1.1515e-02 $\pm$ 1.4e-02 5.0983e+01 $\pm$ 1.1e+00
-EucFW 1.2819e-04 $\pm$ 4.6e-06 7.6259e-04 $\pm$ 8.0e-04 4.3592e+01 $\pm$ 1.0e+00
+BregFW 1.2017e-04 $\pm$ 5.7e-06 1.1515e-02 $\pm$ 1.4e-02 5.0983e+01 $\pm$ 1.1e+$0$
+EucFW 1.2819e-04 $\pm$ 4.6e-06 7.6259e-04 $\pm$ 8.0e-04 4.3592e+01 $\pm$ 1.0e+$0$